A Selected Bibliography of Publications by, and about, J. Robert Oppenheimer

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

05 November 2025 Version 1.62

Title word cross-reference

$1 [Duf46]. $12.95 [Edg91]. $13.50 [Tho03]. $14.00 [Hug07]. $15.95 [Hen81]. $16.00 [RS06]. $16.95 [RS06]. $17.50 [Hen81]. $2.50 [Opp28g]. $20.00 [Hen81, Jor80]. $24.95 [Fra01]. $25.00 [Ger06]. $26.95 [Wol05]. $27.95 [Ger06]. $29.95 [Goo09]. $30.00 [Kev03, Kle07]. $32.50 [Edg91]. $35 [Wol05]. $35.00 [Bed06]. $37.50 [Hug09, Pol07, Dys13]. $39.50 [Edg91]. $39.95 [Bad95]. $8.95 [Edg91]. α [Opp27a, Rut27]. γ [LO34].

-particles [Opp27a]. -rays [Rut27]. -Teilchen [Opp27a].

0-226-79845-3 [Guy07, Hug09]. 0-8014-8661-0 [Tho03]. 0-8047-1713-3 [Edg91]. 0-8047-1714-1 [Edg91]. 0-8047-1721-4 [Edg91]. 0-8047-1722-2 [Edg91]. 0-9672617-3-2 [Bro06, Hug07].

1 [Opp57f]. 109 [Con05, Mur05, Nas07, Sap05a, Wol05, Kru07]. 112 [FW07]. 14.99/$25.00 [Ber04a]. 16 [GHK^{+}96]. 1911 [Meh75]. 1945 [GHK^{+}96, Gow81, Haw61, Bad95, Gol95a, Hew66, See95, She82, HBP94]. 1945-47 [Hew66]. 1950 [Ano50]. 1954 [Ano01b, GM54, SZC54]. 1960s [Sch08a]. 1963 [Kuh63]. 1967 [Bet67a, Bet97, Pun67, RB67]. 1976 [Sag79a, Sag79b]. 1981 [Ano81].

20 [Goe88]. 2005 [Dre07]. 20th [Opp65a, Anoxx, Kai02]. 235U [HAM^{+}21]. 239Pu [HAM^{+}21]. 240pp [Ber04a]. 29 [SZC54].

3 [Opp57h]. 30.00 [Und13].

411-415 [Ber03b]. 45th [Wal96]. 47 [Hew66]. 480pp [Ber04a].

50th [Ano81, Ano93a, Ano93b, Ano93c, Ano93d].

6 [Dom13].

71 [Ber03b, Her03].

8 [Opp57d].

9 [Tho11]. 9.95 [Yox82]. 978 [Dom13, Tho11]. 978-0-224-06262-6 [Und13, Dom13]. 978-0-226-79845-5 [Pol07]. 978-0-674-02828-9 [MW09, Tho11].

= [Kip87].

A&E [Wal96]. A-Bomb [Ano54s, Eas89]. A-Team [Eas89]. A.E.C. [Ano54x, Ano67a, Gin54, Opp54i, Res54]. AA [Opp47e]. Aar [Lau47c]. Abbildungen [Hof96]. Abhandlungen [Bor63a]. Abraham [Wol70, Kle07]. Absorption [HO31b, Opp35a, RAO^{+}49, HO31a, Opp26c]. Absorptionsspektrums [Opp26c]. abstract [CO32a]. abuse [Gre77, Gre23]. Academic [Ano93b]. Accelerator [DOW84]. accelerators [Opp59l]. acceptance [Dir71]. Access [Res54]. Account [Moo21, Nic87]. accuse [AA54a, AA54b]. action [Opp60d]. acts [AS05, BK68]. Adapted [Kip67b, Kip68b]. added [Opp63a, Opp63b]. administrator [Dys06a]. advancement [Hol86, Hol98]. Advances [SFAL21]. adventure [LC68]. adventures [Alv87]. Advisors [Bad77, Dys76, Edg91, Hig76, San77, Wig76, Zac76, Zie91, Yor76, Yor89]. Advisors/Oppenheimer [Zac76]. Advisory [BCD^{+}52, OBC^{+}49, Syl87, Ano52b]. AEC [Ato54, Ano49b, Ano52a, Ano63d, Bet63, BCD^{+}52, Joi54, OBC^{+}49, Opp54e, SZC54, Str83, You05]. Affair [Wit99, BK68, Uda98, Uda94]. Affairs [Ano54j, GR63, Opp51d, Opp55c]. Affection [Aar67]. Affirm [ACU^{+}54]. After [Lag95, Lin81, Wea82, Wyd84, Ano95a, Can09, EPP80, Els81, Els95, EFB02, Pow96]. Afternoon [Mor66, Opp60a]. Afterword [Hol05a]. Again [Lau65, Nue05]. Against [BS01, Tho04, Nic54]. Age [AM86, GR63, Lau47c, Opp50e, Opp61a, Smi70, Zie91, BP90, Kam85, Nel14, Opp45a, Opp50a, Opp61e, Bal91, Boy85, Kun78, Ris06, SFW91, Opp63n]. Agency [Opp47c, Opp47f]. Agent [Gin54]. ages [Ano04a]. ago [Cow03]. Agony [Rho77, Val06b]. Aid [SMSU10]. ailing [Ano67a]. al [Ben14]. al. [See95]. Alamogordo [GHK^{+}96, Ano45, GHK^{+}96, LW71]. Alamos [Arc21, Bad95, Gol95a, Gow81, Haw61, HBP94, Opp45e, Opp80, Per94, Ree17, See95, She82, Ano96, BHB80, Bet91b, Bra80, Bro80, BS97, Con05, CW88, Fer80, Fey80, Gow81, HTS83, Hir80, HHMW93, Joh59, Kis80a, Kun78, Lew19, McM80a, MB21, Mur05, Nas07, Pat05, Sap05a, SR92, Sla24b, Smi70, SW80c, SFAL21, Tho04, Wei67, Wil75b, WS88, WS08, Wol05, Kru07]. Alamos/Examining [Pat05]. alarm [Yor75]. Albert [Opp39b, Opp65e, Opp66d, Opp79a, Opp79b]. Alfred [Bed06, Wol05]. Algae [AO50]. algebraic [Var02]. Alice [Dys83, Gow81, Hen81, Hew66, Jor80, Kar81, Pei80, She82, Kev81]. Alison [Bad95]. alkaline [HS39]. Alleged [Res54]. allgemeines [Opp55i]. Alloy [Moo21]. Alpha [Opp65b]. Alvarez [Alv87]. Always [Bet67b, Mak95]. Am [Her03, Rho77, GHK^{+}96]. Am. [Ber03b]. Ambivalence [Kem83]. Amends [Ano63b]. America [Hew66, FB69, GCD10, Gus04, HKV10, Nic87]. American [Ber04a, Ger06, Opp62e, Pow05, Wan06, AA54b, Ano54a, Bad05, Bed06, BS05b, BS09a, Blu23, Boy85, CHW91, Cas05b, Cas05a, Chi68, Day09b, Dom04, Fin06, Fra76, Gei93, Gei04, Gre69, Har05, Hec05, Her04a, Hol96, Hol02b, Hug89, Hug90, Hug04, Kai05, Kev05a, MP09, McC06, Opp51c, Opp53d, Opp53c, Opp57i, Opp63k, Pow05, Rho05, Sap05b, Sch08a, Sei78, Str06, Val06a, Wer05, WC84a, Dro23b, Fre05a, Pow05, aS06, Wol05, Jul06]. americana [Val06a]. Americans [Ris06]. among [Opp66c]. Analogy [Opp55a, Opp55e, Opp57l]. Analysis [Win00]. anatomía [GL98]. anatomy [GL95b, GL96b, GL97b, GL98, GL11b]. Anchor [Opp64i]. Andrade [Opp64i]. Andrei [Rhe95]. Angular [Opp31a]. Anniversary [Ano81, Ano93a, Ano93b, Ano93c, Ano93d, Opp65a]. Annual [Ano63d, Bet63]. Answered [Lau56]. Antecedent [Rag80]. anthology [MM62b]. antihydrogen [AZ99]. Any [GB89]. Aperiodic [Opp28f]. Apocalypse [Fas93]. Applicability [Opp40]. Application [STY^{+}98, Opp49b]. Applications [BB21]. applied [AZ99]. Appointed [Ano52a, Ano52b]. appreciation [Gol99]. Approach [Lew19]. Approximation [BD54, STY^{+}98, AZ99, BS85, Ess77, GDR04, Jec14, PST07, Sei69, SK17, SNNY99, Woo25, Sce25]. April [Bet97]. APS [Opp47e]. Archive [Hei68]. Archives [CC21]. Arianrhod [RS06]. Arms [Ano54l, Eak05b, Fre05b, Pow05, RS06, Yor95, Eak05a, Kev05b, McM05, Sap05c]. Army [Jon85]. Arrive [Ano93c]. art [Ano95b]. arte [Ano95b]. Arthur [Duf46]. Article [Ano93a, Ano93b, Ano93c, Ano93d]. Articles [GR63]. Arts [Opp55h, Opp64f, Opp64g, Opp64h]. Ask [Whi58]. Asks [Ano54p]. aspects [Bro56, Bro72, Meh75, Opp48e]. Assemblies [KGJ^{+}21]. Assembling [Man74]. Assembly [Bad95, Gol95a, HBP94, KGJ^{+}21, Per94, See95, HHMW93]. Asserts [Ano54x]. Assessing [AM86]. Assessment [SHM^{+}21]. Assistant [Kil77]. association [Wal96, Opp45e, Opp80]. astounding [Haw11]. Astronomical [Kra17]. astrophysics [Sag79a, Sag79b]. Atom [Ano53a, Ano54d, Ano54e, Ano54b, Ano54c, Ano54l, Ano54v, Ano63b, Ano67a, Bor57, Bor63b, Cha96, DM69, Duf46, Lau45, Opp46a, Opp51b, Opp63h, Opp89, CHW91, Hol02b, Kun15, Opp46k, Rut27, SK17, Uda94, Uda98, WC84a, Wit99, Zie91, Opp64i]. Atomalderens [Lau47c]. Atombombe [Hof95, Hof96, Lam66b]. Atomes [LR57, Opp60e]. Atomic [AM86, Ano49d, Ano67a, Ano16, Bal91, BOT^{+}46, Ber95, Ber03a, Bra10, Bur45, Car88, Com56, Die54, Die62, Dri89, Duf46, Fin54, GR63, Hal97, Hec08, HBP94, Kev03, Lag95, Lan94, Lau47c, Lau54, Lau56, Lef95, Lin81, Mas13, Moo21, Mor57, O’R95, Ols62a, Opp45b, Opp45c, Opp46b, Opp46c, Opp46l, O^{+}46, Opp47a, Opp48f, Opp48g, Opp53d, Opp53c, Opp56a, Opp63k, Opp63o, OR00, Pei97, Ste75, Syl87, Val06b, Wal96, Wea82, Wha55, All05, Ano45, Ano95a, Bak76, Bet91b, Chr98, Dee08, EPP80, Els81, Els95, Gam28, GPDW08, Gor09b, Her05b, Hof95, Jon85, Kai05, Kel06, Kug53, Lam66b, Lar88, Lau46, Lau47b, Lau59, Lau72, LW71, MP09, MW46, MW07, McI05, Nel14, Opp45a, Opp46d, Opp46e, OR47, Opp48b]. atomic [Opp48e, Opp50a, Opp51a, Opp57b, Opp61c, Opp61d, O^{+}01, Pat05, Per95, Pow96, PGDW08, Rai02, Rou62, SR92, She72, SK13, Van03, Wai96, AS05, Amu11, Boy85, Edg91, EOA^{+}08, Hev11, HH89, Hun09, Hun14, Kal54, Kal63, Kun78, McF11, Opp57k, SFW91, Wan11]. atomic-bomb [Dee08]. atomique [Rou62]. Atomkernes [Gam28]. Atomkraft [Opp57k]. atomnaja [Rou63, Rou65a]. Atoms [HH89, Lau59, Bal91, LR57]. attempts [Yor75]. Attention [Ken54]. attribute [Opp62a]. August [Wit99, Haw61, Uda94, Uda98]. Auschwitz [Fas93]. Ausgewählte [Bor63a]. author [Dro23b]. Authority [Hec10, Tho02]. Autobiography [Hec10, MS60, Dys18]. Autoelectric [Opp28c]. autour [SB55b]. Autumn [Mor66]. aventure [LC68]. Award [Ano63b, Opp56d, Ano63a, Ano63d, Bet63, Opp54c].

B [Bal91, Con17]. Background [Ada54]. Backs [Ano54h]. Backward [Mea04]. Bad [Mor66]. Badash [Gow81, She82]. Balance [Ano53a]. band [Cas07]. bands [Opp26b]. Banquet [Ano81]. Baptism [Ste75]. Based [Gin54]. Basis [Kip67b, Kip68b]. Baym [Bad95]. BBC [Yox82, Str83]. BBC-2 [Yox82]. Be [Lau56, Ano11, Che59, Che69, Gro62, Gro83, Mor95]. Become [Rho77]. Bedrohung [SB55a]. Bee [Bal91]. Been [Ano67a]. Before [Car21a, Car21b, Wyd84, Kal54, Kal63, OR00, Rig95a, U.S71, Wha55, Rig95b]. beginning [Ano96]. Beginnings [Gon96, Ber87]. Begins [Ano93d]. Behavior [MM39, Opp39c]. Behind [Ano96, Wil70, Dro23c, Neu13, Pat05, SF08]. Belies [Lau45]. Below [Fey80]. Bemerkung [Opp27a]. Bericht [Kip64]. Berkeley [Ano81, Bro06, Hug07, Nis92, You05]. Berlin [Hof96]. Berlin/Heidelberg/New [Hof96]. Bernard [Men84]. Bernstein [Ger06, Lan04, Pow05, Ber03b]. Berry [BKLB92]. Bestrahlung [HS39]. Beta [Opp41d]. Beta-Decay [Opp41d]. Bethe [Aft02, Ano50, Dre07, Edg91, Fra01, Lan00, Olw01, Sar01, Sei01, Ano00a, BWMH81, Bro82, Sch00]. Better [Tel87, Hir88]. Between [Fei60a, LRA21, DD71, Val06a]. Beyond [Bow08, CF21, HO21, SK17, Dom13]. bibliography [SWPS67b]. Bicentenary [Ano54z]. bien [Ano95b]. Big [Hil15, Opp59l]. Bildbiographie [Goo82]. biographical [CC56]. Biographie [BS09b]. Biographies [Ban12]. Biography [Per95, Sch10, Wal96, BS09b, BP90, Dom13, Dri89, Goo82, Guy07, Law88]. Bird [Bed06, Fre05a, Pow05, Wol05, Dro23b]. Birth [Eak05b, Fre05b, Hil15, Lau47c, Pow05, RS06, Eak05a, Kun78, McM05, Sap05c]. Birthday [DPSY64, Opp39b, Opp79b]. Bitter [Sch08a]. Black [ORSSBO^{+}17, Huf05]. Blackett [Wan06]. Blast [BR21, Lau45, Mor21]. Blast-Loaded [Mor21]. Blue [AO50]. Blumberg [Ber78, Zie91]. Board [Wha55, GM54, U.S71]. bodies [Opp26a, Oppxx]. Bohr [Duf46, Opp62e, GH94, Haw94, Opp63p, Opp63f, Opp63g, Opp64k, OBxx, She72]. Bolster [Bro14]. Bomb [Aft02, All56, Ano54e, Ano67a, Ber88, Ber03c, Bet82, Boy85, Car88, Cha96, Duf46, Dys76, Fra01, Gol95c, Gon96, Gus04, Her03, Ken54, Kev03, Lag95, Lan00, Lau45, Lau65, Lef95, Mac54, MD67, Mor57, O’R95, Ols62a, Olw01, Opp46a, Opp51c, OR00, Rhe95, Sar01, Sch03a, Sei01, TB82, Wol09, Ano45, Ano50, Ano95a, Bak76, Ber03b, BS05a, Bro82, Cas05a, Chr98, Coh83, DD71, Dee08, Dom13, Doy02, EPP80, Els81, Els95, GF65, GF67, GPDW08, Han95, Her02, Hof95, Hol02a, Hol02b, Jon85, Kel06, Koj05, Kun15, Lam66b, LC68, Lan05, Lar88, Lau46, Lau47b, Lau72, Law88, Man84, MP09, MW46, MW07, McI05, Mur53, Neu13, Nor02, Nor03, Oli89, Opp45d, Opp46k, Opp61c, Opp61d, Opp65c, O^{+}01, Pat05]. bomb [Per95, Pow96, PGDW08, Rai02, Rou62, SF08, Sch00, SR92, SB54, SB55b, SB55a, SB71, She72, Sti47, Sti76, Str98, Van03, Wai96, WG17, Wol05, Ano54s, Eas89, Kev03, Duf46, Han95, Wal96, Wea82, Lin81]. bomba [Rou63, Rou65a]. Bombardment [OS39a]. Bombe [DD71, LC68, Rou62, Str98]. Bombing [Gol96]. Bombings [Ber95]. Bombs [Opp59i, Die54, Die62]. Book [Bad73, Bad77, Bad95, Bal91, Bed06, Ber03c, Boy87, Bra10, Bro06, Day09a, Dys83, Dys13, Eak05b, Edg91, Fra01, Fre05a, Fre05b, Ger06, Gol95a, Goo09, Gor09a, Han95, Hen81, Hew66, Hig76, Hir88, HBP94, Hor70, Huf69, Hug07, Hug09, Joh13, Jul06, Kev81, Kle07, Lan00, Lan04, MW09, Mas13, Opp28g, Opp63n, Opp64i, Pei80, Per94, Pol07, Pry55, RS06, Rob56, Sar01, Sch03a, Sei01, Set05, aS06, She82, Tho03, Tho11, TS12, Und13, Wan04, Wan06, Wan08, Wei86, Wig76, Wit99, Wol70, Yox82, Zie91, Ber03b, Her03, Hug70]. Books [Ber78, Gol55, Gro56, Ols62a, RS06, Wil69, Wol05, Zac76]. Borderlands [Mas06a, Mas20]. Born [Woo25, Man80, AZ99, BS85, BD54, BKLB92, Bor64, Bow08, DM69, Ess77, GDR04, Jec14, PST07, Röm83, Sce25, Sei69, SK17, STY^{+}98, SNNY99, Tak09, Tap01]. Boron [OS38a]. Bottle [Ano53a]. brain [SF08]. Braziller [Hug70]. Breakdown [BS85]. breakthrough [Opp56c]. Brian [Kev03]. Bridgman [Opp28g]. Brien [Ano50]. Brilliance [RS04]. bring [Yor75]. brings [Ber87]. Britain [Sza95]. British [Moo21]. Broadcast [Ano54r]. Broglie [Var02]. Broida [She82, Gow81]. Brotherhood [Ber03b, Doy02, Her02, Hol02a, Ber03c, Her03, Kev03, Sch03a]. Brown [Kev03]. Bruceton [MB21]. Brunswick [Wit99]. Build [GB89, Hol02b, SR92]. Builders [MD67]. Building [BS05a, Gol95c, Lan05]. Built [Ano54s, Kun15]. bulk [Eck70, Eck73, OBxx, Rabxx, Veb60, von96, vN96]. Bulletin [GR63, AM86, Bro23]. Burchett [BS07]. burdens [Hol86, Hol98]. Burst [KGJ^{+}21]. Bursts [Opp39a]. Bush [Ano52b]. Business [McG02].

C [Ber04a, Ger06, Guy07, Opp64i, Pow05, Wan06]. C.B.S. [Ada54]. ca [Ano00a]. Calculated [Joi54]. Calculation [SFAL21, BS85]. California [Bro06, Hug07, Ano34, Sei78]. Called [Gin54]. Caltech [McC04, McC05]. Cambridge [Bad95, Goo09, Hen81, Per94]. Campaign [Mur53]. can [Gro62, Gro83]. Canada [Inf78b]. Canadian [AAM21]. Cancer [Ano67a, SBB21]. Candid [HH04a]. Cape [Dom13, Und13]. Capture [Opp28d]. career [Opp54a, Wol05, Ano67a]. Carl [Bal91, Eck70, Eck73]. Carlo [SFAL21]. Carson [Bro06, Hug07]. Case [Ano54f, Ano54j, Ano54k, Ano54m, Ano54n, Ano90, Ber90, Edi54, Cos14, Gin54, Gro99, Gro56, Her04b, Hug70, Kal54, Kal63, Ken54, Kev03, Lau54, Lip54, McW54, MT55, New77, OR00, Put63, Rhe95, Rie54, Rie69b, Sch54, Sim54, Str63a, Str64, Tho02, Tri54, WC84b, Ato54, AA54b, Ben14, Bow03, Che65a, Cur55, Gre77, Gre23, Hor70, KL66, Kip67a, Kip68a, Kip05, Put60, Put61, Rie69a, S.54, SG69, SG71, Str63c, Str65, Wha55]. caso [Ben14, Cos14, KL66]. Cassidy [Ber04a, Ger06, Wan06, Pow05]. Catastrophe [Sag83]. Catherine [Bad95, Gol95a, Per94]. Cathryn [Bro06, Hug07]. cause [Kip67a]. Celebrates [Gui04]. Celebration [Ano81, Opp39b, Opp79b]. Censoring [Dee08]. censorship [MM62b]. Censured [Ano67a]. Centenary [Gui04, Fre79]. Centennial [Bro06, Hug07, CH05]. Center [Mon13, Dys13, Joh13, Kru15, Mas13]. Centre [Dom13, Har12, Kev13, TS12, Und13, Mon12]. Century [Hen81, Wan06, Cas05b, Dom04, Har05, Her04a, Hol96, Hug89, Hug90, Hug04, Rho05, Anoxx, Ber04a, Ger06, Har10, Kai02, Kev05b, Pow05, TS01, Wan06]. cerca [Ano95b]. ceremony [Opp62e]. certain [Oppxx]. challenge [Fas93, Opp59c]. changed [Lam66b]. Character [Ano54b, Fla98, Dro23b]. characteristics [Oppxx]. Charge [Opp35c]. Charged [Opp35d]. Charges [Gin54, Nic54, Nic23]. Charisma [TS00, RS04, TS00]. Charles [Dys83, Gow81, Gro56, Hen81, Jor80, Kar81, Kev81, Pei80, Pol07, She82, Hug09]. Chart [Ano93a]. Chase [Opp57j]. Chemical [Tap01]. chemistry [Sce25, Tak09, Tho17, Woo25]. Chevalier [Hug70]. Chicago [Ger06, Guy07, Hug09, Lan04, Opp63n, Pol07, Opp63r]. choice [Mas06b]. Choices [Opp57m]. Christopher [Mec23]. Christy [CC21]. ciencia [Ano95b, Opp55f]. científico [Ano95b]. Cites [Opp54i]. citizen [Rig87b, Rig00b]. City [FOS09, Har09, Kru07, Kun78, Wol05, Con05, Dud80, Mur05, Nas07, Sap05a]. civilian [Blu23]. Civilization [Opp61a, FOS09, Har09]. Claymont [Wal96]. Clearance [Aft03, Ano01b, Tho03, Wan04, Eva54, Pol02]. Climatic [Sag83]. closely [Ano95b]. closer [Har10]. closet [Her05a]. Closing [Hug70]. cloth [Bed06, Fra01, Ger06, Goo09, Kle07]. Cloud [Wei88, Gor09b]. clouds [Hec05]. Club [Opp54c]. Coast [Her09]. Cohen [Men84]. Cold [Bra10, Day09b, Hec08, Wit99, Amu11, Gus96, Hev11, Hun09, Hun14, Kai05, Koj05, Mas06a, Mas20, McF11, Ris06, Uda94, Uda98, Wan11]. collaboration [BB97b]. collapse [IKL99]. Collateral [Blu23, Opp56f]. Colleagues [Lau54, Wil94]. Collected [Kip87]. Collection [SMSU10]. college [Opp46k]. Collisions [Opp40]. Columbia [Wit99, Ano54r, Ano54p, Ano54z]. Comment [Ber03b, Her03, Opp56b, O^{+}65, Sim54, Woo25]. Commentary [Man84]. Comments [Bet82, Opp51b, Opp55b]. Commission [Bal91, HH89, Kal54, Kal63, Syl87, Wha55]. Committee [Syl87, Ano52a, Ano93d, BCD^{+}52, OBC^{+}49]. Common [Gol55, Opp54g, Opp54h, Opp55i, Opp60b, Pry55, Opp55f, Opp64d]. commun [Opp64d]. Communication [Opp63l, Opp63c, Opp63d, Opp63s]. Communications [Szi71]. Communist [Her05a, Ber05, BS05c]. Community [Hab69, Opp59m, Opp59q, Opp76, Opp89, ORSSBO^{+}17, Opp57g, Zie91]. como [Ano95b]. Company [Opp28g]. Comparison [Cha21, LRA21]. Compartmentalization [Gol95c]. competition [SB55b, SB55b]. Complementarity [Opp63e]. Complete [Gol99]. Complex [Gus04, Hil15, TS00]. complicated [Neu13]. Component [CO41]. Comprehension [Opp63l, Opp63s, Opp63c, Opp63d]. Compton [Duf46]. Computer [Gol72]. computers [Opp53a]. Computing [Arc21, Lew21]. común [Opp55f]. Conant [Wol05, Con17]. conception [SB55b]. Concerning [HS39]. Concluding [Opp48a, Opp49a, Opp58a]. Concurring [Mur54]. Conference [Fei60a]. conferences [Meh75]. Confidence [Ano54w]. Confirm [Lau45]. Conflict [Ber78, BO76, DD71, SB55a]. Confronting [Bal91]. Confronts [Nim11]. Congress [Ano74, SMSU10]. connais [Str98]. connection [BKLB92]. conocimeinto [Opp55f]. conocio [Ano95b]. Conquest [GCD10]. Conscience [Car88, Str63c, Str65, Put63]. Conscientious [Rag80]. Constitution [Opp56k]. construction [SB55a]. Containing [Fin06]. contemporanee [Cas14]. Contemporary [Opp48h, Opp51a, Opp59a, Cas14, Opp46i, Opp47a, Opp47d]. Contesting [Lew19]. Continental [Ano54l]. Continued [OS39b, IKL99]. Continuous [Opp26d, Opp26c, Opp27e, Opp27d]. Contraction [OS39b]. Contributions [AAM21, Opp53a, ORSSBO^{+}17, Bad95, Sea67]. Control [BOT^{+}46, Opp46l, O^{+}46, Opp48f, Opp48g, Opp63m, Opp63o, Her05b, Opp46d, Opp46e].

Controversy [Ano54x, Ano67a, San79]. Conventional [Usd09]. conversations [HH04a]. converse [Opp61b]. Conversion [AO50, Opp41a]. Cores [OS38b, OV39]. Cornell [Tho03]. Corrections [AZ99]. Cosmic [CO41, MM39, OS37, Opp39c, Opp44, Opp49a, Opp56c, Ros64, Opp48a]. Cosmic-Ray [OS37, Opp49a]. cosmological [Win00]. Coulomb [Opp29]. Coulombfeld [Opp29]. Counting [Gar08]. coupling [Röm83]. Crane [Hen81]. Craters [Szi71]. Crease [Kle07, Wit99]. Creating [Lew19]. Creation [Opp47b, Chr98]. creative [BB97b]. Creator [Hof95]. Crimes [Ano00b]. Crises [Opp64j, Opp69]. Crisis [Opp45c]. Critical [HHMW93, KGJ^{+}21, Per94, Sag79a, Sag79b, Bad95, Gol95a, HBP94, See95]. Criticality [HAM^{+}21]. Crossing [Opp66a]. cubit [Opp63a, Opp63b]. cultic [Wal13a]. Culture [Boy85, Hol96, Opp58n, Opp59p, Opp60i, Opp61g, Opp61h, Cas14, Opp56e, Opp58d, Opp58e, Opp58k, Opp60d, Opp60g, Opp60h, Opp62d, Opp62c, Pat05]. cultures [Cas14]. Current [Sim54]. Currents [Opp28c, Opp28b]. Curtis [Gro56]. Cyclotron [Hil15]. Czech [Put63].

dai [Opp56m]. damage [Blu23]. Damned [Ano93a]. Damping [BO46]. Danish [Lau47c]. dann [S.54]. dans [Dem15]. Dark [Rum92, Kam85, Han95]. Data [Cha21, Res54]. David [Ano50, Ber04a, Bro06, Ger06, Hug07, OBxx, Pow05, Wan06]. Davis [Hug70, Wil69]. Dawn [Boy85, Car21a, Car21b, Lau46, Lau47a, Lau47b, Lau72, Cah95, Dri89, Gor09b, Lau47d, Opp63n]. Day [EPP80, Els81, Lag95, Lam65, Lam66a, Lam85, Lam89, Mor66, Wea82, Wyd84, Ano95a, Els95, EFB02, HHW99, Pow96, Lam68, Tho16, And66, Hew66, Lin81]. days [McM80a, Opp55d]. Dayton [Tho17]. Deadline [Res91, Res92]. Deak [Chr98]. Death [Ano67c, Rho77]. Debate [Ano54k, Anoxx, Her05a, QJ94]. Debated [Lau54]. Decades [Opp54i]. Decay [EFO48, MKC^{+}21, Opp41d]. Decision [Dys76, GB89, Lef95, Mor57, Opp50g, Opp65c, SZC54, Str63a, Bak76, Bro23, GF65, GF67, Opp50d, Sti47, Sti76, Wai96]. decisions [Str63b]. Declared [Ano54a]. Dedicated [Ano11]. dedication [Opp62e]. Dee [Ber04a, Ger06, Lan04]. Defends [Res54]. Defense [Ano54l, Bro14, Hir88, Ken94, Hol93, Tel87, New77]. degeneracy [Opp27c]. del [GL98]. Delay [Ano54e]. Delete [Dev62]. dell’illuminismo [Cos14]. delphic [Dro23b]. demain [O^{+}65]. Democracy [Tho02]. democratic [Opp63i]. Demring [Lau47d]. Denaturing [ABB^{+}46]. denial [Eva54]. Denied [Ano54g, Res54]. Denken [Opp55i]. Density [OS36, STY^{+}98, SNNY99]. Department [Bir50, Bir65, Bro23, New77]. dependent [PST07]. Depression [Nis92]. Dep’t [ABB^{+}46]. Description [Joh59, Opp55e, Opp57l, Opp58b]. Desert [Ban12, Cah95]. Designing [DOW84]. Despite [Ano54v]. destiny [DD71]. Destroyer [Ban12]. Destruction [Opp47b]. detonated [Ano45]. detonation [GHK^{+}96]. Deuteron [Opp35b]. Deuterons [OP35]. Development [Moo21, Opp47c, Opp47f, Opp54i, Dir71, Har08, Meh75]. Developments [CF21, Opp59a]. Devil [Knu83]. DeWolf [Edg91]. Dexer [Duf46]. día [Lam68]. Dialectics [Cos14]. dialettica [Cos14]. dialogue [Opp64a, Opp64b]. diatomic [Röm83]. Dictionary [DG99, CC56]. Did [Ano94, ORSSBO^{+}17, Dro23a]. Didn’t [Wei94a, Wil94]. Dies [Ano67a]. Difficulties [Ano34]. Dignity [Opp56d]. Dilemma [Wil70]. dilemmas [McC04]. Dimensionally [IKL99]. dimensions [IKL99]. Dirac [Ano34, Ano54g]. directed [Wea82]. direction [Opp27c]. director [Law88, Mec23]. Disarmament [Ano52b]. Disciplining [Tho02]. Disclosure [Lev60]. Discovery [BHT86, Gra64, Opp49b, CHW91, GA71, Lau59, Opp59o, WC84a]. Discuss [Ada54]. Discussion [Opp39c, RAO^{+}49, Ano50, GHK^{+}96]. Disintegration [NNOS37, Opp33, Opp35b, RAO^{+}49]. dispersion [Opp27a]. Dispose [Ano34]. Disputes [Ken54]. disquieted [Lan05]. Dissenters [Lan05]. Distribution [Opp26d]. District [Haw61, HTS83]. Disturbing [She82, Dys79]. Divides [Lau54]. Dixon [Ano54x]. DM [Hof96]. Do [Whi58, Ano50, O^{+}65, Pol19, WS88, WS08]. Doctor [AS05, EOA^{+}08]. Doctors [Bur45]. Documentary [Mer95, CHW91, GA71, SFW91, WC84a]. Documents [Kip67b, Kip68b, U.S71]. Dodd [Hug70]. Does [HO30]. Don’t [Pin15]. Doses [SBB21]. dossier [VK65]. Doubleday [Dys13]. Dr [Ada54, Wha55]. Dr. [Agr65, Ano45, Ano53a, Ano53b, Ano67a, Ano11, Exe54, Goo04, GTS^{+}06, Ken54, Lak66, Mur53, Opp51c, Opp54i, Res54, Rie54, Rie69a, Str63a, Str63c, Str65, Wed67]. Draft [OR47]. drafts [Dee08]. Dragon [KGJ^{+}21]. Drama [Cos14, Cou63, Ken54]. dramatic [Nel14]. dramma [Cos14]. Drawing [Ber03d]. dreams [Haw11]. Drop [Lef95, GF65, GF67, Opp65c, Wai96]. DuBridge [McC04]. Duckworth [Ber04a]. Duckworth/Ivan [Ber04a]. During [Arc21, HBP94, HAM^{+}21, Bad95, Gol95a, HHMW93, Kai05, Mor66, Per94, See95]. Dyson [She82, Ano70, Ano95b].

Each [Pin15]. Earliest [Moo21]. Early [AAM21, Ber88, Boy85, McM80a, ORSSBO^{+}17, Wei94b, Kai05, Ser67, Hec08]. earth [HS39]. East [Con05, Kru07, Mur05, Nas07, Sap05a, Wal96, Wol05]. Eckart [Eck70, Eck73]. ed [Kar81, Tho03]. Edged [Mas13]. Edited [Duf46, Hen81, Pei80, Dys83, Gow81, Jor80, Wha55]. Editor [Sch04, Wan04, BR57, Lev60, Mac54, Ols62b, Opp62f]. Editorial [Ano63c, Edi54]. Editors [Bro06]. eds [Hug07, Per94]. education [Fra59, Fra76, Opp46k]. Edward [Ber78, Ber03c, Ber03b, Her03, Hir88, Kev03, Sch03a, Anoxx, Ano13a, BO76, BP90, Bro82, Cow03, Doy02, Goo04, Har10, Her02, Hol02a, Rhe95, Tel98, Zie91]. Effect [HO31b, Opp28e, Opp63i]. Effects [Opp28f, GHK^{+}96]. Effort [Lew21, Hol02b]. Einstein [MW09, Ano54h, Ano95b, Anoxx, FW07, Fre79, Goo09, Isa07, Kai02, Kol07, Opp39b, Opp56l, Opp65e, Opp66d, Opp79a, Opp79b, Opp79c, RTS^{+}69, RTS^{+}xx, RS06, Sch05, Sch06, Sch08b, Wes14, Tho11, Bey09, Gor09a]. Einzelausgaben [Kip87]. Eisenhower [Bal91, Bir01, HH89, Kil77, LPT^{+}54].

Eleanor [Ano50]. electric [BS85]. Electrodynamics [Gia85, Opp70, Sch58, Sch03c, Var02]. Electromagnetic [Opp31c]. Electron [Ano34, BO46, FO34b, FO34c, Opp29, OP33, Opp34, Opp50f, Opp55e, Opp57l, Opp75, Ano34]. Electronic [Opp28a, Tap01]. Electrons [CO31, CO32b, CO32a, MM39, Opp28d, Opp30b, Opp39c]. Elektronen [Opp29]. élémentaires [Opp58b]. Elementary [Opp39a, Opp58b]. Else [Wea82]. Emerged [Opp54i]. eminent [CC56]. Emission [OS39a]. empire [FOS09, Har09]. Empiricist [Sch86]. Encouragement [Opp50c, Opp51d, Opp50b]. End [Gus96, Ols62a, FW07, FOS09, Gor09b, Har09]. ENDF [Cha21]. Ending [Ano54z]. Enemies [Ano54x]. Energies [KOS37a]. Energy [Ano49d, Ano54c, Bal91, BOT^{+}46, BO76, CO41, Edg91, HH89, Kal54, Kal63, MKC^{+}21, MM39, Moo21, NNOS37, Opp33, Opp35a, Opp39c, Opp46l, O^{+}46, Opp48f, Opp48g, Opp54i, Opp63o, Syl87, Wha55, Her05b, Lau59, LW71, Opp46d, Opp46e, Opp47a, Opp48e, Opp49b, Opp51a, Opp56a, Opp57k, Opp65d, Bro23, Ber78]. engine [Opp56m]. Engineers [Opp57a]. enigma [Ber04a, Ber04c, Cas65, Ger06, Pow05, Ber04a, Lan04, Her07b, Set05]. Enlightenment [Cos14]. Enormoz} [Her09]. Enrico [Gol99]. enthusiasm [Hug89, Hug90, Hug04]. entstehenden [HS39]. environs [Opp57b]. Envisions [Lau56]. epic [Nel14]. Equation [CF21]. equations [Win00]. Equilibrium [HO30]. era [Nel14]. Erdalkalimetalle [HS39]. Ereignis [GHK^{+}96]. Erläuterungen [Gro99]. Ernest [Ber03c, Ber03b, Her03, Kev03, Sch03a, Chi68, Doy02, Her02, Hil15, Hol02a, Opp64i]. Errata [NO34b]. Erratum [Ano49a]. errors [Bro82]. erste [GHK^{+}96]. ersten [Hof95, Hof96, Lam66b]. esemplari [Rou73]. Espionage [Her09, Usd09]. Essay [Dom13, Gow81, Guy07, Edg91]. Essays [Opp89, Zie91, Fra76, Hol86, MS97, Men84]. est [Dem15]. Establishment [Cob72]. Estimated [SBB21]. ethical [Fas93]. Ethics [Mor66]. Ettore [EAW15]. Eugene [Opp63n, HH04b]. Europe [Ell86, FB69]. Even [Wei94a]. event [GHK^{+}96]. Evidence [CC21, Her05a, Wha55]. evolution [MS97]. Evolves [Ano93b]. Ex [QJ94]. Ex-Soviet [QJ94]. examine [Ano45]. Examining [Pat05]. Exchange [Her04b, MT55, Opp54i]. Excitement [Bet67b]. existence [HS39]. experiences [Ano96]. Experimental [Ano34]. Experiments [HAM^{+}21]. Expert [Lau54, Res54]. Experts [Tho02]. Explain [Pin15]. Explaining [Lau56]. Explanations [Gro99]. Exploration [Wit99, Uda94, Uda98]. explore [Wol05]. Explosion [Lam66b, LRA21]. Explosive [BB21]. explosives [Opp46b]. Exposure [SBB21]. Express [Ano54w]. expression [Opp60h]. Extend [Ano34]. extended [Mec23]. extraordinary [SK13]. Eye [Bal91].

F [Dys76, Edg91, Wol70, Zac76, Zie91]. F. [Ano95b]. F.B.I. [Ano54v]. Facility [Arc21]. facing [Ano49c]. facsimiles [Bey49]. Failure [Opp63m]. Faith [ACU^{+}54, Har08]. Fall [Che65a, Put60, Put61, S.54, Bur45, HKV10, Nel14]. Fallout [O^{+}01, Rai02, SBB21]. Fame [Ano54s]. Famed [Ano67a]. Famous [Cou63, HH04a]. fantastic [LC68]. fantastique [LC68]. Farm [Opp63j]. Fashioned [Duf46]. Fast [CO31, CO32b, CO32a, HAM^{+}21]. Fateful [Opp50g, Opp50d]. Father [Ano67a, O’R95, Wal96, Wol05, GPDW08, Per95, PGDW08, Kel06]. Faust [Knu83]. FBI [Ano78, Fed78]. fear [Bur45]. February [Bet97, Ano50]. Feels [Ano54x]. Feldbaum [Bal91]. Fellowships [OBC^{+}49]. fence [McM80b]. fences [Ano96]. Fermi [Ano63a, Ano63d, Bet63, Gol99, Kat21, Wei94a, vH01]. Fermis [Fer80]. Fete [Ano54p]. Feynman [Anoxx, Kai02]. fiction [Cas03]. Field [Opp28c, Opp28b, Opp30a, Opp31c, Opp35c, TO48, Opp29, Opp59a]. Fields [Opp41c]. figure [Cas14, Neu13]. figures [Cas14]. File [Ano54v, Bre94, Fed78, Les94, McM94, Ano78]. Film [Wea82, EG05]. filosofico [Ano95b]. final [Dom08, Wol08]. Finding [SMSU10]. Fire [Opp63n, Fin06, Kun78]. firing [Mac54]. First [Ano50, Ano54x, BS01, Ken54, Opp59i, Bra80, Cas07, Dee08, GHK^{+}96, Hof95, Lam66b, SR92, Kil77]. Fisk [Ano52a]. Fission [BHT86, BW39, Gra64, CHW91, GA71, WC84a]. Fissionable [ABB^{+}46]. Fit [Whi58]. Five [Ano54d, LW71]. Flimsy [McM94]. Fluctuations [Opp35c]. Fluorine [OS39a]. Flying [Opp64j, Opp69]. Focus [Smi70]. Focuses [Ken54]. Fødsel [Lau47c]. Footnote [Oli89]. forbearance [Opp66b]. Force [Opp46a]. forces [Opp65j]. foreknowledge [Opp64c]. Foreword [Duf46, Opp65d, Edg91]. form [Ano95b]. forma [Ano95b]. Formed [Ano93d]. Formula [Cas14, Cas14]. Formulae [Opp35a]. Formulas [LRA21]. Formulation [Gia85]. Forscher [DD71]. fortunate [Pol05]. forum [Ano50]. Foundation [BB21]. Foundations [Bey49, Jor27, Opp47g, Var02]. Four [Ber88]. fox [Pol05]. fra [Val06a]. Fractured [Bor08]. Francis [You05]. Frank [DOW84]. fraternal [Opp64a, Opp64b]. Free [Opp29]. Freedom [Opp59b, Opp62a, Opp57k]. Freely [Kip67b, Kip68b]. Freeman [She82]. freien [Opp29]. Freiheit [Opp57k]. Freis [Ols62a]. French [Dem15, Kip67a, LC68, LR57, Opp58b, Opp64d, O^{+}65, Rou62, SB55b, Str98]. Freundschaft [Che65a]. friends [RTS^{+}69, RTS^{+}xx]. friendship [Che65a, Che65b, Che66, Hug70]. Frontiers [H^{+}58, Wit99, SC98]. Full [Duf46, MW46, MW07]. Function [OP35, SNNY99]. Functional [STY^{+}98, SNNY99]. Functions [Opp47c, Opp47f]. fundamental [Bey49]. Furman [AO44]. Furry [Ano34]. Future [Opp61a, Lau59, Opp48d].

G [Bal91, Zie91]. Gadget [MB21, CC21]. Gaia [MS97]. Gaining [Wes14]. Galileo [De 57, De 68, Wes14]. Galillean [Opp64e]. Gamma [MKC^{+}21, NO33, NO34b, NO34a, HO31a]. Gamma-Rays [NO33, NO34b, NO34a]. gas [Bur45]. Gebunden [Hof96]. Geburtstag [Pun67]. Geburtstag/J [Pun67]. Gen. [Ano45]. genealogy [Wal13b]. General [Ber03a, BCD^{+}52, Opp54i, Dom13, Kun15, Law88, Ano11, Gol96, Law88, Nor02, Nor03, Syl87]. genesis [Hug89, Hug90, Hug04]. Genius [Dys13, MW09, Val06b, BB97b, Chi68, Kun15, Sch08b, Wil83, Goo09, Bey09, Gor09a]. Georg [Pun67]. Gerald [Ber04a]. German [BS09b, BO27, Bor63a, BO63, Che65a, Che69, DD71, Gam28, GHK^{+}96, Goo82, Gro99, HS39, Hof95, Kip64, Kip68a, Kip87, Kip05, Kro81, Lam66b, Opp26c, Opp27a, Opp27e, Opp27d, Opp27c, Opp29, Opp55i, Opp57k, Put60, Put61, S.54, Sch82, SB55a, Wei86]. Gesammelte [Kip87]. Geschichte [Che65a, Lam66b]. Giant [Zie91, BP90]. Giants [MD67, Ris06]. Gita [Hij00]. given [Wha55]. Glance [Mea04]. Glenn [Wol70]. Goal [Lau56]. God [Che59, Che69]. Gödel [FW07]. Goeppert [Goe88]. Golden [Zie91, BP90]. good [Ber97, Opp55d]. Goodchild [Gow81, She82, Yox82]. Gordon [Bad95]. Gott [Che69]. government [Hol02b]. graphs [Oppxx]. Gravitational [OS39b, IKL99]. Great [Ano49b, Hal09, Hug70, Opp46a, Pfa84, Wil70, Bak76, Opp59c, Sza95]. Greatest [Ano54s]. Green [AO50, Hor70, SNNY99]. Gregg [Ber03c, Ber03b, Kev03, Sch03a]. grips [Yor75]. Ground [Ken54]. Group [Ano52b]. Groves [Ano45, Ano11, Ber03a, Gol92a, Gol92b, Gol95b, Gol95c, Gol96, Kun15, Law88, Nor02, Nor03]. growing [Pei80]. Growth [Opp58n, Opp59p, Opp48b, Opp56e]. guerre [SB55b]. Guide [Ano00a]. guru [Hei05]. Gwinn [Ber78].

H [Duf46, Ano54e, Bet82, Bro82, Gon96, Ken54, Mac54, Man84, Mur53, Wei86, Wol09]. H-Bomb [Ano54e, Bet82, Gon96, Ken54, Mac54, Wol09, Bro82, Man84, Mur53]. H. [Hig76]. Haakon [Hug70]. half [Ano49c]. half-length [Ano49c]. Hall [Ber04a]. hands [Kai05]. Hans [Ano50, Edg91, Ano00a, BWMH81, Bro82, Dre07, Pun67]. Hans-Georg [Pun67]. happened [Kam23]. Harald [OR00]. hardback [Dom13, Guy07, Hug09, MW09, Und13]. hardcover [Ber04a, Edg91]. Harold [Hor70]. Harper [Hug70]. Harvard [Goo09, Hen81, Jor80, MW09, Pei80]. having [Ano95b]. Hearing [Ano13b, Lau54, Tho03, Maj71a, Maj71b, Maj83, Pol02, U.S71, Wan04, Aft03]. Hearings [Ano49b, Ano54l, Ano67a, Res54, You05]. heart [Mil05]. heaven [Cas03]. Heidelberg [Hof96]. Heinar [Cos14, Gro99, Wol70, Gro99, Kip87]. Heisenberg [Ano34, Hol84]. Heitler [BO46]. Held [Gin54]. Hells [Cou63]. Help [Ano94]. Hemmelige [Lau47c]. Henrik [OBxx]. Henriksen [Bad95, Gol95a, Per94]. Henry [Edg91, Kev03]. Herbert [Bad77, Dys76, Edg91, Gow81, Ols62a, San77, She82, Wig76, Zac76, Zie91]. Here [Ano93a]. Heresy [Wil70, Hug70]. Heritage [Sch08a]. Herken [Ber03b, Kev03, Ber03c, Sch03a]. Hermitean [Röm83]. Hero [Hec08]. Heroes [RS06]. Hewlett [Bad95, Bal91]. Hidden [Mur53]. High [Ano93c, BB21, CO41, KOS37a, MM39, NNOS37, Opp33, Opp35a, Opp39c, Opp65d]. High-Explosive [BB21]. Him [Joh13]. hindered [Bro82]. Hiroshima [Bur83, Chr98, Fas93, Gol96, TB62, Wyd84]. Hiroshimas [Ano54d]. Hirschfelder [Gow81, She82]. Historic [LRA21]. Historical [Ano13a, Edg91]. Histories [Pei97, BWMH81]. Historiography [Hei68]. History [Ano01b, Bad95, Ber03d, Bet82, Bir50, Bir65, Bro82, Bro06, Gol95a, Haw61, Hei68, HBP94, Hol96, Hug07, Opp62b, Per94, See95, TB82, WP85, Wei88, Bro56, Bro72, CHW91, Dee08, GHK^{+}96, GA71, HTS83, HHMW93, KHFA67, Lam66b, Mer95, Oli89, Opp62e, Opp71, Opp72, Syl87, WC84a]. Hoddeson [Bad95, Gol95a, Per94, See95]. Hole [ORSSBO^{+}17]. holes [Huf05]. Holl [Bal91]. Hollinger [Bro06, Hug07]. Holt [Kev03]. Hommes [LR57, Opp60e]. honor [Gra02, Men84]. Honored [Ano67a]. Honors [Ano53b]. Hope [Hew66, Opp57d, Opp62b, Opp64c, Sch08a, Tho16, Day15, Opp57c, Sch62]. Houghton [Hug70]. hour [Con17]. Hours [Tel83]. House [Duf46, Opp60c]. Human [Duf46, Opp59m, Opp59q, Opp76, Opp56c, Opp57k, Opp57g, Opp61e, Opp62a, Sea67, Yor75]. Humanity [Bor08]. Hundred [Ano54d]. Hungarian [Jon10]. hunt [Hol02b]. Hydride [HAM^{+}21]. Hydrodynamic [MA21]. Hydrogen [All56, DM69, Dys76, HO30, Han95, Ken54, TB82, Ano50, AZ99, SK17, SB54, SB55b, SB55a, SB71, WG17]. Hydrogen-Bomb [Dys76]. hyperpolarizabilities [BS85].

I. [Men84, Rabxx]. IBM [Joh59]. Icons [Hec08]. ideal [Opp59f, She72]. Ignored [Gin54]. Igor [Rhe95]. ihr [DD71]. II [Bad05, Gei93, Gei04, Gol96, HO31b, OW13b, Sag79b, SNNY99]. illustrations [Yox82]. im [DD71, Opp29]. Imaginaries [Sla24a]. imagination [Tho05]. Imagining [RS06]. Impact [Opp27b, Opp35b, LW71]. Impacts [CO32b, Opp28a]. Implications [Sag83]. Implosion [BB21]. Impossible [Opp57m]. Inception [Haw61]. incl [Yox82]. index [Yox82]. indispensable [Nor02, Nor03]. Industrial [Hil15, Usd09]. Industrial-Scale [Usd09]. Infinite [TO48]. Influential [Rog10, Har10]. information [Opp46h]. inner [McM80b]. innocence [FW07]. Inquiry [Ano49b]. Insight [Wei91]. insights [Kel06]. inspiration [Wes14]. Institute [Opp62e, Ano54y]. institutions [Opp63i]. instruments [Oppxx]. Intellect [Hug09, Guy07, Tho06, Guy07, Pol07, Day09a, Wan08]. Intellectual [Hol96, FB69]. Intellectuals [Day09b, Rie69b]. Intensity [Opp26d]. Interaction [Opp30a, OS41]. Interactions [Sch06, Opp58b]. Internal [AO50, Opp41a, Opp41b]. International [BOT^{+}46, Opp46l, Opp46d, Opp46e, O^{+}46, Opp47c, Opp48f, Opp48g, Opp63m, Opp63o, Her05b, Opp57e, She72, Opp47f]. Internationalism [Rhe95]. Interpretation [Opp39a]. interpreting [Ber05]. Intersections [Sch05, Sch06]. Interview [Agr65, Kuh63, Tel98, Mec23, MFO55, Dro23b]. Interviews [DOW84]. intimate [Opp64d]. Introduction [Duf46, Hol98, SFW91]. Invented [CC21]. invention [Hug89, Hug90, Hug04]. inventor [Coh83]. inventory [KHFA67]. Investigated [Erw84]. invitation [Gre99]. Involved [Ano54x]. Inward [Opp58o, Opp58c]. Irony [Str64]. irradiation [HS39]. ISBN [Bro06, Dom13, Edg91, Guy07, Hug07, Hug09, MW09, Pol07, Tho03, Tho11, Und13]. ISBN-13 [Tho11]. issue [Kip87].

Issues [Fra59, Fra76, Ber05]. Italian [Ben14, Cas14, Cos14, Val06a]. Ithaca [Tho03]. IV [HH04a]. Ivan [Ber04a, Ger06, Lan04]. ix [Goo09].

J [Aft03, Ano45, Ano50, Ano67a, Bal91, Bed06, Ber03b, Boy87, Che65a, Cos14, Dom13, Eak05b, Fre05a, Fre05b, Fri67, Gol55, Gro99, Her03, Hij00, Jul06, Kip64, Kip68a, Kip87, Kip05, Kro81, Lin81, Pow05, Pry55, Pun67, RS06, Rob56, Sch82, aS06, Str63a, Str63c, Str65, Tho03, Tho16, TS12, Und13, Wan04, Wea82, Wed67, Wha55, Wol70, Wol05, Yox82, Zie91, Ber04a, Wan06]. J. [All05, AA54b, Amu11, Ano49c, Ano54i, Ano63a, Ano63d, Ano67a, Ano67b, Ano67c, Ano74, Ano78, Ano95a, Anoxx, Ano04a, Ano04b, Ano09, Ano13a, Ban12, Ban16, Bar49, Ben14, Ber82, Ber05, Bet63, Bet67a, BSK67, Bet68, Bet91a, Bet97, BS05b, BS05a, BS09a, BS09b, Bor08, BG09, Cas05b, Cas05a, Che65a, DWS^{+}69, Day15, De 57, De 68, Dir71, Dom08, DPSY64, Eak05a, EPP80, Els81, Els95, Fed78, Feh14, Fla98, GL95a, GL96a, GL97a, GL11a, GNMS54, Ger06, Goo80, Goo81, Goo82, Gow81, Gra02, GDH^{+}04, Gru09, Har12, Har05, Har08, Her04a, Her05b, Hev11, Hof95, Hof96, Hol93, Huf05, Hun09, Hun14, A.I85, Jac81, Kal54, Kal63, Kan85, Kel06, Kem83, Kev05a, Kev05b, Kev13, Kip64, Kip67a, Kip67b, Kip68b, Kip68a, KW69, Kip87, Kip05]. J. [Kle07, Kug53, Kuh63, Lag95, Lan05, Mar97, MM62a, McC06, McF11, McM05, Men03, Mon12, Mor58, Mur08, Nic54, Nue05, O’R95, OBxx, ORSSBO^{+}17, O^{+}01, PC06, Pei55, Pei74, Per95, Pol02, Pow96, Pow05, RB67, Rai02, Rho77, Rho05, Rig95a, RS04, Roy69, San79, Sap05b, Sap05c, SF08, Sch10, Sch01, Sch03b, She82, SW80a, Str06, SMSU10, Sza95, TS00, Tho01, Tho05, Uni54, U.S54, U.S71, Wal13a, Wan11, WS79, Wei86, Wer05, Wil70, Wol05, Wol08, Bra10, Her07a, Wal96]. Jack [Bal91]. James [Opp57d, Opp57f, Opp57h, Con17]. Japan [Ols62b, Opp62f, Ols62a]. Jazz [Ris06]. je [Str98]. Jefferson [Hol86]. Jennet [Wol05]. Jeremy [Ber04a, Lan04, Pow05, Ger06]. Jets [MB21]. Jews [CC56, Hol96]. ji [Opp56m]. jig [Her05b]. Jin [Opp56m]. Jo [Wan06]. Job [Ano54x]. John [Bal91, Wea82, Wol70, Réd05, UKTS68, You05, von96, vN96]. Jonathan [Und13]. Jonathon [Dom13]. Jordan [Var02]. Joseph [Gow81, She82]. journalism [BS07]. journals [Bey49]. Journey [TS01]. Joy [Wei91]. Jr [Hug70]. Judging [Har10]. Juli [GHK^{+}96]. July [Dre07, GHK^{+}96]. June [SZC54, Ano01b]. justice [AA54b, You05].

K-Absorption [HO31b]. Kai [Fre05a, Pow05, Wol05, Bed06, Dro23b]. Kant [Wei86]. Kathlerine [Duf46]. Keepers [Kev03, Van03]. Keeping [Opp60n, Opp59d, Opp59e]. Kept [Bro14]. Kernwaffendetonation [GHK^{+}96]. Kerr [Bad95]. Key [Ken54]. KGB [HKV10]. Khariton [Hol05b]. Killing [McC05]. Kimball [Gow81, Hen81, Hew66, Jor80, Kar81, Kev81, Pei80, She82]. Kipphardt [Gro99, Wol70, Cos14, Gro99, Kip87, Kro81, Sch82]. Kisco [Opp63j]. Klaus [Hof96]. Knew [Cri96]. Knopf [Bed06, Wol05]. know [Str98]. Knowledge [Ano34, Gei93, Gei04, Opp58d, Opp58e, Opp60d, Opp63l, Opp63s, Opp66c, Cas03, Opp58p, Opp59g, Opp59h, Opp60b, Opp63c, Opp63d]. known [Ano95b]. Köhler [Köh76]. Konflikt [SB55a]. Konstruktion [SB55a]. kontinuierlichen [Opp26c, Opp27e]. kontinuierlicher [Opp27d]. Kourtchatov [Rhe95].

L [Pfa84, Wit99]. Lab [Ano93d]. Laboratory [Ano81, Ano93b, DOW84, Gus96, Man80]. Labs [Man74]. Lagrangian [MA21]. Lamont [And66, Hew66]. Language [RS06, Dro23a]. Lansing [And66, Hew66]. Lantern [Oppxx]. L’Arbre [Opp58g]. Larger [Ber04b]. Last [Ano54r, Ano54x]. Later [Ano67a, SHM^{+}21]. laureate [Bro82]. law [Ano34, Gre77, Gre23]. Lawrence [Ano81, Ber03c, Ber03b, DD71, Gow81, Her03, Kev03, Sch03a, She82, Chi68, Dav68, Dav69, DD71, Dav86, Doy02, Her02, Hil15, Hol02a, Hug70, Huf69, Wil69]. layers [Sag79b]. leadership [GL95b, GL96b, GL97b, GL98, GL11b]. Leading [GL95b, GL96b, GL97b, GL98, GL11b, Sei78]. leaflet [Opp56b]. Lecture [Ano54r, Ano54z, Opp47e, Opp57d, Opp57f, Opp57h, Hol86, Opp63e, Opp63h]. Lectures [Opp70, BG09, FOS09, Har09, SR92, Sag79a, Sag79b]. led [Hol02b]. Lee [McC04]. left [Ano49c, Ano50, Inf78b]. legacy [Ano01a, LOD08, TB62]. length [Ano49c]. Leo [O^{+}01, Rai02]. Leona [Hen81]. Leonardo [Rag80]. Leprince [Opp60e]. Leprince-Ringuet [Opp60e]. Leslie [Ano45, Bad95, Law88, Nor02, Nor03, Ber03a]. L’espoir [Opp58h]. lessons [GL95a, GL96a, GL97a, GL11a]. Letter [AO44, BR57, Che54, Che55, Haw94, Lev60, LPT^{+}54, Mac54, Ols62b, Opp49c, Opp54i, Opp62f, Sch04, Wei94a, Dem15, Uni54]. letterature [Cas14]. Letters [And83, Ano80, Cla80, Dys80, Dys83, GH94, Gow81, Hen81, Jor80, Kar81, Kev81, Pei80, Rei80, Sar82, She82, SW80a, SW80b, SW85, SW95, Tay92, Dys18, Réd05, SW80d, SW80e, U.S71]. Level [Guy07, Dom13]. Levels [OS36]. Lewis [Pfa84]. li [Opp56m, Rut27]. Libby [Hen81]. Liberal [Tho02, Tho05]. Liberalism [Sch08a]. Library [Opp62e, Ano74, SMSU10]. liderazgo [GL98]. líderes [GL98]. Lie [Gin54]. Life [Ano54i, Ber78, Ber04b, Bet67b, Dom13, Dys13, Har12, Joh13, Kev13, Kru15, Mas13, Mon13, Opp54i, Pfa84, TS12, Und13, Wit99, Ber87, BO76, BS97, Chi68, Dom08, Isa07, Mon12, Opp62a, PC06, SC98, SK13, Wol08, Kle07, Her07a]. Light [Boy85, GB89, KOS37b, Opp31c, Opp31a, Var02]. Lilienthal [Ano50, Ano54v]. Lillian [Bad95, Gol95a, See95, Per94]. Limitations [Ano34, FO34a]. limits [Cas03]. Line [BS01]. L’intime [Opp64d]. literature [Cas14]. Lithium [Opp33]. Little [Kev03]. live [Opp54d, O^{+}65]. Lives [Ber03c, Her03, Kev03, Opp56f, Sch03a, Ber03b, Doy02, Her02, Hol02a, Hol05b]. Living [Gar08, Che49, Hal97]. Loaded [Mor21]. Logic [Opp28g]. London [Dom13, Goo09, Guy07, Hen81, Hug09, MW09, Tho03, Und13, Yox82]. long [Ano67a, Cow03]. long-ago [Cow03]. Look [Opp58o, Har10, Opp58c]. Looking [Bal91]. Louis [Zie91]. Love [Ano54p, SK13]. Loyalties [Ber03c, Her03, Kev03, Sch03a, Ber03b, Doy02, Her02, Hol02a]. Loyalty [Ano90, Ber90, Bro14, OR00, Ano54t]. lun [Opp56m]. Lunar [Szi71].

M [Bal91, Hor70, Hug70]. machine [McI05]. Machines [Joh59]. MacLeish [GCD10]. Macmillan [Opp28g]. Made [Ano54x, Joh13, Lew21, AZ99, Haw11, Nic87]. Magnetic [CO32b]. Main [Ano34]. Major [AO44]. Majorana [EAW15, Gia85, Var02]. Majority [GM54, SZC54]. Make [Lau65]. Maker [Dys18, She82]. Makers [AG02, Ano52a]. Makes [Ano63b]. Making [Han95, Hec08, Wol09, EOA^{+}08, Koj05, Opp65i, WS88, WS08]. Man [Ano54s, Bor57, Bor63b, Con17, Hol81, McI05, Mor66, Ano04a, Che59, Che69, Dro23c, Hol02b, Nor02, Nor03, Opp66f, Pat05, RTS^{+}69, RTS^{+}xx, Rou64, Rou65b, Opp56d]. Manager [Opp54i]. Mandarin [Opp56m]. Manhattan [AAM21, Arc21, BB97a, Cha21, CF21, Dro23a, Gro62, Gro83, Hal97, Haw61, HTS83, HHW99, HAM^{+}21, Jon85, Kel06, Law88, MFC21, Mas06a, Mas20, Nor02, Nor03, Sla24b, SFAL21, SFW91, Tho17]. Mankind [Ano54p]. Mann [Che69]. Many [Tel55, EOA^{+}08]. March [Ano16, Dre07, Opp49e]. Maria [Goe88]. Marked [Ano67a]. Marks [Aar67]. Marshall [Hen81]. Martin [Bed06, Fre05a, Pow05, Wol05]. Mary [Wan06]. mas [Ano95b]. masks [Bur45]. Mass [Goo09, Hen81, Tho05]. masses [Ort32, Ort57, OM85, Ort93]. Massive [OV39]. Masters [Duf46]. material [Kle07]. Materials [ABB^{+}46]. mathematical [EAW15, Jec14]. Mathematics [Ber03d, RS06]. Matter [Aft03, Ano54b, Ano54c, Ber82, Fla98, GNMS54, Kip67b, Kip68b, Kip87, Kro81, Opp30a, Opp56k, Sch82, Tho03, Wan04, Wei81, Wol70, CH54, DWS^{+}69, Hol93, Kip64, KW69, Man84, MM62a, Nue05, Opp58i, Opp60f, Opp65d, Pei55, Pol02, U.S54, U.S71]. Max [Bor64]. Max-Planck-Medaille [Pun67]. May [GM54]. McMahon [Ano50, Opp49c]. McMillan [Pow05, RS06, Fre05b]. McMullin [Eak05b]. me [Por07]. Mead [Hug70]. Meade [Bad95, Gol95a, Per94]. Meaning [Bey09, Duf46, Goo09, Gor09a, MW09, MW46, MW07, Sch08b]. meanings [Ban16]. Means [Hol84]. meant [She85]. Measurements [MKC^{+}21]. Mechanics [Cob72, Coh78, Opp27c]. Mechanism [AO50, BW39, SB54, SB71, All56]. Medaille [Pun67]. Meeting [Opp47e]. Mein [Che65a]. Melodrama [Str64]. membership [Ber05]. Memoir [Nis92, Kam85, Kil77, Opp63g, Por07, Res91, Res92]. Memoirs [TS01, SSSS95]. Memorial [SWPS67a, Dir71]. memoriam [Pai99]. Memories [Dro23a, McM94]. Men [All56, Kev03, Lau59, LR57, Str63b, Bro56, Bro72, Oli89, Opp66c, SB54, SB71, Van03]. Menace [All56, Bir01, SB54, SB71, Yor75]. menschliche [Opp57k]. Mentes [GL98]. Mercer [FW07]. Merely [Ber01, Ber97]. Mesa [BS97]. Meson [NO42]. Mesons [EFO48, LOW48, Opp47b, RAO^{+}49, Opp66g]. Mesotron [Opp40, Opp41c, Opp41e]. Mesotrons [NO41, OSS40, OS41]. Message [Bor64]. messenger [McC05]. metafora [Cas14]. Metal [HAM^{+}21]. Metallurgy [MFC21]. metals [HS39]. metaphor [Cas14]. method [Ano95b, BKLB92]. Methods [MA21]. metodo [Ano95b]. Mexico [GHK^{+}96, GHK^{+}96, Lau45, Mas06a, Mas20, SBB21]. Michael [Tho16]. Michelmore [Hug70, Bad73]. Microscope [BR85]. Mid [Hol96].

Mid-twentieth-century [Hol96]. Mifflin [Hug70]. migration [FB69]. Milan [Ben14]. Milano [Ben14]. Milestones [Gon96]. Military [Ano93b, Edg91, Hil15, Mur53, Opp51b]. Mind [Opp49f, Opp55g, Opp63q, Rob56]. minds [GL95b, GL96b, GL97b, GL98, GL11b]. Minority [Eva54]. mins [Wal96]. Minutes [Car21a, Car21b]. miracle [Hir80]. miscarriage [AA54b]. Miss [ORSSBO^{+}17]. Mistakes [Ano00b]. mittels [HS39]. Modern [Eak05b, Fre05b, Hol84, Opp28g, Pow05, RS06, Eak05a, McM05, Opp56g, Opp66e, Sap05c, Opp56m]. moderne [Opp58f]. Molecular [Cas07, HO30]. molecules [BO27, BO63, BO98, Röm83]. Molekeln [BO27, BO63]. Möller [Pun67]. moment [Cow03, McI05]. Momentum [Opp31a, Tho04]. mondial [SB55b]. Monk [Dom13, Dys13, Und13, Joh13, Kru15, Mas13, TS12]. monopoly [Gor09b]. Monte [SFAL21]. Month [Ano01b]. Moral [Aft02, Fra01, Lan00, Olw01, Sar01, Sei01, Tho04, Mas06b, Sch00, Zac76]. Morrison [Edg91]. Moscow [Ano94]. Mössbauer [Pun67]. Mössbauer/Max [Pun67]. Mössbauer/Max-Planck-Medaille [Pun67]. Most [Rog10, Har10, Haw11]. Mount [Opp63j]. Mourn [Ano67c]. Movement [Hew66]. Movie [Wal96, Dro23c]. Mrs [Ano50]. Mrs. [RMR^{+}50]. MSS [Goe88]. Multiple [LOW48, NO41]. Multiplicative [CO37]. Multiverse [Opp57f]. Munroe [MB21]. Murrow [Ada54, Ano13a, MFO55]. Museum [DOW84]. Mushrooming [Wei88]. must [Pol19]. Mute [Dev62]. My [Opp54i, Che65a]. Mysteries [Feh14]. mysterious [Dro23b]. mystery [Opp58i, Opp60f]. Myths [Wit99, Uda94, Uda98].

N [Opp64i]. Nachweis [HS39]. Nagasaki [Gol96]. Name [Ano70]. Named [Ano63d, Bet63]. Narrative [Hec10, Com56]. Nation [Opp54i, Wha55]. National [Ano54j, GR63, New77, MM62a]. nations [Opp66b]. NATO [Opp59f]. Natural [MM62a, CC56]. Nature [Day01, Dys79, OS37, Opp64i, Opp65d]. NBC [Opp50d]. near [Ano45, Opp48d]. necessity [Opp59b]. need [Opp59g, Opp59h]. Negative [EFO48]. nel [Cos14]. nelle [Cas14]. Network [Wal96]. Neumann [Gol72, Réd05, UKTS68, vN96]. neutrino [Var02]. Neutron [OS38b, OV39, Coh83, HS39]. Neutronen [HS39]. Neutronics [SFAL21]. Neutrons [CO31, CO32b, CO32a]. Newhouse [Bal91]. News [Ano52a, Ano54l]. News-Makers [Ano52a]. Next [Opp63n]. Neylan [You05]. Nichols [Nic54, Nic23]. Niels [Duf46, Opp62e, Opp63p, Opp63f, Opp63g, Opp64k, OBxx, She72]. Nine [Kev03, Van03]. NM [Ano45]. No [Ano49d, Pfa84, Ano95b]. Nobel [Bro82, Kra17]. Nolan [Mec23]. non [Tak09]. non-Born [Tak09]. nonadiabatic [Röm83]. None [Duf46, MW46, MW07]. Nonlinear [Kap96, Sag79a]. nonspecialist [Guy07]. Norwegian [Lau47d]. Note [Ano13a, EO31, EFO48, KOS37a, KOS37b, Opp30a, Opp31c, Opp31b, Opp35c, Opp35d, OP35, OS37, OS38a, OR47, Opp27a]. Notes [Opp28f, Opp30c, Opp48c, RS06]. Nothing [Mor95]. novel [Che69]. November [Kuh63]. Nuclear [AM86, AAM21, BHT86, Bad05, Bal91, BW39, Cha21, CC21, FBR^{+}41, Gar08, Gra64, Gus96, Gus04, Hec10, Her09, KOS37a, LRA21, Mas06a, Mas20, Mor66, NO33, NO34b, NO34a, OS36, Opp54i, Opp57e, Opp63h, Res54, RAO^{+}49, Rot99, Sag83, SHM^{+}21, SBB21, Sla24a, Tay92, Usd09, Wil70, Bet91b, Bey49, Cah95, CHW91, Ell86, GHK^{+}96, GA71, Kam85, Nic87, Opp49b, Opp64k, Pol19, Syl87, WC84a, Wil75a]. Nuclei [EO31, KOS37b, OS41, Opp31b]. Nucleon [Opp54b]. Nucleon-nucleon [Opp54b]. nucleus [Gam28]. Nuel [Hug70]. Nuell [Wil69]. nullpunkett [Lau47d]. Numbers [Lew19, Lew21]. NY [Opp63j, Wal96]. Nye [Wan06].

O [Gow81, She82]. Obituary [Ano67b, Fri67]. Objector [Rag80]. Obligations [Lau56]. Observations [Ber01]. Obtain [Ano34]. October [Ano81]. odd [IKL99]. Öffentliche [GHK^{+}96]. Office [Bro06, Hug07]. official [Wha55]. Oh [Mil05]. Ohio [Tho17]. old [Opp55d]. Omega [Opp65b]. Once [Gin54]. One [Ano54p, MW46, MW07, SW80b, Wyd84, Ber87, Cas03, Har10, Opp60m, Duf46]. Open [Che54, Che55, LPT^{+}54, Opp49f, Opp55g, Opp63q, Rob56]. opera [AS05]. Opinion [Mur54]. Oppengejmer [Rou63, Rou65a]. Oppenheimer [Aft02, Aft03, Bac72, Bac99, Bad73, Bad77, Bad95, Ber03c, Ber03b, Ber04a, Bet67a, Bet97, Bey09, BS09b, Bra10, Bro06, Che65a, Cos14, DD71, Dys76, Dys83, Dys13, Eak05b, Edg91, Fra01, Fre05b, Ger06, Gol95a, Goo09, Goo82, Gor09a, Gow81, Gro56, Hen81, Her03, Her07a, Hig76, HBP94, Hof95, Hof96, Hor70, Hug70, Hug07, Hug09, Joh13, Jor80, Kar81, Kev81, Kev03, Kip64, Kip05, Kle07, Kru07, Kru15, Lan00, Lin81, MW09, Mas13, Olw01, Pei80, Per94, Pow05, Pun67, RB67, RS06, S.54, Sag79a, Sag79b, San77, Sar01, Sce25, Sch03a, See95, Sei01, She82, Tho03, Wal96, Wan04, Wan06, Wea82, Wig76, Wol05, Woo25, Yox82, Zac76, Zie91, Kro81, Sch82, Ato54, Aar67, Ada54, Agr65, ACU^{+}54, All05, AA54b, Amu11, And83]. Oppenheimer [Ano34, Ano45, Ano49c, Ano49d, Ano50, Ano52b, Ano53a, Ano53b, Ano54f, Ano54h, Ano54i, Ano54j, Ano54k, Ano54m, Ano54n, Ano54l, Ano54t, Ano54o, Ano54q, Ano54r, Ano54p, Ano54s, Ano54u, Ano54w, Ano54y, Ano54z, Ano54-28, Ano63b, Ano63a, Ano63d, Ano67a, Ano67b, Ano67c, Ano70, Ano74, Ano78, Ano80, Ano84, Ano90, Ano93c, Ano94, Ano95b, Ano95a, Anoxx, Ano01b, Ano01a, Ano04a, Ano04b, Ano09, Ano11, Ano13a, Ano13b, Anoxx, AZ99, Ban12, Ban16, Bar49, Bed06, Ben14, Ber82, Ber85, Ber90, Ber97, Ber04a, Ber04c, Ber05, Bet38, Bet63, Bet67b, BSK67, Bet68, Bet91a, BS05b, BS05a, BS05c, BS09a, BS09b, BS85, BD54, Edi54, BKLB92, Bor08, BK68, Bow03, Bow08, BG09, Bre94, Bro82, Bro14, Bro23, Buc63, Can09]. Oppenheimer [Car88, CH05, Cas05b, Cas05a, Cas07, Cas65, Che54, Che55, Che65a, Che65b, Che66, Cla80, CH54, Con05, Coo04, Cou63, Cre05, Cri96, Cro54, Cur55, DOW84, DWS^{+}69, Dav68, Dav69, DD71, DPW^{+}82, Dav86, Day01, Day09b, Day15, De 57, De 68, DM69, Dev62, Dir71, Dom13, Dom04, Dom08, Doy02, Dri89, Dro23c, Dro23b, DPSY64, Dys79, Dys80, Dys06a, Exe54, Eak05a, EPP80, Els81, Els95, EG05, Erw84, Ess77, Eva54, Fed78, Feh14, Fin06, Fla98, Fri67, GDR04, GL95a, GL96a, GL97a, GL11a, GNMS54, Gar54, Ger06, Gia85, Gin54, Gol92a, Gol95b, Gol55, Goo80, Goo81, Goo82, GPDW08, Gow81, Gra02, Gre77, Gre23, GCD10, Gro99, GDH^{+}04, Gru09, Gui04, Har12, Har05, Har08, Haw94, Hec05, Hec08, Hec10, Hei05, Her02, Her04a, Her04b]. Oppenheimer [Her05a, Her05b, Hev11, Hij00, HHMW93, Hof95, Hol02a, Hol02b, Hol93, Hol05b, Hol81, Hol84, Huf05, Hug70, Hun09, Hun14, A.I85, IKL99, Inf78a, Jac81, Jec14, Kai02, Kal54, Kal63, Kam23, Kan85, Kel06, Kem83, Ken67, Ken94, Ken54, Kev03, Kev05a, Kev05b, Kev13, Kip64, KL66, Kip67a, Kip67b, Kip68b, Kip68a, KW69, Kip87, Kip05, Knu83, Kug53, Kuh63, Kun82, Kun15, Lag95, Lak66, Lan05, Lar88, Lau54, Lau56, Lev60, Lif23, Lip54, Mac54, Maj71a, Maj71b, Maj83, Mar97, Mas06b, MM62a, McC05, McC06, McF11, McI05, McM05, McW54, Mea04, Mec23, Men03, MT55, Mic69, Mon12, Mon13, Mor58, Mor66, MS15, Mur53, Mur05, Mur08, MFO55, Nas07, New77, Nic54, Nue05, O’R95, Opp51c, Opp54f, Opp54i, Opp55b, Opp60a, OR69]. Oppenheimer [Opp79c, OBxx, Opp23, ORSSBO^{+}17, O^{+}01, PC06, PST07, Pei55, Pei74, Pei80, Per95, Pol02, Por07, Pow96, Pow05, PGDW08, Put60, Put61, Put63, RSW^{+}69, Rag80, Rai02, Rei80, Res54, Rhe95, Rho77, Rho05, Rie54, Rie69a, Rig87a, Rig95a, Rig95b, Rig00a, RS04, Rob56, Röm83, Rou62, Rou64, Rou65b, Rou73, Roy69, Rum92, S.54, San79, Sap05b, Sap05a, Sap05c, Sar82, SF08, Sch54, Sch10, Sch00, Sch01, Sch03b, Sch05, Sch06, Sch08b, Sei69, SK17, SWPS67a, SWPS67b, STY^{+}98, SNNY99, Sim54, SW80a, SW80b, SW80c, SW80d, SW80e, SW85, SW95, SG69, SG71, Str98, Str63a, SK13, Str63c, Str64, Str65, Str70, Str83, Str06, SMSU10, Sza95, Tak09, Tap01, Tay92, TS00, Tho01, Tho02, Tho05, Tho06, Tri54, Uni54, U.S54, U.S71, Und13, Val06a, Var02]. Oppenheimer [VK65, Wal13a, Wan11, Wed67, WS79, Wei86, Wer05, Wes14, Wha55, Wil94, WC84b, Wil70, Win00, Wol70, Wol05, Wol02, Wol08, Yor76, Yor89, Boy87, Dom13, Fre05a, Pow05, Pry55, Tho11, Tho16, TS12, Wil69, Wol05, Ber04a, Day09a, Gro99, Guy07, Her07b, Huf69, Jul06, Kro81, Lan04, Pol07, Sch82, Set05, aS06, Sla24a, Wan08, Wol70]. Oppenheimer/Unruhe [Pun67]. Oppenheimera [Put63]. Oppie [Dro23a, Lan04, Sch04]. Opportunity [ORSSBO^{+}17]. opposition [Bro82]. oracles [Syl87]. Oral [Wei88, BWMH81]. Order [Lau56, Opp57d, Opp57h, Tho04, Opp47g, Opp57c]. ordinary [Dro23a]. ordre [Opp58h]. O’Regan [Wal96]. Organization [Ano93a, TS00]. Organizing [BB97b]. origin [Rut27]. Original [Ano34]. originally [Bey49]. Origins [Mor21, MA21]. Orlando [Chi68]. Oswald [Veb60]. Other [Pin15, Hol86]. Our [Bai75, Kap96, Man74, Opp56o, Opp64e, Wit99, Bro56, Bro72, Opp56h, Opp61e, Opp20, Uda94, Uda98, Wil70, Wil75a, Wil75b]. ours [Lif23]. Outline [Moo21]. Outpost [Ano93b]. Oversight [Ano93d]. Owens [Ber78]. Oxford [Kle07, RS06, Ano53b]. Oxygen [Opp41b].

P [Gow81, Gro56, HH04b, Hor70, Kle07, Opp28g, She82, Wit99]. Pacific [Ken54, Mac54, Ols62a]. Pact [Knu83]. pages [Bed06, Fra01, Ger06, Goo09, Kle07, Wol05, Wol05]. pains [Pei80]. Pair [NO42, OS39a]. Pairs [Opp35d, Opp41b]. Pais [Kle07, Wol70]. Palace [Kru07, Wol05, Con05, Mur05, Nas07, Sap05a]. Pandora [Kev03, Van03]. panel [GHK^{+}96]. Panos [Zie91]. paper [Cas07, RS06]. Paperback [Bro06, Edg91, Hug07, Tho03]. Papers [Ano74, Ano00a, Goe88, Hug07, Köh76, Sch58, Sch03c, SMSU10, vN96, Eck70, Eck73, Haw11, OBxx, Rabxx, Veb60, von96, Rot99]. Para [Ano95b]. Paradox [Ano54o]. Parallel [Hol05b]. Parsons [Chr98]. Part [SW80b, SW80c, Han95, Yox82, Gol96, HO31b, OW13a, OW13b]. Partial [Ano63b]. Participating [Ano50]. Particle [CW88, DOW84]. Particles [Opp35d, OS37, Opp27a, Opp58b]. Partnership [Gol95b, Kun15]. Parts [Goo05]. Party [Ber05, BS05c]. Pascal [Gol72]. pass [Wil94]. passing [Bro56, Bro72]. Passions [Wei91]. Patents [CC21]. Path [Ano54p, Fer80, Huf05]. patterns [Dys18]. Paul [Bad95, Gol95a, Per94]. Pauli [FW07]. Peace [Bal91, Opp46a, SC98, Fei60a, HH89, Wit99]. Peaceful [Ano54l, Opp56a]. Pegram [Opp63e, Opp63h]. Peierls [Moo21]. Penguin [RS06]. Penneman [Bad95]. pensiero [Rou73]. People [Gus04, Hen81, Lib79, Tel55]. performance [Opp57i]. Performing [Sla24b]. Peril [Hew66]. period [Pai67]. perpetual [Yor75]. Perseus [Rot99]. Persistent [Mur53]. Personal [Ber01, Wit99, Bet91b, Com56, Nic87, Uda94, Uda98]. Personnel [GM54, You05, U.S71, Wha55]. perspective [Bet91b]. Perspectives [Hir88, Opp66e, MP09, Tel87]. Peter [Bad73, Gow81, Hug70, She82, Yox82]. Pharr [Hug70, Wil69]. phenomenological [EAW15]. Philip [Edg91, Hug70, Hor70]. Phillips [Bet38]. Philosophical [Jor27, Ano95b]. Philosophy [Ada54, Mas06b, Opp59j]. Photoeffect [KOS37a]. Photoelectric [HO31a, HO31b, HO31b]. Photosynthesis [Opp41a]. Photosynthetic [AO50]. Phys [Her03]. Phys. [Ber03b]. Physical [Hal09, Opp48d, Oppxx]. Physicist [Ano54r, Ano54-27, Ano67a, Gin54, Gol99, OW13a, OW13b, Wei91, All05, Alv87, Ber87, Dro23a, Yor95]. Physicists [Bad05, Ber88, Day09b, Lau54, Opp56f, Opp64j, Opp69, HH04a, Jon10, Kai05]. Physics [Ano01b, Anoxx, Bir50, Bir65, Hei68, Hol84, Kai02, LL58, Lau56, Opp47d, Opp47e, Opp48h, Opp56n, Opp65f, Opp65g, Opp66f, Sch86, Sei78, Wan06, Yox82, Bey49, BP90, Cas05a, CW88, EAW15, Ess77, FBR^{+}41, GL95a, GL96a, GL97a, GL11a, Haw11, KHFA67, Meh75, Opp56m, Opp62e, Opp66e, Opp71, Opp72, WP85, Dys79, Opp62e, Zie91, Opp28g, Opp65d]. Pi [Ber04a, Ger06]. Piccolo [Ben14]. picture [Goo82]. Pioneer [Ano67a, Mas13, All05]. pioneers [Wil75a]. place [Opp49d]. Plague [Bur45]. Plan [Ber85]. Planck [Pun67]. plasma [Sag79a, Sag79b]. platonic [Cas03]. Play [Buc63, Kip67b, Kip68b, BK68, Kip05]. Playhouse [Tho17]. Plight [Ano54-28]. Pluralistic [Opp57f]. Plus [OS38a, LW71]. Plutonium [MFC21]. Podiumsdiskussion [GHK^{+}96]. poet [Dys06a]. Poison [Ber85, Bur45]. Polarization [Opp27b, Opp54b]. Polenberg [Tho03, Wan04]. policies [CHW91, Nic87, WC84a]. Policy [Mur53, Opp53d, Opp53c, Opp59i, Opp63k, Sag83, DD71]. political [BK68, Bro82, O^{+}01, Rai02, Syl87]. Politics [Ano54-28, Gre67, Gre69, Hab69, Sch08a, Tay92, TS01, Wan06, Ber05, GL95a, GL96a, GL97a, GL11a, Hol93, Inf78b, Kam85, LW71, Opp60j, Sch62, She72]. Politik [DD71]. Polonium [Tho17]. Ponders [Ano53a]. por [Ano95b]. Portrait [Ber04a, Ger06, Her07b, Lan04, Pow05, Set05, Ano49c, Ber04c, Har08]. Portraits [Gus04]. Positive [FO34b, FO34c, OP33, Opp34]. Positives [NO33, NO34b, NO34a]. Positron [FO34a]. Possible [Lew21, Ber05]. Post [Mas06a, Mas20]. Post-Cold [Mas06a, Mas20]. postwar [Tho05]. potere [Val06a]. Potsdam [Fei60a, Fei60b]. Power [Ano53a, Ano16, Hen81, Sla24a, Bet91b, Die54, Die62, OR47, Opp57b, Opp57e, Val06a]. PP [Hen81, Edg91, Hug07, Hug09, MW09, Opp28g, Pei80, Tho03, Und13, Yox82, Bad95, Bro06, Dom13, Duf46, Dys13, Guy07, Hug70, Jor80, Kev03, Pol07]. pp. [Hug70, Lan04, Opp64i, Per94]. practice [Opp48c, Opp57i]. Pragmatism [Zac76]. Preceded [Ken54]. Precision [BB21]. Prelude [Bai75]. Prentice [Ber04a]. Present [Wil70, CHW91, WC84a]. presented [Opp62e]. Presents [Nic54, Nic23]. Presidency [Mer95]. President [Kil77, Ano54v, LPT^{+}54]. Press [Bad95, Ber04a, Edg91, Fra01, Ger06, Goo09, Guy07, Hen81, Hug09, Jor80, Kle07, MW09, Pei80, Per94, Pol07, RS06, Tho03, Wit99, Opp63n, Wit99]. Press/Prentice [Ber04a]. prevent [Pol19]. Price [Opp28g]. primer [CW88, SR92, Ree17]. Prince [Rum92]. Princeton [Fra01, Ano67a, Pai67]. principal [U.S71]. principles [Opp65h]. prípad [Put63]. Priscilla [Eak05b, Fre05b, Pow05, RS06]. Private [Cou63]. prize [Dir71, Ano63d, Ano70, Bet63, Kra17]. Probability [Opp30c]. problem [Opp26a, Opp47a, Opp56c]. problems [Opp48e, Opp51a, Opp61e, Sag79a, Sag79b]. Proceedings [Ano81]. Process [Bet38]. Processes [Tap01]. Produced [Wal96, Wea82]. Production [LOW48, NO33, NO34b, NO34a, NO41, OP33, Opp35d, OSS40, Opp51c]. Productions [Wal96]. profesora [Put63]. professional [Cas07]. Professor [Opp60a, Put63]. Profound [Bet91b]. Program [Gon96, Opp50d]. progress [Opp44]. Project [AAM21, Arc21, BCF^{+}52, BB97a, Cha21, CF21, Dro23a, Ell86, Gro62, Gro83, Hal97, Haw61, HTS83, HHW99, HAM^{+}21, Kel06, Law88, MFC21, Mas06a, Mas20, McC04, McC05, Moo21, Nor02, Nor03, Sla24b, SFAL21, SFW91, Tho17]. Projected [SBB21]. Prometheus [Bed06, Fre05a, Jul06, Pow05, aS06, Wol05, Kev05a, Sap05b, BS05b, BS09a, Fin06, McC06, Str06, Wer05, Dro23b]. prophet [Wal13a]. prophets [Wal13b]. Prospects [Opp55h, Opp64f, Opp64g, Opp64h]. Proteus [Sch01, Sch03b]. Proton [OS38a, OS39a]. Protons [NO41, NNOS37, Opp30b, Opp33]. Proving [Ken54]. Public [Day09b, Duf46, Hec10, Lev60, Opp51d, Sea67, GHK^{+}96, MW46, MW07, Opp55c]. Publications [Yox82]. Pure [Gre67, Mil05]. purpose [Opp65d]. Purposes [Edg91]. pushed [Hol02b]. puzzles [Ber05]. Pyramid [Opp54c].

Quanta [Opp31c, Opp31a]. Quantenmechanik [Opp27c]. Quantentheorie [BO27, BO63, Gam28, Opp27e, Opp27d, Opp26c]. Quantum [Cob72, Gia85, Hei68, Jor27, Opp26d, Opp27b, Opp28c, Opp28d, Opp28a, Opp28b, Opp28e, Opp28f, Opp40, Opp41c, Sch58, Sch03c, TO48, BO27, BO63, BO98, Coh78, Dir71, Gam28, Haw11, KHFA67, Opp26a, Opp26b, Opp27e, Opp27d, Opp27c, Tak09, Var02, Opp26c]. Quest [Lau56, Com56]. Question [Dev62]. Questions [Lau56]. Qui [Dem15]. quien [Ano95b]. Quiet [Mor66]. quotation [Opp56b].

R [Ano13a, Ber03a, Ber04a, Ger06, Hen81, Lan04, Law88, Nor02, Nor03, Pun67, Ano45]. Rabi [Wol70, Day09b, Rabxx, Rig87a, Rig87b, Rig00a, Rig00b]. Rabinowitch [Opp63n]. Race [Ano54l, Duf46, Eak05a, McM05, Sap05c, Fre05b, Pow05, RS06, Eak05b]. Racing [Nor02, Nor03]. radiance [Nel14]. Radiant [Kam85, Mil05]. Radiation [BO46, HO31b, Lau45, MM39, Opp27b, Opp29, Opp39c, SBB21, HO31a]. Radiations [Opp35a]. Radiative [Opp30c]. Radio [BG09]. Radioactive [Ber85, SBB21, Rut27]. Radioactive-Poison [Ber85]. Radiochemistry [HO21]. raison [Opp58f]. Ramsauer [Opp28e]. Ranch [Dud80]. Range [CO31, CO32a, Lau45]. Rate [Opp51c]. rather [Ano95b]. Ray [Dom13, Dys13, Joh13, Kru15, Mas13, OS37, Opp49a, TS12, Opp48a, Und13]. Rays [CO41, LO34, NO33, NO34b, NO34a, Ros64, HO31b, Opp44, Rut27]. Re [Sim54]. Re-Trial [Sim54]. Reach [HO30]. Reaction [BHT86, BO46]. Reactions [OS38a, TO48]. Read [Ano54u]. readership [Dom13]. Reading [Tay92, Tri54]. real [Goo04, GTS^{+}06]. Really [Ano94]. Reappraisal [Dys76]. Reappraising [CH05, Hug07, Bro06]. reason [BG09]. Rebel [BS07, Ano95b, Dys06b]. rebelde [Ano95b]. Receive [Ano63d, Bet63]. receiving [Opp54e]. Recognizing [Ano04b]. Recollections [And83, Ano80, Cla80, Dys80, Hen81, Rei80, Sar82, SW95, Tay92, SW80b, SW80d, SW80e, SW85, She82, Dys83, Gow81, Jor80, Kar81, Kev81, Pei80]. reconnection [Sag79b]. Reconsideration [New77]. Reconsidered [Ano90, Ber90, Ber95]. Reconsidering [Ber03a]. Record [Res54]. recording [KW69]. Recruit [Wil75b]. Red [Gor09b, Kai05, Bir01, Res54]. Redman [Bad95]. reduction [Sce25, Woo25]. Reflected [Wil70]. Reflections [Hug07, Opp59j, Opp60i, Opp61g, Opp61h, Opp62e, Opp62d, Opp71, Opp72, CH05, Inf78b, Mas06b, Bro06]. reformulation [Röm83]. refugee [Jon10]. Register [Ano74]. Regnant [Sch86]. Reins [Gol92b]. Reith [BG09]. relations [Opp57e]. Relativistic [HO31b]. Relativity [Nim11]. Relevant [Gei93, Gei04, Kam23]. remains [Ano45]. remark [Sei69]. Remarks [Opp49a, Opp54c, Opp59k, Opp65h, Opp48a, Opp58a]. Remembering [Coo04, GDH^{+}04]. Remind [Duf46]. reminiscence [Kis80b]. Reminiscences [BHB80, Kis80a, Tel83, Wit99, SC98, Wil75a, Gow81, She82]. renewal [Opp60k, Opp60l, Opp61f]. Repercussions [ABB^{+}46]. Replies [Opp54f, Opp23]. Reply [Opp54i]. Report [ABB^{+}46, BCF^{+}52, Duf46, Eva54, GM54, Opp44, O^{+}46, BCD^{+}52, Kip64, KHFA67, MW46, MW07]. reported [Bey49]. representative [Ris06]. Requests [Gar54]. Research [AAM21, Ano93d, Gei93, Gei04, Opp47c, Opp47f, Sei78, Opp48d, Opp53a]. researchers [DD71]. resembling [You05]. Residents [SBB21]. Resonances [KOS37b, Opp62e, Opp71, Opp72]. Response [Opp50d, Ber03b]. Responsibility [Aft02, Bad05, Fra01, Lan00, Olw01, Sar01, Sei01, Sch00]. resulting [HS39]. Results [Ano34]. retirement [Wei94b]. Retraction [Gin54]. Retrospectives [She85]. Revealed [Usd09]. Reverse [Mur53]. Review [All56, And66, Bad73, Bad77, Bad95, Bed06, Ber03c, Boy87, Bra10, Bro06, Day09a, Dom13, Duf46, Dys76, Dys83, Dys13, Eak05b, Fra01, Fre05a, Fre05b, Gar54, Ger06, Gol95a, Goo09, Gor09a, Gow81, Guy07, Han95, Hen81, Hig76, Hir88, Hor70, Huf69, Hug07, Hug09, Joh13, Jul06, Kev81, Kle07, Lan00, Lan04, MW09, Mas13, Opp28g, Opp63n, Pei80, Per94, Pol07, Pry55, Res54, Rob56, Sar01, Sch03a, Sei01, Set05, aS06, Tho03, Tho11, TS12, Wal96, Wan04, Wan06, Wan08, Wea82, Wei86, Wig76, Wol70, Yox82, Zie91, Aft02, Aft03, Ber03b, Bey09, Her03, Her07a, Her07b, Kar81, Kru07, Kru15, Olw01, Opp64i, San77, See95, Ber04a, Dom13]. Reviews [Bal91, Edg91, Hew66, HBP94, Opp54i, She82, Tho16, Und13, Wit99, Hug70]. Revised [Ano34]. Revisiting [Ree17]. Revoked [Ano01b]. revolt [Ort32, Ort57, OM85, Ort93]. revolution [Opp63i]. Rewriting [Bro82]. Reykjavik [Rho09]. Rezension [Hof96]. rhetoric [Hol93]. rhetorical [Wal13b]. Rhode [Rho09]. Rhodes [Han95]. Richard [Bad95, Bal91, Han95, Opp57j, Wal96, Rho09, Tho03, Wan04]. Richtmeyer [Opp47e]. Richtungsentartung [Opp27c]. Riding [Sch04]. Rigged [Tap01]. right [Ano50]. Ring [Usd09]. Ringuet [Opp60e]. rise [HKV10, Nel14, Sei78]. Risk [Ano54a, Lau54, Rab54, Kun82, SB55b]. Risks [Joi54, SBB21]. risque [SB55b]. Rites [Gus96]. road [Bet91b, Nic87]. Robert [Aft03, All05, AA54b, AO44, Amu11, Ano49c, Ano50, Ano54i, Ano63a, Ano63d, Ano67a, Ano67b, Ano67c, Ano74, Ano78, Ano95a, Anoxx, Ano04a, Ano04b, Ano09, Ano13a, Bad73, Bad95, Ban12, Ban16, Bar49, Bed06, Ben14, Ber82, Ber03c, Ber03b, Ber04a, Ber05, Bet63, Bet67a, BSK67, Bet68, Bet91a, Bet97, BS05b, BS05a, BS09a, BS09b, Bor08, Boy87, Bra10, BG09, Cas05b, Cas05a, Che65a, Cos14, DWS^{+}69, Day15, De 57, De 68, Dir71, Dom13, Dom08, DPSY64, Eak05b, Eak05a, EPP80, Els81, Els95, Fed78, Feh14, Fla98, Fre05a, Fre05b, Fri67, GL95a, GL96a, GL97a, GL11a, GNMS54, Ger06, Gol55, Goo80, Goo81, Goo82, Gow81, Gra02, Gro99, GDH^{+}04, Gru09, Har12, Har05, Har08, Hen81, Her03, Her04a, Her07a, Her05b, Hev11, Hij00, Hof95]. Robert [Hof96, Hol93, Huf05, Hug70, Hun09, Hun14, A.I85, Jac81, Jul06, Kal54, Kal63, Kan85, Kel06, Kem83, Kev03, Kev05a, Kev05b, Kev13, Kip64, Kip67a, Kip67b, Kip68b, Kip68a, KW69, Kip87, Kip05, Kle07, Kro81, Kru07, Kug53, Kuh63, Lag95, Lan05, Lin81, Mar97, MM62a, McC06, McF11, McM05, Men03, Mon12, Mor58, Mur08, Nic54, Nue05, O’R95, ORSSBO^{+}17, O^{+}01, Pei55, Pei74, Per95, Pol02, Pow96, Pow05, Pry55, RB67, Rai02, Rho77, Rho05, Rig95a, RS04, RS06, Rob56, Roy69, San79, Sap05b, Sap05c, SF08, Sch82, Sch10, Sch01, Sch03a, Sch03b, aS06, She82, SW80a, Str63a, Str63c, Str65, Str06, SMSU10, Sza95, TS00, Tho03, Tho05, Tho16, TS12, Uni54, U.S54, U.S71, Und13, Wal96, Wal13a, Wan04, Wan11, Wea82, Wed67]. Robert [WS79, Wei86, Wer05, Wha55, Wil70, Wit99, Wol70, Wol05, Wol08, Yox82, Zie91, And83, Ano54f, Ano80, Bac72, Bac99, Ber97, BS05c, Bro82, Can09, Car88, Cla80, Con05, Cri96, DOW84, Dom04, Doy02, Dri89, Dys80, Fin06, Hec05, Hec08, Hec10, Her02, Her05a, Hol02a, Kev81, Kev03, McC05, Mic69, Mon13, Mur05, Nas07, Opp55b, Pai99, Pei80, Rag80, Rei80, Rhe95, Rig95b, Rou62, Rou63, Rou64, Rou65a, Rou65b, Rum92, Sap05a, Sar82, She82, SW80b, SW80c, SW80d, SW95, SK13, Tay92, Wan06, Dys83, Dys13, Gow81, Joh13, Jor80, Kar81, Kru15, Mas13, Pei80]. Roberts [Wol70]. Robyn [RS06]. Roger [Bad95, Gol95a, Per94]. Role [Ken54, New77, Opp59l, Sce25, Woo25]. Roman [Che69]. Ronald [Bal91]. Roosevelt [Ano50, Ano50, Dem15, RMR^{+}50]. Rosenberg [Usd09]. rotation [Opp26b]. Rough [Mas13, Dee08]. Rough-Edged [Mas13]. Roundup [Ano54l]. Row [Hug70]. Rudolf [Moo21]. Ruin [Eak05b, Fre05b, Pow05, RS06, Eak05a, McM05, Sap05c]. Rules [Opp31a, Opp41d]. Russak [Hen81]. Russell [FW07]. Russian [Gin54]. Rutgers [Wit99]. Rutherford [Opp64i, Wil83, Opp64i].

S [Fra01, Goo09, Kuh63, Lan00, MW09, Sar01, Sei01, Tho11]. S. [Fra01, Sar01]. Sache [Gro99, Kip64, Kip68a, Kip87, Kip05, Kro81, Sch82]. Sacrifice [Pfa84]. Safety [Lew19]. Saga [Sza95]. Saint [Ban12]. Sakharov [Rhe95]. Same [Ano54y]. Sanctifies [Hol84]. say [Wil94]. says [Bro82]. Scale [Usd09]. Scanning [BR85]. Scattering [BO46, LO34, NO42, Opp54b]. scenic [Kip64]. Schauspiel [Kip05]. Scheduling [Tho04]. Schicksal [DD71]. Scholars [Ano34, Nis92, Fra76]. school [Cas05a, Dud80]. Schöpfer [Hof95, Hof96]. Schuster [Hug70, Wol05]. Schweber [Fra01, Goo09, MW09, Sar01, Tho11, Lan00, Sei01]. Science [AG02, Ber01, Bro06, Car88, Cha21, DOW84, Day01, Fin54, Gre67, Gre69, Hab69, Hil15, HBP94, Hol96, Hug07, Kil77, Mor66, MS60, Opp45c, Opp50e, Opp50c, Opp51d, Opp54g, Opp54h, Opp55d, Opp55c, Opp55i, Opp56g, Opp56h, Opp56o, Opp57g, Opp58n, Opp58g, Opp58j, Opp58k, Opp58l, Opp59p, Opp59m, Opp59q, Opp60g, Opp60h, Opp60i, Opp61g, Opp61h, Opp61b, Opp65i, Opp76, Opp80, Opp89, OR00, TS01, Wei88, Zie91, Ano95b, Bet91b, CC56, DD71, Die54, Die62, FW07, HH04a, Hec05, Hol86, Hol98, H^{+}58, Inf78b, Kam85, Koj05, LW71, MM62a, Opp48c, Opp49d, Opp50b, Opp54a, Opp55a, Opp56e, Opp56j, Opp58f, Opp59a, Opp59j, Opp60c, Opp60d, Opp60j, Opp62d, Opp62c, O^{+}01, Rai02]. science [SC98, Str63c, Str65, Val06a, Opp55f, Opp20, Wit99, Gol55, Pry55]. Sciences [Kra17, Opp55h, CC56, Hal09, Men84, Opp59b, Opp64f, Opp64g, Opp64h]. Scientific [Cob72, Hec08, Mea04, Opp46h, Opp47g, Opp61a, Opp63l, Opp63s, ORSSBO^{+}17, Tho02, Bey49, Cas03, Haw11, Hir80, Opp61e, Opp63c, Opp63d, Opp63i, She72, Tho01]. Scientist [Ada54, Ano54a, Ano54v, Ano63b, Hew66, Jor80, Knu83, Lau56, Opp51d, Opp54i, Res54, Ano95b, Bow03, Con17, Dys06a, Dys06b, Opp46i, Opp53b, Rig87b, Rig00b, Sch00, Tho05, Dys79, Fra01, Lan00, Sar01, Sei01, Aft02, Olw01]. Scientists [ACU^{+}54, Ano54w, Ano67c, Ber01, Cou63, DG99, Duf46, GB89, Gol95c, GR63, Hen81, Kev05b, McG02, Rhe95, Rog10, Ste75, Wal13b, Whi58, Cas14, Hal09, Har10, Kil77, Opp57a, GR63, Opp45e, Opp80, AM86]. scienza [Val06a]. scienziati [Cas14]. Scope [Guy07, Dom13]. Scorpions [Ano53a]. screw [Opp46g, Opp46j, Opp07]. Scribner [Hen81]. Seaborg [Wol70]. Searchingly [Mor66]. Seasons [OW13a, OW13b]. seated [Ano49c]. Secondaries [OSS40]. Secrecy [Opp80, Zac76]. Secret [Bro14, Dee08, Fei60b, Kru07, Lau47c, Res54, Wol05, Con05, Dud80, Kai05, Mur05, Nas07, Sap05a, Tho17, WG17]. Secretary [Opp61c, Opp61d]. secrète [SB55b]. Secrets [Rot99, BB97b, Wil94]. sector [Sei78]. Secular [Hol96]. Security [Aft03, Ano54a, Ano90, Ano01b, Ano13b, Ber90, Fed78, GM54, Gre55, Gro56, Hug70, Ken54, Lau54, New77, OR00, Rab54, Res54, Saw54, Tho03, U.S71, Wan04, Ano78, Cur55, Eva54, Hor70, MM62a, Pol02, SG69, SG71, Str63c, Str65, Wha55, Yor75, You05]. Seidel [Bad95]. sein [Che69]. Seiten [Hof96]. Selected [Sch58, Sch03c, Réd05, Bor63a]. Selection [Opp31a, Opp41d]. Selections [AM86]. self [Hol93]. self-defense [Hol93]. Senator [Ano50, Opp49c]. Sense [Ano54o, OMRS84, Cre05, Boy87]. Senselessness [Ano54o]. Sent [Ano54v]. Serber [Wit99, Wol70, Pai99]. serial [Yox82]. Series [Ano54r]. Serve [Ano54y]. service [Hol02b, Sea67]. Set [Ano34, Lau56]. Sets [Ano54p]. Seven [Tel83, Opp63j]. Shadow [Aft02, Fra01, Lan00, Olw01, Sar01, Sei01, Sch00]. Shadows [Bur83]. Shape [Dys13]. shaping [Koj05]. Sharing [Cha96]. Shatterer [Goo81, Gow81, Jac81, She82, Yox82, Goo80]. Sherwin [Bed06, Pow05, Wol05, Fre05a]. Shield [Hir88, Tel87]. shook [Haw11]. Short [Ano93c]. Show [Ada54, Dev62]. Showers [CO37, Opp39a]. shown [Yox82]. Silvan [Goo09, MW09, Tho11]. Simon [Hug70, Wol05]. simple [Wil83]. Sin [Jor80]. Since [Ano63c, Ano67a, Gei93, Gei04, Meh75]. single [Kip87]. Singular [Sag79b]. sino [Ano95b]. Site [Bor08, Lau45]. Sites [Opp57h]. Sixtieth [DPSY64, Opp39b, Opp79b]. Sixty [Gar08]. Slanted [MS97]. slides [Oppxx]. slightly [Ano49c]. Slippery [Shi54]. Slope [Shi54]. Slowly [HO30]. Small [Ben14]. Smith [Dys83, Gow81, Hen81, Jor80, Kar81, Kev81, Pei80, She82, Hew66]. Smyth [Edg91]. Snyder [IKL99]. Social [Bad05, CC56]. society [MP09, Opp46i, Opp53b, Opp56g, Tho05]. sociological [Guy07]. Soft [CO41, OSS40]. Solution [KGJ^{+}21]. Solutions [IKL99]. Solvay [Meh75]. Some [Opp48e, Opp49d, Opp60i, Opp60j, Opp61e, Sag83, Bro56, Bro72].

Something [You05]. Sons [Hen81]. Sorrow [Opp60k, Opp60l, Opp61f]. Sorry [BR57]. Sound [KW69]. Sounders [Yor75]. Sources [KHFA67, Opp49b]. Soviet [Gon96, Her09, Koj05, QJ94, SSSS95, WG17]. spaces [Hal97]. Spanish [Ano95b, GL98, KL66, Lam68, Opp55f]. Speak [Ano54r, Ada54, Dro23a]. Speaking [Opp60m]. speaks [Coh83]. Special [Kil77, SSSS95, GM54, Nim11]. Specific [Bor08]. Spectra [Opp26d, Cas07, Opp26c]. Spectroscopy [MKC^{+}21]. spectrum [Opp27e, Opp27d]. Speech [Opp45e, Opp80, Dir71]. Speeches [Ano81, OW83]. Spektren [Opp27e, Opp27d]. Spencer [Hen81]. Spender [Ber97]. Spies [HKV10, Lan94]. Spin [Opp41e]. Springer [Hof96]. Springer-Verlag [Hof96]. Springs [Opp63j]. Sputnik [Kil77]. spy [QJ94]. spymaster [SSSS95]. St [Wal96]. Stability [OS38b]. Staff [Ano93c]. Stakes [Ano93c]. Stalin [Gor09b]. Standing [WS88, WS08]. Stanford [Edg91]. Stanley [Ber78, Zie91]. Started [Res54]. State [CF21, McG02, Kev05b, ABB^{+}46]. statecraft [Opp58j]. Statement [Ato54, CH54, OBC^{+}49, Bro23]. Statements [Opp49e]. States [Wha55, Opp65j, Cob72, Her09, Sch86]. Statistical [LL58]. Statistics [EO31, Opp31b]. Statues [Ano11]. Stellar [OS38b]. Stephen [Ber97]. Stern [Hor70, Hug70]. Stewart [Wit99]. Still [Men03, Wol03]. Stimson [Opp61c, Opp61d]. Stimulated [EFO48]. stirs [QJ94]. Stood [vH01]. storia [Val06a]. stories [Ano96]. Story [Ano54q, Cro54, Gol95b, Hug70, Mur53, Str83, AA54b, Che65a, Che65b, Che66, Gro62, Gro83, HTS83, Kug53, Lau46, Lau47b, Lau72, Mic69, RTS^{+}69, RTS^{+}xx, Roy69, SK13, Val06a, Hug70, Bad73]. Strahlung [Opp29]. Strange [Kev03]. Strangelove [Goo04, GTS^{+}06]. Strauss [Ano54x, Ano54y, Pfa84]. Street [FW07]. Structure [Opp58n, Opp59p, Wei81, Opp56e, Opp58e, Opp58k, Rut27]. Struggle [Mur53, Wil70]. Strugglers [Ano54p]. Studied [Ano54v]. Studies [Bro06, Hol96, Hug07, Rie69b, Bey49, CH05]. study [Gre77, Gre23, Opp56i, Opp58m]. stuff [Haw11]. style [Cas07]. Subdued [Ols62a, Ols62b, Opp62f]. Subjecting [Lau54]. subversion [Ano95b]. Success [Hol84]. Successful [Ken54]. Sudoplatov [Les94, McM94]. Sul [Ben14, Cos14]. Sum [Goo05]. Sun [Han95, HHW99]. Sunday [Mar97]. Superbomb [Edg91, GB89, Bro82, Yor76, Yor89, Zie91, Zac76, Bad77, Dys76, Hig76, San77, Wig76]. superbombe [SB55b]. Supercomputing [Lew19]. supplemental [Kle07]. Support [Uni54]. survey [H^{+}58]. survivors [Blu23]. Suspended [Res54]. Suspension [Exe54]. suspicion [Men03, Wol02, Wol03]. suspicions [Kai05]. Svedomí [Put63]. Sverdrup [OR00]. Swift [Bad73, Hug70, Mic69]. Sword [Hir88, Tel87]. symbiosis [MS97]. symmetries [Opp65j]. Symmetry [BR21, Opp65h]. Symposium [Ano81, Opp49a, Opp48a]. System [BB21, Gre55, Kra17, AZ99, Cur55, Gro56]. Systems [HAM^{+}21]. szenischer [Kip64]. Szilard [Gol92a, O^{+}01, Rai02].

T [Wol70]. Takes [Gol92b]. taking [Opp56j]. Tales [Bro80, BS97, Lau45]. Talk [Opp57n, Opp59n, Opp63r, Opp63j, Opp62e]. Talks [Mor66]. tall [Ano96]. Taming [MFC21]. Tangled [Ber03c, Cou63, Her03, Kev03, Sch03a, Ber03b, Doy02, Her02, Hol02a]. tanto [Ano95b]. Target [Her09, Mak95, Chr98]. tasks [SSSS95]. Teacher [Lau56, Opp47e]. Teachers [Nis92]. teaching [GL95a, GL96a, GL97a, GL11a]. Team [Eas89]. Tears [Aar67]. Teatro [Ben14]. Technical [Bad95, Gol95a, HBP94, Per94, See95, Bro82, HHMW93]. techniques [SNNY99]. technological [Hir80, Hug89, Hug90, Hug04]. Technology [Bro06, Hug07, LW71, Tel87, Kil77, Hir88]. teflon [Bow03]. Teilchen [Opp27a]. telecast [Ano50]. Television [Yox82]. Tell [Wei94a]. Teller [Bad77, Ber03b, Dys76, Edg91, Hig76, San77, Wig76, Zac76, Zie91, Anoxx, BO76, BP90, Bro82, Cou63, Cow03, Dev62, Doy02, Goo04, Har10, Her02, Hir88, Hol02a, Rhe95, Tel98, Yor76, Yor89, Ber78, Ber03c, Her03, Kev03, Sch03a]. Telling [Str83]. Temper [Sch86]. Ten [Ano63c, Fra76]. tensor [Coh78]. terms [Röm83]. Terrible [Duf46]. Terrorism [BS01]. Test [CC21, Ken54, SHM^{+}21, SBB21, Ano45, Blu23]. testi [Rou73]. Testimony [Ano49b, A.I85, Tri54]. Tests [Lau45, Mac54]. Text [Ano54z, Tay92]. Texts [Opp54i, U.S71]. ThC [LO34]. Theater [Ben14]. Their [HHW99, DD71, Gra02, Ris06]. theorem [Coh78]. Theoretical [Rie69b, Sch86, All05, Cas05a, EAW15, Kai05]. Theoreticians [Ano70]. Theories [NO42, Rou64, Rou65b]. Theory [Ano34, BO46, FO34a, FO34b, FO34c, HO31b, Jor27, Kol07, Opp26d, Opp27b, Opp28c, Opp28d, Opp28a, Opp28b, Opp28e, Opp28f, Opp30a, Opp30b, Opp34, Opp40, Opp41c, Opp50f, Opp57i, Opp75, STY^{+}98, Tap01, TO48, BO27, BO63, BO98, Dir71, Gam28, Opp26c, Opp26a, Opp26b, Opp27e, Opp27d, Opp55e, Opp57l, SNNY99, Var02, HO31b]. There [Bet67b]. therefore [Ano95b]. Thermonuclear [Gon96]. these [Bro56, Bro72]. things [FOS09, Har09]. Thinker [Ano49d]. thinking [Opp56i, Opp58m]. thirteen [Bey49]. Thirty [Car21a, Car21b, Opp66g, Yor75]. Thomas [Hug70, Kuh63]. Thorpe [Guy07, Hug09, Pol07]. Thought [Boy85]. Thoughtful [Mor66]. thoughts [Opp49d, Opp60j]. Threat [Duf46, Fin54, SB55a, Bal91]. Three [BSK67, Oli89, Opp28f, Opp64j, Opp69, BK68]. Tick [Joh13]. Ties [Res54]. Tiger [Bal91]. Time [Ano93c, Ano16, Bai75, Man74, Mor66, Opp49e, Opp61f, Opp63n, Rog10, Tho04, Wil75b, PST07, Ris06, Wil75a]. time-dependent [PST07]. Times [Ano54i, Ber78, Opp56o, BO76, Opp56h, Opp63f, Opp64e, Opp20]. Titanic [Jon10]. Today [Cha21, RMR^{+}50, Opp49d, Opp58l]. Tokyo [Lau45]. told [Gro62, Gro83]. Toll [Lau45]. Tolman [Opp57j]. tomorrow [O^{+}65]. tonight [Opp56n]. Too [Pfa84, GH94, Mor95]. Took [Lau45]. Top [Dee08]. tortured [Wol05]. tower [Ano45]. toy [GDR04]. Tradition [Opp59o, Men84]. tragedian [Guy07]. Tragedy [Bed06, Fre05a, Pow05, Str64, Wil70, Wol05, BS05b, BS09a, Fin06, Jul06, Kev05a, Lif23, McC06, Sap05b, aS06, Str06, Wer05]. Tragic [Day09a, Hug09, Pol07, Wan08, Wit99, Cre05, Guy07, Neu13, Tho06, Uda94, Uda98]. Transcript [Ano54u, Feh14, EG05, U.S71, Wha55]. Transcripts [Bro14]. Transformation [Men84, Tho01]. Transition [Hol84]. Transitions [Opp30c]. Transmission [Cob72]. Transmutation [OP35]. Transmutations [KOS37b]. Trapeze [Opp64j, Opp69]. Traveled [Fei60b]. Treatment [DM69, Jec14]. tree [Opp58p]. Trial [Fla98, Gro56, Lak66, Sim54, Cur55, SG69, SG71, Hug70, Hor70]. Trials [Ano09, Gru09]. Tributes [Aar67, BSK67]. Trinidad [Lam68]. Trinitite [MKC^{+}21]. Trinity [Lin81, Wea82, Ano95a, Bai75, Blu23, BB21, CC21, EPP80, Els81, Els95, EFB02, HO21, Kat21, Kis80b, Lag95, Lam65, Lam66a, Lam68, Lam85, Lam89, Lew21, MB21, Nic87, Opp65a, Pow96, SHM^{+}21, SBB21, And66, Hew66]. Triumph [Bed06, Fre05a, Pow05, Wol05, BS05b, BS09a, Fin06, Jul06, Kev05a, McC06, Sap05b, aS06, Str06, Wer05]. Trouble [She82]. Trouble-Maker [She82]. true [Cas03, Mor95]. Truman [Gor09b, Lef95, Mer95]. truth [Coh83]. truths [MS97]. Tube [Moo21]. Tuesday [Ada54]. Tumult [Hec05]. Tunneling [BR85, Nim11]. Turbulence [Sag79a]. turn [Opp46g, Opp46j, Opp07]. Turned [Ano54-28]. TV [Dev62]. twelve [Wil75a]. Twentieth [Wan06, Hol96, Har10, Hen81, Kev05b, TS01]. Twentieth-Century [TS01]. twenty [LW71]. twenty-five [LW71]. twilight [Dom08, Wol08]. Two [Ano53a, Cou63, Opp30c, SW80c, Wol05, AS05, Opp26a, Ano34].

U [OR00]. U.S. [Lau45]. Udall [Wit99]. UK [Yox82, Moo21]. Ultimate [McI05, Wei81]. ultra [HO31a]. ultra-gamma [HO31a]. Ultragamma [HO31b]. Unbound [Sch03b, Sch01]. Uncommon [OMRS84, Boy87]. Undergraduates [Opp59n, Opp57n]. Understand [Pin15]. Understanding [Opp54g, Opp54h, Opp55i, Opp48b, Opp55f, Opp66f, Pry55, Gol55]. Unified [Kol07]. Unique [Mea04]. Unit [Ano54v]. United [Wha55, Cob72, Her09, Sch86]. unity [Opp59f]. Universe [Lau56, Isa07, She82, Dys79]. Universities [Gei93, Gei04]. University [Bad95, Bro06, Edg91, Fra01, Goo09, Guy07, Hen81, Hug07, Hug09, Jor80, Kle07, MW09, Opp63n, Pei80, Per94, Pol07, RS06, Tho03, Wit99, Fra59, Fra76, San79]. unlikely [Kun15]. Unlocking [Feh14]. unorthodox [Bow03]. Unreason [Opp60n, Opp59d, Opp59e]. Unreliable [Les94]. Unruhe [Pun67]. Unsystematic [Gre55]. until [Haw61]. unwanted [SSSS95]. Upheld [Ano54v]. Uranium [Hen81, Lib79, HS39]. Urans [HS39]. US$14.00 [Bro06]. USA [Opp63j]. USAEC [Opp46h]. Use [Mor57, Sti47, Sti76]. uses [Lau59, Opp56a]. Utopia [Fas93].

Valid [Opp35a]. Validity [BD54]. Value [Opp51b]. Values [Opp59q, Opp57g, Opp57i, Zac76]. VanDeMark [Kev03]. Veblen [Veb60]. verändert [Lam66b]. verdict [Opp54e]. Verhalten [HS39]. Verlag [Hof96]. version [GDR04]. Versions [MM62b]. Versus [Car88, Duf46, Ano00b, Opp57i]. very [Ber97, Dro23b]. Vessels [Mor21]. vibration [Opp26b]. Victim [Ano67a]. Victor [Wol70]. view [Bow03]. viewpoints [Bet91b]. VII [Hof96]. Vinci [Rag80]. virial [Coh78]. Visa [Ano54g]. Vision [Tho16, Day15]. Visit [Mor58]. Vista [McC04, BCF^{+}52, Ell86, McC05]. vita [Rou73]. vivre [O^{+}65]. vocation [Tho01]. Void [Opp89, Zie91]. Vol [Haw61]. Volkoff [Win00]. Vols [Bir50]. Volume [Hug07, Fre79]. Vorgeschichte [GHK^{+}96].

W [Bad95, Gol95a, Hug70, Opp28g, Per94]. War [Bad05, Bal91, Bra10, Gei93, Gei04, Mas06a, Mas20, Mor66, Nis92, Ols62a, Rig95b, Sag83, Uda94, Wan06, Wit99, DD71, Fei60a, HH89, Pol19, SC98, Amu11, Bal91, Can09, Day09b, Gus96, Hec08, Hev11, Hun09, Hun14, Kai05, Koj05, McF11, Rig95a, Ris06, Uda98, Wan11]. Warner [Ano52a]. Warning [Ano54v, Bur45]. warrior [Con17]. Wartime [Man74, Tho04, Kis80a, WS88, WS08]. Was [Ano54v, Bet67b, Her05a, TS00, Ano45, Ano67a, Ano93c, DD71, Her05b, Hol02b, S.54]. Washington [San79]. Wasserstoffbombe [SB55a]. watching [Por07]. Waves [BR21, Sag79a]. Way [Duf46]. weapon [McI05, Opp46f, Opp46g, Opp07]. weaponeers [Wei94b]. Weapons [Bad05, Gar08, Gus96, HO21, Hug70, Opp45c, Opp46c, Opp53d, Opp53c, Opp54i, Opp63k, Wil70, Bet91b, Ell86, Opp45b, Opp64k]. wear [Bur45]. Weart [Hen81]. weekly [Ano50]. Weiner [Gow81, Hen81, Jor80, Kev81, Pei80, Dys83, Kar81, She82]. Weisskopf [Wol70]. well [She85]. Welt [Lam66b]. We’re [BR57, Ano93c, Ano50, Bey49, Nic87]. Werke [Kip87]. West [Amu11, Her09, Hev11, Hun09, Hun14, McF11, Wan11, Bra10]. Westfall [Bad95, Per94, Gol95a]. Where [Bet67b, Lan04, Opp56j, vH01, Whi58]. which [Ano45]. Whitehouse [Mac63]. Whittlesey [Duf46]. Who [Ano54s, CC21, Duf46, TS00, Che59, Che69, Dem15, Hol02b, Dem15]. Widerstreit [DD71]. Wife [Wei94a]. Wigner [Ano52a, HH04b]. Wilfred [BS07]. Will [Ada54, Ano54r]. William [Opp57d, Opp57f, Opp57h]. Wilson [Hug70, DOW84]. Winding [Ano54p]. Winner [Ano70]. Wirkungen [GHK^{+}96]. Wisdom [Ano54-27]. Wissenschaft [DD71, Opp55i]. withdrawn [Kev03]. Without [STY^{+}98, SNNY99]. Witness [Les94, SSSS95]. wizard [LOD08]. Wolfgang [Köh76]. wollte [Che69]. women [HHW99, SK13, WS88, WS08]. Won [Ano54s]. wonderful [Mor95]. Wonders [EOA^{+}08]. Woolwich [MB21]. Work [Moo21, Tel55, Bro82, Bur45, Tho17]. works [Bor63a, Kip87]. World [Ano54p, Ano54-28, Ano67c, Duf46, Dys79, GR63, Kap96, Mea04, Opp48h, RS06, CC56, Haw11, Lam66b, LW71, MW46, MW07, Opp45d, Opp47d, Opp47g, Opp48b, Opp49d, Opp54d, Opp58l, SB55b, Yor75, Bad05, Gei93, Gei04]. world. [Bur45]. Worlds [Ban12, Goo81, Gow81, Jac81, Yox82, Goo80, She82]. Would [Lau65, Che59, Che69]. write [Bur45]. Writer [Gin54]. writings [BS07].

Wrong [GH94]. wu [Opp56m].

x [Duf46, Hen81, HO31b]. X-rays [HO31b]. xi [Hen81]. xii [Bro06, Hug07]. xiii [Bed06, Hen81]. xiv [Edg91, MW09, Opp28g, Und13]. xue [Opp56m]. xv [Per94]. xvi [Bad95, Edg91]. xviii [Fra01]. xx [Hug09, Pol07]. xxii [Kle07]. xxxii [Tho03].

Yates [Ano54x]. Years [Ano63c, Ano84, Bad73, Bad95, Ber88, Bro14, Gar08, Gol95a, HBP94, Hug70, Lau47c, Per94, See95, SHM^{+}21, Bra80, Dom08, HHMW93, Kun82, LW71, Mic69, Opp66g, Ser67, SW80c, Wei67, Wol08, Yor75]. Yield [LRA21, SHM^{+}21]. York [Bad95, Bed06, Duf46, Edg91, Ger06, Hen81, Hig76, Hof96, Opp28g, RS06, Wal96, Wit99, Zie91, Bad77, Dys76, San77, Wig76, Zac76]. Young [Hol81, SW80e, SW85].

zero [Lau46, Lau47d, Lau47a, Lau47b, Lau72]. Zerstreuung [Opp27a]. zum [Pun67]. zur [Opp27a, BO27, BO63, Gam28, Opp27e, Opp27d, Opp27c].

References

Alsop:1954:WA

[AA54a] Joseph Alsop and Stewart Alsop. We accuse! Harper’s Magazine, ??(??):25–45, October 1954. CODEN HAMAA3. ISSN

             1045-7143. URL http://harpers.org/archive/1954/10/we-
             accuse/.

Alsop:1954:WAS

[AA54b] Joseph Alsop and Stewart Alsop. We accuse! The story of the miscarriage of American justice in the case of J. Robert Oppenheimer. Simon and Schuster, New York, NY, USA, 1954. 88 pp. LCCN HD9698.U52 A66.

Andrews:2021:CCM

[AAM21] Stephen A. Andrews, Madison T. Andrews, and Thomas E. Mason. Canadian contributions to the Manhattan Project and early nuclear research. Nuclear Technology, 207(S1):S134–S146, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Aarons:1967:ATM

[Aar67] Leroy F. Aarons. Affection, not tears, marks Oppenheimer tributes. Washington Post, ??(??):??, February 26, 1967. ISSN 0190-8286.

Alvarez:1946:RSD

[ABB^{+}46] L. W. Alvarez, R. F. Bacher, M. Benedict, H. A. Bethe, A. H. Compton, Farrington Daniels, J. R. Oppenheimer, J. R. Ruhoff, G. T. Seaborg, S. H. Spedding, C. A. Thomas, and W. H. Zinn. Repercussions of the State Dep’t report: Denaturing fissionable materials. Bulletin of the Atomic Scientists, 1(9):11, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1954:SAF

[ACU^{+}54] Samuel K. Allison, Edward Condon, Harold C. Urey, Farrington Daniels, F. W. Loomis, Edward Shils, Linus Pauling, S. A. Goudsmit, Julian Schwinger, Albert Einstein, Hugh C. Wolfe, Cyril S. Smith, Victor F. Weisskopf, and Hans Bethe. Scientists affirm faith in Oppenheimer. Bulletin of the Atomic Scientists, 10(5):188–190, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1954:DOS

[Ada54] Val Adams. Dr Oppenheimer to speak on C.B.S.: Scientist will discuss his background, philosophy on Murrow show Tuesday. New York Times, ??(??):20, December 31, 1954. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112948716/.

Aftergood:2002:SBO

[Aft02] Steven Aftergood. In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist (review). Journal of Cold War Studies, 4(1):95–96, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/9218;
http://muse.jhu.edu/article/9218/pdf;   http://muse.
jhu.edu/journals/journal_of_cold_war_studies/v004/4.
1aftergood.html.

Aftergood:2003:MJR

[Aft03] Steven Aftergood. In the Matter of J. Robert Oppenheimer: The Security Clearance Hearing (review). Journal of Cold War Studies, 5(3):122–124, Summer 2003. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:

//muse.jhu.edu/article/45179;   http://muse.jhu.edu/
article/45179/pdf;   http://muse.jhu.edu/journals/journal_
of_cold_war_studies/v005/5.3aftergood.html.

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002. URL http:

//www.gbv.de/dms/goettingen/360423353.pdf;   http://
www.loc.gov/catdir/enhancements/fy0611/2001048396-d.
html.

Agronsky:1965:DOI

[Agr65] Martin Agronsky. Dr. Oppenheimer interview. CBS Evening News with Walter Cronkite (two parts)., August 5–6, 1965.

I:1985:TJR

[A.I85] A.I. ‘The testimony of J. Robert Oppenheimer’. Michigan Quarterly Review, 24(2):368–370, ???? 1985. ISSN 0026-2420.

Allan:1956:RBH

[All56] H. R. Allan. Review of The Hydrogen Bomb: The Men, The Menace, the Mechanism. International Affairs (Royal Institute of International Affairs 1944–), 32(1):132–133, ???? 1956. ISSN 0020-5850 (print), 1468-2346 (electronic). URL http://www.

jstor.org/stable/2607946.

Allman:2005:JRO

[All05] Toney Allman. J. Robert Oppenheimer: theoretical physicist, atomic pioneer. Giants of science. Blackbirch Press, Detroit, MI, USA, 2005. ISBN 1-56711-889-5 (hardcover). 64 pp. LCCN QC16.O62 A77 2005.

Alvarez:1987:AAP

[Alv87] Luis W. Alvarez. Alvarez: adventures of a physicist. The Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-00115-7. xii + 292 + 8 pp. LCCN QC774.A49 A3 1987.

Ackland:1986:ANA

[AM86] Len Ackland and Steven McGuire. Assessing the Nuclear Age: Selections from the Bulletin of the Atomic Scientists. Educational Foundation for Nuclear Science, Chicago, IL, USA, 1986. ISBN 0-941682-07-2, 0-941682-08-0 (paperback). xvii + 382 pp. LCCN U264 .A79 1986.

Amundson:2011:JRO

[Amu11] Michael A. Amundson. J. Robert Oppenheimer, the Cold War, and the Atomic West. Western Historical Quarterly, 42(1):103– 104, Spring 2011. ISSN 0043-3810.

Anderson:1966:RBT

[And66] Oscar E. Anderson, Jr. Review of Day of Trinity, by Lansing Lamont. The American Historical Review, 71(3):1100–1101, April 1966. ISSN 0002-8762 (print), 1937-5239 (electronic). See [Lam65].

Anderson:1983:ORL

[And83] D. L. Anderson. Robert Oppenheimer, letters and recollections. American Journal of Physics, 51(5):477–478, ???? 1983. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Anonymous:1934:ETD

[Ano34] Anonymous. Electron theory of Dirac revised: California scholars dispose of main difficulties but obtain original results. Extend Heisenberg law. Two limitations to experimental knowledge set up by Oppenheimer and Furry. Electron theory of Dirac revised. New York Times, ??(??):n1, February 18, 1934. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1945:GLG

[Ano45] Anonymous. Gen. Leslie Groves (r) and Dr. J. R. Oppenheimer examine the remains of the tower from which a test atomic bomb was detonated near Alamogordo, NM, 1945. URL http://hdl.

loc.gov/loc.pnp/cph.3c2876. Associated Press photograph

number 908D.

Anonymous:1949:E

[Ano49a] Anonymous. Erratum. Bulletin of the Atomic Scientists, 5(4): 128, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:GIT

[Ano49b] Anonymous. The great inquiry: Testimony at AEC hearings. Bulletin of the Atomic Scientists, 5(8–9):221–250, 254, August/ September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This report is a condensation of the 2000-page transcript.

Anonymous:1949:JRO

[Ano49c] Anonymous. J. Robert Oppenheimer, half-length portrait, seated, facing slightly left, May 4, 1949. URL http://hdl.loc. gov/loc.pnp/cph.3c3797. Photograph from Lord and Taylor.

Anonymous:1949:ONT

[Ano49d] Anonymous. Oppenheimer: No. 1 thinker on atomic energy. Life Magazine, 27(15):c1, October 10, 1949. URL https://books.

google.com/books?id=GVIEAAAAMBAJ. Cover photograph and

story.

Anonymous:1950:FTE

[Ano50] Anonymous. First telecast of Eleanor Roosevelt’s weekly forum on February 11, 1950. Participating in a discussion of “what to do with the hydrogen bomb” were, left to right: Senator Brien McMahon, Hans A. Bethe, Mrs. Roosevelt, David E. Lilienthal and J. Robert Oppenheimer, February 11, 1950. URL http:/

/hdl.loc.gov/loc.pnp/cph.3b3842. Photograph by Leonard

McCombe.

Anonymous:1952:NMW

[Ano52a] Anonymous. News-makers: Warner, Wigner, and Fisk appointed to AEC Committee. Chemical & Engineering News Archive, 30(41):4303–4307, 1952. CODEN ???? ISSN

????  URL http://pubs.acs.org/doi/abs/10.1021/cen-
v030n041.p4303.

Anonymous:1952:OBA

[Ano52b] Anonymous. Oppenheimer and Bush appointed to new Advisory Group on Disarmament. Bulletin of the Atomic Scientists, 8(5):

133, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:ABP

[Ano53a] Anonymous. The atom and the balance of power: Dr. Oppenheimer ponders over ‘two scorpions in a bottle’. Fortune, 48(2): 97, 202, August 1953. CODEN FORTAP. ISSN 0015-8259.

Anonymous:1953:OHD

[Ano53b] Anonymous. Oxford honors Dr. Oppenheimer. New York Times, ??(??):10, December 13, 1953. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/112564528/.

Anonymous:1954:ASD

[Ano54a] Anonymous. American scientist declared security risk. Bulletin of the Atomic Scientists, 10(7):243, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:AMC

[Ano54b] Anonymous. The atom: A matter of character. Time, ??(??): ??, June 14, 1954. CODEN TYMEA9. ISSN 0040-781x (print),

2169-1665 (electronic). URL http://content.time.com/time/
magazine/article/0,9171,935123,00.html.

Anonymous:1954:AME

[Ano54c] Anonymous. The atom: A matter of energy. Time, ??(??): ??, September 21, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://content.time.

com/time/magazine/article/0,9171,890631,00.html.

Anonymous:1954:AFH

[Ano54d] Anonymous. The atom: Five hundred Hiroshimas. Time, ?? (??):??, March 22, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://content.time.

com/time/magazine/article/0,9171,819550,00.html.

Anonymous:1954:AHB

[Ano54e] Anonymous. The atom: The H-bomb delay. Time, ??(??):??, November 8, 1954. CODEN TYMEA9. ISSN 0040-781x (print),

2169-1665 (electronic). URL http://content.time.com/time/
magazine/article/0,9171,857625,00.html.

Anonymous:1954:CRO

[Ano54f] Anonymous. The case of Robert Oppenheimer. New Republic, ??(??):8–14, April 26, 1954. ISSN 0028-6583.

Anonymous:1954:DDV

[Ano54g] Anonymous. Dirac denied visa. Physics Today, 7(7):7, July 1954. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1954:EBO

[Ano54h] Anonymous. Einstein backs Oppenheimer. New York Times, ??(??):20–??, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/113110447.

Anonymous:1954:JRO

[Ano54i] Anonymous. J. Robert Oppenheimer: His life and times. Time, ??(??):19–21, April 26, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic).

Anonymous:1954:NAO

[Ano54j] Anonymous. National affairs: The Oppenheimer case. Time, ??(??):??, June 28, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://content.time.

com/time/magazine/article/0,9171,860874,00.html.

Anonymous:1954:NDO

[Ano54k] Anonymous. New debate on the Oppenheimer case. U.S. news and world report, ??(??):86–103, December 24, 1954. CODEN XNWRAV. ISSN 0041-5537.

Anonymous:1954:OH

[Ano54l] Anonymous. News roundup: The Oppenheimer hearings; the arms race; the peaceful atom; continental defense. Bulletin of the Atomic Scientists, 10(6):234–237, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:OCa

[Ano54m] Anonymous. The Oppenheimer case. Newsweek, ??(??):28, April 19, 1954. ISSN 0028-9604.

Anonymous:1954:OCb

[Ano54n] Anonymous. The Oppenheimer case. Nature, 174(4430):569– 573, September 25, 1954. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v174/n4430/pdf/174569a0.pdf.

Anonymous:1954:OPS

[Ano54o] Anonymous. The Oppenheimer paradox: Sense and senselessness. Newsweek, ??(??):28–29, April 26, 1954. ISSN 0028-9604.

Anonymous:1954:OSPb

[Ano54p] Anonymous. Oppenheimer sets path for mankind: Winding up Columbia fete, he asks world strugglers to ‘love one another’. New York Times, ??(??):1, December 27, 1954. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/113003623/.

Anonymous:1954:OS

[Ano54q] Anonymous. The Oppenheimer story. Time, ??(??):69, April 26, 1954. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic).

Anonymous:1954:OSPa

[Ano54r] Anonymous. Oppenheimer to speak: Physicist will broadcast last lecture in Columbia series. New York Times, ??(??):12, December 25, 1954. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/112952570/.

Anonymous:1954:OWG

[Ano54s] Anonymous. Oppenheimer won greatest fame as ‘man who built the A-Bomb’. New York Times, ??(??):16–??, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
docview/113119449.

Anonymous:1954:OL

[Ano54t] Anonymous. [Oppenheimer’s loyalty]. Bulletin of the Atomic Scientists, 10(5):185, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:ROT

[Ano54u] Anonymous. Read the Oppenheimer transcript. Bulletin of the Atomic Scientists, 10(7):258, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:SUD

[Ano54v] Anonymous. Scientist upheld despite F.B.I. file: Warning to Lilienthal in 1947 was studied by atom unit and sent to President. New York Times, ??(??):19–??, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/docview/113116261.

Anonymous:1954:SEC

[Ano54w] Anonymous. Scientists express confidence in Oppenheimer. Bulletin of the Atomic Scientists, 10(7):283, 286, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:SAC

[Ano54x] Anonymous. Strauss asserts A.E.C. job is last: Involved in controversy over Dixon–Yates, he feels he has made first enemies. New York Times, ??(??), December 27, 1954. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/113020693/.

Anonymous:1954:SOS

[Ano54y] Anonymous. Strauss, Oppenheimer serve same Institute. New York Times, ??(??):21–??, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

search.proquest.com/docview/113123872.

Anonymous:1954:TOL

[Ano54z] Anonymous. Text of Oppenheimer lecture ending the Columbia Bicentenary. New York Times, ??(??):10, December 27, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/113033005/.

Anonymous:1954:WP

[Ano54-27] Anonymous. The wisdom of a physicist. New York Times, ?? (??):22, December 28, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/112914520/.

Anonymous:1954:WPT

[Ano54-28] Anonymous. World’s ‘plight’ turned Oppenheimer to politics. New York Times, ??(??):17, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://

search.proquest.com/docview/113111675.

Anonymous:1963:FAJ

[Ano63a] Anonymous. The [1963] Fermi Award [to J. Robert Oppenheimer]. Physics Today, 16(6):21–23, June 1963. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1963:AMP

[Ano63b] Anonymous. Award makes partial amends to atom scientist Oppenheimer. National Observer, ??(??):??, April 8, 1963. ISSN 0027-9803.

Anonymous:1963:EST

[Ano63c] Anonymous. Editorial: Since ten years. Bulletin of the Atomic Scientists, 19(5):2, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Editorial in support of J. Robert Oppenheimer, written before the announcement of the awarding to hime by the Atomic Energy Commission of the Enrico Fermi Prize.

Anonymous:1963:FPJ

[Ano63d] Anonymous. Fermi Prize: J. Robert Oppenheimer named to receive annual AEC Award. Science, 140(3563):161–163, April 12, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1710958.

Anonymous:1967:JROa

[Ano67a] Anonymous. J. Robert Oppenheimer, atom bomb pioneer, dies: Physicist cancer victim censured and later honored by the A.E.C. Dr. J. Robert Oppenheimer, ‘father of the atomic bomb,’ dies in Princeton. Famed physicist was long ailing. Career had been marked with controversy since hearings of 1953. New York Times, ??(??):1, 66, February 19, 1967. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
results/C67ED84E75BB4B2DPQ.

Anonymous:1967:JROb

[Ano67b] Anonymous. [Obituary: J. Robert Oppenheimer]. The Times [London, UK], ??(??):14–??, February 20, 1967. ISSN 0140-0460, 0956-1382.

Anonymous:1967:WSM

[Ano67c] Anonymous. World scientists mourn death of J. Robert Oppenheimer. Washington Post, ??(??), February 20, 1967. ISSN 0190-8286.

Anonymous:1970:TND

[Ano70] Anonymous. Theoreticians name Dyson as winner of Oppenheimer Prize. Physics Today, 23(3):97, 1970. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/23/97/1.

Anonymous:1974:JRO

[Ano74] Anonymous. J. Robert Oppenheimer: A register of his papers in the Library of Congress. Library of Congress, Washington, DC, USA., 1974.

Anonymous:1978:JRO

[Ano78] Anonymous. J. Robert Oppenheimer FBI security file. Scholarly Resources, Wilmington, DE, USA, 1978. LCCN QC16.O62; Microfilm 22,102. 4 microfilm reels.

Anonymous:1980:ORL

[Ano80] Anonymous. Robert Oppenheimer — letters and recollections. Science, Technology and Human Values, ??(32):58, ???? 1980. ISSN 0162-2439 (print),1552-8251 (electronic).

Anonymous:1981:PCA

[Ano81] Anonymous, editor. Proceedings of the Celebration of the 50th Anniversary of the Lawrence Berkeley Laboratory: Symposium and Banquet Speeches October 1981. Lawrence Berkeley Laboratory, National Technical Information Service U.S. Department of Commerce, Springfield, VA 22161, USA, 1981. LCCN ????

URL http://escholarship.org/uc/item/9772r6gq. Report

number LBL-13613.

Anonymous:1984:OY

[Ano84] Anonymous. The Oppenheimer years: 1943–1945. Los Alamos Science, 8:6–25, Winter /Spring 1984. CODEN LASCDI. ISSN

0273-7116. URL http://www.lanl.gov/history/people/
pdf/oppenheimer1.pdf.

Anonymous:1990:OLS

[Ano90] Anonymous. The Oppenheimer loyalty–security case reconsidered. Stanford Law Review, 42(??):1383–1484, July 1990. CO-DEN ???? ISSN 0038-9765 (print), 1939-8581 (electronic).

Anonymous:1993:AAH

[Ano93a] Anonymous. 50th anniversary article: “Here’s Your Damned Organization Chart”. Report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1993. URL http://www.lanl.gov/

history/road/org-chart.shtml.

Anonymous:1993:AAM

[Ano93b] Anonymous. 50th anniversary article: “Military Laboratory” Evolves into Academic Outpost. Report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1993. URL http://www.

lanl.gov/history/road/milit-lab.shtml.

Anonymous:1993:AAOa

[Ano93c] Anonymous. 50th anniversary article: Oppenheimer and his staff arrive: The stakes were high and time was short. Report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1993.

URL http://www.lanl.gov/history/road/oppie_arrives.
shtml.

Anonymous:1993:AAOb

[Ano93d] Anonymous. 50th anniversary article: Oversight Committee formed as Lab begins research. Report, Los Alamos National Laboratory, Los Alamos, NM, USA, 1993. URL http://www.

lanl.gov/history/road/oversight.shtml.

Anonymous:1994:DOR

[Ano94] Anonymous. Did Oppenheimer really help Moscow? Time, 142(21):63–65, May 1994. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic).

Anonymous:1995:DAT

[Ano95a] Anonymous. The day after Trinity: J. Robert Oppenheimer and the atomic bomb. For the record. Voyager, New York, NY, USA, 1995. ISBN 1-55940-685-2. LCCN HAA 0087. 1 CD-ROM.

Anonymous:1995:CCR

[Ano95b] Anonymous. El científico como rebelde — para F. Dyson, quien conocio de cerca a Einstein y a Oppenheimer, la ciencia no es un metodo filosofico sino mas bien una forma de arte y, por tanto, de subversion. (Spanish) [The scientist as rebel — for F. Dyson, having closely known Einstein and Oppenheimer, science is not a philosophical method but rather an art form and therefore subversion]. Quimera, ??(140–141):21–??, 1995. ISSN 0211-3325.

Anonymous:1996:BTF

[Ano96] Anonymous. Behind tall fences: stories and experiences about Los Alamos at its beginning. Los Alamos Historical Society, Los Alamos, NM, USA, 1996. ISBN 0-941232-91-3. xiv + 210 pp. LCCN QC773.3.U5 B43 1996.

Anonymous:19xx:DET

[Anoxx] Anonymous. The debate: Edward Teller and J. Robert Oppenheimer. Report, Los Alamos National Laboratory, Los Alamos, NM, USA, 19xx. URL http://www.lanl.gov/

history/postwar/debate.shtml.

Anonymous:2000:GHB

[Ano00a] Anonymous. Guide to the Hans Bethe papers, ca. 1931–1995. Cornell University Library archives., 2000. URL http://rmc.

library.cornell.edu/EAD/htmldocs/RMA00976.html.

Anonymous:2000:MVC

[Ano00b] Anonymous. Mistakes versus crimes. Physics in Perspective (PIP), 2(1):3–5, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2001:OL

[Ano01a] Anonymous. Oppenheimer’s legacy. Physics World, 14(2):13, February 2001. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://iopscience.iop.org/
article/10.1088/2058-7058/14/2/17.

Anonymous:2001:MPH

[Ano01b] Anonymous. This month in physics history: June 29, 1954: Oppenheimer’s security clearance revoked. APS News, 10(6):

??, June 2001. URL https://www.aps.org/publications/
apsnews/200106/history.cfm.

Anonymous:2004:JRO

[Ano04a] Anonymous. J. Robert Oppenheimer, a man for the ages. Los Alamos News Bulletin, ??(??):22, April 2004. URL http://

www.lanl.gov/orgs/pa/newsbulletin/2004/04/22/text02.

shtml. Today is the 100th birthday of J. Robert Oppenheimer, the Laboratory’s first director.

Anonymous:2004:RJR

[Ano04b] Anonymous. Recognizing J. Robert Oppenheimer. US Congressional Record, 150(89):S7496, June 24, 2004. URL https://

www.congress.gov/congressional-record/2004/6/24/senate-
section/article/s7496-2.

Anonymous:2009:TJR

[Ano09] Anonymous. The trials of J. Robert Oppenheimer. PBS American Experience broadcast (120m)., 2009. URL http://www.

pbs.org/wgbh/americanexperience/features/introduction/
oppenheimer-introduction/;   http://www.shoppbs.org/
product/index.jsp?productId=3444789.

Anonymous:2011:DOG

[Ano11] Anonymous. Dr. Oppenheimer and General Groves statues to be dedicated. Report, Los Alamos National Laboratory, Los Alamos, NM, USA, May 2011. URL http://rosenfest.lanl.

gov/statue%20dedication.shtml. Rosenfest Lectures May

18–20, 2011, Fuller Lodge, Los Alamos, New Mexico USA.

Anonymous:2013:HNE

[Ano13a] Anonymous. Of historical note: Edward R. Murrow with J. Robert Oppenheimer. The Institute Letter, Institute for Advanced Study, ??(??):??, Fall 2013. URL https://www.

ias.edu/ideas/2013/oppenheimer-of-historical-note. In-

cludes 25-minute video of Murrow interview of Oppenheimer.

Anonymous:2013:OSH

[Ano13b] Anonymous. Oppenheimer security hearing. Web document.,

2013.  URL http://www.atomicheritage.org/history/
oppenheimer-security-hearing.

Anonymous:2016:MTA

[Ano16] Anonymous. The march of time: Atomic power! Periscope Film LLC archive, April 17, 2016. URL https://www.youtube.com/

watch?v=iTJJ7tkD5L4.

Anonymous:20xx:EOF

[Anoxx] Anonymous. Einstein, Oppenheimer, Feynman: Physics in the 20th Century. Technical Report MIT OpenCourseWare STS.042J / 8.225J, MIT, Cambridge, MA, USA, 20xx. Free, independent study course that explores the changing roles of physics and physicists during the 20th Century.

Alvarez:1944:LMR

[AO44] Luis Alvarez and J. R. Oppenheimer. Letter to Major Robert Furman. Web document, June 5, 1944. URL http://www.lanl.

gov/history/road/pdf/A%20German%20Atomic%20Bomb.pdf.

Arnold:1950:ICP

[AO50] William Arnold and J. R. Oppenheimer. Internal conversion in the photosynthetic mechanism of blue green algae. Journal of General Physiology, 33(4):423–435, ???? 1950. CODEN JG-PLAD. ISSN 0022-1295 (print), 1540-7748 (electronic).

Archer:2021:ACF

[Arc21] B. J. Archer. The Los Alamos Computing Facility during the Manhattan Project. Nuclear Technology, 207(S1):S190–S203, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Adams:2005:DAO

[AS05] John Adams and Peter Sellars. Doctor Atomic: opera in two acts. Hendon Music, New York, NY, USA, 2005. ISBN 1-4768-1435-X. 1 musical score (556 pages) pp. LCCN M1500.A214 D6 2005.

Settle:2006:BRB

[aS06] Frank a. Settle. Book review: American Prometheus: The triumph and tragedy of J. Robert Oppenheimer. The Journal of Military History, 70(1):205–206, January 2006. ISSN 0899-3718 (print), 1543-7795 (electronic). URL http:

//muse.jhu.edu/journals/journal_of_military_history/
v070/70.1settle.html.

AEC:1954:ASO

[Ato54] Atomic Energy Commission. AEC statement [on Oppenheimer case]. Bulletin of the Atomic Scientists, 10(5):187, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Armour:1999:CBO

[AZ99] E. A. G. Armour and V. Zeman. Corrections to the Born– Oppenheimer approximation as applied to a system made up of hydrogen and antihydrogen. International Journal of Quantum Chemistry, 74(6):645–652, ???? 1999. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=63001781;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=63001781&PLACEBO=IE.pdf. Special Issue: The Role of

Mathematics in Quantum Chemistry: Special Issue in Honor of George G. Hall (Part II of II). Issue Edited by Don Rees, Hiroshi Fujimotoi.

Bacher:1972:RO

[Bac72] Robert F. Bacher. Robert Oppenheimer (1904–1967). Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 116(4):279–293, August 15, 1972. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://www.jstor.org/stable/985898.

Bacher:1999:RO

[Bac99] Robert F. Bacher. Robert Oppenheimer (1904–1967). Los Alamos Historical Society, Los Alamos, NM, USA, 1999. ISBN 0-941232-22-0 (paperback). 55 pp. LCCN QC16.O62 B33 1999.

Badash:1973:BRBb

[Bad73] Lawrence Badash. Book review: The Swift Years. The Robert Oppenheimer Story by Peter Michelmore. Isis, 64(1):138, March 1973. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229903.

Badash:1977:BRBa

[Bad77] Lawrence Badash. Book review: The Advisors. Oppenheimer, Teller, and the Superbomb by Herbert York. Isis, 68(1):165– 167, March 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
230424.

Badash:1995:BRB

[Bad95] Lawrence Badash. Book review: Critical Assembly: A Technical History of Los Alamos during the Oppenheimer Years, 1943–1945. By Lillian Hoddeson, Paul W. Henriksen, Roger A. Meade, Catherine Westfall, [with contributions by] Gordon Baym, Richard Hewlett, Alison Kerr, Robert Penneman, Leslie Redman, and Robert Seidel. (New York: Cambridge University Press, 1993. xvi, 509 pp. $39.95). The Journal of American History, 81(4):1813, 1995. ISSN 0021-8723 (print), 1945-2314 (elec-

tronic). URL http://jah.oxfordjournals.org/content/81/
4/1813.short.

Badash:2005:APN

[Bad05] Lawrence Badash. American physicists, nuclear weapons in World War II, and social responsibility. Physics in Perspective (PIP), 7(2):138–149, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/u2638016081u076h/.

Bainbridge:1975:AOTa

[Bai75] Kenneth T. Bainbridge. “All in our time”: Prelude to Trinity. Bulletin of the Atomic Scientists, 31(4):42–46, April 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baker:1976:ABG

[Bak76] Paul R. Baker, editor. The atomic bomb: the great decision. American problem studies. Dryden Press, Hinsdale, IL, USA, second revised edition, 1976. ISBN 0-03-089873-0. viii + 193 pp. LCCN D842 .B34 1976.

Balogh:1991:BRB

[Bal91] Brian Balogh. Book reviews: Looking the Tiger in the Eye: Confronting the Nuclear Threat. By Carl B. Feldbaum and Ronald J. Bee., Atoms for Peace and War, 1935–1961: Eisenhower and the Atomic Energy Commission. By Richard G. Hewlett and Jack M. Holl, and War and Peace in the Nuclear Age. By John Newhouse. The Journal of American History, 77(4):1419– 1420, 1991. ISSN 0021-8723 (print), 1945-2314 (electronic).

URL http://jah.oxfordjournals.org/content/77/4/1419.
short.

Banco:2012:BJR

[Ban12] Lindsey Michael Banco. The biographies of J. Robert Oppenheimer: Desert saint or destroyer of worlds. Biography (Honolulu), 35(3):492–515, Summer 2012. CODEN BGPYE2. ISSN 0162-4962 (print), 1529-1456 (electronic). URL http://muse.

jhu.edu/journals/biography/v035/35.3.banco.html.

Banco:2016:MJR

[Ban16] Lindsey Michael Banco. The meanings of J. Robert Oppenheimer. The new American canon. University of Iowa Press, Iowa City, IA, USA, 2016. ISBN 1-60938-419-9 (paperback), 1-60938-420-2 (e-book). x + 268 pp. LCCN

QC16.O62 B36 2016. URL http://www.loc.gov/catdir/
enhancements/fy1605/2015033641-b.html;   http://www.
loc.gov/catdir/enhancements/fy1605/2015033641-d.html.

Barnett:1949:JRO

[Bar49] Lincoln Barnett. J. Robert Oppenheimer. Life Magazine, 27 (15):120–125, 131–138, October 10, 1949. URL https://books.

google.com/books?id=GVIEAAAAMBAJ.

Bennis:1997:MP

[BB97a] Warren G. Bennis and Patricia Ward Biederman. The Manhattan Project. In Organizing genius:: the secrets of creative collaboration [BB97b], pages 171–195. ISBN 0-201-33989-7 (paperback). LCCN HD58.9 .B45 1997b.

Bennis:1997:OGS

[BB97b] Warren G. Bennis and Patricia Ward Biederman. Organizing genius:: the secrets of creative collaboration. Perseus Books, Reading, MA, USA, 1997. ISBN 0-201-33989-7 (paperback). xvi + 239 pp. LCCN HD58.9 .B45 1997b.

Brown:2021:THE

[BB21] Eric N. Brown and Dan L. Borovina. The Trinity high-explosive implosion system: The foundation for precision explosive applications. Nuclear Technology, 207(S1):S204–S221, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Buckley:1952:RAG

[BCD^{+}52] Oliver E. Buckley, James B. Conant, Lee A. Dubridge, Willard F. Libby, Eger V. Murphree, John Von Neumann, I. I. Rabi, Walter G. Whitman, and J. R. Oppenheimer. A report by the AEC General Advisory Committee. Physics Today, 5(11): 20–22, November 1952. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bacher:1952:RPV

[BCF^{+}52] Robert F. Bacher, Robert F. Christy, William A. Fowler, Charles C. Lauritsen, J. Robert Oppenheimer, et al. Report on Project Vista. Report ??, California Institute of Technology, Pasadena, CA, USA, February 1952. ???? pp. Little is known about this report on the plans for use of nuclear weapons in Europe in the event of a Soviet invasion. The Air Force strongly opposed the conclusions of this report, and according to [Mon13, page 587], most copies were destroyed on the orders of Thomas Finletter, Secretary of the Air Force, and the remainder were locked up. The Secretary, and General Vandenberg, Air Force Chief of Staff, also ordered that Oppenheimer was henceforth to be denied access to Air Force classified documents, and not to be used as a consultant to the Air Force. According to [McC04, page 340], the two-volume report was partially declassified in 1980, and a redacted version was released. For more on this report, see also [McC05] and [Ell86].

Bludman:1954:VBO

[BD54] S. Bludman and P. B. Daitch. Validity of the Born– Oppenheimer approximation. Physical Review (2), 95(3):823– 830, August 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v95/i3/p823_1.

Bederson:2006:BRK

[Bed06] B. Bederson. Book review: Kai Bird and Martin J. Sherwin, American Prometheus: The Triumph and Tragedy of J. Robert Oppenheimer. New York: Alfred A Knopf, 2005, xiii + 721 pages. $35.00. (cloth). Physics in Perspective (PIP), 8(2):226– 229, May 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bentoglio:2014:SCJ

[Ben14] Alberto Bentoglio. Sul caso di J. Robert Oppenheimer al Piccolo Teatro di Milano. (Italian) [On the case of J. Robert Oppenheimer at the Small Theater of Milan]. In Castellari [Cas14], pages 349–362. ISBN 88-6705-207-1. LCCN ????

Bernstein:1978:BBC

[Ber78] Barton J. Bernstein. Books: Energy and Conflict: The Life and Times of Edward Teller, by Stanley Blumberg and Gwinn Owens. Bulletin of the Atomic Scientists, 34(5):51–53, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BO76].

Bernstein:1982:MJR

[Ber82] Barton J. Bernstein. In the matter of J. Robert Oppenheimer. Historical Studies in the Physical Sciences, 12(2):195–252, ???? 1982. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757496.

Bernstein:1985:ORP

[Ber85] Barton J. Bernstein. Oppenheimer and the radioactive-poison plan. Technology Review, 88(4):14–17, May/June 1985. CODEN TEREAU. ISSN 0040-1692.

Bernstein:1987:LIB

[Ber87] Jeremy Bernstein. The life it brings: one physicist’s beginnings. Ticknor and Fields, New York, NY, USA, 1987. ISBN 0-89919-470-2. xv + 171 + 8 pp. LCCN QC16.B458 A3 1987. US$16.95.

Bernstein:1988:FPB

[Ber88] Barton J. Bernstein. Four physicists and the bomb: The early years, 1945–1950. Historical Studies in the Physical and Biological Sciences, 18(2):231–263, ???? 1988. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic). URL http://

www.jstor.org/stable/27757603.

Bernstein:1990:OLS

[Ber90] B. J. Bernstein. The Oppenheimer loyalty–security case reconsidered. Stanford Law Review, 42(??):1383–1484, July 1990.

Bernstein:1995:ABR

[Ber95] Barton J. Bernstein. The atomic bombings reconsidered. Foreign Affairs, 74(1):135–152, July 1995. CODEN FRNAA3. ISSN

0015-7120. URL http://www.jstor.org/stable/20047025.

Bernstein:1997:MVG

[Ber97] J. Bernstein. The merely very good + Stephen Spender, Robert Oppenheimer. American Scholar, 66(1):31–39, Winter 1997. ISSN 0003-0937 (print), 2162-2892 (electronic).

Bernstein:2001:MPO

[Ber01] Jeremy Bernstein. The Merely Personal: Observations on Science and Scientists. Ivan R. Dee, Chicago, IL, USA, 2001. ISBN 1-56663-344-3. 245 pp. LCCN Q171 .B5375 2001.

Bernstein:2003:RAG

[Ber03a] Barton J. Bernstein. Reconsidering the “Atomic General”: Leslie R. Groves. The Journal of Military History, 67(3):883– 920, July 2003. ISSN 0899-3718 (print), 1543-7795 (electronic).

URL http://www.jstor.org/stable/3397331.

Bernstein:2003:CBR

[Ber03b] J. Bernstein. Comment on book review of “Brotherhood of the bomb: The tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller,” by Gregg Herken [Am. J. Phys. 71(4), 411-415 (2003)] Bernstein’s response ’. American Journal of Physics, 71(7):648, ???? 2003. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bernstein:2003:BRB

[Ber03c] Jeremy Bernstein. Book review: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken. American Journal of Physics, 71(4):411–415, April 2003. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http:

//ajp.aapt.org/resource/1/ajpias/v71/i4/p411_s1. See

comment [Her03].

Bernstein:2003:DWH

[Ber03d] Jeremy Bernstein. The drawing, or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, 2003. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bernstein:2004:EJR

[Ber04a] Barton J. Bernstein. The enigma of J Robert Oppenheimer. [Review of Oppenheimer: Portrait of an Enigma, Jeremy Bernstein, 2004 Gerald Duckworth/Ivan R Dee 240pp, £14.99/$25.00 hardcover, J Robert Oppenheimer: The American Century, David C Cassidy 2004 Pi Press/Prentice Hall 480pp]. Physics World, 12 (12):36–37, December 2004. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://iopscience.

iop.org/pwa/full/pwa-pdf/17/12/phwv17i12a32.pdf.

Bernstein:2004:LTL

[Ber04b] Jeremy Bernstein. Larger than life. Bulletin of the Atomic Scientists, 60(2):35–41, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//bos.sagepub.com/content/60/2/35.abstract;   http:/
/bos.sagepub.com/content/60/2/35.full.pdf+html. Ex-

cerpted from [Ber04c].

Bernstein:2004:OPE

[Ber04c] Jeremy Bernstein. Oppenheimer: portrait of an enigma. Ivan R. Dee, Chicago, IL, USA, 2004. ISBN 1-56663-569-1. xi + 223 pp. LCCN QC16.O62 B43 2004.

Bernstein:2005:PIJ

[Ber05] Barton J. Bernstein. The puzzles of interpreting J. Robert Oppenheimer, his politics, and the issues of his possible Communist Party membership. In Carson and Hollinger [CH05], chapter 5, pages 77–112. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Bethe:1938:OPP

[Bet38] H. A. Bethe. The Oppenheimer–Phillips process. Physical Review, 53(1):39–50, January 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v53/i1/p39_1.

Bethe:1963:FPJ

[Bet63] Hans Bethe. Fermi Prize: J. Robert Oppenheimer named to receive annual AEC Award. Science, 140(3563):161–163, April 12, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bethe:1967:JRO

[Bet67a] Hans A. Bethe. J. Robert Oppenheimer, 1904–1967. Bulletin of the Atomic Scientists, 23(8):3–6, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1967:OWW

[Bet67b] Hans A. Bethe. Oppenheimer: “Where He Was There Was Always Life and Excitement”. Science, 155(3766):1080–1084, March 3, 1967. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/155/3766/1080.full.pdf.

Bethe:1968:JRO

[Bet68] Hans A. Bethe. J. Robert Oppenheimer. 1904–1967. Biographical Memoirs of Fellows of the Royal Society, 14(??):390–416, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769451.

Bethe:1982:CHH

[Bet82] Hans A. Bethe. Comments on the history of the H-bomb. Los Alamos Science, 6:42–53, Fall 1982. CODEN LASCDI.

ISSN 0273-7116. URL http://library.lanl.gov/cgi-bin/

getfile?06-03.pdf. This article corrects a few errors in one book [Goo80, Goo81], and points out numerous errors, fallacies, [SB54]. It also contains a substantial discussion of the role of Edward Teller in the work as Los Alamos, and later, at Lawrence Livermore National Laboratory.

Bethe:1991:JRO

[Bet91a] Hans A. Bethe. J. Robert Oppenheimer. In The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science [Bet91b], pages 221–230. ISBN 0-88318-707-8, 0-671-74012-1. LCCN U264 .B455 1991. US$24.95.

Bethe:1991:RAP

[Bet91b] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Bethe:1997:JRO

[Bet97] Hans A. Bethe. J. Robert Oppenheimer: April 22, 1904– February 18, 1967. Biographical memoirs — National Academy of Sciences of the United States of America, 71:175–218, 1997. CODEN BMNSAC. ISBN 0-309-59031-0. ISSN 0077-2933.

URL http://www.nap.edu/catalog.php?record_id=5737;
http://www.nap.edu/catalog/5737.html.

Beyer:1949:FNP

[Bey49] Robert T. (Robert Thomas) Beyer, editor. Foundations of nuclear physics: facsimiles of thirteen fundamental studies as they were originally reported in the scientific journals. Dover, New York, NY, USA, 1949. 272 pp. LCCN QC173 .B485. See original paper [HS39] and later translation [Gra64].

Beyler:2009:EOM

[Bey09] Richard H. Beyler. Einstein and Oppenheimer: The Meaning of Genius (review). Technology and Culture, 50(3):722– 723, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270547.

Brake:2009:RRR

[BG09] Mark Brake and Martin Griffiths. Radio and reason — the Reith lectures and J. Robert Oppenheimer. Physics Education, 44 (5):474, 2009. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-

6552 (electronic). URL http://iopscience.iop.org/0031-
9120/44/5/003.

Badash:1980:RA

[BHB80] Lawrence Badash, Joseph Oakland Hirschfelder, and Herbert P. Broida, editors. Reminiscences of Los Alamos, 1943–1945, volume 5 of Studies in the history of modern science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1980. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). xxi + 188 pp. LCCN QC791.96 .R44.

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

http://www.jstor.org/stable/987181.

Birge:1950:HPD

[Bir50] Raymond T. Birge. History of the Physics Department. Vols. 3–5 (1928–1950). University of California, Berkeley, Berkeley, CA, USA, 1950. ???? pp.

Birge:1965:HPD

[Bir65] Raymond T. Birge. History of the Physics Department. Department of Physics, University of California at Berkeley, Berkeley, CA, USA, 1965. ???? pp. Five typescript volumes.

Birkner:2001:ERM

[Bir01] Michael J. Birkner. Eisenhower and the Red menace. Prologue: the journal of the National Archives, 33(3):??, Fall 2001. CO-

DEN ???? ISSN 0033-1031. URL http://www.archives.gov/
publications/prologue/2001/fall/eisenhower-and-red-menace-
1.html.

Boskin:1968:OAP

[BK68] Joseph Boskin and Fred Krinsky. The Oppenheimer affair: a political play in three acts. Glencoe Press, Beverly Hills, CA, USA, 1968. ISBN ???? 145 pp. LCCN ????

Bohm:1992:BCB

[BKLB92] A. Bohm, B. Kendrick, Mark E. Loewe, and L. J. Boya. The Berry connection and Born–Oppenheimer method. Journal of Mathematical Physics, 33(3):977–989, March 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/
i3/p977_s1.

Blume:2023:CDA

[Blu23] Lesley M. M. Blume. Collateral damage: American civilian survivors of the 1945 Trinity test. Bulletin of the Atomic Scientists, 79(4):232–237, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2223078.

Born:1927:QMG

[BO27] M. Born and R. Oppenheimer. Zur Quantentheorie der Molekeln. (German) [On the quantum theory of molecules]. Annalen der Physik (1900), 389(20):457–484, 1927. ISSN 1521-

3889. URL http://onlinelibrary.wiley.com/doi/10.1002/

andp.19273892002/abstract. See also English translation [BO98].

Bethe:1946:RRE

[BO46] H. A. Bethe and J. Robert Oppenheimer. Reaction of radiation on electron scattering and Heitler’s theory of radiation damping. Physical Review (2), 70(7–8):451–458, October 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v70/i7-8/p451_1.

Born:1963:QMGb

[BO63] Max Born and J. Robert Oppenheimer. Zur Quantentheorie der Molekeln. (German) [Toward a quantum theory of molecules]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 61, pages 310–337. LCCN ???? Reprinted from Ann. Physik. 84, 457–484 (1927).

Blumberg:1976:ECL

[BO76] Stanley A. Blumberg and Gwinn Owens. Energy and conflict: the life and times of Edward Teller. Putnam, New York, NY, USA, 1976. ISBN 0-399-11551-X. xvii + 492 + 4 pp. LCCN QC16.T37 B58 1976. US$12.95. URL http://alsos.wlu.edu/

information.aspx?id=367.

Born:1998:QTM

[BO98] M. Born and R. Oppenheimer. On the quantum theory of molecules. Rev. Acad. Colombiana Cienc. Exact. Fís. Natur., 22(84):375–391, 1998. CODEN ???? ISSN 0370-3908. Translated from the 1927 German original [BO27] by Arnulfo Poveda, José Luis Villaveces and Gloria Esperanza Moyano.

Born:1957:MA

[Bor57] Max Born. Man and the atom. Bulletin of the Atomic Scientists, 13(6):186–194, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bor63b]. See note in [BR57].

Born:1963:AAGb

[Bor63a] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:MA

[Bor63b] Max Born. Man and the atom. In Grodzins and Rabinowitch [GR63], pages 590–601. LCCN D842 .B78. Reprint of [Bor57].

Born:1964:MMB

[Bor64] Max Born. Message from Max Born. Reviews of Modern Physics, 36(2):509, ???? 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/rmp.aps.org/abstract/RMP/v36/i2/p509_1. Open letter to

Robert Oppenheimer.

Borgwardt:2008:SSF

[Bor08] Elizabeth Borgwardt. Site specific: the fractured humanity of J. Robert Oppenheimer. Modern Intellectual History: MIH, 5(3): 547–571, November 2008. ISSN 1479-2443 (print), 1479-2451 (electronic).

Barnard:1946:ICA

[BOT^{+}46] Chester I. Barnard, J. Robert Oppenheimer, Charles A. Thomas, Harry A. Winne, and David E. Liienthal. International Control of Atomic Energy. US Department of State, Washington, DC, USA, March 16, 1946. xiii + 61 pp. Publication 2498, prepared for the Secretary of State’s Committee on Atomic Energy by a Board of Consultants.

Bowes:2003:TSU

[Bow03] Stanley A. Bowes. The teflon scientist: an unorthodox view of the Oppenheimer case. Xlibris, Philadelphia, PA, USA, 2003. ISBN 1-4134-4246-3 (hardcover), 1-4134-4245-5 (softcover). 446 pp. LCCN QC16.O62 B64 2003.

Bowman:2008:BBO

[Bow08] Joel M. Bowman. Beyond Born–Oppenheimer. Science, 319 (5859):40–41, January 4, 2008. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/319/5859/40.full.pdf.

Boyer:1985:BEL

[Boy85] Paul S. Boyer. By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age. Pantheon Books, New York, NY, USA, 1985. ISBN 0-394-52878-6, 0-394-74767-4. xx + 440 pp. LCCN E169.12 .B684 1985. US$22.50.

Boyer:1987:BRB

[Boy87] Paul Boyer. Book review: Uncommon Sense, by J. Robert Oppenheimer. Physics Today, 40(9):89–90, September 1987. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Blumberg:1990:ETG

[BP90] Stanley A. Blumberg and Louis G. Panos. Edward Teller: giant of the golden age of physics: a biography. Scribner’s, New York, NY, USA, 1990. ISBN 0-684-19042-7. xiv + 306 + 8 pp. LCCN QC16.T37 B57 1990. US$24.95 (CAN$34.95). URL http://

alsos.wlu.edu/information.aspx?id=47.

Born:1957:LEW

[BR57] Max Born and Eugene I. Rabinowitch. Letter to the Editor: We’re sorry. Bulletin of the Atomic Scientists, 13(7):273, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Binnig:1985:STM

[BR85] Gerd Binnig and Heinrich Rohrer. The scanning tunneling microscope. Scientific American, 253(2):50–56, August 1985. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v253/n2/pdf/scientificamerican0885-50.pdf;
http://www.nobelprize.org/nobel_prizes/physics/laureates/
1986/.

Baty:2021:SBW

[BR21] Roy S. Baty and Scott D. Ramsey. On the symmetry of blast waves. Nuclear Technology, 207(S1):S335–S351, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Bradbury:1980:AFY

[Bra80] Norris Bradbury. Los Alamos: the first 25 years. In Badash et al. [BHB80], page ?? ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Brady:2010:BRB

[Bra10] Kevin M. Brady. Book review: J. Robert Oppenheimer, the Cold War, and the Atomic West. Southwestern Historical Quarterly, 114(2):220–221, October 2010. ISSN 0038-478X (print), 1558-9560 (electronic). URL http://muse.jhu.edu/

journals/southwestern_historical_quarterly/v114/114.
2.brady.html.

Breindel:1994:OF

[Bre94] E. Breindel. The Oppenheimer file. National Review, 46(10): 36–40, ???? 1994. ISSN 0028-0038.

Brown:1956:TMS

[Bro56] John Mason Brown. Through these men: some aspects of our passing history. Harper, New York, NY, USA, 1956. ix + 302

pp. LCCN E835 .B7. URL https://archive.org/stream/
throughthesemens006315mb.

Brown:1972:TMS

[Bro72] John Mason Brown. Through these men: some aspects of our passing history. Essay index reprint series. Books for Libraries Press, Freeport, NY, USA, 1972. ISBN 0-8369-2756-7. ix + 302 pp. LCCN E835 .B7 1972.

Brode:1980:TA

[Bro80] Bernice Brode. Tales of Los Alamos. In Badash et al. [BHB80], pages 133–160. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Broad:1982:RHH

[Bro82] William J. Broad. Rewriting the history of the H-bomb: Nobel laureate Hans Bethe says technical errors by Edward Teller, not political opposition by Robert Oppenheimer, hindered work on the superbomb. Science, 218(4574):769–772, November 19, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1689737;   http://www.sciencemag.org/content/
218/4574/769.extract. See response [TB82].

Brown:2006:BRC

[Bro06] Laurie M. Brown. Book review: Cathryn Carson, David A. Hollinger, Editors, Reappraising Oppenheimer, Centennial Studies and Reflections (2005) Office for History of Science and Technology, University of California, Berkeley ISBN 0-9672617-3-2 (xii + 413 pp., US$14.00 Paperback). Studies in History and Philosophy of Modern Physics, 37(4):745–747, December 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000657. See [CH05].

Broad:2014:TKS

[Bro14] William J. Broad. Transcripts kept secret for 60 years bolster defense of Oppenheimer’s loyalty. New York Times, ??(??):A16–??, October 12, 2014. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
docview/1610482802.

Bronson:2023:BSE

[Bro23] Rachel Bronson. Bulletin statement on the Energy Department’s Oppenheimer decision. Bulletin of the Atomic Scientists, 79 (4):265–266, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2223086.

Bishop:1985:BBO

[BS85] D. M. Bishop and S. A. Solunac. Breakdown of the Born– Oppenheimer approximation in the calculation of electric hyperpolarizabilities. Physical Review Letters, 55(19):1986–1988, November 1985. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://adsabs.

harvard.edu/abs/1985PhRvL..55.1986B.

Brode:1997:TAL

[BS97] Bernice Brode and Barbara Storms, editors. Tales of Los Alamos: life on the Mesa, 1943–1945. Los Alamos Historical Society, Los Alamos, NM, USA, 1997. ISBN 0-941232-17-4 (paperback). v + 157 pp. LCCN QC773.3.U5 B76 1997.

Bird:2001:FLA

[BS01] Kai Bird and Martin Sherwin. The first line against terrorism. Washington Post, ??(??):A35–??, December 12, 2001. ISSN

0190-8286. URL http://search.proquest.com/docview/
409200598/.

Bird:2005:BBJ

[BS05a] K. Bird and M. J. Sherwin. Building the bomb (J. Robert Oppenheimer). Smithsonian, 36(5):88–96, August 2005. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Bird:2005:APT

[BS05b] Kai Bird and Martin J. Sherwin. American Prometheus: the triumph and tragedy of J. Robert Oppenheimer. Alfred A. Knopf, New York, NY, USA, 2005. ISBN 0-375-72626-8, 0-375-41202-6. xiii + 721 + 32 pp. LCCN QC16.O62 B57 2005. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://site.ebrary.com/lib/ucsc/Doc?id=10078784.

Bird:2005:ROC

[BS05c] Kai Bird and Martin J. Sherwin. Robert Oppenheimer and the Communist Party. In Carson and Hollinger [CH05], chapter 4, pages 57–76. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Burchett:2007:RJW

[BS07] George Burchett and Nicholas L. Shimmin, editors. Rebel journalism: the writings of Wilfred Burchett. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-71826-0 (paperback). xxi + 314 pp. LCCN PN5516.B87 A25 2007. URL http://

www.loc.gov/catdir/enhancements/fy0803/2007040856-b.
html;   http://www.loc.gov/catdir/enhancements/fy0803/
2007040856-d.html;   http://www.loc.gov/catdir/enhancements/
fy0803/2007040856-t.html.

Bird:2009:APT

[BS09a] Kai Bird and Martin J. Sherwin. American Prometheus: the triumph and tragedy of J. Robert Oppenheimer. Atlantic, London, UK, 2009. ISBN 1-84354-705-8 (paperback). ???? pp. LCCN

???? URL http://catdir.loc.gov/catdir/enhancements/
fy0619/2004061535-d.html;   http://catdir.loc.gov/catdir/
enhancements/fy0619/2004061535-s.html;   http://catdir.
loc.gov/catdir/toc/fy053/2004061535.html.

Bird:2009:JRO

[BS09b] Kai Bird and Martin J. Sherwin. J. Robert Oppenheimer: die Biographie. (German) [J. Robert Oppenheimer: the biography]. Propyläen, Berlin, Germany, third edition, 2009. ISBN 3-549-07358-5. 672 pp. LCCN ????

Bethe:1967:TTJ

[BSK67] Hans A. Bethe, Henry DeWolfe [i.e., Wolf] Smyth, and George Kennan. Three tributes to J. Robert Oppenheimer. Institute for Advanced Study, Princeton, NJ, USA, 1967. 29 pp. LCCN QC16.O62 B45 1967.

Buckley:1963:OP

[Buc63] William F. Buckley. The Oppenheimer play. National Review, ??(??):??, April 23, 1963. ISSN 0028-0038.

Burchett:1945:AP

[Bur45] Wilfred G. Burchett. The atomic plague: ‘I write this as a warning to the world.’ Doctors fall as they work. poison gas fear: all wear masks. Daily Express (London, UK), ??(??):??, September 2, 1945. URL http:/

/www.factsandopinions.com/dispatches/science/people/
wilfred-burchett-a-journalists-warning-to-the-world/.

Burchett:1983:SH

[Bur83] Wilfred G. Burchett. Shadows of Hiroshima. Verso, London, UK, 1983. ISBN 0-86091-783-5 (paperback), 0-86091-080-6 (hardcover). 123 pp. LCCN D767.25.H6 B86 1983.

Bohr:1939:MNF

[BW39] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v56/
i5/p426_1.

Bethe:1981:HAB

[BWMH81] Hans A. (Hans Albrecht) Bethe, Charles Wiener, Jagdish Mehra, and Lillian Hoddeson. Hans Albrecht Bethe oral histories: 1966–1981, 1981.

Cahn:1995:DND

[Cah95] Robert Cahn. Over the desert, a nuclear dawn. The Christian Science Monitor, ??(??):??, July 13, 1995. ISSN 0882-7729 (print), 1540-4617 (electronic). URL http://search.

proquest.com/docview/405745975.

Canaday:2009:ROA

[Can09] John Canaday. Robert Oppenheimer after the War. Hudson Review, 62(1):64, Spring 2009. ISSN 0018-702X.

Carpinterosantamaria:1988:ORA

[Car88] N. Carpinterosantamaria. Robert Oppenheimer and the atomic bomb — science versus conscience. Arbor — Ciencia, Pensamiento, y Cultura, 131(513):9–26, ???? 1988. ISSN 0210-1963.

Carr:2021:TMBa

[Car21a] Alan B. Carr. Thirty minutes before the dawn. arXiv.org, ??

(??):19, March 9, 2021. URL https://arxiv.org/abs/2103.
05722.

Carr:2021:TMBb

[Car21b] Alan B. Carr. Thirty minutes before the dawn. Nuclear Technology, 207(S1):S1–S23, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Castellani:1965:EO

[Cas65] Leandro Castellani. The enigma of Oppenheimer. RAI, Radiotelevisione italiana, Torino, Italia, 1965. v + 77 pp. LCCN PQ4863.A8268 E55 1965.

Casti:2003:OTP

[Cas03] J. L. Casti. The one true platonic heaven: a scientific fiction on the limits of knowledge. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08547-0 (hardcover), 0-309-09510-7 (paperback). xviii + 160 pp. LCCN Q175 .C4339 2003. URL

http://www.loc.gov/catdir/toc/fy037/2003002279.html;
 http://www.nap.edu/catalog.php?record_id=10533.

Cassidy:2005:TPB

[Cas05a] David C. Cassidy. From theoretical physics to the bomb: J. Robert Oppenheimer and the American school of theoretical physics. In Carson and Hollinger [CH05], chapter 1, pages 13– 30. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Cassidy:2005:JRO

[Cas05b] David C. Cassidy. J. Robert Oppenheimer and the American century. Pi Press, New York, NY, USA, 2005. ISBN 0-13-147996-2. xviii + 462 + 16 pp. LCCN QC16.O62 C37 2005. US$27.95.

Cassidy:2007:OFP

[Cas07] David C. Cassidy. Oppenheimer’s first paper: Molecular band spectra and a professional style. Historical Studies in the Physical and Biological Sciences, 37(2):247–270, March 2007. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Castellari:2014:FMF

[Cas14] Marco Castellari, editor. Formula e metafora: figure di scienziati nelle letterature e culture contemporanee. (Italian) [Formula and metaphor: figures of scientists in the literature and contemporary cultures], volume 8 of Di/segni. Ledizioni, Milano, Italy, 2014. ISBN 88-6705-207-1. LCCN ????

Cohen:1956:JWS

[CC56] Harry Cohen and Itzhak J. Carmin, editors. Jews in the world of science; a biographical dictionary of Jews eminent in the natural and social sciences. Monde Publishers, New York, NY, USA, 1956. xxiv + 264 pp. LCCN Q128 .C65. Contributing editors: J. Robert Oppenheimer, Meyer W. Weisgal and Louis Gershenfeld.

Chadwick:2021:WIT

[CC21] Thomas A. Chadwick and M. B. Chadwick. Who invented the Trinity nuclear test’s Christy Gadget? Patents and evidence from the archives. Nuclear Technology, 207(S1):S356– S373, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Crockett:2021:ESM

[CF21] Scott D. Crockett and Franz J. Freibert. Equation of state: Manhattan Project developments and beyond. Nuclear Technology, 207(S1):S286–S294, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Cole:1954:SOM

[CH54] Sterling Cole and Bourke B. Hickenlooper. Statement [on the Oppenheimer matter]. Bulletin of the Atomic Scientists, 10(5): 179, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carson:2005:ROC

[CH05] Cathryn Carson and David A. Hollinger, editors. Reappraising Oppenheimer: centennial studies and reflections, volume 21 of Berkeley papers in history of science. Office for History of Science and Technology, Berkeley, CA, USA, 2005. ISBN 0-9672617-3-2. xii + 413 pp. LCCN QC16.O62 R47 2005.

Chace:1996:SAB

[Cha96] James Chace. Sharing the atom bomb. Foreign Affairs, 75(1): 129–144, January/February 1996. CODEN FRNAA3. ISSN

0015-7120. URL http://www.jstor.org/stable/20047473.

Chadwick:2021:NSM

[Cha21] Mark B. Chadwick. Nuclear science for the Manhattan Project and comparison to today’s ENDF data. Nuclear Technology, 207(S1):S24–S61, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Chevalier:1949:UL

[Che49] Haakon Chevalier. For us the living. Borzoi books. Alfred A. Knopf, New York, NY, USA, 1949. x + 400 pp. LCCN PZ3.C4288 Fo.

Chevalier:1954:OLO

[Che54] Haakon Chevalier. An open letter to Oppenheimer. France Observateur, ??(??):??, September 20, 1954.

Chevalier:1955:OLO

[Che55] Haakon Chevalier. An open letter to Oppenheimer. Frontier, ?? (??):??, March 1955.

Chevalier:1959:MWW

[Che59] Haakon Chevalier. The man who would be God. Putnam, New York, NY, USA, 1959. 449 pp. LCCN PZ3.C4288 Man.

Chevalier:1965:MFJ

[Che65a] Haakon Chevalier. Mein Fall J. Robert Oppenheimer: die Geschichte einer Freundschaft. (German) [My case: J. Robert Oppenheimer: the story of a friendship]. Rütten und Loening, München, Germany, 1965. 223 pp. LCCN ????

Chevalier:1965:OSF

[Che65b] Haakon Chevalier. Oppenheimer: the story of a friendship. George Braziller, New York, NY, USA, 1965. xvi + 219 pp. LCCN QC16.O62 C5.

Chevalier:1966:OSF

[Che66] Haakon Chevalier. Oppenheimer: the story of a friendship. Andre Deutsch, London, UK, 1966. xvi + 219 pp. LCCN QC16.O62 C5 1966.

Chevalier:1969:MGS

[Che69] Haakon Chevalier. Der Mann, der Gott sein wollte: Roman. (German) [The man who would be God: novel]. Aufbau-Verlag, Berlin, West Germany, second edition, 1969. 548 pp. LCCN ???? Translated by Erika Gröger from the American original, with remarks by Friedrich Baadke.

Childs:1968:AGL

[Chi68] Herbert Childs. An American genius: the life of Ernest Orlando Lawrence. Dutton, New York, NY, USA, 1968. 576 pp. LCCN QC16.L36 C5.

Christman:1998:THD

[Chr98] Albert B. Christman. Target Hiroshima: Deak Parsons and the creation of the atomic bomb. Naval Institute Press, Annapolis, MD, USA, 1998. ISBN 1-55750-120-3. xii + 305 pp. LCCN V63.P35 C47 1998.

Cantelon:1991:AAD

[CHW91] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Clark:1980:ORL

[Cla80] R. W. Clark. Robert Oppenheimer — letters and recollections. New York Times Book Review, 85(19):9, ???? 1980. ISSN 0028-7806.

Carlson:1931:RFE

[CO31] J. F. Carlson and J. R. Oppenheimer. On the range of fast electrons and neutrons. Physical Review (2), 38(9):1787–1788, November 1, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v38/i9/p1787_1.

Carlson:1932:RFE

[CO32a] J. F. Carlson and J. R. Oppenheimer. 35. on the range of fast electrons and neutrons [abstract]. Physical Review (2), 35(5): 864–865, March 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v39/i5/p854_1. In Minutes of the New Orleans

Meeting, December 29–30, 1931.

Carlson:1932:IFE

[CO32b] J. F. Carlson and J. R. Oppenheimer. The impacts of fast electrons and magnetic neutrons. Physical Review (2), 41(6):763– 792, September 15, 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v41/i6/p763_1.

Carlson:1937:MS

[CO37] J. F. Carlson and J. R. Oppenheimer. On multiplicative showers. Physical Review (2), 51(4):220–231, February 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v51/i4/p220_1.

Christy:1941:HES

[CO41] R. F. Christy and J. R. Oppenheimer. 7. the high energy soft component of cosmic rays. Physical Review (2), 60(2):159, July 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v60/

i2/p158_1. In Minutes of the Pasadena, California, Meeting June 18–20, 1941.

Coben:1972:SET

[Cob72] Stanley Coben. The scientific establishment and the transmission of quantum mechanics to the United States, 1919–32. The American Historical Review, 76(2):442–466, April 1972. ISSN 0002-8762 (print), 1937-5239 (electronic). URL http://www.

jstor.org/stable/1858707.

Cohen:1978:TVT

[Coh78] Leon Cohen. The tensor virial theorem in quantum mechanics. Journal of Mathematical Physics, 19(9):1838–1840, September 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v19/i9/p1838_s1.

Cohen:1983:TAN

[Coh83] S. T. Cohen. The truth about the neutron bomb: the inventor of the bomb speaks out. Morrow, New York, NY, USA, 1983. ISBN 0-688-01646-4. 226 pp. LCCN UG1282.N48 C64 1983.

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Conant:2005:EPR

[Con05] Jennet Conant. 109 East Palace: Robert Oppenheimer and the secret city of Los Alamos. Simon and Schuster, New York, NY, USA, 2005. ISBN 0-7432-5007-9. xviii + 425 + 16 pp. LCCN QC773.A1 C66 2005.

Conant:2017:MHJ

[Con17] Jennet Conant. Man of the hour: James B. Conant, warrior scientist. Simon and Schuster, New York, NY, USA, 2017. ISBN 1-4767-3088-1 (hardcover), 1-4767-3091-1 (paperback), 1-4767-3092-X (e-book). x + 587 + 16 pp. LCCN CT275.C757 C66 2017.

Cook:2004:RO

[Coo04] Dana Cook. Remembering Oppenheimer. Bulletin of the Atomic Scientists, 60(2):37–41, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Costazza:2014:DDN

[Cos14] Alessandro Costazza. La dialettica dell’illuminismo nel dramma “Sul caso di J. Robert Oppenheimer” di Heinar Kipphardt. (Italian) [The dialectics of enlightenment in the drama “On the Case of J. Robert Oppenheimer” by Heinar Kipphardt]. In Castellari [Cas14], pages 329–348. ISBN 88-6705-207-1. LCCN ????

Coughlan:1963:OTT

[Cou63] Robert Coughlan. Oppenheimer and Teller: The tangled drama and private hells of two famous scientists. Life, ??(??):87A–110,

December 13, 1963. URL http://www.oldlife.net/search.
php?pndate=1963-12-13.

Cowen:2003:LAM

[Cow03] Robert C. Cowen. A long-ago moment with Edward Teller. The Christian Science Monitor, ??(??):11–??, September 12, 2003. ISSN 0882-7729 (print), 1540-4617 (electronic). URL http://

search.proquest.com/docview/405679692.

Crease:2005:OST

[Cre05] Robert P. Crease. Oppenheimer and the sense of the tragic. In Carson and Hollinger [CH05], chapter 15, pages 315–324. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Critchfield:1996:ROK

[Cri96] Charles L. Critchfield. The Robert Oppenheimer I knew. In Behind tall fences: stories and experiences about Los Alamos at its beginning [Ano96], pages 169–177. ISBN 0-941232-91-3. LCCN QC773.3.U5 B43 1996.

Crossman:1954:OS

[Cro54] R. H. S. Crossman. The Oppenheimer story. The New Statesman and Nation, 48(??):525–??, October 23, 1954. ISSN 0952-102X.

Curtis:1955:OCT

[Cur55] Charles P. (Charles Pelham) Curtis. The Oppenheimer case: the trial of a security system. Simon and Schuster, New York, NY, USA, 1955. xiv + 281 pp. LCCN QC16.O62 C8.

Cooper:1988:PPA

[CW88] Necia Grant Cooper and Geoffrey B. West, editors. Particle physics: a Los Alamos primer. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-34542-1, 0-521-34780-7 (paperback). xi + 199 pp. LCCN QC793 .P358

1988. URL http://www.loc.gov/catdir/enhancements/
fy0907/87010858-d.html;   http://www.loc.gov/catdir/
enhancements/fy0907/87010858-t.html.

Davis:1968:LO

[Dav68] Nuel Pharr Davis. Lawrence and Oppenheimer. Simon and Schuster, New York, NY, USA, 1968. 384 pp. LCCN QC16.L36 D38.

Davis:1969:LO

[Dav69] Nuel Pharr Davis. Lawrence and Oppenheimer. Jonathan Cape, London, UK, 1969. 384 pp. LCCN QC16.L36 D38.

Davis:1986:LO

[Dav86] Nuel Pharr Davis. Lawrence and Oppenheimer. The Da Capo series in science; A Da Capo paperback. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1986. ISBN 0-306-80280-5 (paperback). 384 pp. LCCN QC16.L36 D38 1986. US$11.95.

Day:2001:ONS

[Day01] Michael A. Day. Oppenheimer on the nature of science. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 43(2):73–112, April 2001. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Day:2009:BRB

[Day09a] Michael Day. Book review: Oppenheimer: The Tragic Intellect. American Studies, 50(1/2):225–226, Spring/Summer 2009. ISSN 0026-3079 (print), 2153-6856 (electronic). URL http://muse.

jhu.edu/journals/american_studies/v050/50.1.day.html.

Day:2009:ORA

[Day09b] Michael A. Day. Oppenheimer and Rabi: American Cold War physicists as public intellectuals. In Mariner and Piehler [MP09], pages 307–328. ISBN 1-57233-648-X (hardcover). LCCN

QC773.3.U5 A845 2009. URL http://www.loc.gov/catdir/
toc/ecip0825/2008034982.html.

Day:2015:HVJ

[Day15] Michael (Michael A.) Day. The hope and vision of J. Robert Oppenheimer. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2015. ISBN 981-4656-73-9 (hardcover), 981-4656-74-7 (paperback). xvii + 264 pp. LCCN QC16.O62 D38 2015.

Davis:1971:BWI

[DD71] Nuel Phar Davis and N. P. Davis. Die Bombe war ihr Schicksal: die Forscher Oppenheimer und Lawrence im Widerstreit von Wissenschaft und Politik. (German). [The bomb was their destiny: The researchers Oppenheimer and Lawrence in the conflict between science and policy]. Herder, Freiburg im Breisgau, West Germany, 1971. 296 pp. LCCN ????

DeSantillana:1957:GJR

[De 57] Giorgio De Santillana. Galileo and J. Robert Oppenheimer. The Reporter, ??(??):??, December 26, 1957. Reprinted in [De 68].

DeSantillana:1968:GJR

[De 68] Giorgio De Santillana. Galileo and J. Robert Oppenheimer. In Reflections on Men and Ideas, page ?? MIT Press, Cambridge, MA, USA, 1968. LCCN ???? First published in [De 57].

DeepeKeever:2008:TSC

[Dee08] Beverly Ann Deepe Keever. Top secret: Censoring the first rough drafts of atomic-bomb history. Media History, 14(2):

185–204, July 2008. CODEN ???? ISSN 1368-8804 (print),

1469-9729 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/13688800802176805.

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “The letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125– 152, ???? 2015. CODEN RQSCAN. ISSN 0035-2160.

URL https://www.unamur.be/sciences/philosoc/revueqs/
textes-en-ligne/rqs_186_1_demortier.

Devlin:1962:TMO

[Dev62] John C. Devlin. Teller is mute on Oppenheimer; has TV show delete question. New York Times, ??(??):??, April 16, 1962. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp. LCCN Q141

.D52 1994. URL http://www.loc.gov/catdir/enhancements/
fy0637/99488304-d.html;   http://www.loc.gov/catdir/
enhancements/fy0637/99488304-t.html;   http://www.netLibrary.
com/urlapi.asp?action=summary&v=1&bookid=12306.

Dietz:1954:ASB

[Die54] David Dietz. Atomic science, bombs, and power. Dodd, Mead, New York, NY, USA, 1954. 316 pp. LCCN QC773 .D5.

Dietz:1962:ASB

[Die62] David Dietz. Atomic science, bombs, and power. Collier books: science-for-everyone, AS131. Collier Books, New York, NY, USA, new revised edition, 1962. 254 pp. LCCN QC778 .D53 1962; Microfilm 77793 QC.

Dirac:1971:DQT

[Dir71] P. A. M. Dirac. The development of quantum theory: J. Robert Oppenheimer memorial prize acceptance speech. Gordon and Breach Science Publishers, New York, NY, USA, 1971. 66 pp. LCCN ????

Deshpande:1969:BOT

[DM69] V. K. Deshpande and J. Mahanty. Born–Oppenheimer treatment of the hydrogen atom. American Journal of Physics, 37 (8):823–??, August 1969. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v37/i8/p823_s1.

Dominy:2004:ROA

[Dom04] M. F. Dominy. Robert Oppenheimer and the American century. Library Journal, 129(16):108, October 2004. CODEN LIBJA7. ISSN 0363-0277.

Dominy:2008:LTF

[Dom08] Margaret F. Dominy. A life in twilight: The final years of J. Robert Oppenheimer. Library Journal, 133(16):91, October 2008. CODEN LIBJA7. ISSN 0363-0277.

Dombey:2013:ERO

[Dom13] Norman Dombey. Essay review: Oppenheimer, the bomb and beyond, Scope: biography. Level: general readership [Review of Inside the Centre: The Life of J. Robert Oppenheimer, by Ray Monk, London, Jonathon Cape, 2012, 818 pp., £30.00 (hardback), ISBN 978-0-224-06262-6]. Contemporary Physics, 54(5): 252–253, 2013. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Dackman:1984:DSM

[DOW84] L. Dackman, F. Oppenheimer, and R. Wilson. On the designing of a science museum and a particle accelerator laboratory + interviews with Frank Oppenheimer and Robert Wilson. Leonardo (Oxford, England), 17(2):75, ???? 1984. CODEN LEONDP. ISSN 0024-094x (print), 1530-9282 (electronic).

Doyle:2002:BBT

[Doy02] J. Doyle. Brotherhood of the bomb: The tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. Library Journal, 127(12):96, ???? 2002. CODEN LIBJA7. ISSN 0363-0277.

Dyson:1964:JRO

[DPSY64] Freeman Dyson, Abraham Pais, Bengt Stromgren, and Chen Ning Yang. To J. Robert Oppenheimer on his sixtieth birthday. Reviews of Modern Physics, 36(2):507–508, April 1964. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.36.507;   http://rmp.aps.org/
abstract/RMP/v36/i2/p507_1.

Davis:1982:O

[DPW^{+}82] Barry Davis, Peter Prince, Sam Waterston, David Suchet, and Jana Shelden. Oppenheimer, 1982. 1 VHS videocassette.

Drell:2007:HBJ

[Dre07] Sidney D. Drell. Hans Bethe, 2 July 1906–6 March 2005. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 151(3):337–344, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://www.jstor.org/stable/4599076.

Driemen:1989:ADB

[Dri89] J. E. (John Evans) Driemen. Atomic dawn: a biography of Robert Oppenheimer. A People in focus book. Dillon Press, Minneapolis, MN, USA, 1989. ISBN 0-87518-397-2. 160 pp. LCCN QC16.O62 D75 1989.

Drollette:2023:DSO

[Dro23a] Dan Drollette, Jr. “He did not speak the ordinary language”: Memories of Oppie from a Manhattan Project physicist. Bulletin of the Atomic Scientists, 79(4):221–231, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2223075.

Drollette:2023:OVM

[Dro23b] Dan Drollette, Jr. Oppenheimer — “A very mysterious and delphic character.” Interview with Kai Bird, author of American Prometheus. Bulletin of the Atomic Scientists, 79(4):210–215, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2223071.

Drollette:2023:OMB

[Dro23c] Dan Drollette, Jr. Oppenheimer: The man behind the movie. Bulletin of the Atomic Scientists, 79(4):201–202, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2224160.

Dudley:1980:RSS

[Dud80] John H. Dudley. Ranch school to secret city. In Badash et al. [BHB80], pages 1–11. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Duffus:1946:ABV

[Duf46] R. L. Duffus. Atom bomb versus ‘human race’: The scientists who fashioned it remind us of its terrible threat: Review of One World or None. A Report to the Public on the Full Meaning of the Atomic Bomb. Edited by Dexer Masters and Kathlerine Way. Foreword by Niels Bohr. Introduction by Arthur H. Compton. x + 79 pp. New York: Whittlesey House. $1. New York Times, ??(??):BR1–BR2, March 17, 1946. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/107455627.

Davidson:1969:MJR

[DWS^{+}69] Gordon Davidson, Joseph Wiseman, Ruth Speirs, Heinar Kipphardt, and Richard L. (Richard Livingstone) Coe. In the matter of J. Robert Oppenheimer, 1969.

Dyson:1976:HBD

[Dys76] F. J. Dyson. The hydrogen-bomb decision: a reappraisal: Review of The Advisors: Oppenheimer, Teller and The Superbomb, by Herbert F. York. Science, 193(4254):668–669, August 20, 1976. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.sciencemag.org/
content/193/4254/668.full.pdf.

Dyson:1979:BUBa

[Dys79] Freeman Dyson. Disturbing the Universe. Oppenheimer, His Nature and His Physics. 2. World of the Scientist. The New Yorker, 55(26):64, ???? 1979. ISSN 0028-792X.

Dyson:1980:ROL

[Dys80] F. J. Dyson. Robert Oppenheimer — letters and recollections. New Republic, 182(21):37–39, ???? 1980. ISSN 0028-6583.

Dyson:1983:BRBb

[Dys83] Freeman J. Dyson. Book review: Robert Oppenheimer, Letters and Recollections, edited by Alice Smith and Charles Weiner. New Republic, ??(??):??, ???? 1983. ISSN 0028-6583.

Dyson:2006:OSA

[Dys06a] Freeman J. Dyson. Oppenheimer as scientist, administrator, and poet. In The scientist as rebel [Dys06b], pages 229–242. ISBN 1-59017-216-7. LCCN Q158.5 .D977 2006. URL http://

www.loc.gov/catdir/enhancements/fy0665/2006022081-b.
html;   http://www.loc.gov/catdir/enhancements/fy0665/
2006022081-d.html;   http://www.loc.gov/catdir/toc/ecip0616/

2006022081.html. Collection of book reviews.

Dyson:2006:SR

[Dys06b] Freeman J. Dyson. The scientist as rebel. New York Review Books, New York, NY, USA, 2006. ISBN 1-59017-216-

7. xvi + 360 pp. LCCN Q158.5 .D977 2006. URL http://

www.loc.gov/catdir/enhancements/fy0665/2006022081-b.
html;   http://www.loc.gov/catdir/enhancements/fy0665/
2006022081-d.html;   http://www.loc.gov/catdir/toc/ecip0616/

2006022081.html. Collection of book reviews.

Dyson:2013:BRO

[Dys13] Freeman J. Dyson. Book review: Oppenheimer: The shape of genius: Robert Oppenheimer: A Life Inside the Center, by Ray Monk, Doubleday, 825 pp., $37.50. New York Review of Books, 60(13):18–19, August 15, 2013. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2013/aug/15/oppenheimer-shape-genius/.

Dyson:2018:MPA

[Dys18] Freeman J. Dyson. Maker of patterns: an autobiography through letters. Liveright Publishing Corporation, New York, NY, USA, 2018. ISBN 0-87140-386-2 (hardcover). xvi + 400 pp. LCCN QC16.D95 A3 2018.

Eakin:2005:RJR

[Eak05a] H. Eakin. The ruin of J. Robert Oppenheimer and the birth of the modern arms race. New York Times Book Review, ??(??): 12, August 7, 2005. ISSN 0028-7806.

Eakin:2005:BRB

[Eak05b] Hugh Eakin. Book review: The Ruin of J. Robert Oppenheimer: And the Birth of the Modern Arms Race, by Priscilla McMullin. New York Times Book Review, ??(??):??, August 7, 2005. ISSN 0028-7806.

Easton:1989:TB

[Eas89] Nina J. Easton. From ‘A-Team’ to A-Bomb. Los Angeles Times, ??(??):F1–??, November 1, 1989. URL http://search.

proquest.com/docview/420977394.

Esposito:2015:PEM

[EAW15] Salvatore Esposito, E. (Evgeny) Akhmedov, and Frank Wilczek. The physics of Ettore Majorana: phenomenological, theoretical, and mathematical. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-04402-2 (hardcover), 1-316-19108-7 (PDF ebook). xi + 382 pp. LCCN QC19.6 .E87 2015.

Eckart:1970:CEP

[Eck70] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). ????, ????, 1970. 8 (4002413) pp. LCCN Oversize 0113K; 0810D.

URL http://hdl.loc.gov/loc.mss/eadmss.ms001014.

Eckart:1973:CEP

[Eck73] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). US Library of Congress archival manuscript material (collec-

tion)., 1973. URL http://hdl.loc.gov/loc.mss/eadmss.
ms001014.

Edgerton:1991:BRA

[Edg91] David Edgerton. Book reviews: Atomic Energy for Military Purposes, with a new foreword by Philip Morrison and an essay by Henry DeWolf Smyth. Stanford: Stanford University Press, 1990. Pp. xvi + 324. ISBN 0-8047-1721-4, $39.50 (hardcover); 0-8047-1722-2, $12.95 (paperback). Herbert F. York. The Advisors: Oppenheimer, Teller and the Superbomb, with a historical essay by Hans A. Bethe. Stanford: Stanford University Press, 1989. Pp. xiv + 201. ISBN 0-8047-1713-3, $32.50 (hardcover); 0-8047-1714-1, $8.95 (paperback). British Journal for the History of Science, 24(4):476–477, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4027137.

Board:1954:EOC

[Edi54] Editorial Board. Editorial: The Oppenheimer case. Bulletin of the Atomic Scientists, 10(5):173, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Else:2002:DAT

[EFB02] Jon Else, Paul Frees, and Peter Baker. The day after Trinity, 2002. 1 videodisc (89 min.).

Epstein:1948:NSD

[EFO48] S. T. Epstein, R. J. Finkelstein, and J. R. Oppenheimer. Note on stimulated decay of negative mesons. Physical Review (2), 73 (10):1140–1141, May 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v73/i10/p1140_1.

Else:2005:OFT

[EG05] Jon Else and Peter Galison. Oppenheimer in film: a transcript. In Carson and Hollinger [CH05], chapter 18, pages 361–384. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Elliot:1986:PVN

[Ell86] David C. Elliot. Project Vista and nuclear weapons in Europe. International Security, 11(1):163–183, Summer 1986. CODEN ???? ISSN 0162-2889 (print), 1531-4804 (electronic). URL

http://muse.jhu.edu/journals/international_security/
summary/v011/11.1.elliot.html;   http://www.jstor.org/
stable/2538879.

Else:1981:DAT

[Els81] Jon Else. The Day after Trinity: J. Robert Oppenheimer and the atomic bomb. Pyramid Films, Santa Monica, CA, USA, 1981. Video film (88 minutes).

Else:1995:DAT

[Els95] Jon Else. The day after Trinity: J. Robert Oppenheimer and the atomic bomb. For the record. Voyager, New York, NY, USA, 1995. ISBN 1-55940-685-2. LCCN QC16.O62. 1 computer optical disc (89 minutes).

Ehrenfest:1931:NSN

[EO31] P. Ehrenfest and J. R. Oppenheimer. Note on the statistics of nuclei. Physical Review (2), 37(4):333–338, February 15, 1931.

CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v37/i4/
p333_1.

Else:2008:WMH

[EOA^{+}08] Jon Else, Eric Owens, John Adams, Bonni Cohen, and Deborah Hoffmann. Wonders are many: the making of Doctor Atomic, 2008. 1 videodisc (92 min.).

Else:1980:DAT

[EPP80] Jon Else, David Webb Peoples, and Janet Peoples. The day after Trinity, J. Robert Oppenheimer and the atomic bomb, 1980. 1 VHS videocassette (89 min.).

Erwin:1984:OI

[Erw84] R. Erwin. Oppenheimer investigated. The Wilson Quarterly (1976–2012), 18(4):34–46, ???? 1984. ISSN 0363-3276.

Essen:1977:PBO

[Ess77] H. Essén. The physics of the Born–Oppenheimer approximation. International Journal of Quantum Chemistry, 12(4):721– 735, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Evans:1954:MRO

[Eva54] Ward V. Evans. Minority report [on Oppenheimer security clearance denial]. Bulletin of the Atomic Scientists, 10(6):250, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ECFAS:1954:SDO

[Exe54] Executive Committee of the Federation of American Scientists. The suspension of Dr. Oppenheimer. Bulletin of the Atomic Scientists, 10(5):191–192, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fasching:1993:ECA

[Fas93] Darrell J. Fasching. The ethical challenge of Auschwitz and Hiroshima: Apocalypse or Utopia? State University of New York Press, Albany, NY, USA, 1993. ISBN 0-7914-1375-6, 0-7914-1376-4 (paperback). xvi + 366 pp. LCCN BJ1188 .F37 1993.

Fleming:1969:IME

[FB69] Donald Fleming and Bernard Bailyn, editors. The intellectual migration: Europe and America, 1930–1960. Belknap Press of Harvard University Press, Cambridge, MA, USA; London, UK, 1969. ISBN 0-674-33411-6, 0-674-33412-4. 748 pp. LCCN E169.1 .F6.

Fermi:1941:NP

[FBR^{+}41] Enrico Fermi, Gregory Breit, I. I. Rabi, Eugene P. Wigner, and John H. van Vleck, editors. Nuclear physics. University Press of Pennsylvania, Philadelphia, PA, USA, 1941. v + 68 pp.

FBI:1978:JRO

[Fed78] Federal Bureau of Investigation. J. Robert Oppenheimer, FBI security file. Scholarly Resources, Wilmington, DE, USA, 1978.

Fehner:2014:UMJ

[Feh14] Terry Fehner. Unlocking the mysteries of the J. Robert Oppenheimer transcript. Web site., October 3, 2014.

URL http://energy.gov/articles/unlocking-mysteries-
j-robert-oppenheimer-transcript.

Feis:1960:BWP

[Fei60a] Herbert Feis. Between war and peace: the Potsdam Conference. Princeton University Press, Princeton, NJ, USA, 1960. viii + 367 pp. LCCN ????

Feis:1960:STP

[Fei60b] Herbert Feis. The secret that traveled to Potsdam. Foreign Affairs, 38(2):300–317, January 1960. CODEN FRNAA3. ISSN

0015-7120. URL http://www.jstor.org/stable/20029419.

Extract from [Fei60a].

Fermi:1980:FPA

[Fer80] Laura Fermi. The Fermis’ path to Los Alamos. In Badash et al. [BHB80], page ?? ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Feynman:1980:LAB

[Fey80] Richard P. Feynman. Los Alamos from below. In Badash et al. [BHB80], pages 105–132. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44. URL https://www.

chilton-computing.org.uk/harwell/literature/FeynmanLosAlamos.
htm.

Finney:1954:TAS

[Fin54] Nat S. Finney. The threat to atomic science. Bulletin of the Atomic Scientists, 10(7):285–286, 295, September 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finneran:2006:CFA

[Fin06] K. Finneran. Containing the fire — American Prometheus: The triumph and tragedy of Robert Oppenheimer. Issues in Science and Technology, 22(4):85–87, Summer 2006. ISSN 0748-5492 (print), 1938-1557 (electronic).

Flanagan:1998:MJR

[Fla98] Jason Constantine Flanagan. In the matter of J. Robert Oppenheimer: a trial of character. B.A. (Honours History), The University of Queensland, Brisbane, QLD 4072, Australia, Oc-

tober 29, 1998. v + 78 pp. URL http://espace.library.uq.
edu.au/eserv/UQ:266176/Jason_Flanagan.pdf.

Furry:1934:LTP

[FO34a] W. H. Furry and J. R. Oppenheimer. On the limitations of the theory of the positron. Physical Review (2), 45(12):903– 904, June 15, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v45/i12/p903_1.

Furry:1934:TEPa

[FO34b] W. H. Furry and J. R. Oppenheimer. On the theory of the electron and positive. Physical Review (2), 45(5):343–344, March 1, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v45/
i5/p343_1.

Furry:1934:TEPb

[FO34c] W. H. Furry and J. R. Oppenheimer. On the theory of the electron and positive. Physical Review (2), 45(4):245–262, February 15, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v45/i4/p245_1.

Frye:2009:CET

[FOS09] Northrop Frye, J. Robert Oppenheimer, and Edward Togo Salmon. City of the end of things: lectures on civilization and empire. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-543005-0 (paper). 245 pp. LCCN CB5 .F79 2009.

Frankel:1959:IUE

[Fra59] Charles Frankel. Issues in university education. Harper, New York, NY, USA, 1959. 175 pp. LCCN LB2325 .F7.

Frankel:1976:IUE

[Fra76] Charles Frankel, editor. Issues in university education: essays by ten American scholars. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1976. ISBN 0-8371-9353-2. xiv + 175 pp. LCCN LB2325 .F7 1976.

Frauenfelder:2001:BRS

[Fra01] Hans Frauenfelder. Book review: S. S. Schweber, In the Shadow of the Bomb: Bethe, Oppenheimer, and the Moral Responsibility of the Scientist. Princeton: Princeton University Press, 2000, xviii + 260 pages. $24.95 (cloth). Physics in Perspective (PIP), 3(4):493–495, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

French:1979:ECVb

[Fre79] A. P. (Anthony Philip) French, editor. Einstein: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1979. ISBN 0-674-24230-0. xx + 332 pp. LCCN QC16.E5 E37.

Freedman:2005:BRBa

[Fre05a] Lawrence D. Freedman. Book review: American Prometheus: The Triumph and Tragedy of J. Robert Oppenheimer, by Kai Bird and Martin J. Sherwin. Foreign Affairs, 84(3): 136, May/June 2005. CODEN FRNAA3. ISSN 0015-

7120. URL http://www.foreignaffairs.com/articles/
60673/lawrence-d-freedman/american-prometheus-the-triumph-
and-tragedy-of-j-robert-oppenhei;   http://www.jstor.
org/stable/20034367.

Freedman:2005:BRBb

[Fre05b] Lawrence D. Freedman. Book review: The Ruin of J. Robert Oppenheimer, and the Birth of the Modern Arms Race, by Priscilla McMillan. Foreign Affairs, 84(5):174, September/October 2005. CODEN FRNAA3. ISSN 0015-

7120. URL http://www.foreignaffairs.com/articles/
60892/lawrence-d-freedman/the-ruin-of-j-robert-oppenheimer-
and-the-birth-of-the-modern-arm;   http://www.jstor.
org/stable/20031730. See [McM05].

Friendly:1967:OJR

[Fri67] Alfred Friendly. [Obituary: J. Robert Oppenheimer]. US Congressional Record, ??(??):??, ???? 1967.

Feldman:2007:MSE

[FW07] Burton Feldman and Katherine Williams. 112 Mercer Street: Einstein, Russell, Gödel, Pauli, and the end of innocence in science. Arcade Publishing, New York, NY, USA, 2007. ISBN 1-55970-704-6. xx + 243 pp. LCCN Q141 .F345 2007. URL http:

//www.loc.gov/catdir/toc/ecip078/2007001194.html.

Graetzer:1971:DNF

[GA71] Hans G. Graetzer and David L. Anderson. The discovery of nuclear fission: a documentary history, volume 20 of Van Nostrand Reinhold momentum books. Van Nostrand Reinhold, New York, NY, USA, 1971. viii + 120 pp. LCCN QC790 .G68.

Gamow:1928:QAG

[Gam28] George Gamow. Zur Quantentheorie des Atomkernes. (German) [On the quantum theory of the atomic nucleus]. Zeitschrift für Physik, 51(3–4):204–212, March 1928. CODEN ZEPYAA. ISSN

0044-3328. URL http://www.springerlink.com/content/

mw52h8867mr4x185/. Reprinted in [Bey49, pages 77–85].

Garrison:1954:ORR

[Gar54] Lloyd K. Garrison. Oppenheimer requests review. Bulletin of the Atomic Scientists, 10(6):251–254, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter sent to General Nichols by the author, who is the counsel for J. Robert Oppenheimer.

Garwin:2008:LNW

[Gar08] Richard L. Garwin. Living with nuclear weapons: Sixty years and counting. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 152(1):69– 82, March 2008. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://www.jstor.org/stable/
25478469.

Galison:1989:LSD

[GB89] Peter Galison and Barton Bernstein. In any light: Scientists and the decision to build the superbomb, 1952–1954. Historical Studies in the Physical and Biological Sciences, 19(2):267–347, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355

(electronic). URL http://www.jstor.org/stable/27757627.

Grieve-Carlson:2010:MOC

[GCD10] Gary Grieve-Carlson and Michael A. Day. MacLeish, Oppenheimer, and ‘the conquest of America’. Soundings: An Interdisciplinary Journal, 93(??):281–311, Fall/Winter 2010. CODEN ???? ISSN 0038-1861 (print), 2161-6302 (electronic). URL

http://www.jstor.org/stable/41200932.

Gromyko:2004:RJR

[GDH^{+}04] A. Gromyko, F. Dyson, K. C. Hart, P. H. Nitze, A. M. Smith, S. L. Rearden, M. Moore, J. Reston, R. Aron, and A. Mill. Remembering J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 60(2):37–??, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gangopadhyay:2004:BOA

[GDR04] Gautam Gangopadhyay and Binayak Dutta-Roy. The Born– Oppenheimer approximation: a toy version. American Journal of Physics, 72(3):389–??, March 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v72/i3/p389_s1.

Geiger:1993:RRK

[Gei93] Roger L. Geiger. Research and Relevant Knowledge: American Research Universities Since World War II. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-505346-X. xvi + 411 pp. LCCN Q180.U5 G33 1993. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0604/92022148-
d.html;   http://www.gbv.de/dms/bowker/toc/9780195053463.
pdf.

Geiger:2004:RRK

[Gei04] Roger L. Geiger. Research and Relevant Knowledge: American Research Universities Since World War II. Transaction series in higher education. Transaction Publishers, New Brunswick, NJ, USA, 2004. ISBN 0-7658-0569-3. xxi + 411 pp. LCCN Q180.U5G34 2004.

Gerjuoy:2006:BRJ

[Ger06] Edward Gerjuoy. Book review: Jeremy Bernstein, Oppenheimer: Portrait of an Enigma. Chicago: Ivan R. Dee, 2004, 223 pages. $25.00 (cloth). David C. Cassidy, J. Robert Oppenheimer and the American Century. New York: Pi Press, 2005, 462 pages. $27.95 (cloth). Physics in Perspective (PIP), 8(1):109–112, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Giovannitti:1965:DDB

[GF65] Len Giovannitti and Fred Freed. The decision to drop the bomb. Coward-McCann, New York, NY, USA, 1965. 348 pp. LCCN ????

Giovannitti:1967:DDB

[GF67] Len Giovannitti and Fred Freed. The decision to drop the bomb. Methuen and Co. Ltd., London, UK, 1967. 384 pp. LCCN UA23 .G62 1967.

Goldberg:1994:LWA

[GH94] Stanley Goldberg and David Hawkins. Letters: Wrong about Bohr, too. Bulletin of the Atomic Scientists, 50(5):3, 59–60, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1996:EKJ

[GHK^{+}96] Stanley Goldberg, Dieter Hoffmann, Alexei B. Kojewnikow, Fritz Krafft, and Helmut Rechenberg. Die erste Kernwaffendetonation am 16. Juli 1945 in Alamogordo, New Mexico — Vorgeschichte, Ereignis, Wirkungen Öffentliche Podiumsdiskussion. (German) [The first nuclear detonation on 16 July 1945 in Alamogordo, New Mexico — history, event, effects public panel discussion]. Berichte zur Wissenschaftsgeschichte, 19(2–3):157– 182, ???? 1996. CODEN BEWID8. ISSN 0170-6233 (print),

1522-2365 (electronic). URL http://onlinelibrary.wiley.
com/doi/10.1002/bewi.19960190214/abstract.

Gianetto:1985:MOF

[Gia85] Enrico Gianetto. A Majorana–Oppenheimer formulation of quantum electrodynamics. Lettere al Nuovo Cimento Series 2, 44(3):140–144, October 1, 1985. CODEN LNUCAE. ISSN 0375-930X (print), 1827-613X (electronic). URL http://link.

springer.com/article/10.1007/BF02746912.

Giniger:1954:OCH

[Gin54] Henry Giniger. Oppenheimer case held based on lie: Writer, once called a Russian agent by physicist, charges A.E.C. ignored retraction. New York Times, ??(??):17, December 3, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/113042952/.

Gardner:1995:JRO

[GL95a] Howard Gardner and Emma Laskin. J. Robert Oppenheimer: the teaching of physics, the lessons of politics. In Leading minds: an anatomy of leadership [GL95b], pages 89–109. ISBN 0-465-08279-3 (hardcover), 0-465-08280-7 (paperback). LCCN HM141 .G35 1995.

Gardner:1995:LMA

[GL95b] Howard Gardner and Emma Laskin. Leading minds: an anatomy of leadership. Basic Books, New York, NY, USA, 1995. ISBN 0-465-08279-3 (hardcover), 0-465-08280-7 (paperback). xi + 400 pp. LCCN HM141 .G35 1995.

Gardner:1996:JRO

[GL96a] Howard Gardner and Emma Laskin. J. Robert Oppenheimer: the teaching of physics, the lessons of politics. In Leading minds: an anatomy of leadership [GL96b], page ?? ISBN 0-00-255655-3. LCCN ????

Gardner:1996:LMA

[GL96b] Howard Gardner and Emma Laskin. Leading minds: an anatomy of leadership. HarperCollins, London, UK, 1996. ISBN 0-00-255655-3. xi + 400 pp. LCCN ????

Gardner:1997:JRO

[GL97a] Howard Gardner and Emma Laskin. J. Robert Oppenheimer: the teaching of physics, the lessons of politics. In Leading minds: an anatomy of leadership [GL97b], page ?? ISBN 0-00-638123-5 (paperback). LCCN HD57.7.G227.

Gardner:1997:LMA

[GL97b] Howard Gardner and Emma Laskin. Leading minds: an anatomy of leadership. HarperCollins, London, UK, 1997. ISBN 0-00-638123-5 (paperback). xv + 400 pp. LCCN HD57.7.G227.

Gardner:1998:MLA

[GL98] Howard Gardner and Emma Laskin. Mentes líderes: una anatomía del liderazgo. (Spanish) [Leading minds: an anatomy of leadership], volume 11 of Paidós transiciones. Paidós, Barcelona, Spain, 1998. ISBN 84-493-0542-X. 463 pp. LCCN HM141 G3718 1998. Spanish translation by Pedro Tosaus Abadía of [GL97b].

Gardner:2011:JRO

[GL11a] Howard Gardner and Emma Laskin. J. Robert Oppenheimer: the teaching of physics, the lessons of politics. In Leading minds: an anatomy of leadership [GL11b], page ?? ISBN 0-465-02773-3 (paperback), 0-465-02777-6 (e-book). LCCN HM141.

Gardner:2011:LMA

[GL11b] Howard Gardner and Emma Laskin. Leading minds: an anatomy of leadership. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02773-3 (paperback), 0-465-02777-6 (e-book). xxii + 391 pp. LCCN HM141.

Gray:1954:RSP

[GM54] Gordon Gray and Thomas A. Morgan. Report of the Special Personnel Security Board: Majority report, May 27, 1954. Bulletin of the Atomic Scientists, 10(6):243–249, 254, June 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garrison:1954:MJR

[GNMS54] Lloyd K. Garrison, General K. D. Nichols, Thomas E. Murray, and Henry DeWolf Smyth. “In the matter of J. Robert Oppenheimer”. Bulletin of the Atomic Scientists, 10(7):270–274, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

GoeppertMayer:1988:MGM

[Goe88] Maria Goeppert Mayer. Maria Goeppert Mayer papers, 1925– 1973, MSS 20. Collection at the University of California, San

Diego., 1988. URL http://libraries.ucsd.edu/speccoll/
findingaids/mss0020.html.

Goldberger:1955:BBC

[Gol55] Marvin L. Goldberger. Books: Science and the Common Understanding, by J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 11(2):64, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstine:1972:CPN

[Gol72] Herman Heine Goldstine. The Computer: From Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-02367-0 (paperback), 0-691-08104-2. xii + 378 pp. LCCN TK7885.A5 G64 1993. URL http://www.jstor. org/stable/j.ctt7rvp0. Second printing, 1973. Paperback edition 1980. Fifth printing, 1993 with new preface. Reprint 2000 by Books on Demand, Ann Arbor, MI, USA.

Goldberg:1992:GSO

[Gol92a] Stanley Goldberg. Groves, Szilard, and Oppenheimer. Bulletin of the Atomic Scientists, 48(10):37, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:GTR

[Gol92b] Stanley Goldberg. Groves takes the reins. Bulletin of the Atomic Scientists, 48(10):32–39, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1995:BRB

[Gol95a] Stanley Goldberg. Book review: Critical Assembly: A Technical History of Los Alamos during the Oppenheimer Years, 1943– 1945 by Lillian Hoddeson; Paul W. Henriksen; Roger A. Meade; Catherine Westfall. Isis, 86(3):520–522, September 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235091.

Goldberg:1995:GOS

[Gol95b] Stanley Goldberg. Groves and Oppenheimer: The story of a partnership. The Antioch Review, 53(4):482–493, Autumn 1995.

Goldberg:1995:GSC

[Gol95c] Stanley Goldberg. Groves and the scientists: Compartmentalization and the building of the bomb. Physics Today, 48(8):38– 43, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v48/i8/p38_s1.

Goldberg:1996:GGB

[Gol96] Stanley Goldberg. General Groves and the bombing of Hiroshima and Nagasaki (Part II). Berichte zur Wissenschaftsgeschichte, 19(4):207–217, ???? 1996. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Goldberger:1999:AEF

[Gol99] Marvin L. Goldberger. In appreciation: Enrico Fermi (1901– 1954): The complete physicist. Physics in Perspective (PIP), 1(3):328–336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050024;   http:/
/www.springerlink.com/content/y1h02gr3rv7wurrl/.

Goncharov:1996:TMB

[Gon96] German A. Goncharov. Thermonuclear milestones: (2) beginnings of the Soviet H-bomb program. Physics Today, 49(11): 50–54, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goodchild:1980:JRO

[Goo80] Peter Goodchild. J. Robert Oppenheimer: shatterer of worlds. British Broadcasting Corporation, London, UK, 1980. ISBN 0-563-17781-0. 301 pp. LCCN ???? Published in conjunction with the BBC/WGBH television series Oppenheimer, produced by Peter Goodchild and written by Peter Prince. See important corrections [Bet82, page 53].

Goodchild:1981:JRO

[Goo81] Peter Goodchild. J. Robert Oppenheimer: Shatterer of Worlds. Houghton-Mifflin, Boston, MA, USA, 1981. ISBN 0-395-30530-6. 301 pp. LCCN ???? Published in conjunction with the BBC/WGBH television series Oppenheimer, produced by Peter Goodchild and written by Peter Prince. See important corrections [Bet82, page 53].

Goodchild:1982:JRO

[Goo82] Peter Goodchild. J. Robert Oppenheimer: eine Bildbiographie. (German) [J. Robert Oppenheimer: a picture biography].

Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1982. ISBN 3-7643-1292-0. 306 pp. LCCN ????

Goodchild:2004:ETRb

[Goo04] Peter Goodchild. Edward Teller, the real Dr. Strangelove. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01669-6. xxv + 469 + 16 pp. LCCN QC16.T37 G66 2004.

URL http://alsos.wlu.edu/information.aspx?id=2237.

Goodchild:2005:SHP

[Goo05] Peter Goodchild. The sum of his parts. Bulletin of the Atomic Scientists, 61(6):63–66, November 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//bos.sagepub.com/content/61/6/63.abstract. Review of

[BS05b, McM05].

Good:2009:BRS

[Goo09] Gregory A. Good. Book review: Silvan S. Schweber, Einstein and Oppenheimer: The Meaning of Genius. Cambridge, Mass. and London: Harvard University Press, 2008, ix + 412 pages. $29.95 (cloth). Physics in Perspective (PIP), 11(4):467–468, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Gordin:2009:BRB

[Gor09a] Michael D. Gordin. Book review: Einstein and Oppenheimer: The Meaning of Genius. Isis, 100(1):186–188, March 2009. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/599684.

Gordin:2009:RCD

[Gor09b] Michael D. Gordin. Red cloud at dawn: Truman, Stalin, and the end of the atomic monopoly. Farrar, Straus and Giroux, New York, NY, USA, 2009. ISBN 0-374-25682-9. xii + 402 pp. LCCN U264 .G67 2009.

Gowing:1981:ERO

[Gow81] Margaret Gowing. Essay review: Oppenheimer and Los Alamos. Robert Oppenheimer: Letters and Recollections, edited by Alice Kimball Smith and Charles Weiner. J. Robert Oppenheimer, “Shatterer of Worlds”, by Peter Goodchild. Reminiscences of Los Alamos, 1943–1945, edited by Lawrence Badash, Joseph O. Hirschfelder and Herbert P. Broida. Contemporary Physics, 22 (6):681–686, 1981. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Goodchild:2008:OHV

[GPDW08] Peter Goodchild, Peter Prince, Barry Davis, and Sam Waterston. Oppenheimer: the father of the atomic bomb, 2008. ISBN 1-4198-6756-3. 3 videodiscs (ca. 420 min.).

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Graetzer:1964:DNF

[Gra64] Hans G. Graetzer. Discovery of nuclear fission. American Journal of Physics, 32(9):9–15, January 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See [HS39, Bey49].

Gramm:2002:THJ

[Gra02] E. Gramm. With their honor (J. Robert Oppenheimer). New York Times Book Review, ??(??):4, October 6, 2002. ISSN 0028-7806.

Green:1955:USS

[Gre55] Harold P. Green. The unsystematic security system. Bulletin of the Atomic Scientists, 11(4):118–122, 164, April 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Greenberg:1967:PPS

[Gre67] Daniel S. Greenberg. The Politics of Pure Science. World Pub. Co., New York, NY, USA, 1967. xiii + 303 pp. LCCN Q127.U6 G68 1967.

Greenberg:1969:PAS

[Gre69] Daniel S. Greenberg. The Politics of American Science. Penguin, Harmondsworth, UK, 1969. ISBN 0-14-021117-9. 368 pp. LCCN Q127.U6 G68 1969.

Green:1977:OCS

[Gre77] Harold Green. The Oppenheimer case: a study in the abuse of law. Bulletin of the Atomic Scientists, 33(7):12–16, 56–61, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gresser:1999:I

[Gre99] H. Gresser. The invitation. American Heritage, 50(7):40, November 1999. ISSN 0002-8738. Reader’s account of his personal contacts with physicist Robert Oppenheimer.

Green:2023:OCS

[Gre23] Harold P. Green. The Oppenheimer case: a study in the abuse of law. Bulletin of the Atomic Scientists, 79(4):255–264, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2223085.

Grodzins:1956:BBO

[Gro56] Morton Grodzins. Books: The Oppenheimer Case: The Trial of a Security System, by Charles P. Curtis. Bulletin of the Atomic Scientists, 12(9):349, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SG69, SG71].

Groves:1962:NIC

[Gro62] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Harper & Row, New York, NY, USA, 1962. xiv + 464 pp. LCCN QC773.A1 G7.

Groves:1983:NIC

[Gro83] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1983. ISBN 0-306-80189-2, 0-7867-4822-2 (e-book). xiv + 464 + 8 pp. LCCN QC773.A1 G7.

Grobe:1999:EHK

[Gro99] Horst Grobe. Erläuterungen zu Heinar Kipphardt, In der Sache J. Robert Oppenheimer. (German) [Explanations on Heinar Kipphardt, On the Case J. Robert Oppenheimer], volume 160 of Königs Erläuterungen und Materialien. Bange, Hollfeld, Germany, second revised edition, 1999. ISBN 3-8044-1641-1. 104 pp. LCCN ????

Grubin:2009:TJR

[Gru09] David Grubin. The trials of J. Robert Oppenheimer, 2009. ISBN 0-7936-7037-3. 1 videodisc (120 min.).

Goodchild:2006:RDS

[GTS^{+}06] Peter Goodchild, Simon Templeman, Joe Spano, John De Lancie, and Matt August. The real Dr. Strangelove. Audio theatre collection. L. A. Theatre Works, Venice, CA, USA, 2006. ISBN 1-58081-357-7. LCCN RZC 1211. 2 sound discs (122 minutes).

Guinnessy:2004:UCO

[Gui04] Paul Guinnessy. US celebrates Oppenheimer’s centenary. Physics Today, 57(6):35–36, June 2004. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/?PTO/57/35/1;   http://
physicstoday.org/resource/1/phtoad/v57/i6/p35_s1.

Gusterson:1996:NRW

[Gus96] Hugh Gusterson. Nuclear Rites: a Weapons Laboratory at the End of the Cold War. University of California Press, Berkeley, CA, USA, 1996. ISBN 0-520-08147-1 (hardcover), 0-520-21373-4 (paperback). 351 pp. LCCN U264.4.C2 G87 1996.

Gusterson:2004:PBP

[Gus04] Hugh Gusterson. People of the Bomb: Portraits of America’s Nuclear Complex. University of Minnesota Press, Minneapolis, MN, USA, 2004. ISBN 0-8166-3859-4 (hardcover), 0-8166-3860-8 (paperback). xxvi + 312 pp. LCCN U264.3

.G87 2004. URL http://www.loc.gov/catdir/toc/ecip0415/
2004003476.html.

Guy:2007:ERT

[Guy07] Christopher Guy. Essay review: A tragic intellect or a tragedian? [review of Oppenheimer: The Tragic Intellect, by C. Thorpe, The University of Chicago Press, Chicago & London, 2006, pp. 384, £24.00, hardback (ISBN 0-226-79845-3). Scope: sociological biography. Level: nonspecialist]. Contemporary Physics, 48(2):119–121, 2007. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Hutchings:1958:FSS

[H^{+}58] Edward Hutchings et al., editors. Frontiers in science. a survey. Basic Books, New York, NY, USA, 1958. 362 pp. LCCN Q162

.H83. URL http://catalog.hathitrust.org/api/volumes/
oclc/1158177.html.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Hales:1997:ASL

[Hal97] Peter B. (Peter Bacon) Hales. Atomic spaces: living on the Manhattan Project. University of Illinois Press, Urbana, IL, USA, 1997. ISBN 0-252-02296-3 (hardcover). 447 pp. LCCN QC773.3.U5 H35 1997.

Hall:2009:GSP

[Hal09] Derek Hall, editor. Great scientists. Physical sciences. Facts at your fingertips. Brown Bear Books, Redding, CT, USA, 2009. ISBN 1-933834-46-3. 64 pp. LCCN QC15 .P468 2009.

Hutchinson:2021:CEF

[HAM^{+}21] Jesson Hutchinson, Jennifer Alwin, Alexander McSpaden, William Myers, Michael Rising, and Rene Sanchez. Criticality experiments with fast 235U and 239Pu metal and hydride systems during the Manhattan Project. Nuclear Technology, 207(S1):S62–S80, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Hansen:1995:BRB

[Han95] Chuck Hansen. Book review: The bomb, part 2: Dark Sun: The Making of the Hydrogen Bomb, by Richard Rhodes. Bulletin of the Atomic Scientists, 51(5):52–53, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-17290233.
html.

Hart:2005:JRO

[Har05] C. W. Hart. J. Robert Oppenheimer and the American century. Journal of Religion and Health, 44(2):243–245, Summer 2005. CODEN JRHEAT. ISSN 0022-4197 (print), 1573-6571 (electronic).

Hart:2008:JRO

[Har08] Curtis W. Hart. J. Robert Oppenheimer: a faith development portrait. Journal of Religion and Health, 47(1):118–128, March 2008. CODEN JRHEAT. ISSN 0022-4197 (print), 1573-6571 (electronic).

Hart:2009:CET

[Har09] Jonathan Hart, editor. City of the end of things: lectures on civilization and empire. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-543005-0. 245 pp. LCCN CB5 .C53 2009. URL http://www.loc.gov/

catdir/enhancements/fy0914/2009294623-b.htm;   http:
//www.loc.gov/catdir/enhancements/fy0914/2009294623-

d.htm. Lectures from a series given at McMaster University from 1956 to the present.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hargittai:2012:ICL

[Har12] Istvan Hargittai. Inside the centre: The life of J. Robert Oppenheimer. Nature, 491(7426):670, November 29, 2012. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Hawkins:1961:MDH

[Haw61] David Hawkins. Manhattan District history: Project Y: The Los Alamos Project. Vol. 1. Inception until August 1945. Report LAMS-2532, Los Alamos Laboratory, Los Alamos, NM, USA, December 1, 1961. xv + 375 pp. URL http://lib-www.lanl.

gov/cgi-bin/getfile?LAMS-2532.htm.

Hawkings:1994:LOB

[Haw94] David Hawkings. Letter: Oppenheimer and Bohr. Bulletin of the Atomic Scientists, 50(5):60, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Hoddeson:1994:ASB

[HBP94] Lillian Hoddeson, Gordon Baym, and Rudolf Peierls. Atomic science. (book reviews: Critical Assembly. A Technical History of Los Alamos During the Oppenheimer Years, 1943–1945). Science, 263:1162–1163, February 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL https://ui.

adsabs.harvard.edu/#abs/1994Sci...263.1162H.

Hecht:2005:TCR

[Hec05] David K. Hecht. Tumult in the clouds: Robert Oppenheimer and American science. Ph.D. thesis, Yale University, New Haven, CT, USA, May 2005. iii + 298 pp. URL http://search.

proquest.com/docview/305389442.

Hecht:2008:AHR

[Hec08] David K. Hecht. The atomic hero: Robert Oppenheimer and the making of scientific icons in the Early Cold War. Technology and Culture, 49(4):943–966, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

Hecht:2010:NNR

[Hec10] David K. Hecht. A nuclear narrative: Robert Oppenheimer, autobiography, and public authority. Biography (Honolulu), 33(1):167–184, Winter 2010. CODEN BGPYE2. ISSN 0162-4962 (print), 1529-1456 (electronic). URL http://muse.jhu.

edu/journals/biography/v033/33.1.hecht.html;   http:
//www.jstor.org/stable/23541054.

Heilbron:1968:QHA

[Hei68] John L. Heilbron. Quantum historiography and the archive for history of quantum physics. History of Science (UK), 7 (1):90–111, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/7/1/90.full.pdf+html.

Heilbron:2005:OG

[Hei05] J. L. Heilbron. Oppenheimer’s guru. In Carson and Hollinger [CH05], chapter 13, pages 275–292. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Hendry:1981:BRT

[Hen81] John Hendry. Book review: Twentieth Century The Uranium People. By Leona Marshall Libby. New York: Crane Russak; Charles Scribner’s Sons, 1979. PP. x + 341. $15.95. Scientists in Power. By Spencer R. Weart. Cambridge, Mass. and London: Harvard University Press, 1979. PP. xiii + 343. $17.50. Robert Oppenheimer: Letters and Recollections. Edited by Alice Kimball Smith and Charles Weiner. Cambridge, Mass. and London: Harvard University Press, 1980. Pp. xi + 376. $20.00. British Journal for the History of Science, 14(1):97–99, March 1981. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4026084.

Herken:2002:BBT

[Her02] Gregg Herken. Brotherhood of the bomb: the tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. Henry Holt and Company, Incorporated, New York, NY, USA, 2002. ISBN 0-8050-6588-1 (hardcover). xiv + 448 pp. LCCN QC16.O62 H47

2002. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://alsos.wlu.edu/information.aspx?
id=1264;   http://www.loc.gov/catdir/bios/hol051/2002017219.
html;   http://www.loc.gov/catdir/description/hol021/
2002017219.html.

Herken:2003:CBR

[Her03] Gregg Herken. Comment on book review of Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken [Am. J. Phys. 71(4), 411–415 (2003)]. American Journal of Physics, 71(7):647–648, July 2003. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v71/i7/p647_s1;   http://link.aip.

org/link/?AJP/71/647/1. See [Ber03c]. Includes a response by reviewer Jeremy Bernstein.

Herken:2004:JRO

[Her04a] G. Herken. J. Robert Oppenheimer and the American century. Science, 306(5699):1137, November 12, 2004. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Herken:2004:OCE

[Her04b] Gregg Herken. The Oppenheimer case: An exchange. New York Review of Books, 51(5):46–47, March 25, 2004. ISSN 0028-7504 (print), 1944-7744 (electronic). With a reply by Daniel J. Kevles.

Herken:2005:WRO

[Her05a] Gregg Herken. Was Robert Oppenheimer a “closet Communist”? The debate and the evidence. In Carson and Hollinger [CH05], chapter 3, pages 51–56. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Hershberg:2005:JWJ

[Her05b] James G. Hershberg. “The jig was up”: J. Robert Oppenheimer and the international control of atomic energy, 1947–49. In Carson and Hollinger [CH05], chapter 8, pages 149–184. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Herken:2007:JRO

[Her07a] Gregg Herken. J. Robert Oppenheimer: A Life (review). Technology and Culture, 48(4):889–891, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/223361.

Herken:2007:OPE

[Her07b] Gregg Herken. Oppenheimer: Portrait of an Enigma (review). Journal of Cold War Studies, 9(1):98–99, Winter 2007. CODEN ???? ISSN 1520-3972 (print), 1531-3298

(electronic). URL http://muse.jhu.edu/article/210081;
http://muse.jhu.edu/article/210081/pdf;   http://muse.
jhu.edu/journals/journal_of_cold_war_studies/summary/
v009/9.1herken.html;   https://www.math.utah.edu/pub/
bibnet/authors/o/oppenheimer-j-robert.bib.

Herken:2009:TES

[Her09] Gregg Herken. Target Enormoz: Soviet nuclear espionage on the West Coast of the United States, 1942–1950. Journal of Cold War Studies, 11(3):68–90, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://

muse.jhu.edu/article/268391.

Hevly:2011:JRO

[Hev11] Bruce Hevly. J. Robert Oppenheimer, the Cold War, and the Atomic West. New Mexico Historical Review, 86(2):263–265, Spring 2011. ISSN 0028-6206.

Hewlett:1966:BRB

[Hew66] Richard G. Hewlett. Book reviews: A Peril and a Hope: The Scientists’ Movement in America, 1945-47, by Alice Kimball Smith; Day of Trinity, by Lansing Lamont. Technology and Culture, 7(2):250–252, Spring 1966. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

Hewlett:1989:APW

[HH89] Richard G. Hewlett and Jack M. Holl. Atoms for peace and war, 1953–1961: Eisenhower and the Atomic Energy Commission, volume 3 of California studies in the history of science; A history of the United States Atomic Energy Commission. University of California Press, Berkeley, CA, USA, 1989. ISBN 0-520-06018-0. xxix + 696 pp. LCCN HD9698.U52 H55 1989 vol. 3; QC792.7.H48

1989. URL http://energy.gov/sites/prod/files/2013/
08/f2/HewlettandHollAtomsforPeaceandWarComplete.pdf.

With a foreword by Richard S. Kirkendall and an essay on sources by Roger M. Anders.

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711 pp. LCCN QC15 .H295 2004. URL

http://www.worldscibooks.com/physics/p304.html.

Hargittai:2004:EPW

[HH04b] Magdolna Hargittai and István Hargittai. Eugene P. Wigner. In Candid science IV: conversations with famous physicists [HH04a], pages 1–19. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). LCCN QC15 .H295 2004. URL http://www.

worldscibooks.com/physics/p304.html.

Hoddeson:1993:CAT

[HHMW93] Lillian Hoddeson, Paul W. Henriksen, Roger A. Meade, and Catherine L. Westfall, editors. Critical assembly: a technical history of Los Alamos during the Oppenheimer years, 1943–1945. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44132-3 (hardcover). xv + 509 pp. LCCN QC773.3.U5 C75 1993. URL http://www.loc.

gov/catdir/description/cam025/92036611.html;   http:
//www.loc.gov/catdir/toc/cam029/92036611.html. With

contributions by Gordon Baym, Richard Hewlett, Alison Kerr, Robert Penneman, Leslie Redman, and Robert Seidel.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Higinbotham:1976:BRH

[Hig76] William A. Higinbotham. Book review: H. York, The Advisors: Oppenheimer, Teller and the Superbomb. Physics Today, 29(12): 49–50, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hijiya:2000:GJR

[Hij00] James A. Hijiya. The “Gita” of J. Robert Oppenheimer. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 144(2):123–167, June 2000. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://www.amphilsoc.org/
sites/default/files/Hijiya.pdf;   https://www.jstor.
org/stable/1515629.

Hiltzik:2015:BSE

[Hil15] Michael A. Hiltzik. Big Science: Ernest Lawrence, the Cyclotron, and the Birth of the Military–Industrial Complex. Simon and Schuster, New York, NY, USA, 2015. ISBN 1-4516-7575-5 (hardcover), 1-4516-7576-3 (paperback), 1-4516-7603-4 (e-book). x + 512 pp. LCCN QC787.C8 H55 2015.

Hirschfelder:1980:STM

[Hir80] Joseph O. Hirschfelder. The scientific and technological miracle at Los Alamos. In Badash et al. [BHB80], pages 67–88. ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Hirsch:1988:BRB

[Hir88] Daniel Hirsch. Book review: Better a Shield than a Sword: Perspectives on Defense and Technology, by Edward Teller. Bulletin of the Atomic Scientists, 44(4):48–49, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haynes:2010:SRF

[HKV10] John Earl Haynes, Harvey Klehr, and Alexander Vassiliev. Spies: the rise and fall of the KGB in America. Yale University Press, New Haven, CT, USA, 2010. ISBN 0-300-16438-6 (paperback). LCCN UB271.R9 H389 2009.

Hall:1930:WDM

[HO30] Harvey Hall and J. R. Oppenheimer. 1. why does molecular hydrogen reach equilibrium so slowly? Physical Review (2), 35 (1):132–133, January 1, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v35/i1/p132_1. In Minutes of the Stanford

Meeting, December 7, 1929.

Hall:1931:PAU

[HO31a] Harvey Hall and J. R. Oppenheimer. Photoelectric absorption of ultra-gamma radiation. Physical Review (2), 38(3):589, August 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Hall:1931:RTP

[HO31b] Harvey Hall and J. R. Oppenheimer. Relativistic theory of the photoelectric effect. Part I. Theory of the K-absorption of X-rays. Part II. Photoelectric absorption of ultragamma radiation. Physical Review (2), 38(1):57–79, July 1, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v38/i1/p57_1.

Hanson:2021:WRT

[HO21] Susan K. Hanson and Warren J. Oldham. Weapons radiochemistry: Trinity and beyond. Nuclear Technology, 207(S1):S295– S308, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Hoffmann:1995:JRO

[Hof95] Klaus Hoffmann. J. Robert Oppenheimer: Schöpfer der ersten Atombombe. (German) [J. Robert Oppenheimer: Creator of the first atomic bomb]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-642-57815-2, 3-642-63364-1. viii + 307 pp. LCCN ????

Hoffmann:1996:RKH

[Hof96] Dieter Hoffmann. Rezension: Klaus Hoffmann: J. Robert Oppenheimer — Schöpfer der ersten Atombombe. Berlin/ Heidelberg/New York: Springer-Verlag 1995, VII und 307 Seiten, 43 Abbildungen. Gebunden 48,- DM. Berichte zur Wissenschaftsgeschichte, 19(2–3):196, ???? 1996. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Holton:1981:YMO

[Hol81] Gerald Holton. Young man Oppenheimer. Partisan Review, 48 (??):380–388, July 1981. ISSN 0031-2525.

Holton:1984:SSM

[Hol84] Gerald Holton. Success sanctifies the means: Heisenberg, Oppenheimer, and the transition to modern physics. In Mendelsohn [Men84], pages 155–173. ISBN 0-521-26724-

2. LCCN Q126.8 .T73 1984. URL http://www.loc.

gov/catdir/description/cam022/84007832.html;   http:
//www.loc.gov/catdir/toc/cam028/84007832.html.

Holton:1986:ASB

[Hol86] Gerald James Holton. The advancement of science, and its burdens: the Jefferson Lecture and other essays. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25244-X, 0-521-27243-2 (paperback). xi + 351 pp. LCCN Q173

.H734 1986. URL http://www.loc.gov/catdir/description/
cam032/86009722.html.

Holloway:1993:MJR

[Hol93] Rachel L. Holloway. In the matter of J. Robert Oppenheimer: politics, rhetoric, and self-defense. Praeger, Westport, CT, USA, 1993. ISBN 0-275-94429-8. x + 123 pp. LCCN QC16.O62 H65 1993.

Hollinger:1996:SJS

[Hol96] David A. Hollinger. Science, Jews, and Secular Culture: Studies in Mid-twentieth-century American Intellectual History. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01143-5 (hardcover), 0-691-00189-8 (paperback), 1-4008-4774-5 (e-book). xi + 178 pp. LCCN E184.J5 H646 1996. URL http:/

/catdir.loc.gov/catdir/description/prin031/95039843.
html;   http://catdir.loc.gov/catdir/toc/prin031/95039843.
html;   http://www.h-net.org/review/hrev-a0a6h8-aa.

Holton:1998:ASB

[Hol98] Gerald James Holton. The advancement of science, and its burdens: with a new introduction. Harvard University Press, Cambridge, MA, USA, 1998. ISBN 0-674-00530-9 (paperback). xlix + 352 pp. LCCN QK596.C8 Y46.

Hollinger:2002:BBT

[Hol02a] D. A. Hollinger. Brotherhood of the bomb: the tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. New York Times Book Review, ??(??):10, September 15, 2002. ISSN 0028-7806.

Hollinger:2002:HOW

[Hol02b] David A. Hollinger. The hunt for Oppenheimer: Why the man who led the effort to build the American atom bomb was pushed out of government service. New York Times Book Review, ??(??):7.10, September 15, 2002. ISSN 0028-7806. URL

http://search.proquest.com/docview/217302653. Review

of [Her02].

Hollinger:2005:A

[Hol05a] David A. Hollinger. Afterword. In Carson and Hollinger [CH05], pages 385–390. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Holloway:2005:PLO

[Hol05b] David Holloway. Parallel lives? Oppenheimer and Khariton. In Carson and Hollinger [CH05], chapter 6, pages 115–128. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Horton:1970:BRP

[Hor70] Gerald Horton. Book review: Philip M. Stern and Harold P. Green, The Oppenheimer case: security on trial. New York Times, ??(??):??, January 4, 1970. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Hahn:1939:NVB

[HS39] Otto Hahn and Fritz Strassmann. Über den Nachweis und das Verhalten der bei der Bestrahlung des Urans mittels Neutronen entstehenden Erdalkalimetalle. (German) [Concerning the existence of alkaline earth metals resulting from the neutron irradiation of uranium]. Naturwissenschaften, 27(1):11–15, January 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). A facsimile is also available in [Bey49, pages 87– 91] and in [Gra64]. Abridged English translation in [GA71, pages 44–47].

Hawkins:1983:MDH

[HTS83] David Hawkins, Edith C. Truslow, and Ralph Carlisle Smith, editors. Manhattan District history, Project Y, the Los Alamos story, volume 2 of History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-08-0. xxvi + 506 + 9 pp. LCCN QC773.3.U5 M25 1983.

Hufbauer:1969:BRB

[Huf69] Karl Hufbauer. Book review: Lawrence and Oppenheimer. Isis, 60(4):550–551, Winter 1969. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/229114.

Hufbauer:2005:JRO

[Huf05] Karl Hufbauer. J. Robert Oppenheimer’s path to black holes. In Carson and Hollinger [CH05], chapter 2, pages 31–48. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Hughes:1970:COC

[Hug70] H. Stuart Hughes. Closing the Oppenheimer case: book reviews of The Oppenheimer Case: Security on Trial by Philip M. Stern, Harper & Row, 591 pp. The Swift Years: The Robert Oppenheimer Story by Peter Michelmore, Dodd, Mead, 273 pp. The Great Weapons Heresy by Thomas W. Wilson Jr., Houghton, Mifflin, 275 pp. Oppenheimer: The Story of a Friendship by Haakon Chevalier, Braziller, 219 pp. Lawrence and Oppenheimer by Nuel Pharr Davis, Simon & Schuster, 384 pp. New York Review of Books, ??(??): ??, July 2, 1970. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/1970/07/
02/closing-the-oppenheimer-case/. See comments [Str70].

Hughes:1989:AGC

[Hug89] Thomas Parke Hughes. American genesis: a century of invention and technological enthusiasm, 1870–1970. Viking, New York, NY, USA, 1989. ISBN 0-670-81478-4. xii + 529 pp. LCCN T21 .H82 1989. US$24.95.

Hughes:1990:AGC

[Hug90] Thomas Parke Hughes. American genesis: a century of invention and technological enthusiasm, 1870–1970. Penguin, New York, NY, USA, 1990. ISBN 0-14-009741-4. xii + 529 pp. LCCN T21 .H78 1990. US$10.95, CAN$12.95. URL http://www.loc.

gov/catdir/enhancements/fy1209/89039468-b.html.

Hughes:2004:AGC

[Hug04] Thomas Parke Hughes. American genesis: a century of invention and technological enthusiasm, 1870–1970. University of Chicago Press, Chicago, IL, USA and London, UK, 2004. ISBN 0-226-35927-1 (paperback). xxxii + 529 pp. LCCN T21 .H82 2004. URL http://www.

loc.gov/catdir/bios/uchi051/2003067221.html;   http:
//www.loc.gov/catdir/description/uchi052/2003067221.
html;   http://www.loc.gov/catdir/toc/uchi051/2003067221.
html.

Hughes:2007:BRC

[Hug07] Jeff Hughes. Book review: Cathryn Carson and David A. Hollinger (eds.), Reappraising Oppenheimer: Centennial Studies and Reflections. Berkeley Papers in History of Science, Volume 21. Berkeley: Office for History of Science and Technology, University of California, 2005. Pp. xii + 413. ISBN 0-9672617-3-2. $14.00 (paperback). British Journal for the History of Science, 40(3):459–460, September 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4500763.

Hughes:2009:BRC

[Hug09] Jeff Hughes. Book review: Charles Thorpe, Oppenheimer: The Tragic Intellect. Chicago and London: University of Chicago Press, 2006. Pp. xx + 413. ISBN 0-226-79845-3. $37.50, £24.00 (hardback). British Journal for the History of Science, 42(1): 127–128, March 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Hunner:2009:JRO

[Hun09] Jon Hunner. J. Robert Oppenheimer, the Cold War, and the Atomic West, volume 24 of The Oklahoma Western biographies. University of Oklahoma Press, Norman, OK, USA, 2009. ISBN 0-8061-4046-1 (hardcover). xvi + 248 pp. LCCN QC16.O62 H86

Hunner:2014:JRO

[Hun14] Jon Hunner. J. Robert Oppenheimer, the Cold War, and the Atomic West, volume 24 of The Oklahoma Western biographies. University of Oklahoma Press, Norman, OK, USA, 2014. ISBN 0-8061-4046-1 (hardcover), 0-8061-8577-5. xvi + 248 pp. LCCN QC16.O62 H86 2009.

Ilha:1999:DCO

[IKL99] Anderson Ilha, Antares Kleber, and José P. S. Lemos. Dimensionally continued Oppenheimer–Snyder gravitational collapse: Solutions in odd dimensions. Journal of Mathematical Physics, 40(7):3509–3518, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Infeld:1978:O

[Inf78a] Leopold Infeld. Oppenheimer. In Why I left Canada: reflections on science and politics [Inf78b], pages 160–180. ISBN 0-7735-0272-6. LCCN QC16.I6 A3213. Infeld makes a strong rebuttal in support of Albert Einstein against a speech made by J. Robert Oppenheimer on 14 December 1965. He also records harsh criticism by Philip Morrison of Oppenheimer.

Infeld:1978:WLC

[Inf78b] Leopold Infeld. Why I left Canada: reflections on science and politics. McGill-Queen’s University Press, Montréal, Québec, Canada, 1978. ISBN 0-7735-0272-6. xii + 212 pp. LCCN QC16.I6 A3213. Translations from Polish to English by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild (1921–1977).

Isaacson:2007:EHL

[Isa07] Walter Isaacson. Einstein: his life and universe. Simon and Schuster, New York, NY, USA, 2007. ISBN 0-7432-6473-8. xxii + 675 + 16 pp. LCCN QC16.E5 I76 2007. URL http://www.

loc.gov/catdir/enhancements/fy0708/2006051264-d.html.

Jacoby:1981:JRO

[Jac81] T. Jacoby. J. Robert Oppenheimer — shatterer of worlds. New York Times Book Review, 86(28):13–14, ???? 1981. ISSN 0028-

Jecko:2014:MTB

[Jec14] Thierry Jecko. On the mathematical treatment of the Born– Oppenheimer approximation. Journal of Mathematical Physics, 55(5):053504, May 2014. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Johnston:1959:DIM

[Joh59] J. Johnston. Description of the IBM machines at Los Alamos. Report LA-1057, Los Alamos Scientific Laboratory, Los Alamos, NM, June 15, 1959.

Johnson:2013:BRW

[Joh13] George Johnson. Book review: What made him tick: Robert Oppenheimer: A Life Inside the Center, by Ray Monk. New York Times, ??(??):BR.22–??, June 30, 2013. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL

http://search.proquest.com/docview/1399140783/.

JC:1954:CRA

[Joi54] Joint Committee. Calculated risks in the AEC. Bulletin of the Atomic Scientists, 10(7):263, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1985:MAA

[Jon85] Vincent C. Jones. Manhattan, the Army and the atomic bomb. United States Army in World War II. Special studies. Center of Military History, U.S. Army, Washington, DC, USA, 1985. xx + 660 + 4 pp. LCCN QC773.3.U5 J65 1985.

Jones:2010:THR

[Jon10] Derry W. Jones. Titanic Hungarian refugee physicists. Contemporary Physics, 51(3):267–271, May/June 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Jordan:1927:PFQ

[Jor27] Pascual Jordan. Philosophical foundations of quantum theory. Nature, 119(2998):566–569, April 16, 1927. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). From the footnote on first page: “Habilitationsvortrag at the University of Göttingen. Translated by Mr. Robert Oppenheimer. The author is very much obliged to Mr. Oppenheimer for his careful translation.”.

Joravsky:1980:SSB

[Jor80] David Joravsky. Sin and the scientist: Robert Oppenheimer — Letters and Recollections, edited by Alice Kimball Smith, edited by Charles Weiner, Harvard University Press, 376 pp., $20.00. New York Review of Books, 27(12):7– 10, July 17, 1980. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1980/jul/17/sin-and-the-scientist/.

Julian:2006:BRB

[Jul06] T. A. Julian. Book review: American Prometheus: The triumph and tragedy of J. Robert Oppenheimer. The Journal of Military History, 70(1):201–205, January 2006. ISSN 0899-3718 (print), 1543-7795 (electronic).

Kaiser:2002:EOF

[Kai02] David Kaiser. Einstein, Oppenheimer, Feynman: Physics in the 20th Century. Technical Report MIT OpenCourseWare STS.042J / 8.225J, MIT, Cambridge, MA, USA, 2002. Free, independent study course that explores the changing roles of physics and physicists during the 20th Century.

Kaiser:2005:ASR

[Kai05] David Kaiser. The atomic secret in red hands? American suspicions of theoretical physicists during the early Cold War. In Carson and Hollinger [CH05], chapter 9, pages 185–216. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Kalven:1954:CJR

[Kal54] Harry Kalven, Jr. The case of J. Robert Oppenheimer before the Atomic Energy Commission. Bulletin of the Atomic Scientists, 10(7):259–269, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Kal63].

Kalven:1963:CJR

[Kal63] Harry Kalven, Jr. The case of J. Robert Oppenheimer before the Atomic Energy Commission. In Grodzins and Rabinowitch [GR63], pages 442–465. LCCN D842 .B78. Reprint of [Kal54].

Kamen:1985:RSD

[Kam85] Martin David Kamen. Radiant science, dark politics: a memoir of the nuclear age. University of California Press, Berkeley, CA, USA, 1985. ISBN 0-520-04929-2 (hardcover), 0-520-05897-6 (paperback). xii + 348 + 8 pp. LCCN QH31.K277 A37 1985. Foreword by Edwin M. McMillan.

Kammen:2023:WWH

[Kam23] Daniel M. Kammen. Why what happened to Oppenheimer then is relevant now. Bulletin of the Atomic Scientists, 79(4):267–271, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2223087.

Kant:1985:JRO

[Kan85] Horst Kant. J. Robert Oppenheimer. B. G. Teubner, Stuttgart, Germany; Leipzig, Germany, 1985. 176 pp.

Kapitsa:1996:ONW

[Kap96] Serguei Kapitsa. Our nonlinear world. In Behind tall fences: stories and experiences about Los Alamos at its beginning [Ano96], pages 182–196. ISBN 0-941232-91-3. LCCN QC773.3.U5 B43 1996. Twenty-third J. Robert Oppenheimer Lecture, August 1993.

Kargon:1981:ROL

[Kar81] Robert Kargon. Robert Oppenheimer: Letters and Recollections ed. by Alice Kimball Smith, Charles Weiner (review). Technology and Culture, 22(3):649–651, July 1981. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:/

/muse.jhu.edu/pub/1/article/890917/pdf.

Katz:2021:FT

[Kat21] J. I. Katz. Fermi at Trinity. Nuclear Technology, 207(S1):S326– S334, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Kelly:2006:OMP

[Kel06] Cynthia C. Kelly, editor. Oppenheimer and the Manhattan Project: insights into J. Robert Oppenheimer, “Father of the atomic bomb”. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-418-7 (hardcover), 981-256-599-X (paperback). vii + 173 pp. LCCN QC16.O62 O66 2006. URL http://www.loc.

gov/catdir/toc/fy0702/2006283849.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5897.

Kempton:1983:AJR

[Kem83] Murray Kempton. The ambivalence of J. Robert Oppenheimer. Esquire, ??(??):??, December 1983.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of Hbomb at Pacific Proving Ground. New York Times, ??(??):E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/113104238/.

Kennan:1967:O

[Ken67] George Kennan. Oppenheimer. Encounter, 18(??):55–56, ???? 1967. ISSN 0013-7073.

Kennan:1994:DO

[Ken94] George F. Kennan. In defense of Oppenheimer. New York Review of Books, ??(??):??, June 23, 1994. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/1994/06/23/in-defense-of-oppenheimer/.

Kevles:1981:BRA

[Kev81] Daniel J. Kevles. Book review: Alice Kimball Smith and Charles Weiner: Robert Oppenheimer: Letters and Recollections. Isis, 72 (2):330, June 1981. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
231034.

Kevles:2003:SCR

[Kev03] Daniel J. Kevles. The strange case of Robert Oppenheimer: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller by Gregg Herken, Henry Holt, 448 pp., $30.00. Pandora’s Keepers: Nine Men and the Atomic Bomb by Brian VanDeMark Little, Brown, 339 pp. (withdrawn). New York Review of Books, 50(19):37–40, December 4, 2003. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http:

//www.nybooks.com/articles/archives/2003/dec/04/the-
strange-case-of-robert-oppenheimer/.

Kevles:2005:APT

[Kev05a] D. J. Kevles. American prometheus: The triumph and tragedy of J. Robert Oppenheimer. New Republic, 233(16):35–41, October 17, 2005. ISSN 0028-6583.

Kevles:2005:SAS

[Kev05b] Daniel J. Kevles. Scientists, arms, and the state: J. Robert Oppenheimer and the Twentieth Century. In Carson and Hollinger [CH05], chapter 16, pages 327–342. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Kevles:2013:ICL

[Kev13] Daniel J. Kevles. Inside the centre: The life of J. Robert Oppenheimer. TLS — The Times Literary Supplement, ??(5742): 3–5, April 19, 2013. ISSN 0307-661X.

Kimpland:2021:CAD

[KGJ^{+}21] Robert Kimpland, Travis Grove, Peter Jaegers, Richard Malenfant, and William Myers. Critical assemblies: Dragon burst assembly and solution assemblies. Nuclear Technology, 207(S1): S81–S99, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Killian:1977:SSE

[Kil77] James Rhyne Killian, Jr. Sputnik, scientists, and Eisenhower: a memoir of the First Special Assistant to the President for Science and Technology. MIT Press, Cambridge, MA, USA, 1977. ISBN 0-262-11066-0. xix + 315 pp. LCCN Q143.K42 A37.

Kipphardt:1964:SJR

[Kip64] Heinar Kipphardt. In der Sache J. Robert Oppenheimer: ein szenischer Bericht. (German) [In the matter of J. Robert Oppenheimer: a scenic report]. Suhrkamp, Frankfurt am Main, West Germany, 1964. ???? pp.

Kipphardt:1967:CJR

[Kip67a] Heinar Kipphardt. En cause: J. Robert Oppenheimer. (French) [In the case of J. Robert Oppenheimer]. L’Arche, Paris, France, 1967. ISBN 2-85181-130-4. 163 pp. LCCN ????

Kipphardt:1967:MJR

[Kip67b] Heinar Kipphardt. In the Matter of J. Robert Oppenheimer: a Play Freely Adapted on the Basis of Documents. Methuen, London, UK, 1967. 107 pp. LCCN ????

Kipphardt:1968:SJR

[Kip68a] Heinar Kipphardt. In der Sache J. Robert Oppenheimer. (German) [In the case of J. Robert Oppenheimer]. Suhrkamp Verlag, Frankfurt am Main, Germany, sixth edition, 1968. 143 pp. LCCN ????

Kipphardt:1968:MJR

[Kip68b] Heinar Kipphardt. In the Matter of J. Robert Oppenheimer: a Play Freely Adapted on the Basis of Documents. Hill and Wang, New York, NY, USA, 1968. ???? pp. LCCN ????

Kipphardt:1987:GWE

[Kip87] Heinar Kipphardt. Gesammelte Werke in Einzelausgaben = Heinar Kipphardt — In der Sache J. Robert Oppenheimer. (German) [Collected works in single issue = Heinar Kipphardt In the Matter of J. Robert Oppenheimer], volume 3 of Rororo; 12111. Rowohlt, Reinbek bei Hamburg, Germany, 1987. ISBN 3-499-12111-5. 300 pp. LCCN ????

Kipphardt:2005:SJR

[Kip05] Heinar Kipphardt. In der Sache J. Robert Oppenheimer: Schauspiel. (German). [In the case of J. Robert Oppenheimer: a play], volume 58 of Suhrkamp BasisBibliothek. Suhrkamp, Frankfurt am Main, Germany, 2005. ISBN 3-518-18858-5 (paperback). 190 pp. LCCN PT2671.I6 I5 2005.

Kistiakowsky:1980:RWA

[Kis80a] George B. Kistiakowsky. Reminiscences of wartime Los Alamos. In Badash et al. [BHB80], page ?? ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Kistiakowsky:1980:TR

[Kis80b] George B. Kistiakowsky. Trinity — a reminiscence. Bulletin of the Atomic Scientists, 36(6):19–22, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kipphardt:1966:COS

[KL66] Heinar Kipphardt and Adolfo Lozano Borroy. El caso Oppenheimer. (Spanish) [The Oppenheimer case]. Voz e imagen: teatro; 12. Aymá, Barcelona, Spain, 1966. 148 pp. LCCN ????

Kleppner:2007:BRA

[Kle07] Daniel Kleppner. Book review: Abraham Pais, J. Robert Oppenheimer: A Life, with supplemental material by Robert P. Crease. Oxford: Oxford University Press, 2006, xxii + 353 pages. $30.00 (cloth). Physics in Perspective (PIP), 9(4):505– 508, December 2007. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Knust:1983:FOS

[Knu83] Herbert Knust. From Faust to Oppenheimer : the scientist’s pact with the devil. European studies, 13(1–2):122–141, ???? 1983. CODEN ???? ISSN ????

Kohler:1976:WKP

[Köh76] Wolfgang Köhler. Wolfgang Köhler papers 1862–1976. American Philosophical Society Library archive, 1976. URL http://

amphilsoc.org/mole/view?docId=ead/Mss.B.K815-ead.xml.

8.0 Linear feet, Mss.B.K815. From the Web site: “After the dismissal of Jewish professors Köhler took a public stand in April 1933 with “Gespräche in Deutschland.” [(German) [Speech in Germany]] Published in Die Deutsche Allgemeine Zeitung, this was the last anti-Nazi article published openly in Germany, and Köhler expected to be arrested soon afterwards. In 1934 Köhler made his second trip to the United States to teach as the William James Lecturer at Harvard University (1934-1935), and as Visiting Professor at the University of Chicago (1935). At the end of his time in Chicago, Köhler decided to formally emigrate to the United States”.

Kojevnikov:2005:MSB

[Koj05] Alexei Kojevnikov. The making of the Soviet bomb and the shaping of Cold War science. In Carson and Hollinger [CH05], chapter 7, pages 129–146. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Kolb:2007:UTE

[Kol07] Rocky Kolb. A unified theory of Einstein. Bulletin of the Atomic Scientists, 63(6):58–59, November/December 2007. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/63/6/58.abstract;
 http://bos.sagepub.com/content/63/6/58.full.pdf+html.

Review of [Isa07].

Kalckar:1937:NNP

[KOS37a] F. Kalckar, J. R. Oppenheimer, and R. Serber. Note on nuclear photoeffect at high energies. Physical Review (2), 52(4):273– 278, August 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v52/i4/p273_1.

Kalckar:1937:NRT

[KOS37b] F. Kalckar, J. R. Oppenheimer, and R. Serber. Note on resonances in transmutations of light nuclei. Physical Review (2), 52 (4):279–282, August 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v52/i4/p279_1.

Kragh:2017:NPS

[Kra17] Helge Kragh. The Nobel Prize system and the astronomical sciences. Journal for the History of Astronomy, 48(3):257–280, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-

8556 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0021828617721574.

Kroplin:1981:KSJ

[Kro81] W. Kroplin. Kipphardt ‘In der Sache J. Robert Oppenheimer’. (German) [Kipphardt In the Matter of J. Robert Oppenheimer]. Theater der Zeit, 36(11):2, ???? 1981. ISSN 0040-5418.

Krupar:2007:EPR

[Kru07] Jason N. Krupar. 109 East Palace: Robert Oppenheimer and the Secret City of Los Alamos (review). Journal of Cold War Studies, 9(4):143–145, Fall 2007. CODEN ????

ISSN 1520-3972 (print), 1531-3298 (electronic). URL http:/

/muse.jhu.edu/article/223550;   http://muse.jhu.edu/
article/223550/pdf;   http://muse.jhu.edu/journals/journal_
of_cold_war_studies/v009/9.4krupar.html.

Krupar:2015:ROL

[Kru15] Jason Krupar. Robert Oppenheimer: A Life Inside the Center by Ray Monk (review). Technology and Culture, 56(3):774– 775, July 2015. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/593086.

Kugelmass:1953:JRO

[Kug53] J. Alvin Kugelmass. J. Robert Oppenheimer and the atomic story. Julian Messner, New York, NY, USA, 1953. 179 pp. LCCN QC16.O62 K8. Decorations by William Metzig.

Kuhn:1963:IJR

[Kuh63] Thomas S. Kuhn. Interview with J. Robert Oppenheimer by Thomas S. Kuhn, November 18, 1963. Archive for the History of Quantum Physics, ??(??):??, 1963.

Kunetka:1978:CFA

[Kun78] James W. Kunetka. City of fire: Los Alamos and the birth of the Atomic Age, 1943–1945. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ISBN 0-13-134635-0. 234 + 8 pp. LCCN QC792.8.U6 L674 1978. US$10.95.

Kunetka:1982:OYR

[Kun82] James W. Kunetka. Oppenheimer, the years of risk. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-638007-7. xii + 292 + 4 pp. LCCN QC16.O62 K86 1982.

Kunetka:2015:GGG

[Kun15] James W. Kunetka. The general and the genius: Groves and Oppenheimer: the unlikely partnership that built the atom bomb. Regnery Publishing, Washington, DC, USA, 2015. ISBN 1-62157-338-9. xiv + 482 + 16 pp. LCCN QC773.A1 K87 2015.

Kipphardt:1969:MJR

[KW69] Heinar Kipphardt and Joseph Wiseman. In the matter of J. Robert Oppenheimer [Sound recording]. Caedmon TRS 336, ????, 1969. LCCN ????

Laguardia:1995:DAT

[Lag95] C. Laguardia. The day after Trinity — J. Robert Oppenheimer and the atomic bomb. Library Journal, 120(16):126, October 2, 1995. CODEN LIBJA7. ISSN 0363-0277.

Lakoff:1966:TDO

[Lak66] Sanford A. Lakoff. The trial of Dr. Oppenheimer. In Sanford A. Lakoff, editor, Knowledge and Power: Essays on Science and Government, page ?? Free Press, New York, NY, USA, 1966. LCCN Q125 .L28.

Lamont:1965:DT

[Lam65] Lansing Lamont. Day of Trinity. Atheneum, New York, NY, USA, 1965. xi + 333 pp. LCCN QC773.A1 L3.

Lamont:1966:DT

[Lam66a] Lansing Lamont. Day of Trinity. Hutchinson, London, UK, 1966. xii + 333 pp. LCCN QC773.A1 L3 1966b.

Lamont:1966:EVW

[Lam66b] Lansing Lamont. Eine Explosion verändert die Welt die Geschichte der ersten Atombombe. (German) [An explosion that changed the world: the history of the first atomic bomb]. Piper, München, West Germany, 1966. 268 pp. LCCN ????

Lamont:1968:DT

[Lam68] Lansing Lamont. El día de Trinidad. (Spanish) [Day of Trinity]. Diana, México, DF, México, 1a. edition, 1968. 365 + 4 pp. LCCN ????

Lamont:1985:DT

[Lam85] Lansing Lamont. Day of Trinity. Atheneum, New York, NY, USA, 1985. ISBN 0-689-70686-3 (paperback). 363 + 10 pp. LCCN QC773.A1 L3 1985.

Lamont:1989:DT

[Lam89] Lansing Lamont. Day of Trinity. Easton Press, Norwalk, CT, USA, 1989. 363 + 8 pp. LCCN QC773.A1 L36 1985.

Lanouette:1994:AS

[Lan94] William Lanouette. Atomic spies. Debate with the authors of Special Tasks by Soviet spymaster Pavel Sudoplatov [SSSS95], which falsely alleged that Szilard, Niels Bohr, J. Robert Oppenheimer and Enrico Fermi were Soviet agents within the Manhattan Project. MacNeil/Lehrer NewsHour, April 26, 1994.

Lanouette:2000:BRB

[Lan00] William Lanouette. Book review: In The Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist, by S. S. Schweber. The Wilson Quarterly (1976–2012), 24(4):138, Autumn 2000. ISSN 0363-3276. URL http://www.

jstor.org/stable/40260132.

Lanouette:2004:BRW

[Lan04] William Lanouette. Book review: Where’s Oppie? Oppenheimer: Portrait of an Enigma, by Jeremy Bernstein. Chicago: Ivan R. Dee, 2004, 240 pp. Issues in Science and Technology, 21(1):??, Fall 2004. ISSN 0748-5492 (print), 1938-1557 (elec-

tronic). URL http://issues.org/21-1/br_lanouette/.

Lanouette:2005:DDB

[Lan05] W. Lanouette. Dissenters disquieted (‘Building the bomb’, J. Robert Oppenheimer). Smithsonian, 36(7):16, October 2005. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Larsen:1988:OAB

[Lar88] Rebecca Larsen. Oppenheimer and the atomic bomb. Franklin Watts, New York, NY, USA, 1988. ISBN 0-531-10607-1. 192 pp. LCCN QC16.O62 L37 1988.

Laurence:1945:UAB

[Lau45] William Leonard Laurence. U.S. atom bomb site belies Tokyo tales: Tests on New Mexico range confirm that blast, and not radiation, took toll. New York Times, ??(??):1, 4, September 12, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/107140479/.

Laurence:1946:DZS

[Lau46] William Leonard Laurence. Dawn over zero; the story of the atomic bomb. Alfred A. Knopf, New York, NY, USA, 1946. xii + 274 + viii pp. LCCN UF767 .L3.

Laurence:1947:DZ

[Lau47a] William Leonard Laurence. Dawn over zero. Alfred A. Knopf, New York, NY, USA, second edition, 1947. xv + 289 + viii pp. LCCN UF767 .L3 1947.

Laurence:1947:DZS

[Lau47b] William Leonard Laurence. Dawn over zero; the story of the atomic bomb. Museum Press, London, UK, 1947. xvii + 251 pp. LCCN UF767 .L3 1947a.

Laurence:1947:HAA

[Lau47c] William Leonard Laurence. De Hemmelige Aar, Atomalderens Fødsel. (Danish) [The Secret Years: the Birth of the Atomic Age]. J. H. Schultz, København, Danmark, 1947. 243 pp. LCCN UF767 .L312.

Laurence:1947:DN

[Lau47d] William Leonard Laurence. Demring over nullpunkett. (Norwegian) [Dawn over zero]. Bergendahl, Oslo, Norway, 1947. 268 pp. LCCN UF767 .L316.

Laurence:1954:OCD

[Lau54] William L. Laurence. Oppenheimer case divides physicists: Subjecting of atomic expert to ‘security risk’ hearing debated by colleagues. New York Times, ??(??):13, April 30, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/112940304/.

Laurence:1956:PGS

[Lau56] William L. Laurence. Physics goal set by Oppenheimer: Atomic scientist envisions quest for explaining new order in the universe questions to be answered obligations for the teacher. New York Times, ??(??):13, February 3, 1956. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/113908716/.

Laurence:1959:MAD

[Lau59] William Leonard Laurence. Men and atoms; the discovery, the uses, and the future of atomic energy. Simon and Schuster, New York, NY, USA, 1959. 302 pp. LCCN QC773 .L3.

Laurence:1965:WYM

[Lau65] William L. Laurence. Would you make the bomb again? The New York Times Magazine, ??(??):8–??, August 1, 1965. ISSN 0362-1308.

Laurence:1972:DZS

[Lau72] William Leonard Laurence. Dawn over zero; the story of the atomic bomb. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, second edition, 1972. ISBN 0-8371-6064-2. xv + 289 + viii pp. LCCN UF767 .L3 1972.

Lawren:1988:GBB

[Law88] William Lawren. The general and the bomb: a biography of General Leslie R. Groves, director of the Manhattan Project. Dodd, Mead, New York, NY, USA, 1988. ISBN 0-396-08761-2. xii + 324 pp. LCCN UG128.G76 L39 1988.

Lamont:1968:FAB

[LC68] Lansing Lamont and Michel Carrière, Trad. La fantastique aventure de la bombe. (French) [The fantastic adventure of the bomb], volume 229 of L’Air du temps. Gallimard, Paris, France, 1968. 290 pp. LCCN ????

Leffler:1995:TDD

[Lef95] Melvyn P. Leffler. Truman’s decision to drop the atomic bomb. IHJ Bulletin: A Quarterly Publication of the International House of Japan, 15(3):1–7, Summer 1995. ISSN 0285-2608.

Leskov:1994:SFU

[Les94] Sergei Leskov. The Sudoplatov file: An unreliable witness. Bulletin of the Atomic Scientists, 50(4):33–36, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levey:1960:LEO

[Lev60] J. S. Levey. Letter to the Editor: Oppenheimer and public disclosure. Bulletin of the Atomic Scientists, 16(4):145, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:2019:SNC

[Lew19] Nicholas Dale Lewis. Safety in Numbers: Creating and Contesting the Los Alamos Approach to Supercomputing, 1943 to 1980.

Ph.D. dissertation, University of Minnesota, Minneapolis, MN,

USA, 2019. vi + 392 pp. URL https://www.proquest.com/
pqdtglobal/docview/2869179133/.

Lewis:2021:TNC

[Lew21] Nicholas Lewis. Trinity by the numbers: The computing effort that made Trinity possible. Nuclear Technology, 207(S1):S176– S189, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Libby:1979:UP

[Lib79] Leona Marshall Libby. The Uranium People. Crane Russak, New York, NY, USA, 1979. ISBN 0-8448-1300-1 (Crane Russak), 0-684-16242-3 (Scribners). x + 341 + 16 pp. LCCN QC773.3.U5 L52.

Lifton:2023:OTO

[Lif23] Robert Jay Lifton. Oppenheimer’s tragedy — and ours. Bulletin of the Atomic Scientists, 79(4):216–220, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2223073.

Linnett:1981:DAT

[Lin81] R. Linnett. The Day After Trinity, J. Robert Oppenheimer and the Atomic Bomb. Cineaste, 11(2):37, ???? 1981. ISSN 0009-7004.

Lippman:1954:OC

[Lip54] Walter Lippman. The Oppenheimer case. Bulletin of the Atomic Scientists, 10(5):183, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Landau:1958:SP

[LL58] Lev D. Landau and Evgeny M. Lifshitz. Statistical Physics, volume 5 of Course of theoretical physics. Pergamon Press, Oxford, UK, 1958. x + 484 pp. LCCN QC175 .L32 1958. Translation from Russian by Eugenia and Rudolf F. Peierls.

Lauritsen:1934:STC

[LO34] C. C. Lauritsen and J. R. Oppenheimer. On the scattering of the ThC” γ rays. Physical Review (2), 46(1):80–81, July 1, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v46/
i1/p80_1.

Oppenheimer:2008:WL

[LOD08] Frank Lock, Robert Oppenheimer, and Jane Daniels. The wizard’s legacy. Physics Today, 61(7):8, July 2008. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v61/i7/p8/s1.

Lewis:1948:MPM

[LOW48] H. W. Lewis, J. R. Oppenheimer, and S. A. Wouthuysen. The multiple production of mesons. Physical Review (2), 73(2):127– 140, January 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v73/i2/p127_1.

Livingston:1954:OLP

[LPT^{+}54] M. Stanley Livingston, Ernest C. Pollard, Lewis Tonks, Watson Davis, James Franck, Oscar K. Rice, W. A. Higinbotham, David L. Hill, and John S. Toll. Open letter to President Eisenhower. Bulletin of the Atomic Scientists, 10(6):255, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leprince-Ringuet:1957:AHF

[LR57] Louis Leprince-Ringuet. Des Atomes et des Hommes. (French) [Atoms and Men]. Librairie Arthème Fayard, Paris, France, 1957. 184 + 2 pp. LCCN ????

Lestone:2021:CBH

[LRA21] J. P. Lestone, M. D. Rosen, and P. Adsley. Comparison between historic nuclear explosion yield formulas. Nuclear Technology, 207(S1):S352–S355, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Lewis:1971:APT

[LW71] Richard S. Lewis and Jane Wilson, editors. Alamogordo plus twenty-five years; the impact of atomic energy on science, technology, and world politics. Viking, New York, NY, USA, 1971. ISBN 0-670-11151-1. vi + 281 pp. LCCN QC792 .A43 1971. With Eugene Rabinowitch.

Morgan:2021:OLH

[MA21] Nathaniel R. Morgan and Billy J. Archer. On the origins of Lagrangian hydrodynamic methods. Nuclear Technology, 207 (S1):S147–S175, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

MacDougal:1954:LEH

[Mac54] P. S. MacDougal. Letter to the Editor: H-bomb tests in the Pacific and the firing of Oppenheimer. Bulletin of the Atomic Scientists, 10(8):335, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Macdonald:1963:W

[Mac63] Dwight Macdonald. To the Whitehouse. New York Review of Books, ??(??):??, February 1, 1963. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/1963/02/01/to-the-whitehouse/.

Major:1971:OHa

[Maj71a] John Major. The Oppenheimer hearing. Historic trials series. Batsford, London, UK, 1971. ISBN 0-7134-1255-0. 336 + 8 pp. LCCN QC16.O62 M3 1971b.

Major:1971:OHb

[Maj71b] John Major. The Oppenheimer hearing. Historic trials series. Stein and Day, New York, NY, USA, 1971. ISBN 0-8128-1395-2. 336 pp. LCCN QC16.O62 M3.

Major:1983:OH

[Maj83] John Major. The Oppenheimer hearing. Stein and Day, New York, NY, USA, 1983. ISBN 0-8128-6179-5 (paperback). 336 + 4 pp. LCCN QC16.O62 M3 1983. US$9.95 (est.). Reprint of [Maj71b].

Makhijani:1995:AT

[Mak95] Arjun Makhijani. “Always” the target? Bulletin of the Atomic Scientists, 51(3):23–27, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manley:1974:AOT

[Man74] J. H. Manley. “All in our time”: Assembling the wartime labs. Bulletin of the Atomic Scientists, 30(5):42–48, May 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manley:1980:NLB

[Man80] John H. Manley. A new laboratory is born. In Badash et al. [BHB80], page ?? ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

Manley:1984:CMH

[Man84] John H. Manley. Commentary: In the matter of the H-bomb. Bulletin of the Atomic Scientists, 40(1):52–53, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margulis:1997:SJR

[Mar97] Lynn Margulis. Sunday with J. Robert Oppenheimer. In Slanted truths: essays on Gaia, symbiosis, and evolution [MS97], pages 5–28. ISBN 0-387-94927-5 (hardcover). LCCN QH311 .M37 1997. Foreword by Philip and Phylis Morrison.

Masco:2006:NBM

[Mas06a] Joseph Masco. The Nuclear Borderlands: the Manhattan Project in Post-Cold War New Mexico. Short introductions. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12076-5 (hardcover), 0-691-12077-3 (paperback). xiii + 425 pp. LCCN U264.4.N32 .M52 2006.

Mason:2006:OCR

[Mas06b] Richard Mason. Oppenheimer’s choice: reflections from moral philosophy. SUNY series in philosophy. State University of New York Press, Albany, NY, USA, 2006. ISBN 0-7914-6781-3 (hardcover). 183 pp. LCCN BJ59 .M38 2006. URL http://www.loc.

gov/catdir/toc/ecip0516/2005020545.htm.

Maslin:2013:BRR

[Mas13] Jane Maslin. Book review: Rough-edged atomic pioneer Robert Oppenheimer: A Life Inside the Center, by Ray Monk. New York Times, ??(??):??, May 28, 2013. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/1355615864.

Masco:2020:NBM

[Mas20] Joseph Masco. The Nuclear Borderlands: The Manhattan Project in Post-Cold War New Mexico. Princeton University Press, Princeton, NJ, USA, new edition, March 2020. ISBN 0-691-19428-9, 0-691-20217-6. 456 pp. LCCN U264.4.N32 M52 2020.

Moore:2021:WBA

[MB21] Richard Moore and Eric N. Brown. Woolwich, Bruceton, Los Alamos: Munroe jets and the Trinity gadget. Nuclear Technology, 207(S1):S222–S230, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

McCray:2004:PVC

[McC04] W. Patrick McCray. Project Vista, Caltech, and the dilemmas of Lee DuBridge. Historical Studies in the Physical and Biological Sciences, 34(2):339–370, March 2004. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

McCray:2005:KMR

[McC05] W. Patrick McCray. Killing the messenger: Robert Oppenheimer and Caltech’s Project Vista. In Carson and Hollinger [CH05], chapter 11, pages 254–266. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

McCray:2006:APT

[McC06] W. P. McCray. American Prometheus: The triumph and tragedy of J. Robert Oppenheimer. Diplomatic History, 30(2): 321–325, April 2006. ISSN 0145-2096 (print), 1467-7709 (electronic).

McFarlane:2011:JRO

[McF11] Richard A. McFarlane. J. Robert Oppenheimer, the Cold War, and the Atomic West. Journal of the West, 50(1):98–100, Winter 2011. ISSN 0022-5169.

McGrath:2002:SBS

[McG02] Patrick J. (Patrick Joseph) McGrath. Scientists, Business, and the State, 1890–1960. The Luther Hartwell Hodges series on business, society, and the state. University of North Carolina Press, Chapel Hill, NC, USA, 2002. ISBN 0-8078-2655-3 (hardcover). 248 pp. LCCN

Q127.U5 M37 2002. URL http://www.loc.gov/catdir/
description/unc041/2001035149.htm;   http://www.loc.
gov/catdir/toc/fy033/2001035149.htm.

McInernery:2005:MMM

[McI05] Michael McInernery. Man, moment, machine: Ultimate weapon, Oppenheimer and the atomic bomb. 45 minute

video., 2005.  URL http://media.aetv.com;   http:/
/www.aspresolver.com/aspresolver.asp?AHIV;   http://
www.aspresolver.com/aspresolver.asp?VASC.

McMillan:1980:EDA

[McM80a] Edwin M. McMillan. Early days at Los Alamos. In Badash et al. [BHB80], page ?? ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

McMillan:1980:OIF

[McM80b] Elsie McMillan. Outside the inner fence. In Badash et al. [BHB80], page ?? ISBN 90-277-1097-X, 90-277-1098-8 (paperback). LCCN QC791.96 .R44.

McMillan:1994:SFF

[McM94] Priscilla Johnson McMillan. The Sudoplatov file: Flimsy memories. Bulletin of the Atomic Scientists, 50(4):30–33, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMillan:2005:RJR

[McM05] Priscilla Johnson McMillan. The ruin of J. Robert Oppenheimer, and the birth of the modern arms race. Viking, New York, NY, USA, 2005. ISBN 0-670-03422-3. viii + 373 + 16 pp. LCCN QC16.O62 M36 2005.

McWilliams:1954:OC

[McW54] Carey McWilliams. The Oppenheimer case. The Nation, ?? (??):373–374, May 1, 1954. ISSN 0027-8378. URL http://

www.thenation.com/archive;   http://www.unz.org/Pub/
Nation-1954may01.

McDayter:1967:GBB

[MD67] Walt McDayter and Norman Drew. The giants: The bomb builders. Denver Post, ??(??):??, February 3, 1967. URL http:/

/library.ucsd.edu/dc/object/bb0103915g. This is a reason-

ably accurate 83-frame comic strip on the history of the building of the atomic bomb, with Leo Szilard as the central figure of the story.

Meade:2004:BGO

[Mea04] Roger Meade. A backward glance: Oppenheimer: Unique in the scientific world. Nuclear Weapons Journal, ??(??):ii, Spring 2004. CODEN ???? ISSN ???? URL http://

www.lanl.gov/history/road/pdf/Oppenheimer,%20Unique%
20in%20the%20Scientific%20World.pdf. LALP-04-013.

Mecklin:2023:EIC

[Mec23] John Mecklin. An extended interview with Christopher Nolan, director of Oppenheimer. Bulletin of the Atomic Scientists, 79 (4):203–209, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2223070.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mendelsohn:1984:TTS

[Men84] Everett Mendelsohn, editor. Transformation and tradition in the sciences: essays in honor of I. Bernard Cohen. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26724-2. xiv + 577 pp. LCCN Q126.8 .T73 1984. URL http://

www.loc.gov/catdir/description/cam022/84007832.html;
 http://www.loc.gov/catdir/toc/cam028/84007832.html.

Menn:2003:SUS

[Men03] J. M. Menn. Still under suspicion (J. Robert Oppenheimer). American History, 37(6):4, February 2003. ISSN 1076-8866.

Merrill:1995:DHT

[Mer95] Dennis Merrill, editor. Documentary history of the Truman Presidency. University Publications of America, Bethesda, MD, USA, 1995. ISBN 1-55655-567-9 (vol. 1), 1-55655-568-7 (vol. 2), 1-55655-570-9 (vol. 4), 1-55655-577-6 (vol. 11), 1-55655-580-6 (vol. 14), 1-55655-581-4 (vol. 15), 1-55655-697-7 (vol. 28). various pp. LCCN E813 .D56 1995.

Martz:2021:TPP

[MFC21] Joseph C. Martz, Franz J. Freibert, and David L. Clark. The taming of plutonium: Plutonium metallurgy and the Manhattan Project. Nuclear Technology, 207(S1):S266–S285, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Murrow:1955:MOI

[MFO55] Edward R. Murrow, Fred W. Friendly, and J. Robert Oppenheimer. Murrow–Oppenheimer interview, 1955. URL https://

www.youtube.com/watch?v=lVCL3Rnr8xE. 2 film reels of 2 (45

min., ca. 1,620 ft.).

Michelmore:1969:SYR

[Mic69] Peter Michelmore. The swift years: the Robert Oppenheimer story. Dodd, Mead, New York, NY, USA, 1969. viii + 273 pp. LCCN QC16.O62 M5.

Millet:2005:OPR

[Mil05] Lydia Millet. Oh pure and radiant heart. Soft Skull Press, Brooklyn, NY, USA, 2005. ISBN 1-932360-85-9. 489 pp. LCCN

PS3563.I42175 O37 2005. URL http://www.loc.gov/catdir/
toc/ecip056/2005001028.htm.

Mercer:2021:GDE

[MKC^{+}21] David J. Mercer, Katrina E. Koehler, Mark P. Croce, Andrew S. Hoover, Philip A. Hypes, Stosh A. Kozimor, Veronika Mocko, Paul R. J. Saey, Daniel R. Schmidt, and Joel N. Ullom. Gamma and decay energy spectroscopy measurements of trinitite. Nuclear Technology, 207(S1):S309–S320, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Montgomery:1939:BHE

[MM39] C. G. Montgomery and D. D. Montgomery. The behavior of high energy electrons in the cosmic radiation. Reviews of Modern Physics, 11(3–4):255–264, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.11.255;
http://rmp.aps.org/abstract/RMP/v11/i3-4/p255_1. See

discussion [Opp39c].

McCormick:1962:MJR

[MM62a] John McCormick and Mairi MacInnes. In the matter of J. Robert Oppenheimer. natural science and national security. In Versions of censorship: an anthology [MM62b], page ?? LCCN Z657 .M12 1962a.

McCormick:1962:VCA

[MM62b] John McCormick and Mairi MacInnes, editors. Versions of censorship: an anthology. An Aldine library edition. Aldine Publishing Company, Chicago, IL, USA, 1962. 374 pp. LCCN Z657 .M12 1962a.

Monk:2012:ICL

[Mon12] Ray Monk. Inside the centre: the life of J. Robert Oppenheimer. Jonathan Cape, London, UK, 2012. ISBN 0-224-06262-X (hardcover). xiv + 818 + 32 pp. LCCN QC16.O62 M66 2012.

Monk:2013:ROL

[Mon13] Ray Monk. Robert Oppenheimer: a Life Inside the Center. Doubleday, New York, NY, USA, 2013. ISBN 0-385-50407-1 (hardback). xvi + 825 + 32 pp. LCCN QC16.O62 M66 2013.

Moore:2021:RPO

[Moo21] Richard Moore. Rudolf Peierls’s “Outline of the development of the British Tube Alloy Project: a 1945 account of the earliest UK work on atomic energy”. Nuclear Technology, 207(S1):S374– S379, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Morton:1957:DUA

[Mor57] Louis Morton. The decision to use the atomic bomb. Foreign Affairs, 35(2):334–353, January 1957. CODEN FRNAA3. ISSN

0015-7120. URL http://www.jstor.org/stable/20031230.

Morgan:1958:VJR

[Mor58] Thomas B. Morgan. A visit with J. Robert Oppenheimer. Look Magazine, ??(??):??, April 1, 1958.

Morgan:1966:OAD

[Mor66] Thomas B. Morgan. With Oppenheimer, on an autumn day: a thoughtful man talks searchingly about science, ethics and nuclear war on a quiet afternoon during a bad time. Look Magazine, ??(??):61–63, December 27, 1966.

Morrison:1995:NTW

[Mor95] Philip Morrison. Nothing is too wonderful to be true, volume 11 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-363-9 (hardcover). xi + 446 pp. LCCN Q173.

Morgan:2021:OBL

[Mor21] Jonathan E. Morgan. The origins of blast-loaded vessels. Nuclear Technology, 207(S1):S231–S265, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Mariner:2009:ABA

[MP09] Rosemary B. Mariner and G. Kurt Piehler, editors. The atomic bomb and American society: new perspectives. University of Tennessee Press, Knoxville, 2009. ISBN 1-57233-648-X (hardcover). LCCN QC773.3.U5 A845 2009. URL http://www.loc.gov/

catdir/toc/ecip0825/2008034982.html.

Moulton:1960:AS

[MS60] Forest Ray Moulton and Justus Julius Schifferes, editors. The Autobiography of Science. A Science Publications Council book. Doubleday, Garden City, NY, USA, second edition, 1960. 748 pp. LCCN Q111 .M6 1960.

Margulis:1997:STE

[MS97] Lynn Margulis and Dorion Sagan. Slanted truths: essays on Gaia, symbiosis, and evolution. Copernicus, New York, NY, USA, 1997. ISBN 0-387-94927-5 (hardcover). xxiii + 368 pp. LCCN QH311 .M37 1997. Foreword by Philip and Phylis Morrison.

Morton-Smith:2015:O

[MS15] Tom Morton-Smith. Oppenheimer. Oberon Books Ltd., London, UK, 2015. ISBN 1-78319-198-8. ???? pp. LCCN ????

URL http://stratford-upon-avon-theatre.blogspot.com/
2015/01/rsc-oppenheimer-review-stratford-upon.html.

Meyerhoff:1955:OCE

[MT55] H. Meyerhoff and D. Trilling. The Oppenheimer case: An exchange. Partisan Review, 22(??):238–251, ???? 1955.

Murphy:1953:HSH

[Mur53] Charles J. V. Murphy. The hidden struggle for the H-bomb: the story of Dr. Oppenheimer’s persistent campaign to reverse US military policy. Fortune, 47(??):109–110, 230, May 1953. CODEN FORTAP. ISSN 0015-8259.

Murray:1954:CO

[Mur54] Thomas E. Murray. Concurring opinion. Bulletin of the Atomic Scientists, 10(7):277–279, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murphy:2005:EPR

[Mur05] B. J. Murphy. 109 East Palace: Robert Oppenheimer and the secret city of Los Alamos. AMERICAN HISTORY, 40(4):70–71, October 2005. ISSN 1076-8866.

Murray:2008:JRO

[Mur08] William E. Murray. On J. Robert Oppenheimer. Technology Review, 111(1):8, January/February 2008. CODEN TEREAU. ISSN 1099-274X.

Masters:1946:OWN

[MW46] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. Whittlesey House, McGraw-Hill Book Co., Inc., New York, NY, USA, 1946. x + 79 pp. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Masters:2007:OWN

[MW07] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. New Press, New York, NY, USA, 2007. ISBN 1-59558-227-4 (hardcover). xx + 220 pp. LCCN UG1282.A8

O54 2007.  URL http://thenewpress.com/index.php?
option=com_title&task=view_title&metaproductid=1703;
 http://www.loc.gov/catdir/toc/ecip0718/2007020838.htm.

Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46].

Marshall:2009:BRS

[MW09] Trevor Marshall and Max Wallis. Book review: Silvan S. Schweber, Einstein and Oppenheimer: The Meaning of Genius. London: Harvard University Press, 2008. Pp. xiv + 412. ISBN 978-0-674-02828-9. £19.25 (hardback). British Journal for the History of Science, 42(3):471–473, September 2009. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nash:2007:EPR

[Nas07] Philip Nash. 109 East Palace: Robert Oppenheimer and the secret city of Los Alamos. Historian, 69(1):108–109, Spring 2007. ISSN 0018-2370 (print), 1540-6563 (electronic).

Nelson:2014:ARE

[Nel14] Craig Nelson. The age of radiance: the epic rise and dramatic fall of the atomic era. Scribner, New York, NY, USA, 2014. ISBN 1-4516-6043-X (hardcover), 1-4516-6044-8 (paperback), 1-4516-6045-6 (e-book). 4a37 pp. LCCN QC773 .N45 2014.

Neufeld:2013:CTF

[Neu13] Michael J. Neufeld. The complicated, tragic figure behind the bomb. Washington Post, ??(??):B.7–??, July 21,

2013. ISSN 0190-8286. URL http://search.proquest.com/
docview/1404798718.

Newman:1977:OCR

[New77] Steven Leonard Newman. The Oppenheimer Case: a Reconsideration of the Role of the Defense Department and National Security. Ph.D. dissertation, New York University, New York, NY, USA, 1977.

Nichols:1954:NPC

[Nic54] Major General K. D. Nichols. Nichols presents charges [against J. Robert Oppenheimer]. Bulletin of the Atomic Scientists, 10 (5):174–176, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Opp54f].

Nichols:1987:RTP

[Nic87] Major General Kenneth D. (Kenneth David) Nichols. The road to Trinity: a personal account of how America’s nuclear policies were made. Morrow, New York, NY, USA, 1987. ISBN 0-688-06910-X. 401 pp. LCCN QC774.N45 A3 1987.

Nichols:2023:NPC

[Nic23] K. D. Nichols. Nichols presents charges. Bulletin of the Atomic Scientists, 79(4):238–241, 2023. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/00963402.2023.
2223083.

Nimtz:2011:TCS

[Nim11] Günter Nimtz. Tunneling confronts Special Relativity. Foundations of Physics, 41(7):1193–1199, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

springerlink3.metapress.com/content/n67286542k5r4186.

Nisbet:1992:TSM

[Nis92] Robert A. Nisbet. Teachers and Scholars: a Memoir of Berkeley in Depression and War. Transaction Publishers, New Brunswick, NJ, USA, 1992. ISBN 1-56000-034-1. viii + 216 pp. LCCN LD759 .N57 1992.

Nordheim:1937:DHE

[NNOS37] G. Nordheim, L. W. Nordheim, J. R. Oppenheimer, and R. Serber. The disintegration of high energy protons. Physical Review (2), 51(12):1037–1045, June 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v51/i12/p1037_1.

Nedelsky:1933:PPN

[NO33] Leo Nedelsky and J. R. Oppenheimer. The production of positives by nuclear gamma-rays. Physical Review (2), 44(11):948– 949, December 1, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v44/i11/p948_2. See errata [NO34b].

Nedelsky:1934:PPN

[NO34a] Leo Nedelsky and J. R. Oppenheimer. 30. on the production of positives by nuclear gamma-rays. Physical Review (2), 45 (2):136–137, January 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v45/i2/p130_1. In Minutes of the Stanford

Meeting, December 15–16, 1933. See errata [NO34b].

Nedelsky:1934:EPP

[NO34b] Leo Nedelsky and J. R. Oppenheimer. Errata: “The Production of Positives by Nuclear Gamma-Rays”. Physical Review (2), 45 (4):283, February 15, 1934. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v45/i4/p283_1. See [NO33, NO34a].

Nelson:1941:MPM

[NO41] E. C. Nelson and J. R. Oppenheimer. 9. multiple production of mesotrons by protons. Physical Review (2), 60(2):159–160, July 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v60/

i2/p158_1. In Minutes of the Pasadena, California, Meeting June 18–20, 1941.

Nelson:1942:PTM

[NO42] E. C. Nelson and J. R. Oppenheimer. 4. pair theories of meson scattering. Physical Review (2), 61(3–4):202, February 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v61/i3-

4/p202_1. In December 1941 Meeting Held at Stanford University, California.

Norris:2002:RBG

[Nor02] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Steerforth Press, South Royalton, VT, USA, 2002. ISBN 1-58642-039-9. xxi + 722 pp. LCCN UG128.G76 N67 2002. URL

http://www.loc.gov/catdir/toc/fy036/2001057629.html.

Norris:2003:RBG

[Nor03] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Turnaround, London, UK, 2003. ISBN 1-58642-067-4. 228 pp. LCCN UG128.G76 N67 2002. URL http://www.loc.

gov/catdir/toc/fy036/2001057629.html.

Nuechterlein:2005:MAJ

[Nue05] J. Nuechterlein. In the matter (again) of J. Robert Oppenheimer. Commentary, 120(3):54–59, October 2005. ISSN 0010-2601.

Oppenheimer:1946:RIC

[O^{+}46] J. Robert Oppenheimer et al. A report on the international control of atomic energy. Report, U.S. Department of State, Washington, DC, USA, 1946.

Oppenheimer:1965:CVD

[O^{+}65] J. R. Oppenheimer et al. Comment vivre demain?. (French) [How do we live tomorrow?]. In Textes des conférences et des entretiens organisés par les Rencontres internationales de Genève 1964, page ?? La Baconnière, Neuchâtel, Switzerland, 1965.

Ottaviani:2001:FJR

[O^{+}01] Jim Ottaviani et al. Fallout: J. Robert Oppenheimer, Leo Szilard, and the political science of the atomic bomb. G. T. Labs, Ann Arbor, MI, USA, 2001. ISBN 0-9660106-3-9. 239 pp. LCCN QC773.3.U5 O78 2001.

Oppenheimer:19xx:JRO

[OBxx] J. Robert Oppenheimer and Niels Bohr. J. Robert Oppenheimer papers (1799–1980) (bulk 1947–1967) and Niels Henrik David Bohr papers. 19xx. ???? pp. LCCN 0456G; Vault 0202A; Micro-

film 16,646-1P. URL http://hdl.loc.gov/loc.mss/eadmss.
ms998007.

Oppenheimer:1949:AAC

[OBC^{+}49] J. Robert Oppenheimer, Oliver E. Buckley, James B. Conant, Lee A. DuBridge, Enrico Fermi, I. I. Rabi, Hartley Rowe, Glenn T. Seaborg, and Cyril S. Smith. AEC Advisory Committee statement on fellowships. Bulletin of the Atomic Scientists, 5(8–9):210, 254, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliphant:1989:FHT

[Oli89] Mark Oliphant. Footnote to history: Three men and the bomb. Bulletin of the Atomic Scientists, 45(2):41–42, March 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olson:1962:BBS

[Ols62a] James C. Olson. Books: Japan Subdued: The Atomic Bomb and the End of the War in the Pacific, by Herbert Freis. Bulletin of the Atomic Scientists, 18(2):34–35, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Opp62f] and response [Ols62b].

Olson:1962:LEJ

[Ols62b] James C. Olson. Letter to the Editor: Japan subdued. Bulletin of the Atomic Scientists, 18(6):43, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olwell:2001:SBB

[Olw01] Russell B. Olwell. In the Shadow of the Bomb: Bethe, Oppenheimer, and the Moral Responsibility of the Scientist (review). Technology and Culture, 42(4):823–825, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/33834.

OrtegayGasset:1985:RM

[OM85] José Ortega y Gasset and Kenneth Moore, editors. The revolt of the masses. University of Notre Dame Press, Notre Dame, IN, USA, 1985. ISBN 0-268-01609-7. xxxi + 192 pp. LCCN CB103 .O713 1985. US$20.00.

Oppenheimer:1984:US

[OMRS84] J. Robert Oppenheimer, N. (Nicholas) Metropolis, Gian-Carlo Rota, and D. H. (David Howland) Sharp. Uncommon Sense. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1984. ISBN 0-8176-3165-8. ix + 195 pp. LCCN Q127.U6 O66 1984.

Oppenheimer:1933:PPE

[OP33] J. R. Oppenheimer and M. S. Plesset. On the production of the positive electron. Physical Review (2), 44(1):53–55, July 1, 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v44/
i1/p53_2.

Oppenheimer:1935:NTF

[OP35] J. R. Oppenheimer and M. Phillips. Note on the transmutation function for deuterons. Physical Review (2), 48(6):500– 502, September 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v48/i6/p500_1.

Oppenheimer:1926:QTP

[Opp26a] J. R. Oppenheimer. On the quantum theory of the problem of the two bodies. Mathematical Proceedings of the Cambridge Philosophical Society, 23:422–431, 1926. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Oppenheimer:1926:QTV

[Opp26b] J. R. Oppenheimer. On the quantum theory of vibration– rotation bands. Mathematical Proceedings of the Cambridge Philosophical Society, 23:327–335, 1926. CODEN MPCPCO. ISSN 0305-0041 (print), 1469-8064 (electronic).

Oppenheimer:1926:QKA

[Opp26c] J. R. Oppenheimer. Quantentheorie des kontinuierlichen Absorptionsspektrums. (German) [Quantum theory of continuous absorption spectra]. Naturwissenschaften, 14(52):1282, December 24, 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Oppenheimer:1926:QTI

[Opp26d] J. R. Oppenheimer. Quantum theory and intensity distribution in continuous spectra. Nature, 118(2978):771, November 27, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v118/n2978/pdf/118771d0.pdf.

Oppenheimer:1927:BZT

[Opp27a] J. R. Oppenheimer. Bemerkung zur Zerstreuung der α-Teilchen. (German) [Note on the dispersion of α-particles]. Zeitschrift für Physik, 43(5–6):413–415, ???? 1927. CODEN ZEPYAA. ISSN 0044-3328.

Oppenheimer:1927:QTP

[Opp27b] J. R. Oppenheimer. On the quantum theory of the polarization of impact radiation. Proceedings of the National Academy of Sciences of the United States of America, 13(12):800–805, December 15, 1927. CODEN PNASA6. ISSN 0027-8424 (print),

1091-6490 (electronic). URL http://www.jstor.org/stable/
pdfplus/85652.pdf.

Oppenheimer:1927:QRG

[Opp27c] J. R. Oppenheimer. Zur Quantenmechanik der Richtungsentartung (German) [On the quantum mechanics of the direction of degeneracy]. Zeitschrift für Physik, 43(1–2):27–46, January 1927. CODEN ZEPYAA. ISSN 0044-3328.

Oppenheimer:1927:QKSb

[Opp27d] J. R. Oppenheimer. Zur Quantentheorie kontinuierlicher Spektren (German) [On the quantum theory of the continuous spectrum]. Zeitschrift für Physik, 41(8–9):268–293, August 1927. CODEN ZEPYAA. ISSN 0044-3328.

Oppenheimer:1927:QKSa

[Opp27e] Julius Robert Oppenheimer. Zur Quantentheorie der kontinuierlichen Spektren (German) [On the quantum theory of the continuous spectrum]. Dissertation, Georg-August-Universität Göttingen, Göttingen, Germany, 1927. URL http:

//genealogy.math.ndsu.nodak.edu/id.php?id=14001. Pub-

lished in [Opp27d].

Oppenheimer:1928:QTE

[Opp28a] J. R. Oppenheimer. On the quantum theory of electronic impacts. Physical Review (2), 32(3):361–376, September 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v32/i3/
p361_1.

Oppenheimer:1928:QTF

[Opp28b] J. R. Oppenheimer. On the quantum theory of field currents. Physical Review (2), 31(5):914, May 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v31/i5/p912_1.

Oppenheimer:1928:QTA

[Opp28c] J. R. Oppenheimer. On the quantum theory of the autoelectric field currents. Proceedings of the National Academy of Sciences of the United States of America, 14(5):363–365, May 15, 1928. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/pdfplus/
85208.pdf.

Oppenheimer:1928:QTC

[Opp28d] J. R. Oppenheimer. On the quantum theory of the capture of electrons. Physical Review (2), 31(3):349–356, March 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v31/i3/
p349_1.

Oppenheimer:1928:QTR

[Opp28e] J. R. Oppenheimer. On the quantum theory of the Ramsauer effect. Proceedings of the National Academy of Sciences of the United States of America, 14(3):261–262, March 15, 1928. CO-DEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

URL http://www.jstor.org/stable/pdfplus/85421.pdf.

Oppenheimer:1928:TNQ

[Opp28f] J. R. Oppenheimer. Three notes on the quantum theory of aperiodic effects. Physical Review (2), 31(1):66–81, January 1, 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v31/
i1/p66_1.

Oppenheimer:1928:BRL

[Opp28g] J. Robert Oppenheimer. Book review: The Logic of Modern Physics. P. W. Bridgman. Pp. 228 + xiv. The Macmillan Company, New York, 1927. Price $2.50. Physical Review (2), 31(1): 145–146, January 1928. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://journals.aps.

org/pr/abstract/10.1103/PhysRev.31.145.

Oppenheimer:1929:SFE

[Opp29] J. R. Oppenheimer. Über die Strahlung der freien Elektronen im Coulombfeld. (German) [On the radiation of a free electron in a Coulomb field]. Zeitschrift für Physik, 55(11–12):725–737, ???? 1929. CODEN ZEPYAA. ISSN 0044-3328.

Oppenheimer:1930:NTI

[Opp30a] J. R. Oppenheimer. Note on the theory of the interaction of field and matter. Physical Review (2), 35(5):461–477, March 1, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v35/
i5/p461_1.

Oppenheimer:1930:TEP

[Opp30b] J. R. Oppenheimer. On the theory of electrons and protons. Physical Review (2), 35(5):562–563, March 1, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v35/i5/p562_1.

Oppenheimer:1930:TNP

[Opp30c] J. R. Oppenheimer. Two notes on the probability of radiative transitions. Physical Review (2), 35(8):939–947, April 15, 1930. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v35/
i8/p939_1.

Oppenheimer:1931:SRA

[Opp31a] J. R. Oppenheimer. 23. selection rules and the angular momentum of light quanta. Physical Review (2), 37(2):231, January 15, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v37/

i2/p225_1. In Minutes of the Los Angeles Meeting, December 12–13, 1930.

Oppenheimer:1931:NSN

[Opp31b] J. R. Oppenheimer. 28. note on the statistics of nuclei. Physical Review (2), 37(2):232–233, January 15, 1931. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v37/i2/

p225_1. In Minutes of the Los Angeles Meeting, December 12– 13, 1930.

Oppenheimer:1931:NLQ

[Opp31c] J. R. Oppenheimer. Note on light quanta and the electromagnetic field. Physical Review (2), 38(4):725–746, August 15, 1931. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v38/i4/
p725_1.

Oppenheimer:1933:DLP

[Opp33] J. Robert Oppenheimer. The disintegration of lithium by protons of high energy. Physical Review (2), 43(5):380, March 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v43/i5/
p370_1.

Oppenheimer:1934:TEP

[Opp34] J. R. Oppenheimer. 28. the theory of the electron and positive. Physical Review (2), 45(4):290, February 1934. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v45/i4/

p284_1. In Minutes of the Boston Meeting, December 28–30, 1933.

Oppenheimer:1935:FAH

[Opp35a] J. R. Oppenheimer. Are the formulae for the absorption of high energy radiations valid? Physical Review (2), 47(1):44–52, January 1, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v47/i1/p44_1.

Oppenheimer:1935:DDI

[Opp35b] J. R. Oppenheimer. The disintegration of the deuteron by impact. Physical Review (2), 47(11):845–846, June 1, 1935. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v47/i11/p845_1.

Oppenheimer:1935:NCF

[Opp35c] J. R. Oppenheimer. Note on charge and field fluctuations. Physical Review (2), 47(2):144–145, January 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v47/i2/p144_1.

Oppenheimer:1935:NPP

[Opp35d] J. R. Oppenheimer. Note on the production of pairs by charged particles. Physical Review (2), 47(2):146–147, January 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v47/i2/
p146_1.

Oppenheimer:1939:EIS

[Opp39a] J. R. Oppenheimer. 25. on the elementary interpretation of showers and bursts. Physical Review (2), 50(4):389, August 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v50/

i4/p384_1. In Minutes of the Seattle Meeting, June 17–19, 1936.

Oppenheimer:1939:CSB

[Opp39b] J. R. Oppenheimer. Celebration of the sixtieth birthday of Albert Einstein. Science, 89(2311):335–336, April 14, 1939. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.sciencemag.org/content/89/2311/335.2.
extract.

Oppenheimer:1939:DBH

[Opp39c] J. R. Oppenheimer. Discussion [the behavior of high energy electrons in the cosmic radiation]. Reviews of Modern Physics, 11 (3–4):264–266, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.11.264;   http://rmp.
aps.org/abstract/RMP/v11/i3-4/p264_1. See [MM39].

Oppenheimer:1940:AQT

[Opp40] J. R. Oppenheimer. 42. on the applicability of quantum theory to mesotron collisions. Physical Review (2), 57(4):353, February 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v57/

i4/p344_1. In Minutes of the Columbus, Ohio, Meeting, December 28–30, 1939.

Oppenheimer:1941:ICP

[Opp41a] J. R. Oppenheimer. 2. internal conversion in photosynthesis. Physical Review (2), 60(2):158, July 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/prola.aps.org/abstract/PR/v60/i2/p158_1. In Minutes of

the Pasadena, California, Meeting June 18–20, 1941.

Oppenheimer:1941:IPO

[Opp41b] J. R. Oppenheimer. 33. on the internal pairs from oxygen. Physical Review (2), 60(2):164, July 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:/

/prola.aps.org/abstract/PR/v60/i2/p158_1. In Minutes of

the Pasadena, California, Meeting June 18–20, 1941.

Oppenheimer:1941:MQT

[Opp41c] J. R. Oppenheimer. The mesotron and the quantum theory of fields. In Fermi et al. [FBR^{+}41], pages 39–50.

Oppenheimer:1941:SRB

[Opp41d] J. R. Oppenheimer. On the selection rules in beta-decay. Physical Review (2), 59(11):908, June 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v59/i11/p908_1.

Oppenheimer:1941:SM

[Opp41e] J. R. Oppenheimer. On the spin of the mesotron. Physical Review (2), 59(5):462, March 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v59/i5/p462_1.

Oppenheimer:1944:CRR

[Opp44] J. R. Oppenheimer. Cosmic rays: Report of recent progress. Lecture at the University of California, 1944.

Oppenheimer:1945:AA

[Opp45a] J. R. Oppenheimer. The atomic age. Broadcast on the N.Y. Philharmonic Symphony Hour., 1945.

Oppenheimer:1945:AW

[Opp45b] J. R. Oppenheimer. Atomic weapons. Talk at the American Philosophical Society and National Academy of Sciences., 1945.

Oppenheimer:1945:AWC

[Opp45c] J. R. Oppenheimer. Atomic weapons and the crisis in science. Saturday Review of Literature, ??(??):9–11, November 24, 1945.

Oppenheimer:1945:BW

[Opp45d] J. R. Oppenheimer. The bomb and the world. Talk at the National Policy Commission Conference., 1945.

Oppenheimer:1945:SAA

[Opp45e] J. Robert Oppenheimer. Speech to the Association of Los Alamos Scientists. November 2, 1945. URL http://

universityhonors.umd.edu/HONR269J/archive/OppenheimerSpeech.
htm.

Oppenheimer:1946:ABG

[Opp46a] J. R. Oppenheimer. The atom bomb as a great force for peace. The New York Times Magazine, ??(??):59–60, June 9, 1946. ISSN 0362-1308.

Oppenheimer:1946:AE

[Opp46b] J. R. Oppenheimer. Atomic explosives. Talk at the Westinghouse Century Forum, published in the New York Times., 1946.

Oppenheimer:1946:AW

[Opp46c] J. R. Oppenheimer. Atomic weapons. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 90(1):7–10, January 1946. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:/

/www.jstor.org/stable/3301032. Symposium on Atomic En-

ergy and Its Implications (January 1946).

Oppenheimer:1946:ICAb

[Opp46d] J. R. Oppenheimer. International control of atomic energy. Foreign Affairs, ??(??):??, ???? 1946. CODEN FRNAA3. ISSN 0015-7120.

Oppenheimer:1946:ICAc

[Opp46e] J. R. Oppenheimer. International control of atomic energy. In ????, editor, Seven Minutes to Midnight. Basic Books, New York, NY, USA, 1946.

Oppenheimer:1946:NW

[Opp46f] J. R. Oppenheimer. The new weapon. One World or None. (F.A.S.), 1946.

Oppenheimer:1946:NWT

[Opp46g] J. R. Oppenheimer. The new weapon: the turn of the screw. In Masters and Way [MW46], page ?? LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Oppenheimer:1946:SIU

[Opp46h] J. R. Oppenheimer. Scientific information to USAEC. UNAEC, bibliography., 1946.

Oppenheimer:1946:SCS

[Opp46i] J. R. Oppenheimer. The scientist in contemporary society. Princeton University Bicentennial broadcast., 1946.

Oppenheimer:1946:TS

[Opp46j] J. R. Oppenheimer. The turn of the screw. F.A.S. Book, One World or None., 1946.

Oppenheimer:1946:ABC

[Opp46k] J. Robert Oppenheimer. The atom bomb and college education. The General Magazine and Historical Chronicle [University of Pennsylvania General Alumni Society], 48(4):265–??, Summer 1946. Commencement speech at the University of Pennsylvania.

Oppenheimer:1946:ICAa

[Opp46l] J. Robert Oppenheimer. The international control of atomic energy. Bulletin of the Atomic Scientists, 1(12):1–5, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Opp63o].

Oppenheimer:1947:AEC

[Opp47a] J. R. Oppenheimer. Atomic energy as a contemporary problem. Talk at the National War College., 1947.

Oppenheimer:1947:CDM

[Opp47b] J. R. Oppenheimer. Creation and destruction of mesons. Physical Review (2), 71(7):460–489, April 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). Richtmeyer Lecture.

Oppenheimer:1947:FIAa

[Opp47c] J. R. Oppenheimer. Functions of international agency in research and development. Condensed version in [Opp47f]., 1947.

Oppenheimer:1947:PCW

[Opp47d] J. R. Oppenheimer. Physics in the contemporary world. Talk at MIT., 1947.

Oppenheimer:1947:RLA

[Opp47e] J. R. Oppenheimer. Richtmeyer Lecture, APS and AA Physics Teachers’ Meeting. Science Service Wire Report., 1947.

Oppenheimer:1947:FIAb

[Opp47f] J. Robert Oppenheimer. Functions of the International Agency in research and development. Bulletin of the Atomic Scientists, 3 (7):173–176, 197, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1947:SFW

[Opp47g] J. Robert Oppenheimer. Scientific foundations for world order. Denver University, published in pamphlet form and in book, Foundations for world order, University of Denver., 1947.

Oppenheimer:1948:CRC

[Opp48a] J. R. Oppenheimer. Concluding remarks to cosmic ray symposium. Talk at CalTech., 1948.

Oppenheimer:1948:GUA

[Opp48b] J. R. Oppenheimer. The growth of understanding of the atomic world. Talk at Princeton University, 1948.

Oppenheimer:1948:NSP

[Opp48c] J. R. Oppenheimer. Notes on science and practice. Talk at Harvard University, Lawrence Science School., 1948.

Oppenheimer:1948:PRN

[Opp48d] J. R. Oppenheimer. Physical research in the near future. Talk at Cooper Union, NY., 1948.

Oppenheimer:1948:SAP

[Opp48e] J. R. Oppenheimer. Some aspects of the problems of atomic energy. Talk at the New York Bar Association., 1948.

Oppenheimer:1948:ICAa

[Opp48f] Robert Oppenheimer. International control of atomic energy. Foreign Affairs, 26(2):239–252, January 1948. CODEN FR-

NAA3. ISSN 0015-7120. URL http://www.foreignaffairs.
com/articles/70629/j-robert-oppenheimer/international-
control-of-atomic-energy;   http://www.jstor.org/pss/
20030103; http://www.jstor.org/stable/20030103. Reprinted

in [CH05, pp. 149–183].

Oppenheimer:1948:ICAb

[Opp48g] Robert Oppenheimer. International control of atomic energy. Bulletin of the Atomic Scientists, 4(2):39–43, 48, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Opp63m].

Oppenheimer:1948:PCW

[Opp48h] Robert Oppenheimer. Physics in the contemporary world. Bulletin of the Atomic Scientists, 4(3):65–68, 85–86, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1949:CRC

[Opp49a] J. R. Oppenheimer. Concluding remarks to cosmic-ray symposium. Reviews of Modern Physics, 21(1):181–183, January 1, 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Oppenheimer:1949:DAS

[Opp49b] J. R. Oppenheimer. Discovery and application of sources of nuclear energy. Talk at Johns Hopkins University., 1949.

Oppenheimer:1949:LSM

[Opp49c] J. R. Oppenheimer. A letter to Senator McMahon. Bulletin of the Atomic Scientists, 5(6–7):163, 178, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1949:STP

[Opp49d] J. R. Oppenheimer. Some thoughts on the place of science in today’s world. Smith College Lecture., 1949.

Oppenheimer:1949:SMT

[Opp49e] J. R. Oppenheimer. Statements for march of time. Movies., 1949.

Oppenheimer:1949:OM

[Opp49f] J. Robert Oppenheimer. The open mind. Bulletin of the Atomic Scientists, 5(1):3–5, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See minor erratum [Ano49a].

Oppenheimer:1950:AA

[Opp50a] J. R. Oppenheimer. The atomic age. Talk at the National War College., 1950.

Oppenheimer:1950:ESa

[Opp50b] J. R. Oppenheimer. The encouragement of science. Westinghouse Science Talent Search., 1950.

Oppenheimer:1950:ESb

[Opp50c] J. R. Oppenheimer. Encouragement of science. Science, 111 (2885):373–375, April 14, 1950. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Oppenheimer:1950:RFD

[Opp50d] J. R. Oppenheimer. Response. In Fateful decision, NBC Program. Published in the Bulletin of the Atomic Scientists., 1950.

Oppenheimer:1950:AS

[Opp50e] J. Robert Oppenheimer. The age of science: 1900–1950. Scientific American, 183(3):20–23, September 1950. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Oppenheimer:1950:ET

[Opp50f] J. Robert Oppenheimer. Electron theory. In ????, editor, Conseil International de Physique, Solvay, Rapports, pages 1–11. Institut Solvay, Paris, France, 1950. Reprinted in [Sch58, pp. 145–155].

Oppenheimer:1950:FD

[Opp50g] J. Robert Oppenheimer. Fateful decision. Bulletin of the Atomic Scientists, 6(3):74, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1951:CPA

[Opp51a] J. R. Oppenheimer. Contemporary problems of atomic energy. Talk at the New York Bar Association., 1951.

Oppenheimer:1951:CMV

[Opp51b] J. Robert Oppenheimer. Comments on the military value of the atom. Bulletin of the Atomic Scientists, 7(2):43–45, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1951:DOR

[Opp51c] J. Robert Oppenheimer. Dr. Oppenheimer on the rate of American bomb production. Bulletin of the Atomic Scientists, 7(2): 45, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1951:SPA

[Opp51d] J. Robert Oppenheimer. The scientist in public affairs: Encouragement of science. Bulletin of the Atomic Scientists, 7(1):6–8, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Address delivered at the Awards Banquet of the Science Talent Institute in Washington, DC, March 6, 1950.

Oppenheimer:1953:CCR

[Opp53a] J. R. Oppenheimer. Contributions of computers in research. IBM Seminar., 1953.

Oppenheimer:1953:SS

[Opp53b] J. R. Oppenheimer. The scientist in society. Princeton University Graduate Council Talk., 1953.

Oppenheimer:1953:AWAb

[Opp53c] J. Robert Oppenheimer. Atomic weapons and American policy. Bulletin of the Atomic Scientists, 9(6):202–205, July 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Opp63k].

Oppenheimer:1953:AWAa

[Opp53d] Robert Oppenheimer. Atomic weapons and American policy. Foreign Affairs, 31(4):525–535, July 1953. CODEN FRNAA3.

ISSN 0015-7120. URL http://www.foreignaffairs.com/
articles/71043/j-robert-oppenheimer/atomic-weapons-and-
american-policy;   http://www.jstor.org/stable/20030987.

Oppenheimer:1954:CS

[Opp54a] J. R. Oppenheimer. A career in science. Princeton University, Career Forum., 1954.

Oppenheimer:1954:NNS

[Opp54b] J. R. Oppenheimer. Nucleon-nucleon scattering and polarization. In Rochester 1954, High energy nuclear physics, pages 1–26. ????, ????, 1954. Prepared for 4th Annual Rochester Conference on High-Energy and Nuclear Physics, Rochester, New York, 25–27 Jan 1954.

Oppenheimer:1954:RPC

[Opp54c] J. R. Oppenheimer. Remarks at Pyramid Club Award. 1954.

Oppenheimer:1954:WWL

[Opp54d] J. R. Oppenheimer. The world we live in. Life Magazine Radio Broadcast., 1954.

Oppenheimer:1954:RAV

[Opp54e] J. Robert Oppenheimer. On receiving the AEC verdict. Bulletin of the Atomic Scientists, 10(7):258, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1954:OR

[Opp54f] J. Robert Oppenheimer. Oppenheimer replies. Bulletin of the Atomic Scientists, 10(5):177–187, 191, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See charges [Nic54].

Oppenheimer:1954:SCUa

[Opp54g] J. Robert Oppenheimer. Science and the Common Understanding, volume 1953 of BBC Reith lectures. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1954. 127 pp. LCCN Q171 .O6 1954b.

Oppenheimer:1954:SCUb

[Opp54h] J. Robert Oppenheimer. Science and the Common Understanding. The BBC Reith lectures, 1953. Simon and Schuster, New York, NY, USA, 1954. 120 pp. LCCN Q171 .O6.

Oppenheimer:1954:TLC

[Opp54i] J. Robert Oppenheimer. Texts of letter from A.E.C. General Manager to Dr. Oppenheimer and scientist’s reply: Exchange reviews nation’s development of nuclear energy and how weapons emerged scientist cites ‘2 decades of my life’. New York Times, ??(??):16–??, April 13, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/docview/113116087.

Oppenheimer:1955:AS

[Opp55a] J. R. Oppenheimer. Analogy in science. American Psychology Association Meeting., 1955.

Oppenheimer:1955:CRO

[Opp55b] J. R. Oppenheimer. Comments by Robert Oppenheimer. Hiroshima Diary, 1955.

Oppenheimer:1955:SPA

[Opp55c] J. R. Oppenheimer. Science and public affairs. Princeton University, Woodrow Wilson School., 1955.

Oppenheimer:1955:SGO

[Opp55d] J. R. Oppenheimer. Science and the good old days. Princeton Old Guard Talk., 1955.

Oppenheimer:1955:ETD

[Opp55e] J. Robert Oppenheimer. Electron theory: Description and analogy. Report, Department of Physics, Iowa State University, Ames, IA, USA, 1955. Memorial talk about Oppenheimer’s late student, John Franklin Carlson.

Oppenheimer:1955:CCC

[Opp55f] J. Robert Oppenheimer. La ciencia y el conocimeinto común. (Spanish) [Science and the common understanding]. Consejo Superior de Investigaciones Cientificas. Departamento de Filosofía e Historia de la Ciencia, Madrid, Spain, 1955. 67 pp. LCCN Q171 .O618. Spanish translation of [Opp54h].

Oppenheimer:1955:OM

[Opp55g] J. Robert Oppenheimer. The Open Mind. Simon and Schuster, New York, NY, USA, 1955. 145 pp. LCCN QC780 .06.

Oppenheimer:1955:PASa

[Opp55h] J. Robert Oppenheimer. Prospects in the arts and sciences. Bulletin of the Atomic Scientists, 11(2):42–44, 52, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Talk broadcast for the Columbia University Bicentennial, 26 December 1954.

Oppenheimer:1955:WGD

[Opp55i] J. Robert Oppenheimer. Wissenschaft und allgemeines Denken. (German) [Science and the Common Understanding]. Rowohlt, Reinbek bei Hamburg, Germany, 1955. 135 pp. LCCN ???? German translation by Helmtraut Menion of [Opp54g].

Oppenheimer:1956:AEP

[Opp56a] J. R. Oppenheimer. Atomic energy for peaceful uses. Daily Princetonian., 1956.

Oppenheimer:1956:CQL

[Opp56b] J. R. Oppenheimer. Comment for quotation in leaflet. World Universities Service, 1956.

Oppenheimer:1956:CBH

[Opp56c] J. R. Oppenheimer. Cosmic breakthrough and a human problem. Princeton University, Graduate College Forum., 1956.

Oppenheimer:1956:DMA

[Opp56d] J. R. Oppenheimer. Dignity of Man award. Kessler Institute.,

Oppenheimer:1956:GSS

[Opp56e] J. R. Oppenheimer. The growth of science and the structure of culture. American Academy of Arts and Sciences Conference, 1956.

Oppenheimer:1956:CLP

[Opp56f] J. R. Oppenheimer. On the collateral lives of physicists. Science, 123(3193):397, March 9, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Oppenheimer:1956:SMS

[Opp56g] J. R. Oppenheimer. Science and modern society. New Republic, ??(??), ???? 1956. ISSN 0028-6583. Excerpt of [Opp56o].

Oppenheimer:1956:SOTa

[Opp56h] J. R. Oppenheimer. Science and our times. Talk at Roosevelt University. Excerpted in [Opp56g]., 1956.

Oppenheimer:1956:ST

[Opp56i] J. R. Oppenheimer. A study of thinking. Sewanee Review of Bruner Book., 1956.

Oppenheimer:1956:WST

[Opp56j] J. R. Oppenheimer. Where is science taking us? Saturday Review., 1956.

Oppenheimer:1956:CM

[Opp56k] J. Robert Oppenheimer. The Constitution of Matter. Oregon State System of Higher Education, Eugene, OR, USA, 1956. vii + 37 pp. LCCN ???? Condon Lectures 1956 (1955??). Also presented at Goucher College, 1956; Northwestern University, 1956; Naval Research Laboratory, 1956; and Wayne University, 1959.

Oppenheimer:1956:E

[Opp56l] J. Robert Oppenheimer. Einstein. Reviews of Modern Physics, 28(1):1–2, January 1, 1956. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. In memoriam tribute.

Oppenheimer:1956:JDW

[Opp56m] J. Robert Oppenheimer, editor. Jin dai wu li xue lun ji. (Mandarin) [Modern physics engine]. ????, ????, 1956. 1 + 1 + 224 pp. LCCN QC173 .C42.

Oppenheimer:1956:PT

[Opp56n] J. Robert Oppenheimer. Physics tonight. Physics Today, 9(7): 10–13, July 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Oppenheimer:1956:SOTb

[Opp56o] J. Robert Oppenheimer. Science and our times. Bulletin of the Atomic Scientists, 12(7):235–237, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Talk delivered to the Roosevelt University Founders and Friends Dinner in Chicago, 22 May 1956.

Oppenheimer:1957:ES

[Opp57a] J. R. Oppenheimer. Engineers and scientists. Drexel Institute of Technology., 1957.

Oppenheimer:1957:EAP

[Opp57b] J. R. Oppenheimer. The environs of atomic power. American Assembly, Arden House., 1957.

Oppenheimer:1957:HO

[Opp57c] J. R. Oppenheimer. The hope of order. Harvard University, James Lecture., 1957.

Oppenheimer:1957:HOW

[Opp57d] J. R. Oppenheimer. The hope of order [William James Lecture 8]. Box 259, JRO Papers, US Library of Congress, Washington, DC, USA., 1957.

Oppenheimer:1957:NPI

[Opp57e] J. R. Oppenheimer. Nuclear power and international relations. Princeton University, NATO Conference., 1957.

Oppenheimer:1957:PMW

[Opp57f] J. R. Oppenheimer. A pluralistic multiverse [William James Lecture 1]. Box 259, JRO Papers, US Library of Congress, Washington, DC, USA., 1957.

Oppenheimer:1957:SVH

[Opp57g] J. R. Oppenheimer. Science, values and the human community. Fulbright Conference on Higher Education, Sarah Lawrence College., 1957.

Oppenheimer:1957:SOW

[Opp57h] J. R. Oppenheimer. Sites of order [William James Lecture 3]. Box 259, JRO Papers, US Library of Congress, Washington, DC, USA., 1957.

Oppenheimer:1957:TVP

[Opp57i] J. R. Oppenheimer. Theory versus practice in American values and performance. MIT, American Project Conference., 1957.

Oppenheimer:1957:TRC

[Opp57j] J. R. Oppenheimer. Tolman, Richard Chase. In Encyclopaedia Britannica. Encyclopaedia Britannica, Chicago, IL, USA, 1957.

Oppenheimer:1957:AMF

[Opp57k] J. Robert Oppenheimer. Atomkraft und menschliche Freiheit. (German) [Atomic energy and human freedom]. ????, ????, 1957. 126 pp.

Oppenheimer:1957:ETD

[Opp57l] J. Robert Oppenheimer. Electron theory: Description and analogy. Physics Today, 10(7):12–20, July 1957. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). The first John Frank Carlson Lecture.

Oppenheimer:1957:IC

[Opp57m] J. Robert Oppenheimer. Impossible choices. Science, 125(3256): 1021, May 24, 1957. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Oppenheimer:1957:TU

[Opp57n] J. Robert Oppenheimer. Talk to undergraduates. Engineering and Science Monthly [California Institute of Technology], ??(??): ??, ???? 1957.

Oppenheimer:1958:CR

[Opp58a] J. R. Oppenheimer. Concluding remarks. Rochester/CERN Conference., 1958.

Oppenheimer:1958:DPI

[Opp58b] J. R. Oppenheimer. Description des particles et interactions élémentaires. (French) [description of particles and elementary interactions]. L’Université de Paris, France., 1958.

Oppenheimer:1958:ILb

[Opp58c] J. R. Oppenheimer. An inward look. In Kermit Lansner, editor, Second-Rate Brains: [A factual, perceptive report by top scientists, educators, journalists, and their urgent recommendations], page ?? Doubleday’s News Book, New York, NY, USA, 1958. LCCN ????

Oppenheimer:1958:KC

[Opp58d] J. R. Oppenheimer. Knowledge and culture. Hampton Institute., 1958.

Oppenheimer:1958:KSC

[Opp58e] J. R. Oppenheimer. Knowledge and the structure of culture. Helen Kenyon Lecture, Vassar College, Poughkeepsie, NY, 29 October 1958., 1958.

Oppenheimer:1958:SMR

[Opp58f] J. R. Oppenheimer. La science moderne et la raison. Société Française de Philosophie, 1958.

Oppenheimer:1958:LS

[Opp58g] J. R. Oppenheimer. L’arbre de la science. L’Université de Paris, France., 1958.

Oppenheimer:1958:LO

[Opp58h] J. R. Oppenheimer. L’espoir de l’ordre. Science, ??(??):??, ???? ??, 1958. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Oppenheimer:1958:MM

[Opp58i] J. R. Oppenheimer. The mystery of matter. Saturday Evening Post, ??(??), ???? 1958.

Oppenheimer:1958:SS

[Opp58j] J. R. Oppenheimer. Science and statecraft. Talk at the Weizmann Institute., 1958.

Oppenheimer:1958:SSC

[Opp58k] J. R. Oppenheimer. Science and the structure of culture. Talk at Rutgers University., 1958.

Oppenheimer:1958:SWT

[Opp58l] J. R. Oppenheimer. Science and the world today. Talk at Princeton Theological Seminary., 1958.

Oppenheimer:1958:ST

[Opp58m] J. R. Oppenheimer. A study of thinking. The Sewanee Review, 66:481–490, 1958. ISSN 0037-3052 (print), 1934-421X (electronic). Sewanee Review of Bruner Book.

Oppenheimer:1958:GSS

[Opp58n] J. Robert Oppenheimer. The growth of science and the structure of culture. Daedulus, ??(??):??, Winter 1958.

Oppenheimer:1958:ILa

[Opp58o] Robert Oppenheimer. An inward look. Foreign Affairs, 36 (3):209–220, January 1958. CODEN FRNAA3. ISSN 0015-

7120. URL http://www.foreignaffairs.com/articles/
71376/robert-oppenheimer/an-inward-look;   http://www.
jstor.org/stable/20029277. Reprinted in [Opp58c].

Oppenheimer:1958:TK

[Opp58p] Robert Oppenheimer. The tree of knowledge. Harper’s Magazine, ??(??):55–60, October 1958. CODEN HAMAA3. ISSN 1045-7143. Presented at the International Press Institute, and later as a lecture at the University of Wisconsin, 10 May 1959.

Oppenheimer:1959:CDF

[Opp59a] J. R. Oppenheimer. Contemporary developments in the field of science. Lawrenceville Herodotus Club., 1959.

Oppenheimer:1959:FNS

[Opp59b] J. R. Oppenheimer. Freedom and necessity in the sciences. Dartmouth College., 1959.

Oppenheimer:1959:GC

[Opp59c] J. R. Oppenheimer. The great challenge. CBS/TV., 1959.

Oppenheimer:1959:KUa

[Opp59d] J. R. Oppenheimer. In the keeping of unreason. Congress for

cultural freedom, published in Prospective., 1959.

Oppenheimer:1959:KUb

[Opp59e] J. R. Oppenheimer. In the keeping of unreason. Prospective, ??

(??):??, ???? 1959.

Oppenheimer:1959:NIU

[Opp59f] J. R. Oppenheimer. NATO and the ideal of unity. Princeton

University, NATO Conference., 1959.

Oppenheimer:1959:NNKa

[Opp59g] J. R. Oppenheimer. The need for new knowledge. In Dad Wolfie, editor, Symposium on Basic Research, pages 1–15. American Association for the Advancement of Science, Washington, DC, 1959. LCCN ????

Oppenheimer:1959:NNKb

[Opp59h] J. R. Oppenheimer. The need for new knowledge. Weaver Symposium. Published in translation in Revista de Occidente, March 1963., 1959.

Oppenheimer:1959:PFB

[Opp59i] J. R. Oppenheimer. Policy and the first bombs. Science, 130 (3388):1530–1592, December 4, 1959. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Oppenheimer:1959:RSP

[Opp59j] J. R. Oppenheimer. Reflections on science and philosophy. Yale

University, Hoyt Lecture., 1959.

Oppenheimer:1959:R

[Opp59k] J. R. Oppenheimer. Remarks. Dinner for Harold Taylor., 1959. Oppenheimer:1959:RBA

[Opp59l] J. R. Oppenheimer. The role of the big accelerators. Think

magazine, ??(??):??, ???? 1959.

Oppenheimer:1959:SHC

[Opp59m] J. R. Oppenheimer. Science and the human community. In Issues in university education [Fra59], pages 48–62. LCCN LB2325 .F7.

Oppenheimer:1959:TU

[Opp59n] J. R. Oppenheimer. Talk to undergraduates. Kansas Teacher, ??(??):18–20, 39–40, February 1959.

Oppenheimer:1959:TD

[Opp59o] J. R. Oppenheimer. Tradition and discovery. ACLS, Rochester., 1959.

Oppenheimer:1959:GSS

[Opp59p] J. Robert Oppenheimer. The growth of science and the structure of culture. Bulletin of the Atomic Scientists, 15(2):76, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Two-paragraph excerpt from the author’s article of the same name in Daedalus, winter 1958.

Oppenheimer:1959:SVH

[Opp59q] J. Robert Oppenheimer. Science, values, and the human community. Bulletin of the Atomic Scientists, 15(1):7, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1960:APO

[Opp60a] J. R. Oppenheimer. An afternoon with Professor Oppenheimer. Society of Science and Man, Tokyo, Japan., 1960.

Oppenheimer:1960:CK

[Opp60b] J. R. Oppenheimer. Common knowledge. Talk at Reed College., 1960.

Oppenheimer:1960:HS

[Opp60c] J. R. Oppenheimer. The house of science. American Institute of Architects., 1960.

Oppenheimer:1960:KSA

[Opp60d] J. R. Oppenheimer. Knowledge as science, action, culture. Talk at Queen’s University, Kingston, ON, Canada., 1960.

Oppenheimer:1960:LRA

[Opp60e] J. R. Oppenheimer. Leprince-Ringuet’s ’Des Atomes et des Hommes’. University of Chicago Press. See [LR57]., 1960.

Oppenheimer:1960:MM

[Opp60f] J. R. Oppenheimer. The mystery of matter. In Adventures of the mind, page ?? Vintage Books, New York, NY, USA, 1960. Reprint of [Opp58i].

Oppenheimer:1960:SC

[Opp60g] J. R. Oppenheimer. Science and culture. International House of Japan Bulletin, ??(??):2–10, October 1960. CODEN ???? ISSN 0285-2608. Variation on [Opp62c].

Oppenheimer:1960:SCE

[Opp60h] J. R. Oppenheimer. Science, culture et expression. Prospective, 5(??):??, ???? 1960. Translated abbreviated version of [Opp59e].

Oppenheimer:1960:SRS

[Opp60i] J. R. Oppenheimer. Some Reflections on Science and Culture. The John Calvin McNair lectures, 1959. University of North Carolina, Chapel Hill, NC, USA, 1960. 28 pp. Talk at University of North Carolina, Chapel Hill (1959).

Oppenheimer:1960:STS

[Opp60j] J. R. Oppenheimer. Some thoughts on science and politics. Princeton University Woodrow Wilson School, 1960.

Oppenheimer:1960:SRa

[Opp60k] J. R. Oppenheimer. Sorrow and renewal. Speech at Congress for Cultural Freedom, Berlin. Published in [Opp60l]., 1960.

Oppenheimer:1960:SRb

[Opp60l] J. R. Oppenheimer. Sorrow and renewal. Encounter, ??(??):??, ???? 1960. Originally given in [Opp60k].

Oppenheimer:1960:SOA

[Opp60m] J. R. Oppenheimer. Speaking to one another. University of Pennsylvania, Franklin Lecture., 1960.

Oppenheimer:1960:KU

[Opp60n] J. Robert Oppenheimer. In the keeping of unreason. Bulletin of the Atomic Scientists, 16(1):18–22, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This article was given at a seminar on “Industrial Society and the Western Political Dialogue”, help by the Congress for Cultural Freedom at Rheinfelden, Switzerland, September 20–27, 1959.

Oppenheimer:1961:FCS

[Opp61a] J. R. Oppenheimer. The future of civilization in the scientific age. France-Asie, ??(??):1807–1815, March/April 1961.

Oppenheimer:1961:SC

[Opp61b] J. R. Oppenheimer. Science and converse. Princeton University Graduate College Forum., 1961.

Oppenheimer:1961:SSAa

[Opp61c] J. R. Oppenheimer. Secretary Stimson and the atomic bomb. Phillips Academy, Andover. Published in [Opp61d]., 1961.

Oppenheimer:1961:SSAb

[Opp61d] J. R. Oppenheimer. Secretary Stimson and the atomic bomb. Andover Bulletin, ??(??):??, Spring 1961. Original given as talk [Opp61c].

Oppenheimer:1961:SHP

[Opp61e] J. R. Oppenheimer. Some human problems of our scientific age. Tribune Libre Universitaire, Brussels. Text reprinted in review of Tribune Libre Universitaire., 1961.

Oppenheimer:1961:TSR

[Opp61f] J. R. Oppenheimer. A time of sorrow and renewal. Encounter, 16(??):71–??, February 1961. Originally given in [Opp60k].

Oppenheimer:1961:RSCa

[Opp61g] J. Robert Oppenheimer. Reflections on science and culture. Colorado Quarterly, 10(2):101–111, Autumn 1961. CODEN ???? ISSN 0010-1710. Reprinted in [Opp62d].

Oppenheimer:1961:RSCb

[Opp61h] J. Robert Oppenheimer. Reflections on science and culture. University of Alberta, Edmonton, AB, Canada., 1961.

Oppenheimer:1962:FAH

[Opp62a] J. R. Oppenheimer. Freedom as an attribute of human life. Talk at the Congress for Cultural Freedom. Published in [Opp62b].,

Oppenheimer:1962:HH

[Opp62b] J. R. Oppenheimer. History and Hope. ????, London, UK, 1962. ???? pp. LCCN ???? Original talk in [Opp62a].

Oppenheimer:1962:SC

[Opp62c] J. R. Oppenheimer. On science and culture. Encounter, 19 (4):3–11, October 1962. Variations entitled Some reflections on science and culture widely published.

Oppenheimer:1962:RSC

[Opp62d] J. R. Oppenheimer. Reflections on science and culture. The Mexico Quarterly Review, ??(??):??, Spring 1962. Reprint of [Opp61g].

Oppenheimer:1962:RRP

[Opp62e] J. R. Oppenheimer. Reflections on the resonances of physics history: talk presented at the dedication ceremony of the Niels Bohr Library of the American Institute of Physics, 1962. American Institute of Physics collection, 1962. 6 pp.

Oppenheimer:1962:LEJ

[Opp62f] J. Robert Oppenheimer. Letter to the Editor: Japan subdued. Bulletin of the Atomic Scientists, 18(6):41, 43, June 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ols62a].

Oppenheimer:1963:ACa

[Opp63a] J. R. Oppenheimer. The added cubit. Talk at the National Book Awards. Published in [Opp63b]., 1963.

Oppenheimer:1963:ACb

[Opp63b] J. R. Oppenheimer. The added cubit. Encounter, ??(??):??, ???? 1963. Original talk in [Opp63a].

Oppenheimer:1963:CCSb

[Opp63c] J. R. Oppenheimer. Communication and comprehension of scientific knowledge. National Academy of Sciences Centennial. Published in [Opp63l, Opp63d]., 1963.

Oppenheimer:1963:CCSc

[Opp63d] J. R. Oppenheimer. Communication and comprehension of scientific knowledge. In Melvin Calvin, editor, The Scientific Endeavour: centennial celebration of the National Academy of Sciences, page ?? Rockefeller Institute Press, New York, NY, USA, 1963. Addresses presented at the centennial celebration of the National Academy of Sciences.

Oppenheimer:1963:CPL

[Opp63e] J. R. Oppenheimer. Complementarity [Pegram Lecture 2]. Pegram Lectures, Brookhaven., 1963.

Oppenheimer:1963:NBH

[Opp63f] J. R. Oppenheimer. Niels Bohr and his times. Pegram Lectures, Brookhaven., 1963.

Oppenheimer:1963:NBM

[Opp63g] J. R. Oppenheimer. Niels Bohr memoir. In Year Book, American Philosophical Society, page ?? American Philosophical Society, Philadelphia, PA, USA, 1963.

Oppenheimer:1963:NAP

[Opp63h] J. R. Oppenheimer. The nuclear atom [Pegram Lecture 1]. Pegram Lectures, Brookhaven., 1963.

Oppenheimer:1963:SRE

[Opp63i] J. R. Oppenheimer. The scientific revolution and its effect on democratic institutions. Talk at Fund for the Republic, 10th Anniversary. Published in [Opp63r]., 1963.

Oppenheimer:1963:TSS

[Opp63j] J. R. Oppenheimer. Talk at Seven Springs Farm, Mount Kisco, NY, USA. June 1963.

Oppenheimer:1963:AWA

[Opp63k] J. Robert Oppenheimer. Atomic weapons and American policy. In Grodzins and Rabinowitch [GR63], pages 188–196. LCCN D842 .B78. Reprint of [Opp53c].

Oppenheimer:1963:CCSa

[Opp63l] J. Robert Oppenheimer. Communication and comprehension of scientific knowledge. Proceedings of the National Academy of Sciences of the United States of America, 50(6):1194–1200, ???? 1963. CODEN PNASA6. ISSN 0027-8424 (print), 1091-

6490 (electronic). URL http://www.pnas.org/content/50/6/
1194.full.pdf+html?.

Oppenheimer:1963:FIC

[Opp63m] J. Robert Oppenheimer. The failure of international control. In Grodzins and Rabinowitch [GR63], pages 64–75. LCCN D842 .B78. Reprint of [Opp48g].

Oppenheimer:1963:FNT

[Opp63n] J. Robert Oppenheimer. The fire next time: Book review: Dawn of a new age, by Eugene Rabinowitch, University of Chicago Press. New York Review of Books, ??(??):??, September 26,

1963. URL http://www.nybooks.com/articles/1963/09/26/
the-fire-next-time/.

Oppenheimer:1963:ICA

[Opp63o] J. Robert Oppenheimer. The international control of atomic energy. In Grodzins and Rabinowitch [GR63], pages 53–63. LCCN D842 .B78. Reprint of [Opp46l].

Oppenheimer:1963:NB

[Opp63p] J. Robert Oppenheimer. Niels Bohr. Bulletin of the Atomic Scientists, 19(7):9, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1963:OM

[Opp63q] J. Robert Oppenheimer. The Open Mind. Simon and Schuster, New York, NY, USA, 1963. viii + 152 pp. LCCN QC780.

Oppenheimer:1963:TC

[Opp63r] J. Robert Oppenheimer. A talk in Chicago. Bulletin of the Atomic Scientists, 19(8):4–6, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). First presented as a talk, “The Scientific Revolution and Its Effects on Democratic Institutions”, in Chicago on June 13, 1963, as part of the symposium on “Challenges to Democracy”, sponsored by the Fund for the Republic.

Oppenheimer:1963:CCSd

[Opp63s] R. Oppenheimer. Communication and comprehension of scientific knowledge. Science, 142(3596):1143–1146, November 29, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Oppenheimer:1964:FDa

[Opp64a] J. R. Oppenheimer. The fraternal dialogue. Talk at l’Université de Paix. Published in [Opp64b]., 1964.

Oppenheimer:1964:FDb

[Opp64b] J. R. Oppenheimer. The fraternal dialogue. From Heart to Heart, page ??, 1964.

Oppenheimer:1964:HF

[Opp64c] J. R. Oppenheimer. Hope and foreknowledge. California Institute of Technology., 1964.

Oppenheimer:1964:LCF

[Opp64d] J. R. Oppenheimer. L’intime et le commun. (French) [The intimate and the common]. Talk at Rencontres Internationales de Genève., 1964.

Oppenheimer:1964:OTG

[Opp64e] J. R. Oppenheimer. Our times as Galillean times. In Carlo Maccagni, editor, Saggi su Galileo Galilei. (Italian) [Essays on Galileo Galilei], volume 3, page ?? G. Barbèra, Firenze, Italy, 1964. Pubblicazioni del Comitato nazionale per le manifestazioni celebrative del 4. centenario della nascita di Galileo Galilei.

Oppenheimer:1964:PASa

[Opp64f] J. R. Oppenheimer. Prospects in the arts and sciences. Talk at the Columbia University Bicentennial. Reprinted in [Opp64g, Opp64h]., 1964.

Oppenheimer:1964:PASb

[Opp64g] J. R. Oppenheimer. Prospects in the arts and sciences. In Mark van Doren, editor, Man’s Right to Knowledge: An international symposium presented in honor of the two-hundredth anniversary of Columbia University, 1754–1954, page ?? Muschel, New York, NY, USA, 1964. LCCN ???? Second series. With an introduction by Richard R. Powell.

Oppenheimer:1964:PASc

[Opp64h] J. R. Oppenheimer. Prospects in the arts and sciences. In Huberman Elizabeth and Huberman Edward, editors, Fifty Great Essays, page ?? Bantam Books, New York, NY, USA, 1964. LCCN ????

Oppenheimer:1964:ERB

[Opp64i] J. Robert Oppenheimer. Ernest Rutherford: Book review of Rutherford and the Nature of the Atom, by E. N. da C. Andrade, Anchor, 213 pp. New York Review of Books, ??(??): ??, May 14, 1964. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/1964/05/
14/ernest-rutherford/.

Oppenheimer:1964:FTT

[Opp64j] J. Robert Oppenheimer. The Flying Trapeze: Three Crises for Physicists. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1964. x + 65 pp. LCCN ???? MacMaster University Whidden lectures for 1962.

Oppenheimer:1964:NBN

[Opp64k] J. Robert Oppenheimer. Niels Bohr and nuclear weapons. New York Review of Books, 3(9):6–8, December 17, 1964. CODEN ???? ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1964/dec/17/niels-bohr-and-atomic-weapons/.

Oppenheimer:1965:AT

[Opp65a] J. R. Oppenheimer. The 20th anniversary of Trinity. Washington Post, Outlook Section, ??(??):??, ???? 1965.

Oppenheimer:1965:AO

[Opp65b] J. R. Oppenheimer. Alpha or omega. Washington Post, ??(??): ??, ???? 1965. ISSN 0190-8286.

Oppenheimer:1965:DDB

[Opp65c] J. R. Oppenheimer. Decision to drop the bomb. NBC White Paper, Books by Coward McCann, 1965.

Oppenheimer:1965:FBM

[Opp65d] J. R. Oppenheimer. Foreword to Nature of matter — purpose of high energy physics. Current foreword, published by Brookhaven National Laboratory., 1965.

Oppenheimer:1965:AE

[Opp65e] J. R. Oppenheimer. On Albert Einstein. UNESCO, Paris. Published in [Opp66d]., 1965.

Oppenheimer:1965:Pa

[Opp65f] J. R. Oppenheimer. Physics. Published in [Opp65g]., 1965.

Oppenheimer:1965:Pb

[Opp65g] J. R. Oppenheimer. Physics. In Albert Love and James Saxon Childers, editors, Listen to leaders in science, pages 48–?? Tuve and Love, Inc., Atlanta, GA, USA, 1965. LCCN ????

Oppenheimer:1965:RSP

[Opp65h] J. R. Oppenheimer. Remarks on symmetry principles. In Behram Kursunoğlu, Arnold Perlmutter, and Ismail Sakmar, editors, Coral Gables conferences on symmetry principles at high energy: second conference, January 20–22, 1965 University of Miami, page ?? W. H. Freeman, New York, NY, USA, 1965. LCCN ????

Oppenheimer:1965:SM

[Opp65i] J. R. Oppenheimer. Science in the making. In Principal addresses of the 1965 U.S. Army National Junior Science and Humanities Symposium, 29–30 April 1965. Ft. Belvoir Defense Technical Information Center, Ft. Belvoir, VA, USA, 1965.

Oppenheimer:1965:SFS

[Opp65j] J. R. Oppenheimer. The symmetries of forces and states. In Amos de Shalit, Herman Feshbach, and Léon van Hove, editors, Preludes in Theoretical Physics in honor of V. F. Weisskopf, page ?? North-Holland Publishing Co., Amsterdam, The Netherlands, 1965. LCCN QC71 .S39 1966.

Oppenheimer:1966:C

[Opp66a] J. R. Oppenheimer. Crossing. Hound and Horn, 1(??):335–??, ???? 1966.

Oppenheimer:1966:FN

[Opp66b] J. R. Oppenheimer. The forbearance of nations. Herald Tribune Paris — Washington Post, Paris Edition, ??(??):16, December 6, 1966.

Oppenheimer:1966:KAM

[Opp66c] J. R. Oppenheimer. Knowledge among men. Encounter, ??(??): ??, ???? 1966. Original talk in [Opp66f].

Oppenheimer:1966:AE

[Opp66d] J. R. Oppenheimer. On Albert Einstein. New York Review of Books, 6(4):4–5, March 17, 1966. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1966/mar/17/on-albert-einstein/. Lecture deliv-

ered at UNESCO House in Paris on December 13, 1965.

Oppenheimer:1966:PMP

[Opp66e] J. R. Oppenheimer. Perspectives in modern physics. In Robert E. Marshak, editor, Perspectives in modern physics: Essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July, 1966, page ?? Interscience Publishers, New York, NY, USA, 1966. LCCN QC774.B4 M3. With the assistance of J. Warren Blaker.

Oppenheimer:1966:PMU

[Opp66f] J. R. Oppenheimer. Physics and man’s understanding. Talk at Smithsonian Institution Bicentennial. Published in [Opp66c]., 1966.

Oppenheimer:1966:TYM

[Opp66g] Robert Oppenheimer. Thirty years of mesons. Physics Today, 19(11):51–58, November 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v19/i11/p51_s1.

Oppenheimer:1969:FTT

[Opp69] J. Robert Oppenheimer. The Flying trapeze: Three Crises for Physicists. Harper & Row, New York, NY, USA, 1969. ix + 69 pp. LCCN ???? MacMaster University Whidden lectures for 1962.

Oppenheimer:1970:LE

[Opp70] J. Robert Oppenheimer. Lectures on Electrodynamics. Gordon and Breach, New York, 1970. ???? pp. Notes compiled at University of California Berkeley by S. Kusaka with S. Frankel and E. Nelson. Edited by B. S. DeWitt.

Oppenheimer:1971:RRP

[Opp71] J. R. Oppenheimer. Reflections on the resonances of physics history. Niels Bohr Library, New York, NY, USA, 1971. 7 pp.

Oppenheimer:1972:RRP

[Opp72] J. R. Oppenheimer. Reflections on the resonances of physics history. American Institute of Physics, Woodbury, NY, USA, 1972. 8 pp. Publication R-239.1.

Oppenheimer:1975:ET

[Opp75] J. Robert Oppenheimer. 5.2. electron theory. In Mehra [Meh75], chapter 10, pages 257–259. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Oppenheimer:1976:SHC

[Opp76] J. R. Oppenheimer. Science and the human community. In Issues in university education [Fra59], page ?? LCCN LB2325 .F7.

Oppenheimer:1979:AE

[Opp79a] J. R. Oppenheimer. On Albert Einstein. In French [Fre79], page ?? ISBN 0-674-24230-0. LCCN QC16.E5 E37.

Oppenheimer:1979:CSB

[Opp79b] J. Robert Oppenheimer. Celebration of the sixtieth birthday of Albert Einstein. Bulletin of the Atomic Scientists, 35(3):39, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Oppenheimer:1979:OE

[Opp79c] J. Robert Oppenheimer. Oppenheimer on Einstein. Bulletin of the Atomic Scientists, 35(3):36–38, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Oppenheimer:1980:SSS

[Opp80] J. R. Oppenheimer. Secrecy and science + speech to Association of Los Alamos Scientists 1945. New York Times Book Review, 85(19):9, ???? 1980. ISSN 0028-7806.

Oppenheimer:1989:AVE

[Opp89] J. Robert Oppenheimer. Atom and Void: Essays on Science and Community. Princeton science library. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-08547-1, 0-691-02434-0 (paperback). xi + 155 pp. LCCN Q175

.O65 1989. URL http://www.loc.gov/catdir/description/

prin031/89010413.html. Preface by Freeman J. Dyson.

Oppenheimer:19xx:LSG

[Oppxx] J. Robert Oppenheimer. Lantern slides of graphs, instruments, and characteristics of certain physical bodies, 19xx. URL http:

//lcweb2.loc.gov/pp/cphquery.htm. 35 slides.

Oppenheimer:2007:NWT

[Opp07]  J. R. Oppenheimer.  The new weapon: the turn of

the screw. In Masters and Way [MW07], pages 53– 60. ISBN 1-59558-227-4 (hardcover). LCCN UG1282.A8

O54 2007.  URL http://thenewpress.com/index.php?
option=com_title&task=view_title&metaproductid=1703;
 http://www.loc.gov/catdir/toc/ecip0718/2007020838.htm.

Foreword by Niels Bohr. Introduction by Arthur H. Compton. Reprint of [MW46].

Oppenheimer:2020:SOT

[Opp20] J. Robert Oppenheimer. 1956: Science and our times. Bulletin of the Atomic Scientists, 76(6):356–358, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847481.

Oppenheimer:2023:OR

[Opp23] J. R. Oppenheimer. Oppenheimer replies. Bulletin of the Atomic Scientists, 79(4):242–254, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:/

/www.tandfonline.com/doi/full/10.1080/00963402.2023.
2223084.

Oppenheimer:1947:DNA

[OR47] J. Robert Oppenheimer and I. I. Rabi. Draft note on atomic power. In archive of Oppenheimer papers., October 23, 1947.

Oppenheimer:1969:OCR

[OR69] J. Robert Oppenheimer and I. I. (Isidor Isaac) Rabi, editors. Oppenheimer. Scribner, New York, NY, USA, 1969. x + 90 pp. LCCN QC16.O62 O6.

ORegan:1995:JRO

[O’R95] Richard O’Regan. J. Robert Oppenheimer: Father of the atomic bomb. Produced by Claymont Productions in association with A&E Network, 1995. 50 mins. (A&E Biography, 235 East 45th St., New York, NY 10017)., 1995.

Oreskes:2000:SSB

[OR00] Naomi Oreskes and Ronald Rainger. Science and security before the atomic bomb: The loyalty case of Harald U. Sverdrup. Studies in History and Philosophy of Modern Physics, 31(3):309– 369, September 2000. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219800000198.

Ortega-Rodriguez:2017:ESC

[ORSSBO^{+}17] M. Ortega-Rodríguez, H. Solís-Sánchez, E. Boza-Oviedo, K. Chaves-Cruz, M. Guevara-Bertsch, M. Quirós-Rojas, S. Vargas-Hernández, and A. Venegas-Li. The early scientific contributions of J. Robert Oppenheimer: Why did the scientific community miss the black hole opportunity? Physics in Perspective (PIP), 19(1):60–75, April 2017. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-017-0195-6.

OrtegayGasset:1932:RM

[Ort32] José Ortega y Gasset. The revolt of the masses. G. Allen and Unwin, Ltd., London, UK, 1932. 204 + 4 pp. LCCN CB103 .O72 1932a.

OrtegayGasset:1957:RM

[Ort57] José Ortega y Gasset. The revolt of the masses. W. W. Norton & Co., New York, NY, USA, 25th anniversary edition, 1957. 190 pp. LCCN CB103 .O72 1957. Authorized translation from the Spanish.

OrtegayGasset:1993:RM

[Ort93] José Ortega y Gasset. The revolt of the masses. W. W. Norton & Co., New York, NY, USA, 1993. ISBN 0-393-31095-7 (paperback). 190 pp. LCCN CB103 .O713 1993. US$6.95.

Oppenheimer:1936:DNL

[OS36] J. R. Oppenheimer and R. Serber. 35. the density of nuclear levels. Physical Review (2), 50(4):391, August 1936. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v50/i4/

p384_1. In Minutes of the Seattle Meeting, June 17–19, 1936.

Oppenheimer:1937:NNC

[OS37] J. R. Oppenheimer and R. Serber. Note on the nature of cosmicray particles. Physical Review (2), 51(12):1113, June 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v51/i12/
p1113_1.

Oppenheimer:1938:NBP

[OS38a] J. R. Oppenheimer and R. Serber. Note on boron plus proton reactions. Physical Review (2), 53(8):636–638, April 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v53/
i8/p636_1.

Oppenheimer:1938:SSN

[OS38b] J. R. Oppenheimer and Robert Serber. On the stability of stellar neutron cores. Physical Review (2), 54(7):540, October 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v54/
i7/p540_1.

Oppenheimer:1939:PEP

[OS39a] J. R. Oppenheimer and J. S. Schwinger. On pair emission in the proton bombardment of fluorine. Physical Review (2), 56 (10):1066–1067, November 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v56/i10/p1066_1.

Oppenheimer:1939:CGC

[OS39b] J. R. Oppenheimer and H. Snyder. On continued gravitational contraction. Physical Review (2), 56(5):455–459, September 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v56/
i5/p455_1.

Oppenheimer:1941:IMN

[OS41] J. R. Oppenheimer and Julian Schwinger. On the interaction of mesotrons and nuclei. Physical Review (2), 60(2):150–152, July 15, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v60/i2/p150_1.

Oppenheimer:1940:PSS

[OSS40] J. R. Oppenheimer, H. Snyder, and R. Serber. The production of soft secondaries by mesotrons. Physical Review (2), 57(2): 75–81, January 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v57/i2/p75_1.

Oppenheimer:1939:MNC

[OV39] J. R. Oppenheimer and G. M. Volkoff. On massive neutron cores. Physical Review (2), 55(4):374–381, February 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v55/i4/p374_1.

Oppenheimer:1983:S

[OW83] J. Robert Oppenheimer and Thomas J. Watson, Jr. Speeches. Annals of the History of Computing, 5(2):157–160, April/June 1983. CODEN AHCOE5. ISSN 0164-1239. URL http://dlib.

computer.org/an/books/an1983/pdf/a2157.pdf;   http:/
/www.computer.org/annals/an1983/a2157abs.htm.

Oppenheimer:2013:PASa

[OW13a] Frank Oppenheimer and Charles Weiner. A physicist for all seasons: Part I. Physics in Perspective (PIP), 15(1):33–91, March 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0009-6.

Oppenheimer:2013:PASb

[OW13b] Frank Oppenheimer and Charles Weiner. A physicist for all seasons: Part II. Physics in Perspective (PIP), 15(2):178–240, June 2013. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-009-0010-0.

Pais:1967:PP

[Pai67] Abraham Pais. The Princeton period. Physics Today, 20(10):42– 48, October 1967. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v20/i10/p42_s1.

Pais:1999:MRS

[Pai99] Abraham Pais. In memoriam: Robert Serber (1909–1997). Physics in Perspective (PIP), 1(1):105–109, March 1999. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s000160050008.

Patsilelis:2005:IAE

[Pat05] Chris Patsilelis. Inside Los Alamos/examining the man and the culture behind the atomic bomb. Houston Chronicle, ??(??): Z.20–??, July 10, 2005. ISSN 1074-7109. URL http://search.

proquest.com/docview/396202146.

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN

QC16.O62 P35 2006. URL http://www.loc.gov/catdir/
enhancements/fy0636/2005002173-d.html;   http://www.
loc.gov/catdir/enhancements/fy0723/2005002173-b.html;
 http://www.loc.gov/catdir/toc/ecip056/2005002173.html.

Peierls:1955:MJR

[Pei55] Rudolf Peierls. In the matter of J. Robert Oppenheimer. Atomic Scientists’ Journal, 4(??):145–154, ???? 1955.

Peierls:1974:JRO

[Pei74] Rudolf Peierls. J. Robert Oppenheimer. In Charles Coulston Gillispie, editor, Dictionary of Scientific Biography: Volume 10: S. G. Navashin–W. Piso, pages 213–218. Charles Scribner’s Sons, New York, NY, USA, 1974. ISBN 0-684-10121-1. LCCN Q141D5 1970.

Peierls:1980:BRG

[Pei80] Sir Rudolf Peierls. Book review: The growing pains of Robert Oppenheimer. Robert Oppenheimer: Letters and Recollections. Edited by Alice Kimball Smith and Charles Weiner. Pp. 250. (Harvard University Press: 1980). Nature, 286(5775):827–828, August 21, 1980. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL https://ui.adsabs.harvard.
edu/#abs/1980Natur.286..827P.

Peierls:1997:AH

[Pei97] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Peruzzi:1994:BRL

[Per94] Giulio Peruzzi. Book review: Lillian Hoddeson, Paul W. Henriksen, Roger A. Meade, Catherine Westfall (eds.), Critical Assembly: A Technical History of Los Alamos during the Oppenheimer Years, 1943–1945, Cambridge, Cambridge University Press, 1993, xv + 509 pp. Nuncius, 9(2):888–892, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01332.

Perkins:1995:BJR

[Per95] Jack Perkins. Biography. J. Robert Oppenheimer, father of the atomic bomb, 1995. ISBN 1-56501-601-7. 1 videocassette of 1 (vhs) (ca. 50 min.) :1 videodisc of 1 (dvd) (ca. 50 min.).

Pfau:1984:NST

[Pfa84] Richard Pfau. No Sacrifice Too Great: the Life of Lewis L. Strauss. University Press of Virginia, Charlottesville, VA, USA, 1984. ISBN 0-8139-1038-2. x + 314 pp. LCCN QC774.S77 P43 1984.

Prince:2008:JRO

[PGDW08] Peter Prince, Peter Goodchild, Barry Davis, and Sam Waterston. Oppenheimer : the father of the atomic bomb. BBC Worldwide film series (420 minutes on three DVDs)., 2008. ISBN 1-4198-6756-3.

Pines:2015:WWD

[Pin15] David Pines. What we don’t understand, we explain to each other. The Physics Teacher, 53(9):526–531, December 2015. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Polenberg:2002:MJR

[Pol02] Richard Polenberg, editor. In the matter of J. Robert Oppenheimer: the security clearance hearing. Cornell paperbacks. Cornell University Press, Ithaca, NY, USA, 2002. ISBN 0-8014-8661-0 (paperback), 0-8014-3783-0 (cloth). xxxii + 409 pp. LCCN QC16.O62 I5 2002. URL http://www.h-net.org/

review/hrev-a0c5p6-aa.

Polenberg:2005:FF

[Pol05] Richard Polenberg. The fortunate fox. In Carson and Hollinger [CH05], chapter 12, pages 267–272. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Polenberg:2007:BTI

[Pol07] Richard Polenberg. Book review: Oppenheimer: The Tragic Intellect. By Charles Thorpe. (Chicago: University of Chicago Press, 2006. xx, 413 pp. $37.50, ISBN 978-0-226-79845-5). The Journal of American History, 94(2):634–635, 2007. ISSN 0021-8723 (print), 1945-2314 (electronic). URL http://jah.

oxfordjournals.org/content/94/2/634.short.

Polanyi:2019:WMD

[Pol19] John Polanyi. We must do more to prevent nuclear war. Times Higher Education, ??(??):??, December 9, 2019.

URL https://www.timeshighereducation.com/features/
we-must-do-more-prevent-nuclear-war.

Porter:2007:OWM

[Por07] Jeffrey Lyn Porter. Oppenheimer is watching me: a memoir. Sightline books. University of Iowa Press, Iowa City, IA, USA, 2007. ISBN 1-58729-616-0 (hardcover). xiv + 132 pp. LCCN E169.Z8 P5924 2007. URL http://site.ebrary.com/

lib/stanford/Doc?id=1035469.

Powell:1996:DAT

[Pow96] C. S. Powell. The day after Trinity: J. Robert Oppenheimer and the atomic bomb. Scientific American, 274(5):105, May 1996. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Powers:2005:ATB

[Pow05] Thomas Powers. An American tragedy: American Prometheus: The Triumph and Tragedy of J. Robert Oppenheimer by Kai Bird and Martin J. Sherwin. The Ruin of J. Robert Oppenheimer and the Birth of the Modern Arms Race by Priscilla J. McMillan. Oppenheimer: Portrait of an Enigma by Jeremy Bernstein. J. Robert Oppenheimer and the American Century by David C. Cassidy. New York Review of Books, 52(14):73–79, September 22, 2005. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2005/sep/22/an-american-tragedy/.

Pryce:1955:BRB

[Pry55] M. H. L. Pryce. Book review: Science and the Common Understanding, by J. Robert Oppenheimer. Physics Today, 8(8): 16, August 1955. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Panati:2007:TDB

[PST07] Gianlucaa Panati, Herberta Spohn, and Stefana Teufel. The time-dependent Born–Oppenheimer approximation. Mathematical modelling and numerical analysis = Modelisation mathématique et analyse numérique: M^{2}AN, 41(??):297–314, March 2007. CODEN RMMAEV. ISSN 0764-583X (print), 1290-3841 (electronic).

Pungs:1967:HGM

[Pun67] L. Pungs. Hans-Georg Möller zum 85. Geburtstag/J. R. Oppenheimer/Unruhe um Mössbauer/Max-Planck-Medaille 1967. Physikalische Blätter, 23(5):223–227, 1967. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Putik:1960:FOG

[Put60] Jaroslav Putík. Der Fall Oppenheimer. (German) [The Oppenheimer case]. Domowina Verlag, Bautzen, Germany, 1960. 243 pp. LCCN MLCS 86/03442 (H). Translation to German from Czech by Ben Budar.

Putik:1961:FOG

[Put61] Jaroslav Putík. Der Fall Oppenheimer. (German) [The Oppenheimer case]. Domowina Verlag, Bautzen, Germany, second edition, 1961. 244 pp. LCCN MLCS 86/03442 (H). Translation to German from Czech by Ben Budar.

Putik:1963:SPP

[Put63] Jaroslav Putík. Svědomí; případ profesora Oppenheimera. (Czech) [Conscience: The Case of Professor Oppenheimer]. Mladá fronta, Praha, Czechoslovakia, third edition, 1963. 161 pp. LCCN QC16.O62 P8 1963.

Quinn-Judge:1994:ESS

[QJ94] Paul Quinn-Judge. Ex-Soviet spy stirs debate. Boston Globe, ??(??):25–??, May 8, 1994. ISSN 0743-1791. URL http://

search.proquest.com/docview/403615251/.

Rabinowitch:1954:WSR

[Rab54] Eugene Rabinowitch. What is a security risk? Bulletin of the Atomic Scientists, 10(7):241, 256, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabi:19xx:RPB

[Rabxx] I. I. (Isidor Isaac) Rabi. I. I. Rabi papers, 1899–1989 (bulk 1945– 1968). 19xx. ???? pp. LCCN 0639F; LCA (Unprocessed ac. 22,752); Vault 0202A; Oversize 6:10. URL http://hdl.loc.

gov/loc.mss/eadmss.ms998009.

Ragghianti:1980:LVC

[Rag80] C. Ragghianti. Leonardo da Vinci, conscientious objector, as an antecedent of Robert Oppenheimer. Critica d’Arte, 45(172–??): 72, ???? 1980. ISSN 0011-1511.

Raiteri:2002:FJR

[Rai02] S. Raiteri. Fallout: J. Robert Oppenheimer, Leo Szilard, and the political science of the atomic bomb. Library Journal, 127 (14):153, September 2002. CODEN LIBJA7. ISSN 0363-0277.

Rossi:1949:DDN

[RAO^{+}49] B. Rossi, C. D. Anderson, J. R. Oppenheimer, G. E. Valley, and R. D. Sard. Discussion on the disintegration and nuclear absorption of mesons. Reviews of Modern Physics, 21 (1):31–37, January 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.21.31;   http:
//rmp.aps.org/abstract/RMP/v21/i1/p31_1. Oppenheimer:

Remark on μ-Decay (pages 34–35).

Rabinowitch:1967:JRO

[RB67] Eugene Rabinowitch and Hans Bethe. J. Robert Oppenheimer, 1904–1967. Bulletin of the Atomic Scientists, 23(8):2–6, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Reed:2017:RAP

[Ree17] B. Cameron Reed. Revisiting The Los Alamos Primer. Physics Today, 70(9):42–49, September 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Reingold:1980:ORL

[Rei80] N. Reingold. Robert Oppenheimer — letters and recollections. Reviews in American History, 8(4):541–546, ???? 1980. ISSN 0048-7511 (print), 1080-6628 (electronic).

Reston:1954:DOS

[Res54] James Reston. Dr. Oppenheimer suspended by A.E.C. in security review; scientist defends record: Hearings started. access to secret data denied nuclear expert — Red ties alleged. New York Times, ??(??):1, 15, April 13, 1954. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/113112494/D7D8E4644FBF404EPQ/1.

Reston:1991:DM

[Res91] James Reston. Deadline: a memoir. Random House, New York, NY, USA, 1991. ISBN 0-394-58558-5. xvii + 525 + 16 pp. LCCN PN4874.R447 A3 1991.

Reston:1992:DM

[Res92] James Reston. Deadline: a memoir. Times Books, New York, NY, USA, 1992. ISBN 0-8129-2071-6. xvii + 539 pp. LCCN PN4874.R447 A3 1992. US$14.00 (US$17.50 Can.).

Rheaume:1995:ISB

[Rhe95] C. Rheaume. Internationalism of scientists and the bomb — the case of Robert Oppenheimer, Edward Teller, Igor Kourtchatov, and Andrei Sakharov. Social Science Information = Information sur les sciences sociales, 34(2):269–285, ???? 1995. ISSN 0539-0184 (print), 1461-7412 (electronic).

Rhodes:1977:BDA

[Rho77] Richard Rhodes. ‘I Am Become Death’: The agony of J. Robert Oppenheimer. American Heritage, 20(6):72–82, October 1977.

Rhodes:2005:JRO

[Rho05] R. Rhodes. J. Robert Oppenheimer and the American century. New York Times Book Review, ??(??):7–8, May 15, 2005. ISSN 0028-7806.

Rhodes:2009:RRB

[Rho09] Richard Rhodes. Richard Rhodes’ Reykjavik. Bulletin of the Atomic Scientists, 65(5):71–81, September 2009. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/65/5/71.abstract;
 http://bos.sagepub.com/content/65/5/71.full.pdf+html.

Rieff:1954:CDO

[Rie54] Philip Rieff. The case of Dr. Oppenheimer. The Twentieth Century, 156(??):113–232, September 1954.

Rieff:1969:CDO

[Rie69a] Philip Rieff. The case of Dr. Oppenheimer. In On Intellectuals; Theoretical Studies, Case Studies [Rie69b], pages 314–340. LCCN HM213 .R53.

Rieff:1969:ITS

[Rie69b] Philip Rieff. On Intellectuals; Theoretical Studies, Case Studies. Doubleday, Garden City, NY, USA, 1969. x + 347 pp. LCCN HM213 .R53.

Rigden:1987:RO

[Rig87a] John S. Rigden. Rabi and Oppenheimer. In Rabi, scientist and citizen [Rig87b], page ?? ISBN 0-465-06792-1. LCCN QC16.R2 R54 1987. US$22.95.

Rigden:1987:RSC

[Rig87b] John S. Rigden. Rabi, scientist and citizen. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-06792-1. xiii + 302 + 8 pp. LCCN QC16.R2 R54 1987. US$22.95.

Rigden:1995:JRO

[Rig95a] John S. J. Rigden. J. Robert Oppenheimer: before the War. Scientific American, 273(1):76–81, July 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v273/n1/pdf/scientificamerican0795-76.pdf. Al-

though Oppenheimer is now best remembered for his influence during World War II, he made many important contributions to theoretical physics in the 1930s.

Rigden:1995:ROB

[Rig95b] John S. J. Rigden. Robert Oppenheimer: Before the war. Scientific American, 273(1):68–??, July 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Rigden:2000:RO

[Rig00a] John S. Rigden. Rabi and Oppenheimer. In Rabi, scientist and citizen [Rig00b], pages 211–231. ISBN 0-674-00435-3. LCCN QC16.R2 R54 2000.

Rigden:2000:RSC

[Rig00b] John S. Rigden. Rabi, scientist and citizen. Harvard University Press, Cambridge, MA, USA, 2000. ISBN 0-674-00435-3. xxix + 302 pp. LCCN QC16.R2 R54 2000.

Risjord:2006:GTT

[Ris06] Norman K. Risjord. Giants in their time: representative Americans from the Jazz Age to the Cold War. Representative Americans. Rowman and Littlefield, Lanham, MD, USA, 2006. ISBN 0-7425-2784-0 (hardcover), 0-7425-2785-9 (paperback). x + 259 pp. LCCN CT220 .R57 2006. URL http://catdir.loc.gov/

catdir/toc/ecip0516/2005017389.html.

Rodgers:1950:TMR

[RMR^{+}50] Douglas Rodgers, Roger Muir, Eleanor Roosevelt, Albert Einstein, David Eli Lilienthal, J. Robert Oppenheimer, Brien McMahon, Hans A. (Hans Albrecht) Bethe, Harry Winne, Allan Kline, and Detlev W. (Detlev Wulf) Bronk. Today with Mrs. Roosevelt. NBC Television Network, United States, February 12, 1950. LCCN FSA 1920 (arch neg); FSA 1922 (mag trk); FCB 4264 (ref print); FTA 3029 (pos trk); FTA 3028 (masterpos); FCB 4367 (ref print, copy 2); FSA 1921 (neg trk). Seven tape reels.

Roberts:1956:BRB

[Rob56] Henry L. Roberts. Book review: The Open Mind, by J. Robert Oppenheimer. Foreign Affairs, 35(1):151–166, October 1956. CODEN FRNAA3. ISSN 0015-7120. URL http://www.

foreignaffairs.com/articles/129043/j-robert-oppenheimer/
the-open-mind;   http://www.jstor.org/stable/20031213.

Rogers:2010:MIS

[Rog10] Kara Rogers, editor. The 100 Most Influential Scientists of All Time. The Britannica guide to the world’s most influential people. Britannica Educational Publishers, in association with Rosen Educational Services, New York, NY, USA, 2010. ISBN 1-61530-002-3 (library binding). 360 pp. LCCN Q162 .A15 2010.

Romelt:1983:HRB

[Röm83] Joachim Römelt. A Hermitean reformulation of the Born– Oppenheimer nonadiabatic coupling terms for diatomic molecules. International Journal of Quantum Chemistry, 24(6):627–631, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rossi:1964:CR

[Ros64] Bruno Rossi. Cosmic Rays. McGraw-Hill, New York, NY, USA, 1964. viii + 268 pp. LCCN ????

Rothstein:1999:NSP

[Rot99] Linda Rothstein. Nuclear secrets: The Perseus Papers. Bulletin of the Atomic Scientists, 55(4):17–19, July 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/4/17.full.

Rouze:1962:ROB

[Rou62] Michel Rouzé. Robert Oppenheimer et la bombe atomique. (French) [Robert Oppenheimer and the atomic bomb], volume 7 of Savants du monde entier. Seghers, Paris, France, 1962. 223 pp. LCCN ????

Rouze:1963:ROA

[Rou63] Michel Rouzé. Robert Oppengejmer i atomnaja bomba. Gosatomizdat, Moskva, USSR, 1963. 147 pp. LCCN ????

Rouze:1964:ROM

[Rou64] Michel Rouzé. Robert Oppenheimer, the man and his theories. A Profile in science. Souvenir Press, London, UK, 1964. 189 pp. LCCN QC16.O62 R613 1964.

Rouze:1965:ROA

[Rou65a] Michel Rouzé. Robert Oppengejmer i atomnaja bomba. Atomizdat, Moskva, USSR, second edition, 1965. 147 pp. LCCN QC16.O62 R62217 1965x.

Rouze:1965:ROM

[Rou65b] Michel Rouzé. Robert Oppenheimer, the man and his theories. A Profile in science. P. S. Eriksson, New York, NY, USA, 1965. 192 pp. LCCN QC16.O62 R613 1965.

Rouze:1973:OVP

[Rou73] Michel Rouzé. Oppenheimer: la vita, il pensiero, i testi esemplari, volume 45 of I Memorabili. Edizioni Accademia, Milano, Italias, 1973. 222 pp. LCCN ????

Royal:1969:SJR

[Roy69] Denise Royal. The story of J. Robert Oppenheimer. St. Martin’s Press, New York, NY, USA, 1969. x + 196 pp. LCCN QC16.O62 R62.

Rigden:2004:JRO

[RS04] J. S. Rigden and R. H. Stuewer. J. Robert Oppenheimer — brilliance and charisma. Physics in Perspective (PIP), 6(4):369–371, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Rigden:2006:BNR

[RS06] John S. Rigden and Roger H. Stuewer. Book notes [Robyn Arianrhod, Einstein’s Heroes: Imagining the World Through the Language of Mathematics (New York: Oxford University Press, 2006), $16.95 (paper); Priscilla J. McMillan, The Ruin of J. Robert Oppenheimer and the Birth of the Modern Arms Race (New York: Penguin Books, 2005), $16.00 (paper)]. Physics in Perspective (PIP), 8(4):466–469, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-006-0324-0.

Rabi:1969:O

[RSW^{+}69] I. I. (Isidor Isaac) Rabi, Robert Serber, Victor F. Weisskopf, Abraham Pais, and Glenn T. Seaborg. Oppenheimer. Scribner, New York, NY, USA, 1969. x + 90 pp. LCCN QC16.O62 O6.

Reid:1969:ESM

[RTS^{+}69] R. W. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. 42-minute black-and-white motion picture from Time-Life Films, Paramus, NJ, USA., 1969. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Reid:20xx:ESM

[RTS^{+}xx] R. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. Motion picture, 20xx. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Rummel:1992:TOD

[Rum92] Jack Rummel. Robert Oppenheimer: Dark Prince. Makers of modern science. Facts On File, New York, NY, USA, 1992. ISBN 0-8160-2598-3. iv + 140 pp. LCCN QC16.O62 R85 1992.

Rutherford:1927:LSR

[Rut27] Sir Ernest Rutherford, O.M., P.R.S. LI. Structure of the radioactive atom and origin of the α-rays. Philosophical Magazine (7), 4(22):580–605, September 1927. CO-DEN PHMAA4. ISSN 1941-5982 (print), 1941-5990 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.

1080/14786440908564361. Cited in [Wil83, page 441] as ‘a great paper’. Wilson (page 559) later notes that this paper inspired George Gamow to his prediction of the quantum tunneling effect in 1929 (credit also goes to Edward Condon and Ronald Gurney who wrote two papers in 1928 on that idea, and to Robert Oppenheimer, who published a paper on that topic five months before those of Condon and Gurney).

S:1954:FOW

[S.54] H. S. Der Fall Oppenheimer - und was dann?. (German) [The Oppenheimer case — and then what?]. Sozialdemokratischer Pressedienst, 97:3–??, April 28, 1954. URL http://library.

fes.de/spdpd/1954/540428.pdf.

Sagdeev:1979:OLCa

[Sag79a] Roald Z. Sagdeev. The 1976 Oppenheimer Lectures: Critical problems in plasma astrophysics. I. Turbulence and nonlinear waves. Reviews of Modern Physics, 51(1):1–9, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.51.1;   http://rmp.aps.org/abstract/
RMP/v51/i1/p1_1.

Sagdeev:1979:OLCb

[Sag79b] Roald Z. Sagdeev. The 1976 Oppenheimer Lectures: Critical problems in plasma astrophysics. II. Singular layers and reconnection. Reviews of Modern Physics, 51(1):11–20, January 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.51.11;   http://rmp.aps.org/abstract/RMP/
v51/i1/p11_1.

Sagan:1983:NWC

[Sag83] Carl Sagan. Nuclear war and climatic catastrophe: Some policy implications. Foreign Affairs, 62(2):257–292, Winter 1983. CODEN FRNAA3. ISSN 0015-7120. URL http://www.jstor.

org/stable/20041818.

Sanders:1977:AOT

[San77] Ralph Sanders. The Advisors: Oppenheimer, Teller and the Superbomb by Herbert York (review). Technology and Culture, 18(1):129–131, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/891895/pdf.

Sanders:1979:UWC

[San79] Jane A. Sanders. The University of Washington and the controversy over J. Robert Oppenheimer. Pacific Northwest Quarterly, 70(1):8–19, January 1979. CODEN ???? ISSN 0030-8803.

Sapp:2005:EPR

[Sap05a] G. Sapp. 109 East Palace: Robert Oppenheimer and the secret city of Los Alamos. Library Journal, 130(10):168, June 2005. CODEN LIBJA7. ISSN 0363-0277.

Sapp:2005:APT

[Sap05b] G. Sapp. American prometheus: The triumph and tragedy of J. Robert Oppenheimer. Library Journal, 130(7):115, April 15, 2005. CODEN LIBJA7. ISSN 0363-0277.

Sapp:2005:RJR

[Sap05c] G. Sapp. The ruin of J. Robert Oppenheimer and the birth of the modern arms race. Library Journal, 130(12):117, July 2005. CODEN LIBJA7. ISSN 0363-0277.

Sargent:1982:ORL

[Sar82] B. W. Sargent. Robert Oppenheimer, letters and recollections. Queens Quarterly, 89(1):174–177, ???? 1982. ISSN 0033-6041.

Sartori:2001:BRS

[Sar01] Leo Sartori. Book review: S. S. Schweber, In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist. American Journal of Physics, 69(3):394–??, March 2001. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/ajpias/
v69/i3/p394/s2.

Sawyer:1954:MTS

[Saw54] Roland Sawyer. More than security. Bulletin of the Atomic Scientists, 10(7):284, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the author’s article in The Christian Science Monitor.

Shepley:1954:HBM

[SB54] James R. Shepley and Clay Blair, Jr. The hydrogen bomb: the men, the menace, the mechanism. Jarrolds, London, UK, 1954. 216 pp. LCCN UF767 .S4 1955. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56], and a longer paper [Tel55].

Shepley:1955:WKB

[SB55a] James R. Shepley and Clay Blair, Jr. Die Wasserstoffbombe: der Konflikt, die Bedrohung, die Konstruktion. (German) [The hydrogen bomb: the conflict, the threat construction]. Steingrüben Verl., Stuttgart, Germany, 1955. 288 pp. LCCN ???? Translation of [SB54] to German by Hans Dieter Müller. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56].

Shepley:1955:GSA

[SB55b] James R. Shepley and Clay Blair, Jr. La guerre secrète autour de la superbombe: la conception, la compétition, le risque mondial. (French) [The hydrogen bomb: conception, competition, and world risk]. Plon, Paris, France, 1955. x + 272 pp. LCCN ???? Translation of [SB54] to French by Anne Frauger. Preface by Yves Rocard. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56].

Shepley:1971:HBM

[SB71] James R. Shepley and Clay Blair, Jr. The hydrogen bomb: the men, the menace, the mechanism. Jarrolds, London, UK, 1971. ISBN 0-8371-5235-6. vii + 244 pp. LCCN UF767 .S4 1955. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56].

Simon:2021:ERD

[SBB21] Steven L. Simon, André Bouville, and Harold L. Beck. Estimated radiation doses and projected cancer risks for New Mexico residents from exposure to radioactive fallout from the Trinity nuclear test. Nuclear Technology, 207(S1):S380–S396, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Serber:1998:PWR

[SC98] R. (Robert) Serber and Robert P. Crease. Peace and war: reminiscences of a life on the frontiers of science. The George B. Pegram lecture series. Columbia University Press, New York, NY, USA, 1998. ISBN 0-231-10546-0 (hardcover). xxiii + 241 pp. LCCN QC16.S46 A3 1998. URL http://www.h-net.org/

review/hrev-a0b1d3-aa.

Scerri:2025:BOA

[Sce25] Eric Scerri. The Born–Oppenheimer Approximation and its role in the reduction of chemistry. Foundations of Chemistry, 27(2):183–197, August 2025. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://link.

springer.com/article/10.1007/s10698-025-09543-3. See

comment [Woo25].

Schlesinger:1954:OC

[Sch54] A. M. Schlesinger. The Oppenheimer case. The Atlantic, 194 (??):24–46, October 1954.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schlesinger:1962:PH

[Sch62] Arthur M. (Arthur Meier) Schlesinger, Jr. The politics of hope. Houghton-Mifflin, Boston, MA, USA, 1962. 298 pp. LCCN E183 .S3. Reprinted in [Sch08a].

Scholz:1982:KSJ

[Sch82] G. Scholz. Kipphardt ‘In der Sache J. Robert Oppenheimer’. (German) [Kipphardt In the Matter of J. Robert Oppenheimer]. Theater Heute, ??(5):58, ???? 1982. ISSN 0040-5507.

Schweber:1986:ETR

[Sch86] S. S. Schweber. The empiricist temper regnant: Theoretical physics in the United States 1920–1950. Historical Studies in the Physical and Biological Sciences, 17(1):55–98, ???? 1986. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.jstor.org/stable/27757575.

Schweber:2000:SBB

[Sch00] S. S. (Silvan S.) Schweber. In the shadow of the bomb: Bethe, Oppenheimer, and the moral responsibility of the scientist. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-04989-0 (hardcover). xviii + 260 + 8 pp. LCCN QC774.O56 S32 2000.

Schweber:2001:JRO

[Sch01] S. S. Schweber. J. Robert Oppenheimer: Proteus unbound. Physics World, 14(2):39–46, February 2001. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL

http://iopscience.iop.org/article/10.1088/2058-7058/
14/2/31.

Schweber:2003:BRB

[Sch03a] Silvan Schweber. Book review: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken. Physics Today, 56(5):59–60, May 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

org/resource/1/phtoad/v56/i5/p59_s1.

Schweber:2003:JRO

[Sch03b] Silvan S. Schweber. J. Robert Oppenheimer: Proteus unbound. Science in Context, 16(1–2):219–242, March 2003. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schwinger:2003:SPQ

[Sch03c] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schoke:2004:LER

[Sch04] James A. Schoke. Letter to the editor: Riding with Oppie. Bulletin of the Atomic Scientists, 60(4):4–5, July 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://www.highbeam.com/doc/1G1-119224463.html.

Schweber:2005:IOE

[Sch05] S. S. Schweber. Intersections: Oppenheimer and Einstein. In Carson and Hollinger [CH05], chapter 17, pages 343–360. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Schweber:2006:EOI

[Sch06] Silvan S. Schweber. Einstein and Oppenheimer: Interactions and intersections. Science in Context, 19(4):513–559, December 2006. CODEN SCCOEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Schlesinger:2008:PHB

[Sch08a] Arthur M. (Arthur Meier) Schlesinger, Jr. The Politics of Hope: and, The Bitter Heritage: American Liberalism in the 1960s. The James Madison library in American politics. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13475-8 (paperback). xxxv + 539 pp. LCCN E183 .S33 2008. With a new foreword by Sean Wilentz.

Schweber:2008:EOM

[Sch08b] S. S. (Silvan S.) Schweber. Einstein and Oppenheimer: the meaning of genius. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02828-7. xiv + 412 pp. LCCN QC16.E5

S3285 2008. URL http://www.loc.gov/catdir/toc/ecip082/
2007043108.html.

Schwarz:2010:JRO

[Sch10] Siegfried Schwarz. J. Robert Oppenheimer. the biography. Zeitschrift für Geschichtswissenschaft, 58(12):1028–1030, ???? 2010. ISSN 0044-2828.

Seaborg:1967:PSH

[Sea67] Glenn T. Seaborg. Public service and human contributions. Physics Today, 20(10):48–51, October 1967. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v20/i10/p48_s1.

Seely:1995:CAT

[See95] Bruce E. Seely. Critical Assembly: a Technical History of Los Alamos during the Oppenheimer Years, 1943–1945 by Lillian Hoddeson, et al. (review). Technology and Culture, 36(4): 1053–1055, October 1995. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/887781/pdf.

Seiler:1969:RBO

[Sei69] Rudolf Seiler. A remark on the Born–Oppenheimer approximation. International Journal of Quantum Chemistry, 3(1):25–32, January 1969. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seidel:1978:PRC

[Sei78] Robert Wayne Seidel. Physics Research in California: the rise of a leading sector in American physics. Ph.D. thesis, University of California, Berkeley, Berkeley, CA, USA, June 1978. vii

+ 596 pp. URL http://search.proquest.com/pqdtglobal/
docview/302900022.

Seidel:2001:BRBb

[Sei01] Robert W. Seidel. Book review: In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist by S. S. Schweber. Isis, 92(2):422–423, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/3080703.

Serber:1967:EY

[Ser67] Robert Serber. The early years. Physics Today, 20(10):35–39, October 1967. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v20/i10/p35_s1.

Settle:2005:BRB

[Set05] Frank A. Settle. Book review: Oppenheimer: Portrait of an Enigma. The Journal of Military History, 69(1):278– 279, January 2005. ISSN 0899-3718 (print), 1543-7795 (elec-

tronic). URL http://muse.jhu.edu/journals/journal_of_
military_history/v069/69.1settle.html.

Scherer:2008:JRO

[SF08] Glenn Scherer and Marty Fletcher. J. Robert Oppenheimer: the brain behind the bomb. MyReportLinks.com, Berkeley Heights, NJ, USA, 2008. ISBN ???? 128 (est.) pp. LCCN ????

Sood:2021:NCA

[SFAL21] Avneet Sood, R. Arthur Forster, B. J. Archer, and R. C. Little. Neutronics calculation advances at Los Alamos: Manhattan Project to Monte Carlo. Nuclear Technology, 207(S1):S100– S133, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Stoff:1991:MPD

[SFW91] Michael B. Stoff, Jonathan F. Fanton, and R. Hal (Richard Hal) Williams, editors. The Manhattan Project: a documentary introduction to the Atomic Age. Temple University Press, Philadelphia, PA, USA, 1991. ISBN 0-87722-787-X. xxi + 290 pp. LCCN QC773.A1 M36 1991.

Stern:1969:OCS

[SG69] Philip M. Stern and Harold P. Green. The Oppenheimer case: security on trial. Harper & Row, New York, NY, USA, 1969. xii + 591 pp. LCCN QC16.O62 S69. With a special commentary by Lloyd K. Garrison.

Stern:1971:OCS

[SG71] Philip M. Stern and Harold P. Green. The Oppenheimer case: security on trial. Hart-Davis, London, UK, 1971. ISBN 0-246-64035-9. xiv + 591 pp. LCCN QC16.O62 S69 1971. Foreword by Sir Solly Zuckerman.

Sherwin:1972:NBA

[She72] Martin J. Sherwin. Niels Bohr and the atomic bomb: the scientific ideal and international politics, 1943–1944. In Proceedings of the History of the 20th Century Physics conference, Varenna, Italy, 31 July–12 August, 1972, page ?? ????, ????, 1972.

Sherman:1982:OWT

[She82] Michael Sherman. Oppenheimer: What a trouble-maker!. book reviews: Reminiscences of Los Alamos, 1943–1945 by Lawrence Badash, Joseph O. Hirschfelder, and Herbert P. Broida. Disturbing the Universe by Freeman Dyson. J. Robert Oppenheimer, Shatterer of Worlds by Peter Goodchild. Robert Oppenheimer, Letters and Recollections by Alice Kimball Smith and Charles Weiner. The Public Historian, 4(4):97–117, Autumn 1982. ISSN 0272-3433 (print), 1533-8576 (electronic). URL http://www.

jstor.org/stable/3377050.

Sherwin:1985:RHW

[She85] Martin J. Sherwin. Retrospectives: How well they meant. Bulletin of the Atomic Scientists, 41(7):9–15, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1954:SS

[Shi54] Edward Shils. A slippery slope. Bulletin of the Atomic Scientists, 10(6):242, 256, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Selby:2021:NYA

[SHM^{+}21] Hugh D. Selby, Susan K. Hanson, Daniel Meininger, Warren J. Oldham, William S. Kinman, Jeffrey L. Miller, Sean D. Reilly, Allison M. Wende, Jennifer L. Berger, Jeremy Inglis, Anthony D. Pollington, Christopher R. Waidmann, Roger A. Meade, Kevin L. Buescher, James R. Gattiker, Scott A. Vander Wiel, and Peter W. Marcy. A new yield assessment for the Trinity nuclear test, 75 years later. Nuclear Technology, 207(S1):

321–325, November 2021. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Simpson:1954:CCR

[Sim54] Alan Simpson. Current comment: The re-trial of the Oppenheimer case. Bulletin of the Atomic Scientists, 10(10):387–388, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Streshinsky:2013:ALS

[SK13] Shirley Streshinsky and Patricia Klaus. An atomic love story: the extraordinary women in Robert Oppenheimer’s life. Turner Publishing Company, New York, NY, 2013. ISBN 1-61858-019-1 (hardcover). xxv + 369 pp. LCCN QC16.O62 S696 2013.

Sellier:2017:HAB

[SK17] Jean Michel Sellier and Kristina G. Kapanova. On the hydrogen atom beyond the Born–Oppenheimer approximation. International Journal of Quantum Chemistry, 117(21):??, November 5, 2017. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slaughter:2024:NIP

[Sla24a] Aimee Slaughter. Nuclear imaginaries and power in Oppenheimer. Technology and Culture, 65(1):319–332, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/920526.

Slaughter:2024:PMP

[Sla24b] Aimee Slaughter. Performing the Manhattan Project in Los Alamos. History of Science (UK), 62(2):305–325, June 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.
1177/00732753231187011.

Smith:1970:AFA

[Smi70] Alice Kimball Smith. Los Alamos: Focus of an age. Bulletin of the Atomic Scientists, 26(6):15–20, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sung:2010:JRO

[SMSU10] Carolyn H. Sung, David Mathisen, Michael Spangler, and Stephen Urgola. J. Robert Oppenheimer papers: A finding aid to the collection in the Library of Congress. Report, Manuscript Division, Library of Congress, Washington, DC, USA, April

2010. 75 pp. URL http://memory.loc.gov/service/mss/
eadxmlmss/eadpdfmss/1998/ms998007.pdf.

Shigeta:1999:DFT

[SNNY99] Y. Shigeta, H. Nagao, K. Nishikawa, and K. Yamaguchi. Density functional theory without the Born–Oppenheimer approximation. II. Green function techniques. International Journal of Quantum Chemistry, 75(4–5):875–883, November 15, 1999. CODEN IJQCB2. ISSN 0020-7608 (print),

1097-461X (electronic). URL http://www3.interscience.
wiley.com/cgi-bin/abstract/66004983/START;   http://
www3.interscience.wiley.com/cgi-bin/fulltext?ID=66004983&
PLACEBO=IE.pdf.

Serber:1992:APF

[SR92] R. (Robert) Serber and Richard Rhodes, editors. The Los Alamos primer: the first lectures on how to build an atomic bomb. University of California Press, Berkeley, CA, USA, 1992. ISBN 0-520-07576-5. xxxiii + 98 + 8 pp. LCCN QC773.A1 S47 1992. URL http://www.loc.gov/

catdir/bios/ucal051/91014068.html;   http://www.loc.
gov/catdir/description/ucal041/91014068.html. Anno-

tated by Robert Serber. Edited with an introduction by Richard Rhodes.

Sudoplatov:1995:STM

[SSSS95] Pavel Sudoplatov, Anatolii Pavlovich Sudoplatov, Jerrold L. Schecter, and Leona Schecter. Special tasks: the memoirs of an unwanted witness, a Soviet spymaster. Little, Brown and Co., Boston, MA, USA, updated edition, 1995. ISBN 0-316-82115-2. xxxi + 527 pp. LCCN JN6529.I6 S83 1995.

Steiner:1975:BAS

[Ste75] A. Steiner. Baptism of atomic scientists. Bulletin of the Atomic Scientists, 31(2):21–28, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stimson:1947:DUB

[Sti47] H. L. Stimson. The decision to use the bomb. Harper’s Magazine, 194(2):97–107, February 1947. CODEN HAMAA3.

ISSN 1045-7143. URL http://www.harpers.org/archive/

1947/02/0032863. Reprinted in [Sti76]. See also [Lef95] for a contrary view based on a half-century of historical scholarship, and access to previously-secret US documents.

Stimson:1976:DUB

[Sti76] H. L. Stimson. The decision to use the bomb. In Baker [Bak76], pages 14–28. ISBN 0-03-089873-0. LCCN ???? Reprinted from [Sti47]. See also [Lef95] for a contrary view based on a halfcentury of historical scholarship, and access to previously-secret US documents.

Strauss:1963:DCD

[Str63a] Lewis L. Strauss. Decision in the case of Dr. J. Robert Oppenheimer. In Men and decisions [Str63b], chapter 14, pages 267–295. LCCN E741 .S78.

Strauss:1963:MDa

[Str63b] Lewis L. Strauss. Men and decisions. Doubleday, Garden City, NY, USA, 1963. viii + 468 + 16 pp. LCCN E741 .S78.

Strout:1963:CSS

[Str63c] Cushing Strout, editor. Conscience, science and security: the case of Dr. J. Robert Oppenheimer. Berkeley series in American history. Rand McNally, Chicago, IL, USA, 1963. 58 pp. LCCN QC16.O62 S7.

Strout:1964:OCM

[Str64] Cushing Strout. The Oppenheimer case: Melodrama, tragedy, and irony. The Virginia Quarterly Review, 40(1):268–280, Winter 1964.

Strout:1965:CSS

[Str65] Cushing Strout, editor. Conscience, science and security: the case of Dr. J. Robert Oppenheimer. Berkeley series in American history. Rand McNally, Chicago, IL, USA, 1965. 58 pp. LCCN QC16.O62 S7.

Strout:1970:O

[Str70] Cushing Strout. Oppenheimer. New York Review of Books, ?? (??):??, September 24, 1970. ISSN 0028-7504 (print), 1944-7744

(electronic). URL http://www.nybooks.com/articles/1970/
09/24/oppenheimer/.

Strout:1983:TOS

[Str83] Cushing Strout. Telling the Oppenheimer story: From the AEC to the BBC. The Yale Review, 73(1):122–130, October 1983.

Strathern:1998:OBJ

[Str98] Paul Strathern. Oppenheimer et la bombe, je connais!. (French) [Oppenheimer and the bomb, I know!]. Mallard, Paris, France, 1998. ISBN 2-84372-013-3. 93 pp. LCCN ????

Strout:2006:APT

[Str06] Cushing Strout. American Prometheus: The triumph and tragedy of J. Robert Oppenheimer. Reviews in American History, 34(1):86–92, March 2006. ISSN 0048-7511 (print), 1080-6628 (electronic).

Shigeta:1998:DFT

[STY^{+}98] Y. Shigeta, H. Takahashi, S. Yamanaka, M. Mitani, H. Nagao, and K. Yamaguchi. Density functional theory without the Born– Oppenheimer approximation and its application. International Journal of Quantum Chemistry, 70(4–5):659–669, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=75044;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=75044&PLACEBO=IE.pdf.

Smith:1980:LOJ

[SW80a] A. K. Smith and C. Weiner. Letters of J. Robert Oppenheimer. The Sciences (New York), 20(3):6, ???? 1980. CODEN SCN-CAD. ISSN 0036-861X.

Smith:1980:POR

[SW80b] Alice Kimball Smith and Charles Weiner. Part one: Robert Oppenheimer: Letters and recollections. Bulletin of the Atomic Scientists, 36(5):19–27, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1980:PTR

[SW80c] Alice Kimball Smith and Charles Weiner. Part two: Robert Oppenheimer: the Los Alamos years. Bulletin of the Atomic Scientists, 36(6):11–17, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1980:ROL

[SW80d] Alice Kimball Smith and Charles Weiner, editors. Robert Oppenheimer, letters and recollections. Harvard University Press, Cambridge, MA, USA, 1980. ISBN 0-674-77605-4. xvi + 376 + 5 pp. LCCN QC16.O62 A4 1980.

Smith:1980:YOL

[SW80e] Alice Kimball Smith and Charles Weiner. The young Oppenheimer: letters and recollections. Physics Today, 33(4):24– 33, April 1980. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v33/i4/p24_s1. Reprinted in [SW85].

Smith:1985:YOL

[SW85] Alice Kimball Smith and Charles Weiner. The young Oppenheimer: Letters and recollections. In Weart and Phillips [WP85], pages 221–227. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [SW80e].

Smith:1995:ROL

[SW95] Alice Kimball Smith and Charles Weiner. Robert Oppenheimer: Letters and Recollections. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1995. ISBN 0-8047-2620-5. xxii + 376 pp. LCCN QC16.O62 A4 1995. URL http://

www.loc.gov/catdir/description/cam027/95067508.html.

Serber:1967:MO

[SWPS67a] Robert Serber, Victor F. Weisskopf, Abraham Pais, and Glenn T. Seaborg. A memorial to Oppenheimer. Physics Today, 20(10):34, October 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/PHTOAD/v20/i10.

Serber:1967:OB

[SWPS67b] Robert Serber, Victor F. Weisskopf, Abraham Pais, and Glenn T. Seaborg. Oppenheimer bibliography. Physics Today, 20(10):52–53, October 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v20/i10/p52_s1.

Sylves:1987:NOP

[Syl87] Richard Terry Sylves. The nuclear oracles: a political history of the General Advisory Committee of the Atomic Energy Commission, 1947–1977. Iowa State University Press, Ames, IA, USA, 1987. ISBN 0-8138-0062-5. xviii + 319 + 16 pp. LCCN QC773.3.U5 S95 1987. With an introduction by Anthony A. Tomei.

Szasz:1995:GBS

[Sza95] Ferenc Morton Szasz. Great Britain and the saga of J. Robert Oppenheimer. War in History, 2(3):320–333, ???? 1995. CO-DEN ???? ISSN 0968-3445 (print), 1477-0385 (electronic).

Strauss:1954:AMD

[SZC54] Lewis L. Strauss, Eugene M. Zuckert, and Joseph Campbell. AEC majority decision: 29 June 1954. Bulletin of the Atomic Scientists, 10(7):275–277, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1971:CLC

[Szi71] Gertrud Weiss Szilard. Communications: Lunar craters. Bulletin of the Atomic Scientists, 27(4):4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Takatsuka:2009:TNB

[Tak09] Kazuo Takatsuka. Toward non-Born–Oppenheimer quantum chemistry. International Journal of Quantum Chemistry, 109 (10):2131–2142, ???? 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tapia:2001:TRB

[Tap01] O. Tapia. Towards a rigged Born–Oppenheimer electronic theory of chemical processes. Advances in Quantum Chemistry, 40: 103–131, 2001. CODEN AQCHA9. ISSN 0065-3276. URL http:

//www.sciencedirect.com/science/article/pii/S0065327601400116.

Taylor:1992:PNT

[Tay92] Bryan C. Taylor. The politics of the nuclear text: Reading Robert Oppenheimer’s letters and recollections. Quarterly Journal of Speech, 78(4):429–449, November 1992. CODEN ???? ISSN 0033-5630 (print), 1479-5779 (electronic).

Teller:1962:LH

[TB62] Edward Teller and Allen Brown. The legacy of Hiroshima. Doubleday, Garden City, NY, 1962. 325 pp. LCCN UF767 .T4 (LC); UF767 .T4.

Teller:1982:HBH

[TB82] Edward Teller and Hans A. Bethe. Hydrogen bomb history. Science, 218(4579):1270, December 24, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1690175. Response to [Bro82].

Teller:1955:WMP

[Tel55] Edward Teller. The work of many people. Science, 121(3139): 267–275, February 25, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1681529.

Teller:1983:SHR

[Tel83] Edward Teller. Seven hours of reminiscences. Los Alamos Science, 7:190–195, Winter/Spring 1983. CODEN LASCDI.

ISSN 0273-7116. URL http://library.lanl.gov/cgi-bin/
getfile?07-22.pdf.

Teller:1987:BST

[Tel87] Edward Teller. Better a shield than a sword: perspectives on defense and technology. Free Press, New York, NY, USA, 1987. ISBN 0-02-932461-0. xiv + 257 pp. LCCN UA23 .T39 1987.

URL http://alsos.wlu.edu/information.aspx?id=2449.

Teller:1998:IET

[Tel98] Edward Teller. Interview with Edward Teller. Web transcript., November 15, 1998. URL http://alsos.wlu.

edu/information.aspx?id=3425;   http://www.gwu.edu/~
nsarchiv/coldwar/interviews/episode-8/teller1.html.

Thorpe:2001:JRO

[Tho01] Charles Robert Thorpe. J. Robert Oppenheimer and the transformation of the scientific vocation. Ph.D. dissertation, University of California, San Diego, San Diego, CA, USA, 2001.

xiii + 700 pp. URL http://search.proquest.com/docview/

252126285. Listed in Dissertation Abstracts International. A, 62, (2001), 2254–2254.

Thorpe:2002:DES

[Tho02] Charles Thorpe. Disciplining experts: Scientific authority and liberal democracy in the Oppenheimer case. Social Studies of Science, 32(4):525–562, August 1, 2002. CO-DEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (elec-

tronic). URL https://journals.sagepub.com/doi/pdf/10.
1177/0306312702032004002.

Thorpe:2003:BRR

[Tho03] Charles Thorpe. Book review: Richard Polenberg (ed.), In the Matter of J. Robert Oppenheimer: The Security Clearance Hearing. Ithaca and London: Cornell University Press, 2002. Pp. xxxii + 409. ISBN 0-8014-8661-0. $13.50 (paperback). British Journal for the History of Science, 36(3):378–379, September 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4028170.

Thorpe:2004:ATS

[Tho04] Charles Thorpe. Against time: Scheduling, momentum, and moral order at wartime Los Alamos. Journal of Historical Sociology, 17(1):31–55, ???? 2004.

Thorpe:2005:SMS

[Tho05] Charles Thorpe. The scientist in mass society: J. Robert Oppenheimer and the postwar liberal imagination. In Carson and Hollinger [CH05], chapter 14, pages 293–314. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Thorpe:2006:OTI

[Tho06] Charles Thorpe. Oppenheimer: the tragic intellect. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-79845-3 (hardcover), 0-226-79846-1 (paperback). xx + 413 + 12 pp. LCCN QC16.O62 T56 2006. URL http://

www.loc.gov/catdir/enhancements/fy0666/2006015223-b.
html;   http://www.loc.gov/catdir/enhancements/fy0666/
2006015223-d.html;   http://www.loc.gov/catdir/enhancements/
fy0666/2006015223-t.html.

Thorpe:2011:BRS

[Tho11] Charles Thorpe. Book review: Silvan S. Schweber, Einstein and Oppenheimer, ISBN-13 978-0-674-02828-9. Annals of Science, 68 (4):558–561, 2011. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Thorpe:2016:RHV

[Tho16] Charles Thorpe. Reviews: The Hope and Vision of J. Robert Oppenheimer — by Michael A. Day. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 58(4):315–316, November 2016. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Thomas:2017:PPM

[Tho17] Linda Carrick Thomas. Polonium in the Playhouse: the Manhattan Project’s secret chemistry work in Dayton, Ohio. Trillium, an imprint of The Ohio State University Press, Columbus, OH, USA, 2017. ISBN 0-8142-1338-3 (hardcover). xxi + 247 pp. LCCN QC773.3.U5 T46 2017.

Tomonaga:1948:IFR

[TO48] Sin-Itiro Tomonaga and J. R. Oppenheimer. On infinite field reactions in quantum field theory. Physical Review (2), 74(2): 224–225, July 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v74/i2/p224_1.

Trilling:1954:OCR

[Tri54] Diana Trilling. The Oppenheimer case: a reading of the testimony. Partisan Review, 21(6):604–635, ???? 1954. CODEN ???? ISSN 0031-2525.

Thorpe:2000:WWJ

[TS00] Charles Thorpe and Steven Shapin. Who was J. Robert Oppenheimer? Charisma and complex organization: Charisma and complex organization. Social Studies of Science, 30 (4):545–590, August 1, 2000. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.

jstor.org/stable/285781;   https://journals.sagepub.
com/doi/pdf/10.1177/030631200030004003.

Teller:2001:MTC

[TS01] Edward Teller and Judith L. Shoolery. Memoirs: a Twentieth-Century Journey in Science and Politics. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0532-X, 1-903985-12-9. xii + 628 pp. LCCN QC16.T37 M55 2001. US$24.99.

URL http://alsos.wlu.edu/information.aspx?id=784;
http://www.gbv.de/dms/bowker/toc/9781903985120;   http:
//www.loc.gov/catdir/enhancements/fy0833/2001097880-
d.html;   http://www.loc.gov/catdir/toc/fy02/2001097880.
html.

Turney:2012:BRB

[TS12] Jon Turney and Karen Shook. Book review: Inside the Centre: The Life of J. Robert Oppenheimer, by Ray Monk. The Times Higher Education Supplement: THE, ??(2079):48–??, De-

cember 6, 2012. URL http://search.proquest.com/docview/
1283226145.

Udall:1994:MAP

[Uda94] Stewart L. Udall. The myths of August: a personal exploration of our tragic Cold War affair with the atom. Pantheon Books, New York, NY, USA, 1994. ISBN 0-679-43364-3. xii + 399 pp. LCCN E840 .U33 1994. US$25.00; CAN$33.50.

Udall:1998:MAP

[Uda98] Stewart L. Udall. The myths of August: a personal exploration of our tragic Cold War affair with the atom. Rutgers University Press, New Brunswick, NJ, 1998. ISBN 0-8135-2546-2. xii + 399 pp. LCCN E840 .U33 1998.

Ulam:1968:JN

[UKTS68] S. Ulam, H. W. Kuhn, A. W. Tucker, and Claude E. Shannon. John von Neumann, 1903–1957. Perspectives in American History, 2(??):235–269, ???? 1968. CODEN ???? ISSN 0079-0990.

Underwood:2013:BRR

[Und13] Martin Underwood. Book reviews: Ray Monk, Inside the Centre: The Life of J. Robert Oppenheimer. London: Jonathan Cape. 2012. Pp. xiv + 818. ISBN 978-0-224-06262-6. £30.00 (hardback). British Journal for the History of Science, 46(1): 172–173, March 2013. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

UIP:1954:LSJ

[Uni54] University of Illinois Physicists. [Letter of support for J. Robert Oppenheimer]. Bulletin of the Atomic Scientists, 10(5):191, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter signed by 27 individuals.

USAEC:1954:MJR

[U.S54] U.S.Atomic Energy Commission. In the matter of J. Robert Oppenheimer. United States Government Printing Office, Washington, DC, USA, 1954. 992 pp. LCCN QC16.O62 U54 1954.

Transcript of hearing before Personnel Security Board, Washington, DC, April 12, 1954 and through May 6, 1954.

USAEC:1971:MJR

[U.S71] U.S.Atomic Energy Commission. In the matter of J. Robert Oppenheimer: transcript of hearing before Personnel Security Board and texts of principal documents and letters. MIT Press, Cambridge, MA, USA, 1971. ISBN 0-262-71002-1 (paperback). xii + 1084 pp. LCCN QC16.O62 U54 1971. MIT 161. Reprint of [U.S54].

Usdin:2009:RRR

[Usd09] Steven T. Usdin. The Rosenberg ring revealed: Industrialscale conventional and nuclear espionage. Journal of Cold War Studies, 11(3):91–143, Summer 2009. ISSN 1520-3972 (print),

1531-3298 (electronic). URL http://muse.jhu.edu/article/
268392.

Valdevit:2006:OFS

[Val06a] Giampaolo Valdevit. Oppenheimer fra scienza e potere. Una storia americana. (Italian) [Oppenheimer between science and power. An American story]. Studi Storici, 47(1):115–142, January/March 2006. ISSN 0039-3037 (print), 2036-458X (elec-

tronic). URL http://www.jstor.org/stable/20567340.

Valiunas:2006:AAG

[Val06b] Algis Valiunas. The agony of atomic genius. The New Atlantis: a Journal of Technology and Society, 14(??):85–104, Fall 2006. CODEN ???? ISSN 1543-1215 (print), 1555-

5569 (electronic). URL http://www.thenewatlantis.com/
publications/the-agony-of-atomic-genius.

VanDeMark:2003:PKN

[Van03] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2

V36 2003. URL http://www.loc.gov/catdir/toc/fy045/
2002043646.html.

Varlamov:2002:AAF

[Var02] V. V. Varlamov. About algebraic foundations of Majorana– Oppenheimer quantum electrodynamics and de Broglie–Jordan neutrino theory of light. Annales de la Fondation Louis de Broglie, 27(2):273–286, 2002. ISSN 0182-4295.

Veblen:1960:OVP

[Veb60] Oswald Veblen. Oswald Veblen papers, 1881–1960 (bulk 1920– 1960). US Library of Congress archival manuscript material (col-

lection)., 1960. URL http://hdl.loc.gov/loc.mss/eadmss.
ms011049;   http://hdl.loc.gov/loc.mss/eadmss.ms011049.
3.

vonHippel:2001:WFS

[vH01] Frank N. von Hippel. Where Fermi stood. Bulletin of the Atomic Scientists, 57(5):26–29, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/
57/5/26.full;   http://bos.sagepub.com/content/57/5/
26.full.pdf+html;   http://www.highbeam.com/doc/1G1-
78334537.html.

Vilar:1965:DO

[VK65] Jean Vilar and Heinar Kipphardt. Le dossier Oppenheimer. Éditions Gonthier, Genève, Switzerland, 1965. 125 pp. LCCN QC16.O62 V5.

vonNeumann:1996:PJN

[vN96] John von Neumann. Papers of John von Neumann, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. 11,660 items. 34 containers plus 1 vault container. 13.4 linear feet. Manuscript number MSS44180. Correspondence, memoranda, journals, speeches, article and book drafts, notes, charts, graphs, patent, biographical material, family papers, printed materials, newspaper clippings, photographs, and other materials pertaining primarily to von Neumann’s career as professor of mathematics at the Institute for Advanced Study including his directorship of the Electronic Computer Project; adviser and commissioner on the U.S. Atomic Energy Commission; scientific consultant to government and private concerns, including the Los Alamos Scientific Laboratory, Los Alamos, New Mexico, and the U.S. Army Ballistic Research Laboratory, Aberdeen, Maryland; and author of works on ballistic research, computers, continuous geometries, logic, operator theory, quantum mechanics, and the theory of games. Includes evaluations of his work written after his death by colleagues including Herman Heine Goldstine, Paul R. Halmos, and Abraham H. Taub. Of special interest are an Albert Einstein letter and report on theoretical physics (1937). Also includes a small amount of material pertaining to Eva and Peter Aldor. Correspondents include Eva Aldor, Frank Aydelotte, Hans Albrecht Bethe, Garrett Birkhoff, S. Chandrasekhar, George Bernard Dantzig, P. A. M. Dirac, Carl Eckart, Enrico Fermi, Abraham Flexner, George Gamow, Kurt Gödel, Herman Heine Goldstine, Werner Heisenberg, L. van Hove, Cuthbert Corwin Hurd, Pascual Jordan, R. H. Kent, George B. Kistiakowsky, Oskar Morgenstern, J. Robert Oppenheimer, Rudolf Ortvay, Wolfgang Pauli, Marshall H. Stone, Lewis L. Strauss, Abraham Haskel Taub, Edward Teller, Stanislaw M. Ulam, Oswald Veblen, Klara Dan Von Neumann, Warren Weaver, Hermann Weyl, Norbert Wiener, and Eugene Paul Wigner. Gift, Marina Von Neumann Whitman, 1974–1975. Gift, Nicholas A. Vonneuman, 1993.

vonNeumann:1996:JNP

[von96] John von Neumann. John von Neumann papers, 1912–1996 (bulk 1935–1957). US Library of Congress archival manuscript material (collection)., 1996. URL http://hdl.loc.gov/loc.

mss/eadmss.ms996003;   http://hdl.loc.gov/loc.mss/eadmss.
ms996003.3.

Wainstock:1996:DDA

[Wai96] Dennis D. Wainstock. The decision to drop the atomic bomb. Praeger, Westport, CT, USA, 1996. ISBN 0-275-95475-7. 180 pp. LCCN ????

Walker:1996:MRB

[Wal96] J. Samuel Walker. Movie review: J. Robert Oppenheimer: Father of the Atomic Bomb. Produced by Richard O’Regan. Claymont Productions in association with A&E Network, 1995. 50 mins. (A&E Biography, 235 East 45th St., New York, NY 10017). The Journal of American History, 83 (3):1129, 1996. ISSN 0021-8723 (print), 1945-2314 (elec-

tronic). URL http://jah.oxfordjournals.org/content/83/
3/1129.1.full.pdf+html;   http://jah.oxfordjournals.
org/content/83/3/1129.1.short.

Walsh:2013:JRO

[Wal13a] Lynda Walsh. J. Robert Oppenheimer: cultic prophet. In Scientists as prophets: a rhetorical genealogy [Wal13b], pages 100–118. ISBN 0-19-985709-1 (hardcover), 0-19-985711-3 (paperback). LCCN Q175.5 .W35 2013.

Walsh:2013:SPR

[Wal13b] Lynda Walsh. Scientists as prophets: a rhetorical genealogy. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-985709-1 (hardcover), 0-19-985711-3 (paperback). x + 264 + 1 pp. LCCN Q175.5 .W35 2013.

Wang:2004:BRR

[Wan04] Zuoyue Wang. Book review: Richard Polenberg (Editor): In the Matter of J. Robert Oppenheimer: The Security Clearance Hearing. Isis, 95(4):750–751, December 2004. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/432357.

Wang:2006:BRB

[Wan06] Jessica Wang. Book review: Blackett: Physics, War, and Politics in the Twentieth Century J, by Mary Jo Nye; David C. Cassidy. Robert Oppenheimer and the American Century. Isis, 97(1):144–147, March 2006. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/501108.

Wang:2008:BRBa

[Wan08] Zuoyue Wang. Book review: Oppenheimer: The Tragic Intellect. Isis, 99(1):228–229, March 2008. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/589401.

Wang:2011:JRO

[Wan11] Zuoyue Wang. J. Robert Oppenheimer, the Cold War, and the Atomic West. Pacific Historical Review, 80(1):161–162, February 2011. ISSN 0030-8684 (print), 1533-8584 (electronic).

Williams:1984:AAD

[WC84a] Robert Chadwell Williams and Philip L. (Philip Louis) Cantelon, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present, 1939–1984. University of Pennsylvania Press, Philadelphia, 1984. ISBN 0-585-17297-8 (e-book), 0-8122-7920-4, 0-8122-1169-3 (paperback). xv + 333 pp. LCCN U264 .A44 1984.

Williams:1984:OC

[WC84b] Robert Chadwell Williams and Philip L. (Philip Louis) Cantelon. The Oppenheimer case. In The American atom: a documentary history of nuclear policies from the discovery of fission to the present, 1939–1984 [WC84a], pages 141–175. ISBN 0-585-17297-8 (e-book), 0-8122-7920-4, 0-8122-1169-3 (paperback). LCCN U264 .A44 1984.

Weart:1982:FRB

[Wea82] Spencer R. Weart. Film review: The Day After Trinity: J. Robert Oppenheimer and the Atomic Bomb, produced and directed by John Else. Bulletin of the Atomic Scientists, 38(7): 41–42, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wedgwood:1967:DJR

[Wed67] C. V. Wedgwood. Dr. J. Robert Oppenheimer. The Times [London, UK], ??(??):14–??, February 23, 1967. ISSN 0140-0460, 0956-1382.

Weisskopf:1967:AY

[Wei67] Victor F. Weisskopf. The Los Alamos years. Physics Today, 20(10):39–42, October 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v20/i10/p39_s1.

Weinberg:1981:USM

[Wei81] Steven Weinberg. The ultimate structure of matter. In Anonymous [Ano81], pages 66–88. LCCN ???? URL http://

escholarship.org/uc/item/9772r6gq. Report number LBL-

13613.

Weiss:1986:BRJ

[Wei86] B. Weiss. Book review: J. Robert Oppenheimer — German — H Kant. Annals of Science, 43(6):570–571, ???? 1986. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Weiner:1988:OHS

[Wei88] Charles Weiner. Oral history of science: A mushrooming cloud? The Journal of American History, 75(2):548–559, September 1988. ISSN 0021-8723 (print), 1945-2314 (electronic). URL

http://www.jstor.org/stable/1887871.

Weisskopf:1991:JIP

[Wei91] Victor Frederick Weisskopf. The Joy of Insight: Passions of a Physicist. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1991. ISBN 0-465-03678-3, 0-465-03677-5. xiv + 336 + 8 pp. LCCN QC16.W516 A3 1991.

Weiner:1994:LFD

[Wei94a] Nella Fermi Weiner. Letter: Fermi: Didn’t even tell his wife. Bulletin of the Atomic Scientists, 50(4):3, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisman:1994:ERW

[Wei94b] Jonathan Weisman. Early retirement for weaponeers? Bulletin of the Atomic Scientists, 50(4):16–22, July 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://www.highbeam.com/doc/1G1-15587904.html.

Werner:2005:APT

[Wer05] A. Werner. American Prometheus: The triumph and tragedy of J. Robert Oppenheimer. American Journal of Psychiatry, 162 (12):2402, December 2005. CODEN AJPSAO. ISSN 0002-953X (print), 1535-7228 (electronic).

Westfall:2014:GIG

[Wes14] Catherine Westfall. Gaining inspiration from Galileo, Einstein and Oppenheimer. History of Physics Newsletter, 12(5):5, 7, 9, Fall 2014. CODEN ???? ISSN ???? URL http://www.aps.

org/units/fhp/newsletters/fall2014/index.cfm.

Wellerstein:2017:SSH

[WG17] Alex Wellerstein and Edward Geist. The secret of the Soviet hydrogen bomb. Physics Today, 70(4):40–47, April 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wharton:1955:NSC

[Wha55] Michael Wharton. A nation’s security: the case of Dr J. Robert Oppenheimer: edited from the official transcript of evidence given before the Personnel Security Board of the United States Atomic Energy Commission. Secker and Warburg, London, UK, 1955. 398 pp. LCCN ????

White:1958:WDW

[Whi58] Theodore White. “Where do we fit in?” the scientists ask. The New York Times Magazine, ??(??):??, May 18, 1958. ISSN 0362-1308.

Wigner:1976:BRB

[Wig76] Eugene P. Wigner. Book review: The Advisors: Oppenheimer, Teller, and the Superbomb, by Herbert York. American Scientist, 64(5):561, September 1976. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.

jstor.org/stable/27847469.

Wilson:1969:BBO

[Wil69] Jane Wilson. Books: Lawrence and Oppenheimer, by Nuell Pharr Davis. Bulletin of the Atomic Scientists, 25(1):31–32, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav68, Dav86].

Wilson:1970:GWH

[Wil70] Thomas Williams Wilson, Jr. The Great Weapons Heresy: The Struggle Behind our Present Nuclear Dilemma as Reflected in the Tragedy of J. Robert Oppenheimer. Houghton-Mifflin, Boston, MA, USA, 1970. xxii + 275 pp. LCCN QC16.O62 W5.

Wilson:1975:AOT

[Wil75a] Jane Wilson, editor. All in our time: the reminiscences of twelve nuclear pioneers. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1975. 236 pp. LCCN QC773.A1 A44. Reprinted by the Educational Foundations for Nuclear Science, Chicago, IL, USA (1975).

Wilson:1975:AOTa

[Wil75b] Robert R. Wilson. “All in our time”: a recruit for Los Alamos. Bulletin of the Atomic Scientists, 31(3):41–47, March 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1983:RSG

[Wil83] David Wilson. Rutherford, simple genius. Hodder and Stoughton, London, UK, 1983. ISBN 0-340-23805-4. 638 + 8 pp. LCCN QC16.R8 W54 1983. US$14.95. URL http://catalog.

hathitrust.org/api/volumes/oclc/10560773.html.

Wildavsky:1994:OCS

[Wil94] Ben Wildavsky. Oppenheimer’s colleagues say he didn’t pass secrets. San Francisco Chronicle, ??(??):A3–??, April 19,

1994. ISSN 1932-8672. URL http://search.proquest.com/
docview/411086158.

Winter:2000:ACO

[Win00] Dale Winter. Analysis of the cosmological Oppenheimer–Volkoff equations. Journal of Mathematical Physics, 41(8):5582–5597, August 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wittner:1999:BRB

[Wit99] Lawrence S. Wittner. Book reviews: Peace & War: Reminiscences of a Life on the Frontiers of Science. By Robert Serber and Robert P. Crease. New York: Columbia University Press and The Myths of August: A Personal Exploration of Our Tragic Cold War Affair with the Atom. By Stewart L. Udall. New Brunswick: Rutgers University Press. The Journal of American History, 86(1):298–299, June 1999. ISSN 0021-8723 (print), 1945-2314 (electronic). See [SC98, Uda94, Uda98].

Wolfenstein:1970:BRB

[Wol70] Lincoln Wolfenstein. Book review: Oppenheimer by Abraham Pais; Glenn T. Seaborg; Robert Serber; Victor F. Weisskopf; I. I. Rabi; In the Matter of J. Robert Oppenheimer by Heinar Kipphardt; John Roberts. Isis, 61(2):288–289, Summer 1970. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230002.

Wolverton:2002:OUS

[Wol02] M. Wolverton. Oppenheimer under suspicion. American History, 37(3):36–42, August 2002. ISSN 1076-8866. J. Robert Oppenheimer became the subject of an investigation that question his loyalty to the United States during the Cold War.

Wolverton:2003:SUS

[Wol03] M. Wolverton. Still under suspicion. American History, 37(6):4, February 2003. ISSN 1076-8866. Reply to Jonathan M. Menn’s comments concerning J. Robert Oppenheimer.

Wolff:2005:BFB

[Wol05] Carlo Wolff. Books: Father of the bomb two books explore tortured career of J. Robert Oppenheimer. American Prometheus: The Triumph and Tragedy of J. Robert Oppenheimer. By Kai Bird and Martin J. Sherwin. Alfred A. Knopf. $35. 719 pages. 109 East Palace: Robert Oppenheimer and the Secret City of Los Alamos. By Jennet Conant. Simon & Schuster. $26.95. 424 pages. The Atlanta Journal–Constitution, ??(??):L4–??, June 5,

2005. ISSN 1539-7459. URL http://search.proquest.com/
docview/337235829.

Wolverton:2008:LTF

[Wol08] Mark Wolverton. A life in twilight: the final years of J. Robert Oppenheimer. St. Martin’s Press, New York, NY, USA, 2008. ISBN 0-312-37440-2 (hardcover). x + 339 pp. LCCN QC16.O62 W65 2008. URL http://www.loc.gov/

catdir/enhancements/fy0838/2008025766-b.htm;   http:
//www.loc.gov/catdir/enhancements/fy0838/2008025766-
d.htm;   http://www.loc.gov/catdir/toc/fy1002/2008025766.
htm.

Wolk:2009:MHB

[Wol09] Herman S. Wolk. Making the H-bomb. Air Force magazine, 95(3):66–69, March 2009. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.airforce-

magazine.com/MagazineArchive/Pages/2009/March%202009/
0309H-Bomb.aspx.

Woolley:2025:CBO

[Woo25] R. Guy Woolley. Comment on The Born–Oppenheimer approximation and its role in the reduction of chemistry. Foundations of Chemistry, 27(3):373–380, December 2025. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (electronic). URL https://

link.springer.com/article/10.1007/s10698-025-09552-2.

See [Sce25].

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Weiner:1979:JRO

[WS79] Charles Weiner and Alice Kimball Smith. J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 35(1):57, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Request for additional Oppenheimer letters.

Wilson:1988:SMD

[WS88] Jane Wilson and Charlotte Serber, editors. Standing by and making do: women of wartime Los Alamos. Los Alamos Historical Society, Los Alamos, NM, USA, 1988. ISBN 0-941232-08-5 (paperback). xi + 130 pp. LCCN QC773.A1 S82 1988. US$8.95.

Wilson:2008:SMD

[WS08] Jane Wilson and Charlotte Serber, editors. Standing by and making do: women of wartime Los Alamos. Los Alamos Historical Society, Los Alamos, NM, USA, second edition, 2008. ISBN 0-941232-08-5 (paperback). 152 pp. LCCN QC773.A1

S82 2008. URL http://www.loc.gov/catdir/toc/ecip0822/
2008029781.html.

Wyden:1984:DOB

[Wyd84] Peter Wyden. Day One: Before Hiroshima and After. Simon and Schuster, New York, NY, USA, 1984. ISBN 0-671-46142-7 (hardcover). xxxii + 412 pp. LCCN D767.25.H6 W93 1984.

York:1975:SAT

[Yor75] Herbert F. York. Sounders of the alarm: Thirty years of attempts to bring the world to grips with the perpetual menace to human security. Bulletin of the Atomic Scientists, 31(10):43– 45, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1976:AOT

[Yor76] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. W. H. Freeman, New York, NY, USA, 1976. ISBN 0-7167-0718-7. x + 175 pp. LCCN UG1282.A8 Y67.

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67 1989. URL http://www.loc.

gov/catdir/description/cam023/88062671.html;   http:
//www.loc.gov/catdir/toc/cam027/88062671.html.

York:1995:AP

[Yor95] Herbert F. (Herbert Frank) York. Arms and the physicist, volume 12 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-099-0. xiii + 294 pp. LCCN QC16.Y67 A3 1995.

Young:2005:SRJ

[You05] Stephanie Young. Something resembling justice: John Francis Neylan and the AEC personnel security hearings at Berkeley, 1948–49. In Carson and Hollinger [CH05], chapter 10, pages 217–253. ISBN 0-9672617-3-2. LCCN QC16.O62 R47 2005.

Yoxen:1982:BRN

[Yox82] Edward Yoxen. Book review: The New Physics: Television review of ‘Oppenheimer’, 7 part serial, shown on BBC-2 in the UK in 1980 and of Peter Goodchild, J. Robert Oppenheimer ‘Shatterer of Worlds’. London: BBC Publications, 1980. Pp 301 incl. index and illustrations. £9.95. British Journal for the History of Science, 15(2):204–207, July 1982. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.

jstor.org/stable/4025976.

Zacharias:1976:BPS

[Zac76] Jerrold R. Zacharias. Books: Pragmatism, secrecy and moral values: The Advisors/Oppenheimer, Teller, and the superbomb, by Herbert F. York. Bulletin of the Atomic Scientists, 32(10):57– 59, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor76, Yor89].

Ziegler:1991:BRS

[Zie91] Charles Ziegler. Book review: Stanley A. Blumberg and Louis G. Panos: Edward Teller: Giant of the Golden Age of Physics. Herbert F. York: The Advisors: Oppenheimer, Teller, and the Superbomb. J. Robert Oppenheimer: Atom and Void: Essays on Science and Community. Isis, 82(3):589–590, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233287.